- 1

A Bibliography of ACM SIGMOD Record

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

04 November 2023 Version 1.79

Title word cross-reference

$20M [2219]. 3 [1652, 2532, 698]. 6 [495]. 8 [2185]. ^{∗} [880, 924]. ^{+} [2008, 1095, 687, 1190, 1748, 1258]. ^{3} [932]. nd [2004]. ^{th} [1931, 2047, 1930, 2046, 2137, 2214]. _{2} [755, 735, 1199, 744]. ϵ [2659, 2660]. G+ [605]. k [2578, 2579, 2675, 1495]. n [1110]. ¬1 [384]. O_{2} [486, 900]. R [598, 1570, 256]. ρ [1971].

* [598]. *-tree [598].

-dimensional [1110]. -Nearest [1495]. -Shape [2578, 2579]. -tree [1570, 1190]. -trees [2008, 1095, 1748, 256, 687, 1258]. -valued [698].

.KBMS [924].

/GraphLog [605]. /information [160].

0 [2123]. 0-8020-8860-0 [2123]. ’00 [1764, 1791]. 007 [848]. 03 [2049].

1 [37, 2088]. 1-55860-570-3 [2088]. 1.0 [2192]. 10th [2203, 2354, 2458]. 13th [2394, 1609]. ’17 [2626]. 17th [2751]. 18 [1930]. 19 [2724, 2137]. 1970 [14]. 1971 [20]. 1972 [38, 34]. 1975 [66]. 1980 [204]. 1988 [553]. 1995 [1215]. 1996 [1362]. 1998 [1546]. 1999 [2044, 1617]. 1st [2153].

2 [60]. 2-variable [1631]. 2.0 [2288, 2810]. 2000 [1861]. 2001 [1945]. 2002 [2027, 1977, 2043, 2045]. 2003 [2067, 2108]. 2005 [2190, 2213]. 2006 [2263]. 2007 [2288]. 2008 [2363, 2341, 2286]. 2009 [2364, 2362, 2371]. 2010 [2379, 2394, 2387]. 2011 [2411, 2447, 2449, 2436]. 2012 [2493, 2458, 2460, 2445]. 2014 [2530]. 2016 [2643]. 2020 [2724, 2718]. 2021 [2766]. 21st [1238]. ’22 [2813]. 2nd [2027, 2215].

3 [1380, 1527, 608, 2088]. 3.0 [2480]. 30540 [2122]. 30540-44008-9 [2122]. 3D [1343, 1344]. 3Dynamic [1605]. 3rd [2139, 2107].

400 [951]. 4th [1455].

5 [1931]. 5th [1569, 2372].

6.0 [770]. 6th [2485].

7 [1937, 2214]. 7.0 [1553]. 72 [42]. 7th [2216].

’84 [2860, 2861]. 8th [2555, 2426, 2255].

9 [2122]. ’93 [2868, 1036, 1064]. ’94 [1099, 1071]. ’96 [1444]. ’97 [1457, 1466]. ’98 [1481, 1609]. ’99 [1676, 1669]. 9th [2106, 2323].

Aachen [1100]. Aalborg [2455]. Abadi [2453]. Aberdeen [196]. Abiteboul [2044, 2284]. Abridged [2593]. Abstract [777, 85, 80, 72, 371, 182, 89, 194, 321, 1278, 236, 1538, 1149, 776, 69, 1420, 1537, 77, 320, 884]. abstract-object [1278]. abstracting [2415]. Abstraction [141, 329, 188, 157, 2183, 75, 710, 170, 174, 143, 187, 79, 191, 192]. abstractions [74, 166, 2265]. abstracts [41, 42]. academia [1932, 1926, 2267, 2275]. Academic [1616, 2793, 1908, 2441]. Academics [2846]. Accelerated [2509]. Accelerating [2763, 2735, 2736, 2793]. acceleration [1869]. Access [2370, 13, 38, 1530, 1620, 1555, 2106, 1643, 1410, 602, 1672, 1033, 531, 601, 803, 1532, 2647, 1531, 949, 2204, 1455, 657, 598, 1263, 1086, 749, 409, 1969, 319, 2119, 450, 872, 1808, 171, 1559, 238, 821, 812, 1209, 1198, 694, 126, 1357, 883, 2139]. accesses [406]. Accessing [1295, 2584, 127, 515]. Accommodating [626]. account [1889]. accounting [679]. accrued [2385]. Accurate [2838, 2578, 2579, 276, 2837, 855]. achieve [858]. Achievements [1238, 619, 627]. ACID [1688]. ACM [38, 2139, 2216, 14, 2106, 2290, 2291, 2530, 42, 2239, 2445, 13, 2863, 2866, 2867, 2868, 2870, 2871, 1580, 2563, 1099, 2869, 2130, 2865, 1946, 1697, 2213, 2029, 2143, 2319, 2397, 2815, 2262]. ACM-SIGFIDET [38, 14]. Acquisition [288]. Across [1645, 2847, 2393, 1936]. ACT [1322]. ACT-NET [1322]. ACTA [586]. action [761]. Actions [1329]. activation [851]. Active [1322, 1239, 970, 833, 1471, 272, 1574, 1407, 1071, 1321, 1236, 1174, 420, 1171, 546, 1914, 1178, 1470, 766, 1098, 521, 1228, 1444, 1126, 1065]. Activities [2797, 2848, 136, 984, 216, 156, 1965, 587, 67, 748, 820]. activity [1311, 196]. Actual [211]. actually [2474]. Ad [1394, 2819, 2820, 2001, 2056, 3]. ad-hoc [2001]. Ada [249, 458]. ADAM [1897]. adapt [2406]. Adaptable [1807]. adaptation [2113, 2328, 1209]. Adapting [1121, 254]. Adaptive [1730, 1675, 1002, 1642, 1024, 1826, 1112, 1737, 304, 1282, 307, 1809, 567]. ADBIS’2001 [1931]. ADBMS [1321, 1322]. adding [463]. Addison [2043, 66]. Addison-Wesley [66]. Address [1148, 749, 1147, 642]. Addressing [1199]. ADEPT [1588]. Aditya [2534]. Administering [221, 1303]. Administration [1478, 53]. Ado [1317]. ADO.NET [2236]. ADS [1071]. Advanced [1881, 952, 740, 2557, 1765, 2405, 1062, 678, 924, 1177]. Advancements [2105]. Advances [1227, 1931]. advantage [1163]. adventure [2310]. adversarial [2337]. Adversarially [2726, 2725]. Advice [2718]. Affiliation [2402]. after [1121]. again [1875]. against [471, 1939]. Agenda [1040].

Agent [1386, 1607, 2047, 1588, 1919, 1198, 2065]. Agent-Based [1386, 1607, 1588, 1198]. Agent-Oriented [2047]. Agents [1592, 1969, 1605]. Aggregate [1747, 1514, 1503, 1821, 2141, 508, 1830, 1333]. Aggregates [1634, 1383, 1797, 2133, 2209]. Aggregation [1641, 1384, 2607, 1407, 2606, 259, 1848, 1112, 2038]. Agnes [2045]. Agnostic [2688]. Agora [2720]. ahead [2102, 780, 805]. AI [2710, 2049, 497, 239]. AI/DB [497]. aided [358, 1867]. Ailamaki [2689]. Air [990]. AJAX [2078]. Alan [1944]. Alberto [2179, 2409]. alerters [1687]. Algebra [1516, 1628, 2609, 2608, 2633, 2632, 1287, 2438, 302, 527, 463, 412, 2259]. Algebraic [1571, 660, 2199]. Algebras [2334, 885, 699, 1284, 282, 195]. Algorithm [307, 1488, 593, 2836, 1650, 1761, 1067, 1383, 1270, 1828, 254, 1117, 381, 462, 327, 2073, 543, 877, 1118, 1077, 438]. Algorithmic [2698, 2699]. Algorithms [2626, 2615, 1623, 2438, 2726, 1290, 2725, 1028, 1502, 2714, 2835, 882, 1618, 804, 1744, 1729, 1484, 2720, 2522, 236, 637, 364, 1806, 1249, 684, 879, 597, 2496, 2487, 788, 407, 512, 1903, 484, 1112, 686, 1337, 346]. Algres [1171]. Alistair [2088]. Allocation [2857, 2056, 1825, 854, 683, 685]. Almaden [704]. Alonso [2122]. alphanumeric [1274]. AlphaSort [1008]. Already [1388]. Altaïr [641]. Alternative [1503, 2352, 2398, 759, 1911]. Alternatives [1511, 290, 432]. amateur [326, 2297]. ambiguity [2316, 2272]. ambiguous [2296]. AMDB [1559]. America [2275]. Amherst [748]. Amicalola [1980]. Among [1440, 1012, 1605, 150, 2053, 1011, 724, 1812]. AMS [1311]. analog [932]. Analysis [1706, 2686, 1990, 311, 2745, 409, 780, 2594, 2657, 2658, 2011, 2684, 2789, 2522, 2402, 1895, 2478, 1513, 470, 772, 478, 759, 410, 471, 1891, 661, 308, 1772, 395, 2186, 451, 636, 2251, 1250, 2222, 2157, 2221]. Analytical [2147, 1591, 1480]. Analytics [2763, 2854, 2454, 2537, 2746, 2557, 2505, 2545, 2488]. Analyzing [2488, 2075]. Anand [2491]. Anatomy [1762, 710]. Andreas [2483]. Andrew [2653]. Andy [2535]. AnHai [2319]. annealing [373]. annotated [2443, 1097, 746, 2083, 2451, 1098, 828, 1214]. Annotating [738, 2162]. Annotation [2228, 2074]. Announcement [1582, 29, 1069]. Announcements [846, 949, 1204]. Annual [57, 2861, 2860, 200]. anomaly [2095]. Anonymity [2339]. Anonymous [1592]. ANSI [62, 55, 1103, 222, 1989]. ANSI/X3/SPARC [62, 222]. ANSI/X3/SPARC/study [55]. answer [1341]. Answering [1640, 1749, 1793, 243, 2567, 2566, 2680, 1717, 2218, 1821, 1779, 2048, 2175]. Answers [2799, 1510, 1066, 2831, 517, 738]. Anthill [2546]. any [1874, 239]. Anything [1208]. AOSE [2047]. Apache [2854]. API [2087, 2365]. apparatus [203]. APPC [911]. append [800]. append-only [800]. apples [2188]. apples-to-apples [2188]. applicable [367, 370]. Application [151, 941, 1030, 1521, 1600, 1048, 1875, 1539, 424, 1090, 940, 913, 311, 1139, 1975, 459, 1527, 1923, 1906, 208, 731, 187, 2045]. Application-centric [941, 940]. application-oriented [1923]. Applications [1490, 2666, 1670, 1469, 1022, 1441, 2665, 1446, 2153, 2721, 1562, 1994, 1784, 1948, 1529, 735, 157, 1840, 1914, 120, 334, 1210, 2425, 397, 673, 1144, 1780, 905, 398, 314, 2073, 1456, 121, 1183, 1786, 338, 1312, 1167, 678, 924, 518, 1967, 343, 419, 2229]. Applied [2717, 94]. Applying [1124, 1342, 1834]. Approach [1732, 1603, 1588, 858, 676, 1988, 1821, 349, 351, 1839, 164, 526, 867, 1122, 2329, 174, 1923, 767, 204, 2199, 2161, 180, 565, 1999, 1943, 682, 338, 550, 289, 242, 557, 644, 870, 1987, 352, 1952]. Approaches [941, 2516, 2612, 101, 940, 1917, 1291, 430, 909, 1168, 2118, 2313, 417]. Approximate [1640, 1749, 1510, 1518, 1634, 1821, 1830, 1826]. Approximating [1747]. Approximations [1439]. April [2859, 1362]. Arabic [568]. ARANEUS [1550]. arbitrary [2468]. Archimedes [2613]. Architectural [1092]. Architecture [1324, 2338, 1658, 388, 2532, 1589, 138, 906, 1038, 2653, 1158, 1968, 285, 128, 2119, 380, 108, 1763, 205, 582, 2054, 614, 390, 521, 125, 894, 126, 202, 681]. Architectures [2424, 1414, 1044, 993, 756, 1140, 292, 1308, 1921, 680, 1307, 415, 1155, 1827, 1254, 1255, 796, 700]. Archival [2795, 30]. Archive [1682]. archiver [285]. Archives [1746, 2428, 1888]. archiving [861]. ARCS [1065]. area [1791, 1599, 1926, 2230]. Areas [2712, 2562]. ARIADNE [1556, 1601]. Ariel [545, 761]. ARIES [805, 993, 2120]. ARIES/CSA [993]. ARIES/IM [805]. ARISE [2174]. ARISE/NISR [2174]. arithmetic [1800]. Arizona [1244]. Arlington [1995]. arms [2419]. ARPA [973, 1222]. ARPA/NSF [973]. Array [1383]. Array-Based [1383]. Arrays [1644, 1501, 879, 1269, 442, 660, 685]. art [2430, 1300, 2058, 2503]. ARTDB [1239]. ARTDB-95 [1239]. Artemis [2170]. article [1331]. Artifacts [2798, 2809]. artificial [140]. Arulraj [2756]. AS/400 [951]. ASCII [46]. Asia [2460, 2460]. Asilomar [1594]. Ask [2647]. ASKALON [2168]. asked [2256]. Asking [2597]. aspect [1195, 2207]. aspect-oriented [2207]. Aspects [1091, 674, 2083, 2451, 429, 310]. assembly [659]. Asserting [1442]. assertions [716]. Assessment [1584, 1075, 213]. asset [1156, 992, 2352]. Assets [1614, 1157]. assigning [520]. Assignment [2778, 2777]. assist [2187]. assistant [2230, 2268]. assisted [541]. associate [113, 2474]. Associating [173]. Association [1392, 1393, 1630, 1472, 1408, 1511, 1484, 1482, 868, 1249, 1310, 2198, 1118, 1776, 1900, 1836, 1248, 2169]. Association-Rule [1482]. Associations [1483, 1971]. Associative [105, 132, 238, 774]. Associative/parallel [105]. assumptions [1125]. assurance [1369]. Astronomy [430, 1746]. asymmetric [1120]. Asynchronous [1719, 682]. AT&T [845]. Athena [2435]. Athens [2217, 1219, 1220]. Atomic [503, 865, 2138, 2030]. Atomicity [994]. Attack [965]. Attacking [2566]. attention [2085]. Attribute [607, 667, 759, 821]. Attributed [2775, 2776]. Attributes [1637, 1747, 520, 599]. audit [1899]. Australia [2648]. author [2439, 1901, 1858]. Author-X [1858]. authorities [1912]. Authority [1592, 1539]. authorship [2011]. AutoAdmin [1513]. Automata [1958]. automated [925, 51]. Automatic [1490, 1811, 1657, 2079, 2513, 2110, 2533, 1752, 1717, 1282, 2439, 2132, 1860, 2224]. Automatically [1506]. Automating [1753]. Automation [70, 468, 1346, 1167, 1345]. automization [135]. autonomic [1960]. Autonomous [1047, 2053]. Autoplex [2116]. auxiliary [2008]. Availability [39, 2798, 1547, 1042, 1728, 787, 785, 511, 1144, 769, 418]. available [2344]. Avi [1926]. aviation [1845]. avoid [773, 2267]. Award [2302, 2243, 2268, 2445, 2530, 2319, 2397, 2815]. awarding [2187]. Awards [1447, 1598, 1674, 1695, 1207, 1226, 1330, 1353]. aware [2061, 1814, 2430, 2054]. Awareness [2370, 2549, 2514]. Axiomatic [176, 186]. Azure [2857].

B [101, 255, 2008, 1110, 2202, 2246, 1095, 428, 2827, 687, 1572, 1889, 1190, 1748, 688, 686, 1258]. B-tree [1110, 2202, 1572, 1889, 686]. B-trees [255, 2246, 428, 688]. B2B [1878, 1921, 1917, 1916]. Back [2858, 1450, 2583, 2477, 690, 2487, 278]. back-end [278]. Backed [1699]. backend [120, 121]. Backup [1711, 788]. Backward [1554]. bad [2294]. Baden [1359]. Baden-Württemberg [1359]. Baeza [2275]. Baeza-Yates [2275]. Bag [2784, 2783]. Bailis [2649]. Balancing [1385, 2712, 1013, 1123, 332, 1128, 2369, 1864]. Bandwagon [1074]. bandwidth [1135]. BANG [394]. banks [2343]. Bao [2768]. Base [32, 33, 62, 2576, 37, 1550, 2577, 594, 221, 101, 154, 95, 547, 2061, 120, 1221, 47, 361, 102, 2212, 347, 54, 2754, 100, 172, 608, 739, 99, 73, 98, 861, 403, 63, 121, 218, 49, 50, 482, 138, 58, 678, 79, 632, 924, 42, 115, 193, 91, 198]. Based [1386, 1400, 1716, 1435, 1487, 1499, 1555, 1046, 1607, 1510, 1424, 1588, 1725, 1503, 1050, 1033, 954, 1421, 1557, 1520, 957, 1707, 1371, 596, 2612, 1509, 1493, 1750, 1383, 1863, 1784, 1308, 1820, 583, 669, 561, 2283, 2381, 2638, 1129, 1584, 2033, 589, 2464, 1866, 1821, 349, 677, 2079, 2838, 1284, 1307, 919, 1917, 1135, 1813, 1587, 1788, 335, 342, 386, 903, 996, 2084, 526, 1181, 2059, 1311, 435, 308, 2037, 1652, 1916, 2448, 896, 582, 1358, 460, 1049, 443, 1993, 614, 767, 824, 1888, 2410, 248, 1999, 2019, 683, 2118, 1254, 2022, 1118, 2055, 1272, 760, 1198, 1287, 280]. based [289, 2187, 315, 190, 801, 1896, 2837, 1967, 2257, 2280, 2169, 1765]. Bases [2613, 1318, 182, 525, 153, 105, 649, 489, 74, 602, 670, 173, 175, 1102, 59, 455, 371, 183, 51, 191, 644]. basic [94]. Basket [1651, 1391]. Baskets [1392]. battlefield [916]. Bay [2313]. Bayan [568]. BCNF [206, 461, 574]. BCS [32]. be [145, 495, 1925, 1940, 2243, 2230, 2689, 402]. beat [827]. Beckman [2519]. Become [2762]. Becomes [2753]. been [224, 2067]. Beeri [2433]. beginning [2463]. Behavior [762, 586]. Behaviour [147, 359]. behavioural [162]. Being [2624, 2085, 2244, 2267, 2472]. Belief [1627, 834]. Beliefs [1442, 1939]. Believing [2815]. Bell [845, 2088]. Bellcore [617]. below [2060]. Benchmark [2498, 1381, 476, 2271, 2388, 1691, 827, 849, 1884, 847, 848]. Benchmarking [2729, 2730, 405, 1751, 2498, 2156]. benchmarks [1097, 754, 1782]. benefits [2502]. Beng [2614]. Berkeley [2299, 632]. Berlin [2654]. Bernstein [2100]. Bertino [2385]. BeSS [1324, 1313]. Best [2780, 2779, 2547, 2364, 1184, 2227]. better [1268, 2268]. Between [2589, 868, 150, 2428, 1296, 464, 909, 49, 768, 1297, 1978]. BEWEB [2449]. Beyond [2626, 2615, 1392, 1755, 2375, 2721, 371, 1707, 2856, 1659, 1169, 2330, 2311, 2219, 2656]. BeyondMR [2626]. BeyondMR’16 [2615]. Bhowmick [2749]. Bias [2622]. Biased [2672, 1715, 2671]. Bibliography [841, 962, 745, 638, 639, 365, 1097, 746, 2083, 2451, 1098, 2235, 828, 1237, 1214]. bidirectional [2404]. Bifocal [1273]. Big [2554, 2746, 2527, 2549, 2640, 2839, 2545, 2840, 2651, 2559, 2427, 2465, 2348]. BIGDATA [2650]. BigDAWG [2539]. biggest [2293]. Bildschirmtext [457]. Bilkent [2171]. billion [1009]. billion-record [1009]. billions [1870]. bin [1699]. binary [87]. binding [1218]. Bindings [1597]. bio [2080, 2078]. BIO-AJAX [2078]. bio-molecular [2080]. BioFast [2081]. bioinformatics [1974, 1791, 2073]. biological [2079, 2078, 2102]. biologists [2401]. Bipartite [2778, 2777]. BIRCH [1257]. BIRTE [2528]. BIRTE’14 [2555]. Bit [1800, 84]. Bit-sliced [1800]. Bitmap [1512, 1635, 2008]. bitmapped [1191]. Bitmaps [1636]. blind [2211, 2222, 2256, 2221]. block [231]. Blockchain [2706, 2707]. blocking [636]. Bluetooth [2033]. Bluetooth-based [2033]. BNCOD [1930]. BNR [895]. board [322, 2195, 2220, 2245, 2529, 2580, 2610]. Boards [2818]. boat [1517, 1632]. Bob [966]. body [2002]. Bombay [2253, 2464, 1243]. Book [1902, 1922, 1941, 1955, 1976, 2003, 2042, 2121, 2754, 2309, 2252, 2463]. books [97]. Boolean [277]. Boon [2268]. boosting [1803]. bootstrap [2332]. Borkin [204]. born [610, 1875]. born-again [1875]. Boston [2860]. Both [2780, 2779, 574]. Bottom [1647, 650]. Bottom-Up [1647, 650]. bought [2409]. bound [2489]. Boundary [1656]. bounding [2234, 1111]. box [1443]. BP [2227, 2283]. BP-Mon [2283]. BP-UML’05 [2227]. BPEL [2283]. Brain [2544]. Brain-Inspired [2544]. Branding [2090]. Brands [2560]. Brazil [2409]. Brazilian [2137, 1609]. Break [2759]. breakdown [708]. Breaking [1494, 1443]. Brick [1043]. bricks [2398]. Bricolage [2573]. Bridger [2304]. bridges [2486]. Bridging [2657, 2658, 921]. brief [2439, 1947, 1599]. Bringing [54, 1668, 1933, 2720]. briskly [2219]. British [1930, 118, 2353]. Broad [2666, 2665]. Broadcast [1120, 1385, 1625, 273]. Broadening [2568]. brokering [1127, 1909]. Browsing [1084, 270, 1743, 1707, 751, 1566, 1181]. Bruce [2156]. BSR [1135]. Btx [457]. Bubba [441]. Bubbles [133, 1803]. bucket [2212]. BUCKY [1381]. buddy [2234]. budget [1247]. Buffer [1492, 1379, 1280, 1825, 683]. buffering [534, 877]. Buffers [1003]. Build [2677, 2524]. builders [768, 324]. Building [1633, 2760, 735, 1585, 2327, 732, 2635, 1398, 2623, 2634, 2157, 2828, 2614, 2312, 876, 770, 1945, 1907, 2284, 2649]. bulk [1377]. bullet [1977]. Buneman [2044, 2353]. Bunge [2123]. burned [2219]. bushy [661, 1251]. Business [2800, 2485, 2555, 1545, 2384, 1753, 1588, 1764, 2449, 2528, 1918, 1754, 70, 1863, 2283, 1914, 913, 2425, 1944, 1916, 1843, 2244, 208]. Business-to-Business [1764, 1918]. businesses [730]. Buyer [2801]. BX [2404]. byte [406]. Byzantine [2499].

C [101, 100, 2088, 66, 2374, 2388, 849, 1163, 2232]. C-SPARQL [2374]. C. [2120]. Cache [1748, 681, 1778, 1374, 1560]. cached [1826]. caches [579]. Caching [1540, 988, 1504, 1572, 594, 860, 1261, 1260, 1841, 680, 1869, 1286, 2043]. CAD [666, 321, 1433, 334, 397, 908, 693, 194]. CAD/CAM [397]. Cadlab [1179]. CAiSE*00 [1764]. CALANDA [1187]. Calculus [1702, 562, 555, 1107, 884, 267, 2259]. Calendar [950, 2169]. California [2163]. Call [13, 1582, 2419, 2442, 2501, 2142]. Called [1329]. Calls [27, 1898]. CAM [397]. came [2120]. Camps [1331]. Can [2748, 1974, 1989, 2231, 2071, 827, 768, 2176, 2401]. Canada [2859]. Candidate [1708]. Candidates [2130]. canonical [715]. Capabilities [1654, 740]. Capability [1311, 1557]. CapBasED [1311]. CapBasED-AMS [1311]. capture [560, 494]. carbohydrate [2075]. Cardinality [2838, 2258, 2837]. Care [2819, 2820]. Career [2803, 1894, 2812, 2846, 2401, 2463]. Career-enhancing [1894]. Carlo [2591]. Carlson [101]. Carnot [914, 1151]. Cartesian [1251]. Casati [2122]. Case [1217, 2790, 1575, 1414, 2546, 1368, 2234, 1939, 442, 467, 2115, 1412, 554, 346]. Catalog [1844, 135]. Catalogs [1585]. Catching [1517]. Categorization [1508, 259]. categorizing [1802]. Category [2807]. Catriel [2433]. Cattell [2552]. CDB’04 [2153]. Center [705, 2163, 2151, 2542, 2465, 2336, 2435, 1325, 704, 2455, 963, 1202]. Centered [1678]. centers [2501]. Centralized [706, 254]. Centre [196]. Centric [2513, 941, 940, 2727, 2728, 2484]. Century [1238]. Ceri [2252]. certification [315]. Cgi [1699]. Cgi-bin [1699]. chain [1920, 1780]. chair [1959, 2655, 2386]. Chairman [11]. Challenge [2549, 887, 558, 1142, 1687]. Challenges [2540, 2745, 2457, 1753, 2639, 2558, 1072, 2516, 2543, 2747, 2582, 991, 2502, 1351, 1350, 2427, 2081, 2147, 2197, 891, 1413, 2325, 2430, 1882, 2013, 2102, 2145, 1899, 2019, 889, 1980, 2360, 2348]. Chamberlain [99]. Champaign [1964]. chance [2437]. Chang [1943]. Change [1292, 1372, 1571, 1698, 846, 832, 1926, 2176]. Changed [2812]. Changes [2529, 2580, 2195, 2220, 2245, 2764, 1538]. Changing [1487, 1065, 2664, 2663, 1044]. character [104]. characteristics [794]. Characterization [1072, 461, 235, 563]. characterizations [1849, 2149]. Characterizing [2805, 2388]. charging [1710]. Chase [1731, 266]. chasing [2294]. Chaudhuri [1417, 2335]. cheap [423]. Checking [2692, 2693, 510, 1953, 294, 1288]. Chenggang [2815]. Chicago [2228, 1965]. Chien [2653]. children [2408]. Chiller [2727, 2728]. CHIMERA [1171]. Chin [2614]. choose [2085]. choosing [2230]. chopping [795]. Christopher [2397]. Christos [2193]. Chunks [1504]. CICS [1246]. CIM [1168]. circuit [322, 1908]. Citation [2186, 2187]. citation-based [2187]. Citations [2588]. Cite [2588]. CITRI [829]. City [1983]. CiVeDi [2096]. Claremont [2315]. Class [939, 1678, 487, 714, 1252, 1115, 1268]. class-division [1115]. Classes [2703, 2611, 1094, 2702, 156, 718, 719, 1804, 89]. CLASSIC [507]. Classification [1683, 1484, 474, 1898, 109]. Classifier [2629, 2628]. classify [1802]. classifying [1804]. classroom [501]. Claude [2316]. Clean [2254]. CleanDB [2254]. Cleaning [2559, 2078]. Cleanix [2559]. cleanly [2165]. CLI [1218]. Client [1521, 1646, 993, 1039, 1374, 1140, 680, 1210, 779, 1262, 1089, 700, 681]. Client-Server [1521, 993, 1374, 680, 779, 1262, 700]. Client-Site [1646]. Client/Server [1039, 1210, 1089, 681]. Clio [1860, 1795]. clips [932]. clocks [1105]. Close [1093, 1189]. Closer [2595]. Closest [1724]. Closing [1997, 1071]. Closure [1028, 401, 1689, 571]. closures [613]. Cloud [2511, 2380, 2467, 2493, 2493]. Cloud-I [2493]. CloudDB [2511, 2380, 2467]. Clouds [2618]. CloudViews [2717]. cloudy [2437]. Cluster [1938]. Clustering [1623, 1490, 1622, 1318, 1488, 1953, 1665, 2578, 1621, 1761, 1715, 2579, 2456, 1803, 534, 645, 307, 2132, 599, 2055, 644, 775, 1257]. Clusters [1714, 2573, 1798]. CLV [470]. CMM [2178]. CNN [2705, 2704]. Co [66, 858, 2011, 2357]. co-authorship [2011]. co-existence [858]. Coalesce [1701]. Coast [2298]. coauthors [2385]. CoBase [920]. cocktail [362]. CODASYL [101, 32, 136, 311, 67]. CODAYSL [319]. Codd [495, 699, 2028, 2007]. code [2156]. coder [1940]. Codes [1728, 469, 344, 106]. Cognitive [2857]. cohabitation [909]. coherency [772]. Collaborating [2712]. collaboration [2413]. Collaborative [2318, 2711, 2505, 2814, 953]. Collection [1021, 1396, 1282, 503, 865]. Collections [1738, 1865, 2161]. collective [2142]. Columbia [1578, 1896]. column [1902, 1922, 1941, 1955, 1976, 2003, 2042, 2121]. columnar [2246]. Columns [2748]. Combinatorial [998, 536]. Combined [1732, 858]. Combining [1733, 264, 1950, 420, 1184, 536]. Come [2806, 2833]. COMFORT [925]. coming [2184]. Commanding [2546]. commentary [370, 1064, 544]. Comments [65, 64]. Commerce [1584, 1592, 1762, 1844, 1586, 1661, 1526, 1977, 1910, 1944, 1780, 1782, 1918, 1765, 1580]. Commercial [2715, 903, 1144, 727, 744]. Commit [1411, 652]. Commitment [480, 771]. Committed [2825, 2826]. Committee [136, 2655, 3, 33]. Committees [1447, 1598, 1674, 1695, 1207, 1226, 1330, 1353]. Common [1499, 2667, 2668, 49, 642, 1939, 239, 1771, 1035]. Commonsense [2627]. Communication [2721, 1836, 1120, 2001, 169]. Communication-efficient [1836]. communications [2053]. Communities [2563, 2457, 2176]. Community [2551, 2163, 2135, 1940, 2176]. commute [2275, 2385]. compact [255]. companion [1990]. Comparative [1706, 1772]. Comparing [879]. Comparison [1637, 600, 1316, 101, 255, 2188, 2388, 511, 287, 783, 237, 269, 113, 1102]. Comparisons [2829, 2830]. compatible [249]. Compensation [801]. Compensation-based [801]. competition [2364]. compiled [497]. compiler [382]. Compiling [522, 473]. complementary [265]. complete [2027, 699, 2398, 438]. complete-story [2398]. completed [2365]. completeness [741]. completion [2192, 2480]. Complex [2644, 1320, 1500, 2289, 954, 1024, 1717, 1781, 1129, 448, 645, 522, 363, 733, 1796, 790, 659, 827, 314, 2358, 2163, 1302, 928, 660]. complex-object [790]. Complexity [2784, 2709, 2783, 1979, 515, 571]. compliant [1766]. Component [1584, 1158, 1919, 1428]. Component-based [1584]. Componentized [1589]. Composite [2723, 533, 666, 2079, 753, 2158, 299]. Composition [2359]. compound [371]. Comprehension [977]. Comprehensive [2853]. Compressed [2609, 2738, 2608, 2737, 1819, 1984, 1770]. compressing [2088]. Compression [942, 1820, 873, 1364]. Compressor [1704, 1721]. computability [447]. Computation [1647, 2700, 2701, 1638, 1634, 2534, 266, 2382, 1801, 1796, 2656, 1798]. computational [1692, 2392].

Computations [2659, 1406, 2660].

Computer [2136, 2287, 937, 836, 1460, 118, 57, 4, 41, 541, 129, 108, 358, 1867, 2232, 209, 125, 116, 260, 904, 2264, 2120, 1446]. computer-aided [358]. computer-assisted [541]. Computers [989, 642]. Computing [2778, 2336, 2454, 2540, 837, 196, 2777, 2538, 1678, 613, 1654, 886, 1213, 1208, 1210, 1797, 102, 2030, 113, 1960, 371, 2025, 1211, 2166, 1778]. concept [917]. ConceptBase [549]. Concepts [392, 975, 2817, 321, 1942, 2197, 974, 818, 358, 171, 1444, 484, 494, 193]. Conceptual [751, 1916, 1970, 1048, 501, 2236, 1968, 172, 179, 181, 190, 191]. concerns [1777]. Concurrency [1620, 539, 956, 905, 2808, 582, 1375, 422, 1625, 958, 959, 2750, 1105, 290, 540, 583, 254, 236, 513, 654, 655, 778, 444, 1230, 1229, 799, 826, 803, 794, 805, 219, 1903, 686, 706, 1232, 681]. Concurrent [2694, 428, 2695, 1374, 1872, 127, 126]. Condition [1206, 761, 276, 402]. conditions [1126]. Conference [32, 2862, 2863, 2866, 2867, 2868, 2870, 2871, 1481, 1582, 2818, 2724, 2766, 1793, 2869, 1931, 2338, 2751, 2865, 1930, 950, 2371, 2460, 1466, 1994, 1099, 2152, 215, 2321, 200, 2469, 2048, 60]. Conferences [2797, 2848, 2100, 2440, 2213]. confidence [1999]. configuration [1412]. Conflict [520]. Conflicts [716]. Confluence [2692, 2693, 762]. Congratulations [2762]. Congressional [1749]. conjectures [2373]. Conjunction [2371, 2213]. Conjunctive [2611, 2531, 2829, 2830]. Connected [2563, 2818, 910]. connecting [2751]. connection [768]. connectivity [2117]. Conscious [1748, 2687, 867]. Consensus [975, 974]. considerations [164, 2360]. Considered [1572]. Consistency [1523, 149, 2784, 2783, 1374, 777, 2174, 2343, 363, 1233, 1289, 681]. Consistent [2218, 2823, 2567, 2566, 2824, 245, 1198]. Consortium [1322]. Constant [2697, 2696, 2531, 461]. Constant-Delay [2697, 2696]. constant-time-maintainability [461]. constitutes [2230]. Constrained [1631, 1483, 2181, 415, 672]. Constraint [2153, 1769, 1106, 1849, 896, 1956, 1689, 1288]. constraint-like [1849]. Constraints [2682, 1789, 2604, 163, 1631, 807, 420, 162, 541, 2206, 363, 852, 508, 295, 1126]. Constructing [1556, 971, 333]. Construction [1519, 1632, 2576, 2577, 88, 2163, 2224]. constructive [186, 194]. constructs [649, 463, 389]. Consul [177]. Consumer [1033, 1363]. Containment [2141, 2066, 1832]. contemporary [2360]. Content [1658, 1699, 1843, 1173, 1308, 1857, 1841, 1307, 1869, 1254, 1901]. content-based [1308, 1307, 1254]. Contention [2727, 2728, 467]. Contention-centric [2727, 2728]. contents [1240, 2000]. contest [2387, 2448]. Contests [2849]. Context [2370, 938, 723, 2061, 2273, 311, 2023, 1357, 1422]. context-aware [2061]. contexts [436]. Contextualizing [2512, 1604]. continent [2275]. continual [1797]. Continuous [1558, 1893, 1740, 2740, 2739, 2617, 800, 2094, 2461, 1255]. Continuously [1730, 1736]. contra [817]. Contracting [1915]. contradictions [487]. contrasts [164]. contribution [2303]. Contributions [2641, 2307, 2306]. contributors [2399]. Control [1320, 1620, 2106, 2780, 956, 2779, 1410, 1625, 2647, 959, 1105, 290, 540, 440, 583, 254, 236, 513, 1282, 772, 163, 444, 539, 679, 1230, 1229, 799, 826, 905, 2808, 582, 2201, 794, 682, 219, 1903, 422, 686, 706, 958, 1232, 681, 2085, 1558, 13, 38]. controlled [789]. controller [127]. controlling [483]. controls [679]. Convergence [2834, 2123]. Conversion [1497, 213, 211, 311]. Converting [1365]. Cooperation [1616, 2059]. Cooperative [1481, 604, 922].

CoopIS [1481]. CoopWARE [1608]. Coordinating [126]. Coordination [2504, 2021, 2375]. Copy [1137]. CORAL [926, 864]. Corfu [2371]. Corpora [2517]. Corporation [1203]. CoRR [2586]. correct [2307, 637]. Correctness [2504, 85, 2616, 1122, 481, 542]. Correlated [1485, 1797, 2056]. correlation [1863]. Correlations [1392]. correspondence [1331, 716]. Corrigenda [40]. COSIMA [1852, 1853]. cosine [2257, 2280, 29]. COSS [1035]. Cost [1677, 1318, 1743, 1287, 1509, 1493, 1010, 1750, 2345, 867, 789]. Cost-Based [1509, 1750, 1287]. cost-controlled [789]. Cost-effective [1743]. Cost/Performance [1010]. Costly [1406]. Costs [2778, 2777]. coterie [786]. Cougar [2036, 1952]. Council [2498]. count [2200, 1802, 672]. Counting [2703, 2799, 1391, 2786, 2702, 1750, 2785, 376, 2267]. countries [2502]. country [2409]. Course [29, 2677, 61]. CourseRank [2368]. Cover [1157]. covers [488, 637, 615]. Covid [2724]. Covid-19 [2724]. Cow [2754]. CPR [2694, 2695]. CPU [406]. CPUs [2750]. CQ [1551]. Cracking [2598]. Crash [779, 1119]. Crawlers [1660]. creating [804, 2483]. Creation [1431, 2793]. Credit [476]. Crete [2492]. Crisis [2803, 1040, 2321]. criteria [200]. Critical [1670, 1075, 2271]. critique [1103, 2316, 1062, 2272, 2316]. critiquing [2316, 2272]. Cross [2682, 2823, 2824, 2255]. cross-cutting [2255]. Cross-Engine [2823, 2824]. Cross-Row [2682]. crossing [2486]. Crossroads [1417, 1416]. Crowd [2521, 2558]. CS [1940, 1959, 2275]. CSA [993]. CSCW [1675]. CSCW-98 [1675]. Cube [1752, 1798, 1377]. Cubes [1439, 1376, 1378, 1796, 1267, 1647]. Cubetree [1377]. Cubetrees [1503]. cultural [135]. curated [2353]. curation [2413]. CURE [1488]. Current [1584, 2758, 819, 825, 2653, 55, 2102, 634, 48, 52]. Curricula [2793]. curriculum [2437, 1940].

Curse [1494]. curve [1785]. Customer [1539, 1149]. Customizable [1475]. Customized [117, 2096]. cut [2385]. cutting [2255]. CVDB [2136, 2287]. CWI [2424]. cyclic [234]. Cymbal [1673].

D [100, 99, 1351, 1350, 1380, 1652, 2532, 305]. Dagstuhl [2322, 2723, 2068, 2404, 2405, 2651]. DAI [2204]. Daisy [2455]. DaMoN’06 [2238]. Dan [2044, 2624]. dangers [1264]. Daniel [2453]. DANTE [2645]. DARE [1862]. Dark [2588]. Darmstadt [2189, 388]. DARPA [1352, 730, 822]. DART [1444]. DASlab [2536]. Data [2654, 2237, 446, 2593, 1633, 1385, 1382, 1506, 1651, 1405, 1490, 1745, 2238, 2813, 2571, 2760, 2521, 2498, 13, 32, 38, 33, 62, 2862, 2863, 2866, 2867, 2868, 2870, 2871, 1463, 2748, 1423, 1450, 2800, 2278, 2139, 1734, 2854, 2526, 95, 1606, 1076, 1942, 2044, 2645, 1263, 1530, 2216, 2821, 156, 112, 941, 1803, 1391, 157, 1047, 2189, 2869, 2646, 2285, 680, 1611, 2276, 1664, 2464, 2354, 1475, 720, 1465, 1415, 1372, 2554, 2492, 2537, 2172, 1497, 161, 2229, 441, 2758, 2847, 2686, 2603, 2788, 1653, 2347, 257, 1031, 1910, 2746, 1221, 53]. Data [2527, 833, 2548, 1610, 2727, 2794, 361, 1735, 1536, 1360, 2822, 1249, 2323, 2865, 1399, 2068, 2790, 996, 1944, 2740, 2416, 2215, 2715, 74, 1301, 1430, 1393, 1093, 2541, 2640, 169, 1376, 989, 1891, 837, 452, 1642, 2318, 2754, 1621, 2667, 1671, 2706, 37, 2722, 1548, 2801, 2602, 2839, 1672, 2685, 2713, 2757, 1446, 119, 1648, 2807, 1431, 1546, 1033, 1727, 954, 1549, 2570, 1704, 1721, 1728, 2721, 2455, 2511, 2759, 2668, 7, 1451, 1445, 1024, 2380, 2467, 1847, 2850, 1408, 2, 2505, 1378, 2490, 371, 1454, 1507, 1539, 2544, 2739, 453, 1661, 1715, 2545, 1655, 1682, 2786, 2840]. Data [2291, 2466, 2494, 2553, 2676, 2546, 2814, 2639, 2789, 1371, 1744, 1532, 183, 1709, 1698, 1726, 1203, 1478, 1468, 143, 2707, 2625, 2589, 187, 2214, 550, 2217, 1434, 2650, 79, 1966, 1153, 1896, 632, 594, 1666, 191, 2770, 2239, 2520, 1634, 998, 2559, 721, 1531, 2586, 2683, 1924, 2785, 1837, 1563, 2728, 2484, 2522, 2279, 2562, 2793, 2064, 1881, 622, 1189, 1120, 1799, 1784, 525, 505, 653, 2390, 153, 440, 454, 71, 1212, 101, 154, 2502, 2113, 2027, 2204, 2075, 2021, 1893, 1820, 1897, 1895, 2346, 2443, 469, 2381, 2314, 2481, 85, 2174, 105, 2427, 1828]. data [2061, 2273, 507, 869, 1981, 940, 81, 158, 1145, 1293, 1789, 2413, 2197, 589, 2331, 2428, 1905, 485, 891, 1840, 2396, 2478, 2482, 1367, 1856, 1086, 1872, 2079, 1859, 65, 2343, 907, 120, 797, 873, 426, 1065, 1851, 1813, 163, 1368, 1839, 2325, 749, 2430, 1117, 2425, 1296, 2001, 47, 1241, 691, 2146, 834, 1686, 128, 1797, 697, 2099, 2196, 1849, 1996, 9, 2542, 225, 2184, 86, 2437, 102, 2071, 875, 80, 489, 526, 2212, 2013, 1946, 2077, 1885, 2056, 166, 1267, 2167, 2119, 2018, 2078, 2127, 2249, 783, 2004, 54, 75, 90, 1954, 1291, 2403, 673, 737, 509]. data [435, 648, 2198, 72, 790, 308, 769, 100, 2333, 732, 2405, 2145, 658, 2375, 172, 1867, 460, 2391, 1612, 443, 739, 296, 173, 1947, 1873, 1785, 174, 486, 1899, 2421, 2046, 2133, 2489, 69, 390, 531, 1888, 439, 2032, 915, 99, 2422, 178, 2349, 1102, 1794, 180, 98, 59, 78, 861, 3, 1842, 1845, 395, 1453, 226, 1833, 382, 1209, 68, 1, 63, 2022, 756, 1934, 121, 218, 1773, 2446, 182, 2055, 2452, 2501, 1305, 1918, 49, 50, 1333, 482, 1359, 51, 8, 220, 109, 1874, 1377, 138, 715, 1297, 2462, 411, 2251, 1250, 1907, 77, 188, 58, 723]. data [2164, 908, 493, 271, 2465, 1448, 2038, 2157, 1337, 2020, 42, 910, 2173, 1357, 202, 115, 278, 2497, 717, 2058, 1771, 574, 2407, 192, 193, 91, 250, 1926, 2193, 2275, 2335, 2385, 2415, 2423, 2414, 2491, 1778, 1257, 2224, 2165, 204, 33, 2450, 2454, 2536, 963, 221, 194]. data-base [101, 47, 100, 99, 98, 50]. Data-based [2464]. data-bases [59]. Data-centric [941, 940]. Data-Driven [2676, 2586, 1837, 2375]. Data-Intensive [1611, 1446, 2721, 2455, 440, 2167, 2403, 1907, 963]. data-mining [1117]. data-oriented [2273]. data-structure-set [65]. Data/knowledge [721, 482]. Database [2859, 2519, 2690, 505, 1055, 886, 2370, 2797, 2848, 676, 843, 982, 1322, 964, 2795, 1423, 842, 293, 1670, 1561, 1433, 1239, 137, 224, 1594, 2694, 2843, 1323, 2203, 1061, 216, 336, 1677, 2473, 1047, 2540, 1029, 1616, 1995, 1578, 1475, 1073, 1718, 2270, 1555, 1326, 1457, 1021, 705, 1418, 1585, 223, 2206, 2513, 1200, 1534, 2745, 1702, 1380, 833, 258, 1046, 1577, 457, 984, 1409, 2698, 1425, 1515, 946, 999, 754, 956, 1684, 1411, 569, 704, 593, 64, 1059, 845, 2072, 634, 2701, 950, 975, 553, 2595, 1574, 2518, 609, 2601, 1037, 962, 196, 2804, 1699]. Database [1541, 1727, 2620, 2721, 1579, 629, 2633, 963, 1445, 2410, 1048, 993, 1663, 1407, 997, 1609, 728, 747, 506, 2600, 889, 948, 2151, 830, 1243, 2695, 955, 943, 1202, 1726, 942, 829, 1053, 1027, 2856, 2699, 2516, 1511, 2677, 2217, 1467, 2089, 2647, 617, 2360, 189, 619, 1238, 1443, 116, 972, 260, 2623, 1052, 959, 820, 1564, 835, 2755, 1379, 947, 844, 2171, 1244, 1040, 2582, 1543, 2264, 1071, 953, 846, 987, 1964, 2525, 2535, 2598, 2614, 2619, 2756, 1374, 831, 1090, 1321, 1136, 312, 1124, 292, 2315, 331, 991, 762, 2014, 350, 1236, 142, 1220]. database [2402, 55, 2419, 2500, 623, 755, 131, 340, 1529, 2188, 1099, 1988, 1227, 317, 791, 354, 1070, 1216, 1303, 268, 735, 1890, 112, 157, 222, 158, 325, 2413, 1469, 408, 616, 2002, 510, 1841, 236, 513, 624, 1413, 1840, 436, 160, 1171, 249, 1538, 133, 273, 677, 1819, 2034, 798, 306, 161, 522, 2208, 281, 1965, 1327, 1566, 426, 887, 252, 478, 631, 2307, 311, 862, 1192, 1309, 456, 974, 888, 288, 203, 1139, 2425, 2001, 274, 2345, 733, 261, 621, 1155, 903, 466, 983, 1199, 339, 434, 766, 918, 971, 103, 753, 564, 165, 649, 1294, 1146, 1979, 2132]. database [1458, 2050, 378, 471, 1687, 94, 397, 701, 1342, 2167, 168, 380, 461, 499, 1122, 108, 414, 826, 2135, 372, 435, 818, 780, 752, 1188, 742, 459, 1340, 633, 1527, 708, 316, 2808, 552, 590, 205, 967, 568, 1082, 237, 1937, 460, 295, 1231, 234, 174, 1304, 345, 381, 614, 767, 1960, 824, 359, 204, 690, 383, 43, 177, 2096, 521, 491, 179, 2092, 925, 1104, 403, 1289, 270, 626, 584, 738, 933, 1525, 455, 559, 1842, 543, 656, 2393, 981, 1984, 877, 2357, 353, 810, 313, 1183, 388, 467, 745, 646, 675, 815, 61, 465, 1179, 2186]. database [926, 864, 748, 480, 1777, 389, 828, 1089, 1874, 185, 240, 731, 405, 413, 694, 620, 197, 219, 1101, 951, 318, 396, 87, 1883, 424, 572, 663, 280, 1980, 289, 242, 678, 300, 2207, 1348, 2080, 417, 315, 1126, 744, 493, 1234, 126, 1886, 518, 2361, 557, 429, 640, 702, 819, 958, 860, 904, 1237, 1219, 1214, 883, 898, 494, 639, 192, 343, 2226, 66, 1987, 548, 592, 1936, 1926, 1940, 2085, 2100, 2205, 2244, 2335, 2401, 2432, 1151, 229, 916, 230, 247, 890, 1832, 419, 1462, 2424]. Database-Backed [1699]. Database-Centric [2513]. Database-Driven [1585, 1841]. database-intensive [157]. Database-like [1555]. Databases [1619, 841, 2288, 2858, 2136, 2287, 1439, 1087, 1489, 1370, 1479, 1626, 1657, 1056, 1664, 1740, 1440, 1499, 1724, 2394, 1662, 1025, 1486, 2794, 1735, 834, 1930, 1732, 2616, 1488, 1057, 1480, 2152, 1665, 1502, 2277, 1063, 1081, 1627, 2486, 1496, 1092, 1299, 1652, 2334, 1446, 2510, 2561, 1414, 2153, 1472, 1713, 2254, 1923, 2565, 2742, 2372, 2426, 979, 980, 2460, 1650, 957, 1743, 2581, 2851, 938, 2592, 2621, 2853, 2191, 1710, 1402, 2741, 2543, 2816, 1432, 1759, 2680, 2142, 2591, 498, 1994, 1259, 540, 771, 868, 666, 377, 784, 365, 857, 251, 1455, 131, 399, 669, 207]. databases [2082, 1988, 2218, 1125, 2281, 1145, 2363, 2076, 1931, 420, 2103, 720, 1097, 645, 302, 1528, 367, 370, 1282, 765, 243, 1096, 1065, 1990, 541, 363, 1276, 504, 2458, 1208, 1165, 1365, 2183, 630, 103, 127, 319, 1009, 625, 662, 852, 1306, 799, 468, 168, 1114, 2127, 783, 672, 515, 244, 348, 1802, 1098, 892, 1780, 1997, 2049, 1364, 170, 358, 1180, 265, 921, 1809, 827, 809, 500, 582, 1197, 647, 1956, 922, 437, 1182, 1689, 2137, 418, 722, 978, 241, 2499, 2026, 141, 750, 856, 741, 893, 430, 1295, 529, 973, 1198, 1935, 635, 1228, 1444]. databases [1366, 793, 814, 2091, 384, 284, 715, 502, 490, 1884, 422, 2160, 1150, 1880, 301, 628, 638, 716, 1279, 538, 800, 2125, 729, 320, 627, 825, 2282, 1769, 1694, 2223, 310, 693, 901, 870, 1185, 1770, 2045, 2219, 2284, 2335, 2353, 352, 2434, 712, 227, 474, 724, 1812, 438, 698, 890, 1257, 612, 1568, 1569, 1215]. DataBlade [1245, 1038]. DataBlitz [1670]. Databse [1328]. datacycle [380]. DataEd [2813]. Dataflow [1087, 440, 1907, 419]. datalog [446, 504, 447, 2268, 2686, 581, 2787, 2834]. DataMine [1170]. Datasets [2664, 1394, 1395, 1638, 2663, 2720, 1117, 1420, 1768, 2512]. Dataspaces [2584, 2183]. DataSplash [1552]. DataSpot [1566]. Date [66, 102, 73, 2272, 839, 840]. David [1940, 2432]. Davidson [2712]. day [2013]. days [1915, 2156]. Daytona [1673]. DB [2190, 1263, 332, 2148, 497, 2157, 763, 1940, 2176, 2524]. DB&IR [2322]. DB/IR [2190]. DB2 [1160, 1161, 1667, 1533, 1960, 1039, 1224]. DBA [1990]. DBAs [2156]. DBGlobe [2025]. DBLearn [1057]. DBLP [2148]. DBLP-DB [2148]. DBMiner [1306]. DBMS [323, 2864, 496, 680, 546, 1435, 534, 356, 1054, 1091, 930, 1178, 425, 545, 1514, 205, 241, 263, 1451, 248, 902, 1398, 700, 484, 1434, 1659, 532, 239, 554, 611, 899, 1232, 2060, 681, 2366]. DBMSs [1036, 2572, 2573, 689]. DBPL [2203]. DBRank’07 [2277]. DBS [222]. DBS-SG [222]. DBSim [1304]. DBTG [81]. DBTG-data [81]. dbTrento [2446]. DC [2869]. DDB [908]. DDL’ [60]. DDT [1046]. DDTS [203]. Dead [2754, 2476]. Deadlock [423, 943, 479, 1067, 364]. Dean [2472]. debate [65]. Debit [476]. debugging [1559]. DEC [1031]. decade [1529, 2184]. DECdtm [896]. December [1580, 2108, 620, 2263]. Deciding [2792]. Decision [1041, 2526, 1415, 1042, 1632, 1600, 1509, 1154, 1205, 1782, 1268]. Decision-Making [1600]. Declarative [2504, 2700, 2686, 1753, 2701, 2577, 2352, 2373, 2375, 2324]. declustering [759]. Deco [2452]. decomposing [504]. decomposition [303, 206, 232]. decompositions [207]. decompression [1984]. DEDALE [1500]. deduction [746]. Deductive [1627, 1053, 1027, 627, 377, 662, 348, 552, 590, 827, 614, 664, 675, 926, 864, 572, 663, 729, 898, 2219, 474]. DeeDS [1236]. Deep [2575, 2574, 2705, 2704, 2582, 661, 2376]. DeepDive [2576, 2577]. Deepening [2568]. Default [159]. deferred [510, 1290]. Defined [1677, 753, 89]. Defining [2659, 426, 2660]. definition [9, 73, 68, 1, 1333, 1069, 250]. degradation [1265]. degrees [2148]. DEIS’10 [2405]. DELAB [452]. Delaunay [1418]. Delay [2697, 2696, 2531]. delayed [771]. Delays [1493]. deleting [346]. Deletion [1095, 1190]. Delivering [1547]. delta [1959]. demand [1133, 1825]. demo [2180, 1420]. demonstration [1858, 931, 2364, 2386]. Dennis [2401]. denotational [2094, 813]. dense [346]. Densest [2732, 2731]. Density [1716, 1715]. Density-Based [1716]. department [1959, 2385]. dependence [679, 502]. dependencies [244, 232, 234, 462, 565, 1289, 914]. dependency [1863, 311, 1813, 462]. dependency-based [1813]. deploying [1783]. Dept. [1446]. depth [433, 60]. Derivability [2343]. derivation [439]. Derived [1382]. Describing [1905, 1879]. Description [13, 38, 869, 7, 180, 8]. descriptive [2, 3]. Descriptors [1722]. Design [1590, 1611, 1512, 1030, 862, 1192, 1046, 1686, 946, 1269, 1088, 2019, 407, 187, 253, 1177, 2522, 153, 454, 154, 2419, 293, 1134, 322, 2061, 2271, 273, 667, 798, 920, 363, 288, 287, 545, 397, 701, 468, 2093, 673, 358, 552, 1867, 1194, 2046, 491, 455, 543, 1934, 50, 384, 185, 413, 894, 424, 2360, 242, 2207, 417, 908, 355, 557, 1911, 230, 304, 2069]. designers [324, 404]. Designing [2766, 245, 263, 2675, 1746, 1784, 157, 51]. designs [574]. desirable [75]. Desktop [1437]. detail [2221]. Detection [1898, 1372, 2551, 2516, 1067, 1799, 1897, 1895, 1137, 1292, 364, 423, 2040, 1279]. detectors [1896]. determine [2035]. Determining [724]. determinism [762]. deterministic [377, 663]. developed [2502]. Developing [1544, 2210, 2077, 2409]. Development [1325, 913, 80, 459, 2349, 208, 928, 731, 1967, 2223, 2408]. Developmental [2253]. Developments [2029, 2143, 2487, 729]. deviation [2040]. DEVise [1420, 1395]. DeWitt [1940]. DFI [2770]. Diagram [1056]. DIAM [87]. DIAMetrics [2729, 2730]. Dick [18]. dictionaries [169]. Dictionary [1090]. did [2302]. Diego [2804]. Dieter [1977]. Differential [2852, 327]. Differentially [2659, 2841, 2842, 2660]. Differentially-Private [2659, 2660]. diffusion [2470]. Digital [1542, 1213, 2751, 1697, 1078, 2341, 2371, 2691, 2544, 1604, 2853, 1746, 2311, 1787, 1137, 2428, 1788, 2378, 1888, 299, 932, 1201]. Digree [2623]. DILS’05 [2215]. Diluting [1688]. dimension [1689, 263]. Dimensional [1714, 1490, 1400, 1475, 1399, 1742, 1401, 2063, 1505, 1799, 1784, 2027, 1529, 1828, 1813, 1110, 1249, 1281, 1747, 658, 1829, 1785, 433, 1934]. Dimensionality [1494, 1496, 1671, 1809]. DIPS [550]. Direct [1672, 284, 2110, 608, 123, 750]. Directed [2732, 2731]. Directions [2593, 2864, 1698, 2650, 95, 2082, 496, 361, 558, 630, 662, 64, 2091, 77, 628, 115, 627]. Directories [1629]. Directory [910]. DIRECTV [1142]. Disaggregated [2827, 2828]. Disambiguation [1000, 2439]. disc [530, 406]. Disco [1428]. Discovering [1591, 2471, 1971, 1879]. Discovery [1657, 1656, 999, 2509, 1057, 2732, 2731, 998, 1432, 1170, 2329, 1773, 1900, 1185]. discrepancies [647]. discs [530, 122]. Discussion [2858, 317, 1908]. disjunction [519]. Disjunctions [1019]. Disjunctive [1018, 816]. Disk [1501, 955, 285, 879, 877, 688, 579, 685]. Disks [1120, 1575, 470, 442]. disparate [1839]. display [365, 733]. Displays [970, 1344, 1343]. disruptive [2005]. Dissemination [1851, 2024, 2201]. Dissertation [2319, 2397, 2815, 2562, 2243, 2268, 2445]. Distance [1400, 1502, 1737, 1784, 2489]. Distance-Based [1400, 1784]. Distinct [2786, 2785]. Distinguished [2205, 2243]. distorted [878]. Distributed [1464, 764, 2281, 836, 379, 1013, 1411, 2701, 2771, 2827, 1728, 838, 2040, 1682, 2546, 1681, 943, 1726, 220, 2214, 1614, 2623, 557, 1010, 475, 580, 1067, 2772, 1261, 1260, 440, 1236, 2461, 317, 332, 513, 249, 273, 907, 504, 364, 203, 630, 2387, 127, 852, 2167, 2373, 881, 652, 418, 880, 135, 1842, 371, 313, 646, 928, 682, 480, 479, 576, 219, 1836, 422, 906, 1348, 2080, 315, 493, 126, 689, 860, 901, 221, 250, 712, 230, 1428]. Distributing [1012, 1011, 229]. Distribution [785, 2570, 308, 769, 449, 502, 1412, 763]. Distributions [2740, 2739]. Distributor [1031]. diversification [2377]. Diversity [2797, 2848, 2255]. Divesh [2415]. division [1115]. DLFM [1756]. DM [2157]. DMBS [468]. DMDCON [2201]. DMDW’03 [2046]. DMDW’99 [1686]. DMSN [2278]. DNA [2076, 1871]. DNA-miner [1871]. Do [2667, 2588, 2792, 2668, 2778, 2474, 1155, 2231, 2777, 768, 2853, 1940, 2409]. Doan [2319]. Doctoral [2520]. Document [2697, 1722, 2696, 2517, 1707, 1014, 1432, 1543, 1679, 751, 1887, 285, 415, 1185].

Documents [1506, 1080, 1016, 2788, 1656, 1722, 1085, 2590, 1858, 1079, 1303, 2147, 1137, 2059, 2088, 1302, 1883, 2282, 1812]. does [2474, 2335]. doesn’t [1940]. doing [1089]. DOLAP [1946]. DOLAP’05 [2239]. DOLAP’07 [2291]. Domain [2328, 2688, 2650, 2170, 73, 248, 455, 709]. Domain- [2688]. Domain-Specific [2650]. Domains [1315, 1499, 2847, 1332, 1624, 173, 1331]. Dominates [2778, 2777]. Domino [1663]. Domino/Notes [1663]. Donald [99]. Don’t [2647, 2476]. DOODLE [765]. double [2211, 2222, 2256, 2221]. double-blind [2211, 2222, 2256]. Doubly [878, 402]. Down [2759, 240]. Dr. [1064]. Drawing [2749]. Dresden [2620]. drinking [2035]. Drive [1478]. Driven [1585, 833, 2342, 2676, 1841, 1687, 1311, 2375, 2349, 2223, 2586, 1837, 2522]. drives [2401]. DSQ [1732]. DSS [1155]. DTL [1566]. Dublin [1983]. due [2724]. duplicates [1131]. durable [2448]. during [1963, 1522, 1265, 42]. duties [2472]. Duty [957]. DyDa [1872]. DynaMat [1648]. Dynamic [2200, 155, 1386, 1391, 1001, 1869, 1127, 1658, 1013, 970, 535, 1336, 391, 1496, 2839, 1648, 1699, 1825, 2744, 1764, 492, 2840, 1589, 2743, 2709, 685, 1857, 1841, 1851, 256, 298, 658, 2201, 359, 774, 1337]. Dynamically [2664, 2663, 1451, 753]. Dynamically-Fetched [1451]. Dynamics [2107, 220, 1978].

e-business [1914, 1843]. E-commerce [1765, 1944, 1780, 1584, 1762]. e-mail [1901]. E-salesperson [1852, 1853]. e-science [2164]. e-services [1774, 1879]. each [1152, 2071]. early [1982, 2087]. earth [2237, 749, 1360, 910]. Earthquake [2163]. East [1931, 2172, 1325]. Eastern [17]. easy [1566, 366]. eat [2748]. EBCDIC [46]. eBusiness [1915]. ebXML [1916]. eCo [1766]. Ecobase [1890]. Economy [1586].

Ecosystem [2720]. ECRC [1188]. ECSA’2007 [2338]. EDBT [2379, 2724, 2125]. EDBT/ICDT [2379, 2724]. EDBT’02 [1981]. EDBT’04 [2092]. Eddies [1730, 2057]. Edgar [2028]. Edge [2618]. edit [2489]. Editing [121]. Edition [1160, 2027, 1161, 2004]. editor [1881, 2015, 1910, 2159, 1943, 1223, 31, 1335, 2474, 19, 12, 35]. Editorial [2529, 2580, 2610, 2195, 2220, 2245]. Education [2813, 2858, 2754, 2804, 2850, 557]. Educational [1601]. EDUCE [357]. EFDBS [1457]. Efendi [1179]. effect [354, 369]. Effective [645, 1742, 534, 1107, 447, 1889, 1283, 1743, 1118, 1812]. Effectiveness [999]. efficiency [2501]. Efficient [1105, 525, 1405, 2703, 2666, 575, 2823, 871, 1060, 1620, 1635, 2613, 1720, 2847, 1610, 1805, 1817, 2246, 319, 1488, 1796, 2541, 2664, 990, 2578, 2665, 1491, 1652, 659, 1712, 1721, 1521, 1132, 2732, 1638, 2663, 773, 523, 1743, 2579, 2702, 1744, 1729, 1027, 1625, 2731, 2169, 267, 2765, 1812, 1823, 2468, 598, 787, 519, 2076, 2382, 637, 1801, 1818, 1954, 1993, 2358, 2133, 1277, 1283, 805, 861, 1776, 1836, 2038, 2497, 574, 2407, 1257]. Efficiently [1382, 1489, 328, 1394, 396, 2824, 1267]. effort [2498]. EFIS [1676]. Eighth [2239]. EIHA [1783]. Elasticity [2525]. elections [2130]. Electronic [1526, 1580, 1590, 1593, 1585, 1587, 1592, 2290, 1586, 1661, 1589, 1090, 1116, 1766, 1977, 1910, 1918]. Elements [2786, 2785]. Eliminating [1406]. Elisa [2385]. Elsewhere [846, 1222]. embedded [656]. Embedding [2575, 71, 2775, 2574, 2776]. embodied [462]. EMC [1672]. Emergence [2123]. Emergent [1973]. Emerging [2554, 2393, 2532, 1166]. Empirical [255]. Empiricism [2625]. employment [1967]. Empowering [2618]. Enable [1753, 1869]. enabled [1968, 1791]. enabler [1159].

Enabling [1841, 1586]. Enactment [2459, 2495]. Encapsulation [577, 91]. Encina [906]. Encoding [1635, 2150]. Encrypted [2589, 2282]. encryption [2360]. encyclopedias [2440]. End [2760, 136, 569, 67, 2688, 1908, 279, 208, 706, 278]. End-to-End [2760, 2688, 1908]. End-User [569, 208]. EnDM [2466, 2494, 2553]. endogamy [2469]. Energy [990, 2037, 2466, 2494, 2553, 2369, 2501, 2497, 2407, 2522]. energy-efficient [2497, 2407]. enforce [363]. enforcement [280]. Enforcing [1410]. Engine [2823, 2605, 1027, 2677, 2623, 2824, 1846, 2387]. engineer [971]. Engineering [2390, 2139, 1676, 2152, 2629, 2342, 324, 2628, 1298, 541, 2047, 2050, 2093, 398, 1456, 1966]. Engines [2729, 1547, 2730, 2459, 1755, 1921, 2444, 2495]. England [15]. Enhanced [1508, 1570, 1434, 2065, 2385]. enhancement [532]. Enhancing [1122, 418, 1233, 915, 2197, 1894]. ENORM [2506]. enough [1519, 1940, 1388]. enriched [2097]. Enrichment [2790]. ensemble [2224]. Ensuring [994]. Enterprise [2336, 2555, 1143, 1159, 1044, 2528, 2817, 1531, 1436, 2340, 901, 2485]. Enterprise-Wide [1436, 901]. Enterprises [2850, 2384, 1980]. Entity [2062, 2849, 1441, 2661, 2635, 1452, 2688, 2634, 2662, 2236, 817, 1193, 542]. Entity-Relationship [2062, 1441, 542]. entries [428]. Enumerating [1750]. Enumeration [2697, 2703, 2696, 2702, 2531]. Environment [1056, 1437, 1700, 1338, 1212, 158, 160, 907, 1139, 2096, 928, 512, 484, 855, 458, 2168, 230, 1130, 505, 2163]. Environmental [1354, 1600, 1356, 2035, 1890, 1368, 1361, 1355, 1358, 1359, 1357]. Environments [988, 1386, 1625, 1605, 1120, 2181, 293, 1872, 414, 2378, 1089, 916]. EOS [1058]. ephemeral [866]. Epsilon [1828]. Equi [433]. Equi-depth [433]. Equity [2848]. equivalence [2456, 713]. equivalent [402]. Era [2640]. ERIC [2526]. Erich [2441]. Error [2782, 672, 2661, 2781, 2662]. Error-constrained [672]. Errors [137, 2550, 671, 197]. Essential [2506, 636]. Establish [2667, 2668]. Establishing [210]. establishment [2052]. Estimating [231, 1274, 1003, 2257, 2280, 433]. Estimation [1619, 1677, 1002, 2838, 1520, 2837, 2550, 1273, 1835, 802, 1123, 2489, 824, 567, 276, 1275, 240, 2258, 855, 1834]. Estimators [1637]. ETH [2336, 982, 981]. Ethics [2749]. EUFC [136]. EUG’91 [743]. Europe [2244, 2252, 2284]. European [1931, 2371, 2338]. evaluate [246, 2409]. Evaluating [1285, 2059, 502]. Evaluation [2682, 1512, 1001, 2841, 2852, 2842, 339, 872, 1492, 1398, 788, 596, 1067, 1505, 331, 2571, 583, 259, 1125, 124, 357, 2411, 268, 519, 546, 258, 1805, 307, 535, 672, 866, 614, 2098, 2133, 330, 2362, 523, 656, 796, 407, 512, 1807, 650, 60, 1911, 337, 1113, 651, 438, 98]. Evanston [620]. Even [2647, 492]. Event [2644, 2723, 2203, 1435, 766, 2263, 1863, 2356, 363, 1311]. Event-Based [1435]. event-driven [1311]. event/trigger [363]. Every [2706, 2707, 1149]. everything [1296, 1297]. Evolution [939, 209, 1698, 2113, 399, 2083, 2451, 472, 1889, 2235, 1777, 828, 814, 709]. Evolving [1041, 1402]. example [557]. Exchange [2174, 2346, 2170, 1145, 2405]. exchanged [2170]. Exclusive [1523]. executing [2165]. Execution [2605, 2727, 1004, 1642, 1491, 1750, 2728, 653, 469, 1138, 1127, 757, 1108, 761, 2167, 1286, 1818, 2465, 123, 250, 556]. Executions [1406]. executive [361]. existence [858]. existing [1990, 2357, 1366]. EXODUS [485, 779, 385, 389]. expected [201]. Expensive [1018, 874, 1286]. Experience [2261, 2126, 2850, 792, 735, 241]. Experiences [876, 2724, 1465, 2849, 1845, 382, 953, 211, 2478, 2110, 1517, 2373, 135, 1907]. Experiencing [1608]. Experiment [1437, 2286]. experimental [203, 397, 237, 482]. experiments [737, 2089]. Expert [316, 2163, 218]. Explanations [2637, 2636, 2705, 2704]. explicit [717]. Exploiting [1439, 156, 534, 2196, 1849, 758, 132, 1232, 2175]. Exploration [2760, 2541, 1395, 1338, 1566, 1420, 1250]. Exploratory [2540, 2510, 2561, 1483, 2592]. ExploreDB [2510, 2561, 2592]. Exploring [2314, 2400, 2527, 2564, 1897, 2081, 1856, 2486]. Exposing [2014]. Expressing [1733, 1754, 2165]. Expressions [1000, 266, 1285, 267]. Expressive [2708, 2735, 2736, 2669, 2670, 499, 515, 296, 572]. Extend [1624]. Extended [85, 992, 1441, 1564, 1759, 356, 1193, 679, 2004, 447, 542, 796, 277, 318, 884]. extender [690]. extendible [552, 253]. Extending [842, 510, 806, 2604, 585, 412, 560, 572, 2207, 674]. Extensibility [1515]. Extensible [340, 624, 537, 1704, 1562, 760, 1726, 1729, 1473, 1038, 128, 2078, 1058]. Extensible/rule [760]. extension [390, 814, 494, 192]. Extensions [1034, 1646]. External [297, 1522, 1451, 1138, 1233, 1794, 3, 576]. Extracting [1506, 654, 655, 1722, 1452, 1507, 1865]. Extraction [2556, 1366, 2326, 2331, 2325, 287, 2084, 2328, 2327, 2324, 1947, 1153, 2332]. Extremely [2460]. Eyes [2505].

F [2028, 100, 2122, 514]. F-logic [514]. Fabric [2538]. Face [1536]. FaceKit [732]. Facilities [136, 1016, 257, 75, 872]. Facility [67, 425, 750]. Factorized [2581, 2851]. factors [433]. facts [2418, 2212]. faculty [2010, 2301]. fad [631]. Fagin [839, 2619, 840]. fails [1205]. Fair [2734, 2733]. Fairness [2698, 2699]. Faloutsos [2193]. False [2676]. Family [2712, 2204, 906]. fan [2293]. Far [1731]. Farallones [2313]. Fast [1623, 784, 1370, 1025, 2570, 1761, 2307, 1806, 1117, 397, 246, 291, 1196, 1850]. Fast-Start [1850]. Faster [2831, 2705, 2704]. FastMap [1117]. FastSort [576]. Fault [1134, 1255, 1864, 1948, 1850, 2499, 786, 1150]. Fault-tolerant [1255, 1864, 1948, 2499, 1150]. faults [1898]. favorite [2433]. FDM [319]. FDT [23]. feasibility [2502]. Features [1322, 1039, 1321, 647, 2447, 902, 1184, 1232, 2134]. Federated [1457, 1676, 1059, 2759, 1604, 1994, 1775, 1179, 715, 870]. Federation [1414]. Feedback [1002, 404]. feel [2476]. fellow [2120, 2156]. Fensel [1977]. Fetched [1451]. Fetches [1003]. fever [1926]. Fewer [2831]. field [2255, 2294]. fields [106]. fifth [2048]. figures [2418]. File [1010, 1140, 1136, 124, 41, 394, 449, 524, 1184, 685, 45]. FileNet [1543]. Files [1015, 287, 450, 872, 882, 880, 603, 299, 346]. Filling [2533, 1785]. Filtering [1806, 2023, 2065]. Filters [2673, 2674, 2200]. Final [38, 222, 621, 1069]. Finally [1044]. financial [1691]. find [1145, 311]. find-statements [311]. Finding [2778, 1714, 1738, 2527, 2777, 2247, 1201]. Fine [1020]. Fine-Grained [1020]. Finite [5, 6, 1003]. FinTime [1691]. Firm [1409]. First [722, 2599, 1678, 1994, 2338, 2493, 2127, 1993, 2126, 2191, 2227, 2813, 2136, 2475, 1239, 1323, 2270, 2537, 2459, 2289, 2277, 2685, 2510, 2254, 2341, 2512, 2380, 2410, 2460, 2656, 2070, 2684, 2342, 2466, 2213, 2279]. First-Class [1678]. fits [1155]. Five [1706, 1460, 397]. Five-Minute [1460]. fix [1205, 2100]. fixed [1810]. fixed-size [1810]. Flash [2410, 2448, 2420]. Flash-based [2410, 2448]. FlashDB [2410]. flat [356]. flawed [2320]. flaws [1279]. Flexible [2810, 1793, 2538, 683, 994, 2381, 773, 861, 2048]. floating [106]. Flocks [1482]. Florida [1202]. Flow [2229, 2780, 2779, 2770, 128, 1786].

FlowBack [1554]. flows [905]. Fly [2541]. following [2319]. food [1780]. Footprint [1659]. FORAL [87]. Force [29]. forces [2176]. forecast [2437]. Foreign [2841, 2842]. Forensics [2853, 1901]. Foreword [1448, 1335]. Forgetful [2544]. form [1140, 722, 384, 366]. Formal [1046, 742, 481, 2594, 360, 1882, 2182]. formalization [2479]. format [3]. formats [1145, 2208]. formerly [1617]. Formidable [2803]. Forms [2533, 242, 4]. formulas [474]. Forth [1450]. FORUM [2381, 23]. Forward [2583, 590, 2115]. foundation [2182]. FoundationDB [2771, 2772]. Foundations [2521, 1463, 2723, 1957, 2680, 2346, 76]. Four [564, 2427, 2138, 512]. Fourth [2428, 1673, 57, 1685, 1946, 2626, 2278, 1793, 2467]. Fourth-Generation [1673]. FoW [2710]. FQAS [1793, 2048]. FQP [2538]. fractals [2193]. Fragmentation [2515, 2335]. Fragments [228]. Framework [1733, 2854, 2726, 2659, 2725, 2556, 581, 1389, 1410, 1092, 2629, 1159, 1855, 2628, 2660, 1972, 2170, 2236, 586, 808, 2078, 1291, 1339, 447, 2349, 180, 1198, 337, 458]. Frameworks [2746, 731, 693]. Francisco [2313]. Franklin [2683]. freaks [2085]. Free [2795, 1549, 1709, 1283]. freely [595]. freely-reorderable [595]. Freire [2764]. Frequent [1708, 1631]. Frequently [2256]. Frequently-asked [2256]. Freshness [1718]. friendly [606]. fringe [2294]. front [2477, 279]. front-end [279]. Frontier [2557]. Fry [98]. Full [763, 2429]. full-text [2429]. Fully [687, 1872, 1989]. Function [670, 339]. Functional [2780, 129, 2779, 234, 741, 676, 1108, 282, 462, 432, 635]. Functionality [1471, 858, 1914, 1470]. Functions [1514, 2601, 2600, 446, 1784, 436, 110]. Fundamental [1253, 470]. Fundamentals [2004]. Funding [730, 846, 949, 987, 1204, 1247, 1940, 2244, 2441]. further [636].

Fusionplex [2116]. Future

[26, 2864, 496, 1665, 2754, 949, 2406, 631, 64, 138, 627, 1926, 2193]. fuzziness [2207]. Fuzzy [1472, 1950]. FZI [1446].

G [2122]. Gaia [458]. Gain

[2775, 2776, 2133]. gains [683]. Games [2742, 2741, 2264, 1546]. Gamma [478]. ’Gap [2304, 2252]. Garbage [1021, 1396, 1282, 503, 865]. Garcia [1959]. Garcia-Molina [1959]. Garlic [1314]. gathering [2038]. GDPR [2719]. GEM [278, 247]. GenAlg [2077]. gene [2074]. genealogical [2414]. general [71, 807, 519, 2413, 808, 1110, 1252, 2249, 291, 1883, 698, 2248]. general-purpose [2413, 2249, 291]. Generalization [1482]. Generalized [1714, 1375, 93, 1484, 1285, 466, 241]. Generalizing [1392]. Generating [1823, 1699, 1768, 1009]. Generation [1328, 2703, 1449, 1673, 1708, 1522, 2355, 2702, 616, 2079, 1327, 1869, 887, 1685, 733, 1916, 1159, 1166, 1525, 1776, 1089]. generation/OODBMS [727]. generations [2116]. generator [385, 2249]. generic [1987]. Genesis [313]. GENESYS [1060]. genomic [2077]. Geo [2070, 1398]. Geo-Spatial [1398]. Geographic [1441, 703, 1344, 1343, 888, 1935, 712]. Geometry [1652]. Geometry-based [1652]. GeoMiner [1430]. geophysical [241]. Georg [2260]. George [1461, 1943]. Geospatial [1440]. Gerhard [2348]. German [457]. Germany [2348]. Gerome [2243]. GESTALT [499]. Get [1560, 2085, 2230, 2275]. getting [1959]. GHM [2248]. GI [2405]. GI-Dagstuhl [2405]. gigabytes [2088]. Ginsburg [2128]. GIOSS [999]. GIS [1361, 931, 979, 980, 1183, 2045]. gist [739]. GitHub [2789]. GIUKU [739]. give [2071]. giving [2180]. Glass [2789]. Global [1581, 1815, 1600, 341, 1912, 852, 2232, 823, 889, 2025]. globally [245]. Glossary [975, 974, 818]. Glue [862, 675]. Glue-Nail [862, 675]. go [1873, 2348]. Goal [1280]. Goal-oriented [1280]. GOAT [2074]. goes [936]. Goetz [2716]. Going [1450, 1755, 224]. golden [1834]. Good [346, 2476, 1813, 1949, 2180, 1940, 2294, 918, 734]. Google [2285, 2409]. gossiping [1972]. got [2284]. Gottlob [2260]. GPU [2532]. graduate [61, 2397, 2409]. Graefe [2716]. Grained [1020]. Grammar [432, 526]. grammar-based [526]. Grammar-like [432]. grand [2348]. Graph [2708, 2652, 2821, 2603, 2822, 569, 2602, 954, 2496, 2623, 2816, 2508, 2498, 918, 565, 2011, 2434]. Graph- [2603, 2602]. Graph-Based [954]. Graph-Oriented [569, 918]. Graphical [952, 1056, 737, 570, 606, 1986, 400, 261, 608, 739, 543]. Graphics [2508, 260]. GraphLog [605]. Graphs [2838, 2289, 2844, 2551, 2523, 2837, 498, 595]. Gray [2397, 2815, 2307, 344, 2299, 2305, 2306, 2302, 2303, 1991, 2294, 2445]. great [2243]. Greece [2371, 2648]. GREO [903]. Grid [2108, 1828, 394, 450, 882, 2026, 2055, 2168, 2166, 2017, 1975]. grid-based [2055]. Ground [2564, 890]. Group [1749, 2644, 2336, 32, 62, 982, 2518, 2587, 2524, 2614, 2689, 1220, 55, 2500, 222, 1362, 444, 559, 981, 2446, 1101, 1219, 2226, 2284, 2348, 2454, 1462, 2424, 67, 2620, 1579]. Group-By [1749]. groupware [539]. Growing [1012, 1011]. Guarantee [1066]. Guarantees [2661, 2662]. Guard [2298]. Guest [2015, 2159, 1881, 1910, 1335]. GUI [1294]. guidance [795]. Guide [2594, 2675, 2027, 2074]. guided [604]. Guidelines [1316]. Gulf [2313].

H [100, 98, 2088, 2122, 774, 2586]. H-trees [774]. Haas [2408]. Hadoop [2677]. Hagen [1101]. Han [1942, 2423]. HANA [2425]. Handling [697, 2780, 378, 2779, 2444, 1759, 2092, 1206, 538, 698]. handwritten [1116]. handyman [971]. Hank [2463]. Happening [1667, 37]. Hard [573, 1751, 2524, 2764, 508, 2230]. HardBD [2554]. harder [1959]. Hardware [2238, 2554, 2572, 2543, 2687, 2522, 903, 291]. Hardware-Conscious [2687]. Hardware-driven [2522]. harnessing [2005]. Harvard [2536]. Hash [2738, 1004, 1271, 2744, 2743, 2737, 854, 603, 853, 1118]. hash-based [1118]. Hash-Joins [1271]. Hashing [451, 344, 298, 297, 880, 299, 262]. Haystack [2527, 1804, 2351]. Healey [1943]. Health [2685, 2713, 2757, 2262]. healthcare [2170]. heap [503, 865]. Hector [1959]. Heisenbugs [2156]. held [1362, 620]. Helix [2527]. help [1974, 2401]. helped [2401]. Helping [937]. her [2385]. heritage [135]. Heterogeneity [944, 1795, 717, 709, 721, 712]. Heterogeneous [1423, 1645, 1499, 836, 1092, 170, 2573, 1498, 1532, 1614, 1455, 720, 713, 1182, 646, 1198, 620, 1176, 689, 2210, 724]. heuristics [536]. HG [1612]. HG-trees [1612]. Hidden [2533, 1802]. hiding [1265]. Hierarchical [2248, 1803, 1989, 127, 526, 90, 395, 2051, 100]. hierarchically [1292]. Hierarchies [1700, 156, 487, 159, 356]. Hierarchy [1484, 509]. High [1714, 1490, 1390, 1529, 477, 1400, 1547, 1144, 769, 1401, 1548, 1728, 1711, 73, 1544, 1562, 467, 1202, 2770, 2255, 1023, 1799, 1784, 85, 787, 1828, 511, 1813, 2202, 86, 799, 380, 90, 170, 2448, 1829, 381, 238, 805, 1934, 678, 706, 310, 1870, 2409]. High-Availability [1728, 511]. High-Dimensional [1400, 1401, 1529, 1828, 1813, 1829, 1934]. high-end [706]. high-impact [2409]. high-level [90, 678, 310]. High-Performance [1390, 1562, 381]. High-Scale [1548]. High-Speed [2770]. high-throughput [2448]. higher [514]. higher-order [514]. Highlights [2657]. Highly [1374, 1842]. HIKM [2262].

Hilbert [1785]. him [2335, 2401]. Hints [2747]. HiPAC [420]. Hipp [2678]. Histogram [1519, 1814, 1813, 1123]. Histograms [1633, 1520, 1815, 433, 1275, 2244]. historical [302, 528]. History [1476, 2244]. Hoc [1394, 2819, 2820, 2001, 2056, 3]. HODFA [1092]. hold [891]. Holding [2724]. holistic [2111]. home [2294]. Homogenizing [1092]. Homoiconicity [2814]. Honey [1659]. Hong [2796, 1221]. Hongjun [2144]. hope [622]. Hopkins [2390]. Horizontal [206, 232]. horizontally [289]. hostage [2267]. Hot [1265, 2091]. HP [728]. HPCC [965, 986]. HPI [2587, 2817]. Huanchen [2765]. hub [1913]. huh [2812]. Hull [2597]. Human [368, 2534]. Human-Powered [2534]. Humans [2710, 2667, 2668]. husband [2230]. Hy [919]. Hybrid [2766, 2349]. HYDRO [646]. Hygraph [919]. Hygraph-based [919]. HyperBase [953]. Hypergraph [1129]. hyperinflation [2409]. Hyperion [2021]. Hyperlinks [1508]. HyperStorM [1303]. hypertabular [1341]. Hypertext [1508, 1664, 953, 969, 334]. Hypothetical [1389].

I/O [2805, 2388, 1266]. I3 [1362]. I3/POB [1362]. Ian [2088]. IBM [704, 2509, 2305, 902, 951, 2120, 2156]. ICDE [2240]. ICDT [2379, 2724]. Iceberg [1647, 1796, 1798]. Iceberg-cube [1798]. icons [335, 1175]. icons-based [335]. Idea [2806, 1565, 1250]. ideas [86, 2294]. identification [2079, 563, 490, 1901]. Identify [1622]. Identifying [1716, 2399]. identity [516, 699, 711, 660]. idf} [2257, 2280]. IDISKs [1575]. IDLOG [572]. Idreos [2598]. IDS [1896]. IEEE [2229, 1466, 2129, 2191]. if [1513]. IFCIS [1481]. IFIP [60, 2340]. IFO [1189, 1093]. ignore [2353]. II [1953, 690, 87]. IIT [2253, 2464]. iiWAS [2263]. IIWeb [2049].

IIWeb-03 [2049]. Illinois [1965, 620, 1964]. Illustra [1245]. IM [805]. Image [1082, 1421, 1650, 1173, 1866, 1364]. Images [1083, 1758, 969, 2088, 2080]. imaginary [106]. Imminent [2520]. immortality [2308]. Impact [2211, 2221, 2653, 1070, 1847, 2409]. Imperative [2780, 2779, 2373]. Implementation [2682, 807, 1590, 2261, 252, 1091, 427, 954, 2126, 997, 382, 218, 1398, 864, 484, 278, 320, 104, 399, 448, 798, 920, 120, 862, 1806, 552, 1269, 2232, 1190, 121, 388, 389, 1287, 253, 429, 1177, 1770]. Implementations [2595, 1216, 1924]. Implementing [578, 1669, 1389, 1267, 1095, 1119, 547, 82, 595]. Implication [1391, 565, 518]. Implications [1030, 2532, 1511, 290, 1154, 2103, 661, 119, 462, 241]. Implicit [2575, 2788, 2574]. importance [2085, 2230, 2244, 2415, 2433]. imprecision [626]. IMPRESS [1046]. Impressions [14]. Improve [1718, 777, 206, 1232]. Improved [2643, 1373, 1715, 1275]. Improving [1816, 1510, 2399, 786, 1234, 1357]. IMS [51]. In-context [2023]. in-depth [60]. In-memory [2543, 2448]. in-network [1952]. INA [607]. inapplicable [367, 370]. Inaugural [2807]. Inc. [1064]. include [2284]. including [656]. Inclusion [2797, 2848, 565]. Inclusive [2761]. Incomplete [244, 673, 1000, 612, 348, 437]. Incompleteness [175]. Inconsistent [2680]. Incorporating [509, 1812]. Incremental [1702, 1131, 2576, 1502, 2656, 1719, 1337, 1014, 651, 890, 865, 1377]. incrementally [863]. Independence [33, 1813, 2018, 235, 1689, 7]. independent [575, 265, 674]. Index [851, 1512, 1553, 1401, 2827, 1757, 1003, 1473, 1259, 1861, 1529, 1810, 1513, 1816, 2008, 770, 256, 1230, 1229, 1808, 2448, 269, 1829, 2420, 774, 805, 1800]. indexed [1785]. Indexes [1373, 2675, 2828, 2202, 658, 2098, 804, 554]. Indexing [1651, 1758, 1400, 1741, 1165, 1603, 990, 1742, 1671, 1743, 1736, 1402, 1316, 1116, 2468, 1117, 1877, 1809, 658, 2358, 2088, 1115]. India [2120]. Indian [1243]. Indiana [705, 2151]. indicator [2469]. Indices [1402, 1374, 1191]. indirect [2110]. induction [1804]. Industrial [2114, 1616, 1070, 1926, 2100]. Industrial-strength [2114]. Industrial/Academic [1616]. Industry [2204, 1459, 2816, 2498, 2097, 917, 912, 2303, 2409, 2441]. INDY [120, 121]. INEX [2233]. inexpensive [442]. inference [679, 403, 506]. Inferring [1454, 1453]. infinite [517]. Influence [1725]. Influencing [983]. Influential [2845, 2855, 1703, 2802, 2791, 2811, 1573, 1595, 1680, 1615, 1790, 1892, 1929, 1961, 1992, 2124, 2086, 2101, 2140, 2155, 2177, 2194, 1474]. InfoHarness [1176]. infological [87]. Infomaster [1426]. informatics [2253, 2457, 2226]. Information [1235, 2181, 38, 1581, 1481, 2435, 1386, 2645, 2216, 1334, 2354, 1645, 1593, 1676, 836, 10, 1462, 2325, 2583, 2751, 2556, 558, 2323, 1608, 1426, 2470, 1181, 1904, 945, 1461, 1100, 1492, 2049, 1672, 2523, 1521, 934, 1600, 1451, 1044, 1452, 2587, 2544, 1604, 2622, 1072, 57, 2213, 1602, 2006, 1605, 1428, 2262, 2339, 1201, 1356, 1338, 935, 1455, 1920, 286, 703, 995, 217, 1890, 81, 158, 1974, 1931, 1792, 160, 1986, 1292, 367, 370, 1965, 1368, 134, 1210, 1685, 1950, 1969, 47, 361, 2183, 309, 2084, 1355, 1354, 2328, 2077, 1775, 2127, 244, 348, 1336, 2327]. information [472, 2486, 710, 1339, 708, 692, 314, 2405, 2024, 1358, 2324, 1993, 437, 176, 2444, 915, 2092, 2116, 889, 1943, 2023, 2446, 1156, 1776, 1786, 1369, 1359, 1983, 324, 413, 277, 1903, 1346, 1345, 1909, 1980, 1176, 2065, 910, 860, 1357, 1771, 2255, 2332, 897, 1765, 304, 698, 612, 2069]. Informix [770, 1162, 1759]. InfoSleuth [1386, 1427, 1607, 1151]. Infosphere [1786]. Infrastructure [2810, 972, 1919, 1945, 973].

Infrastructures [1764, 1920]. INGRES [806, 427, 2417]. Inheritance [325, 159, 534, 514, 664, 660]. Initial [1493, 2057, 545, 2462]. initiative [2237]. ink [1116]. Inktomi [1547]. Innovative [2475]. INODE [2760]. input [110, 576, 571]. input-output [110]. input/output [571]. Inputs [1618]. INRIA [622]. ins [1296, 1297]. inserting [346]. Insight [2735, 2736]. insights [2448]. Inspired [2544]. Installations [2623]. Instance [2841, 2842, 575]. Instance-optimal [2841, 2842]. Instant [2542, 855]. instantiating [1888]. Institute [2390, 1243, 2226, 2435, 2454]. institutional [1330, 1353]. Institutions [2793]. Instrumental [928]. integrate [1989]. Integrated [460, 1395, 1543, 153, 356, 545, 314, 1420, 491, 50]. Integrating [1356, 1240, 2810, 589, 160, 1085, 165, 1451, 1228, 1511, 2160, 935, 1158, 2077, 2378]. Integration [154, 1386, 1606, 1720, 1499, 1426, 2215, 2102, 2417, 1642, 1600, 1618, 1168, 2817, 2322, 2204, 2021, 1920, 2381, 2481, 2174, 360, 477, 1981, 1878, 1921, 1917, 2347, 2109, 1846, 2430, 558, 834, 2013, 1291, 2049, 1182, 2116, 2118, 2115, 917, 951, 1843, 897, 2353]. Integrity [294, 510, 541, 852, 464, 1288, 280, 253, 1126]. Intel [2274, 2465]. IntelligencE [2449, 1591, 2485, 2555, 2528, 2627, 140, 2493, 1944, 2384]. Intelligent [1455, 1084, 1575, 897, 510, 739, 106, 122]. Intensive [2390, 1611, 1446, 2721, 2455, 550, 440, 157, 2167, 2403, 1907, 855, 963]. inter [758, 915]. inter-operability [915]. inter-operation [758]. Interaction [1492, 2096]. Interactive [1306, 1033, 1250, 153, 365, 1313, 158, 41, 5, 1170, 739, 6, 260]. InterBase [923, 895]. interchange [723, 1422]. interdependencies [912]. Interdisciplinary [2684]. Interests [36]. Interface [952, 1421, 2770, 1136, 570, 606, 604, 1192, 261, 427, 309, 608, 922, 49, 310]. Interfaces [2638, 1399, 569, 2721, 350, 735, 737, 2376, 1102, 2207, 1337]. interfering [653]. Interim [62]. internal [1284, 3]. internals [2014, 933]. International [2238, 2813, 2370, 2136, 2287, 2862, 2863, 2866, 2867, 2868, 2870, 2871, 1568, 2278, 1239, 1323, 2203, 2216, 1569, 1793, 2869, 2261, 2485, 2555, 2354, 2270, 2537, 2229, 2394, 2323, 2865, 2215, 2289, 2277, 2685, 2713, 2757, 2510, 2561, 2153, 2254, 2372, 2511, 2599, 2426, 2380, 2410, 2467, 2656, 2291, 2466, 2494, 2528, 2553, 1202, 2592, 2213, 1215, 2191, 2239, 1765, 2340, 1994, 2139, 2493, 2458, 1685, 1946, 1993, 2046, 2048, 2227, 2228, 2554, 2449, 838, 2154, 910, 2339]. Internet [1994, 2537, 1556, 1449, 1591, 1760, 1740, 1387, 936]. Interoperability [1581, 618, 1440, 711, 1601, 1607, 1608, 1425, 911, 2340, 2170, 913, 647, 812]. Interoperable [1200, 1309, 716]. Interoperation [944]. interpolation [656]. interpreted [497]. interpreted-compiled [497]. interpreter [483]. Interpreting [884]. Interrupted [1712]. INtersect [917]. intertask [914]. Interval [1408, 658]. intervals [2150]. interview [1908, 2142]. Interviewing [1963]. Introducing [2083, 2451]. Introduction [2109, 1583, 1956, 1881, 2015, 1223, 326, 1910, 2159, 1599, 1769, 66, 1196]. Intrusion [2516, 1897, 1895, 1896]. intrusions [1898]. invalidation [1778]. Invariant [2692, 2693]. Invariants [1485]. inversion [2359, 124]. Inverted [1014, 41]. Investigating [1936]. Investigation [1379]. invitational [973]. involved [2284]. Ioannidis [2244]. IOT [2537]. IP [1926]. IPAW’06 [2228]. IPSI [747]. IQIS [2213]. IR [2190]. IRCDL2021 [2751]. Iris [609]. ISBN [2123, 2088, 2122]. ISDO [1764]. ISIS [309]. isn’t [1940]. isolation [1103, 2095, 2375]. ISS [559]. issue [1895, 2109, 1197]. Issues [285, 1457, 1996, 1603, 1319, 2093, 892, 909, 1661, 1698, 185, 417, 1543, 1071, 1765, 991, 886, 2502, 1351, 2082, 1350, 445, 797, 1917, 1065, 1910, 2001, 625, 2132, 634, 2198, 2145, 626, 2019, 2091, 2266, 1926, 2423, 2463]. Italian [2751, 111]. ITASCA [764]. items [868, 2056]. Itemset [1391]. iterative [342]. Iteratively [2564]. ITISA [2684]. ITRI [1073]. IWEI [2340].

J [98, 1943, 566, 66]. Jagadish [2586]. James [1943]. Jan [1190]. Jannink [1190]. Jasmin [274]. Jason [1943]. Java [2365, 1469, 1347, 1528, 1693, 2087, 1907, 1531, 1924]. JedAI [2688]. Jeff [2472]. Jeffrey [2085]. Jennifer [2230]. Jiawei [1942, 2423]. Jim [2397, 2445, 2815, 2307, 2299, 2301, 2305, 2306, 2302, 2303, 1991, 2294]. JISBD [2152]. Job [1094, 1963]. John [2027, 2129, 2176]. Johns [2390]. Johnson [2445]. Join [1640, 1138, 1723, 1502, 2714, 671, 2839, 1397, 2594, 2607, 1618, 2199, 2840, 1719, 1737, 2571, 1129, 1828, 266, 231, 740, 2274, 1273, 597, 369, 462, 2098, 2489, 2487, 1191, 1272, 407, 201, 512, 859, 1077, 431, 536, 725, 1251, 2503, 1113, 2176]. join-order [1251]. Joins [1005, 1639, 1641, 1004, 2570, 1006, 1271, 596, 871, 510, 93, 854, 1984, 1191, 853, 855]. Journal [2588, 2012]. journals [2188, 2440]. Joy [2756]. Juggling [2601, 2600]. Juliana [2764]. Jumping [1074]. June [2860]. Jungle [1593]. just [2307, 2812, 1855].

K-d [305]. Kabiria [751]. Kaleidoscope [604]. Kamber [1942]. Kanata [2113]. Kardamom [419]. Karlsruhe [1446]. Karpinski [18]. Kauffman [2088]. Kaufmann [2044, 2045]. KBMS [1177]. Kenneth [2641]. kernel [120, 257, 121, 388]. Key [2771, 1678, 2772, 634, 1997, 269, 238, 84, 2052]. Keynote [749, 1147, 1148, 642]. Keys [2841, 2842, 2567, 1810, 299]. Keyword [2590, 2507, 2399]. Kian [2524]. Kian-Lee [2524]. Kim [1064]. Kimball [2027]. kind [394]. Kleisli [1567]. know [1149, 2176]. Knowledge [1569, 2613, 2576, 1057, 2300, 2844, 2123, 2517, 1773, 2577, 190, 2627, 2262, 525, 1455, 1977, 547, 2002, 677, 1221, 288, 1788, 1306, 347, 472, 2486, 2329, 171, 608, 175, 1888, 403, 584, 455, 482, 280, 289, 678, 924, 1967, 1177, 721, 198, 139, 2217]. knowledge-base [403]. Knowledge-based [190, 1788, 1888, 289]. knowledgeable [739]. Known [1597, 2846, 1617, 2397]. Kong [1221, 2796]. Kooi [966]. Korea [2267]. Korth [2463]. Koutris [2642]. KRDB’98 [1569]. KRISYS [1177]. KSR1 [904]. KTELO [2659, 2660]. Kuno [2122]. Kùzu [2856]. Kweelt [1855]. Kyu [2267]. Kyu-Young [2267].

L [1944, 1943]. Lab [1565, 2217, 1909, 2796, 2509]. Laboratories [845]. Laboratory [2536]. Labs [728, 1926, 2100, 2416]. Laender [2409]. Land [2060]. Language [2575, 2858, 2708, 607, 1657, 2686, 2637, 1597, 1438, 1046, 1673, 2735, 2574, 2636, 2667, 647, 1050, 2565, 2668, 2564, 2736, 976, 2755, 969, 516, 505, 71, 144, 561, 1106, 1293, 1469, 485, 436, 1986, 400, 765, 426, 483, 1693, 335, 983, 165, 2077, 90, 72, 514, 967, 1049, 922, 1269, 664, 208, 3, 506, 283, 2452, 1935, 814, 143, 87, 663, 678, 1069, 310, 494, 1957, 91, 813, 247, 70, 551]. Languages [1463, 2203, 1706, 1720, 2752, 1425, 2640, 154, 2233, 676, 2438, 85, 325, 1362, 1879, 1284, 2383, 1107, 1108, 9, 86, 5, 1772, 767, 78, 2, 1, 49, 8, 1363, 185, 6, 318, 572, 320, 2353, 2434]. Large [2858, 2788, 1533, 1488, 1480, 1665, 1394, 2355, 1395, 1396, 2341, 2371, 1638, 2460, 1743, 1744, 1710, 1602, 1759, 2689, 1865, 525, 868, 1076, 791, 105, 354, 547, 2428, 2343, 541, 2039, 103, 468, 2111, 1114, 783, 597, 1180, 1809, 865, 1420, 804, 430, 1692, 2112, 484, 1248, 431, 536, 1257]. Large-Scale [1602, 2039, 1692]. lasting [1847]. Later [1460]. Latin [2275]. Laura [2408]. Laws [1723, 2193]. Layer [2551]. Layered [1658]. layers [2250]. Lazy [881]. LCE [459]. Learned [2724, 2849, 2846, 2768, 2767, 2110]. Learning [2805, 2774, 2609, 2013, 2608, 2701, 2773, 2639, 2809, 2582, 1751, 1839, 1924, 2224]. least [2502, 2319]. led [2335]. Lee [2524]. Left [661, 523]. left- [523]. Left-deep [661]. Legacy [1092, 909, 2173]. Leis [2750]. length [428]. Lens [2838, 2837]. less [492]. Lessons [2724, 2849, 261, 2717, 1412, 2110, 1692]. Letter [18, 566]. level [85, 2236, 477, 1990, 86, 1306, 679, 90, 170, 781, 73, 180, 2330, 2677, 87, 678, 643, 310]. levels [1103]. Leveraging [1156]. LH [880]. LH* [1728]. LH*/RS [1728]. Libraries [1593, 2751, 2341, 2371, 2691, 1604, 1787, 2428, 1788, 1361]. Library [2751, 1213, 1201]. LID [2363]. lies [2219]. Life [2071, 2215, 2081, 1891, 2072, 2309, 792, 1959, 2120, 2156, 2268, 2335]. Lifecycle [2639, 2809]. Lifestreams [1241]. light [1186]. lightweight [1305]. like [1555, 971, 1849, 432, 1932, 2546]. Limit [2686]. Limited [1643, 1518]. Lindsay [2156]. Line [1558, 1592, 1480, 1024, 1259, 1361, 783, 1713, 1711, 1403, 801, 1258]. Linear [1741, 2609, 2608, 599, 2633, 2632, 880, 523, 402]. link [1246]. linkage [498]. linkage-graphs [498]. linkages [1812]. linked [2081, 2498]. Linking [2512]. LISP [123]. Lists [1014]. literature [48, 2222, 52, 2353]. Live [2724]. living [261]. LiXQuery [2182]. LLO [664]. Load [332, 1013, 2369, 1864]. load-balancing [1864]. Loading [1087, 869, 882]. Loads [1712]. Local [1716, 852, 330]. Locally [1809, 332]. Locating [2507]. Lochovsky [100]. lock [772, 2059, 479]. lock-based [2059]. Locking [826, 1094, 771, 237, 773]. LOF [1716]. Log [2657, 1711, 2658, 780].

Logging [1624, 379, 866, 805]. Logic [2638, 2363, 2084, 2591, 362, 2316, 339, 564, 282, 2373, 827, 514, 585, 465, 2272, 277, 573, 650, 475, 698]. Logic-based [2638, 2084]. logic-programming [362]. logic-programming/object-oriented [362]. Logical [2346, 1624, 1711, 411, 377, 2412, 1990, 167, 296, 856, 183, 320, 337, 1694, 901]. LOGRES [929]. logs [2432]. Logspace [2703, 2702]. Lomet [2432]. London [2108]. Long [2795, 1489, 2544, 587]. long-running [587]. Long-Term [2795, 2544]. Loo [2268]. Look [2583, 1093, 2595, 1189, 2271, 2307, 1982, 2087, 59]. Looking [1633, 2583, 2789, 2415]. loophole [1997]. loops [767]. loosely [1105, 270]. LORE [1305, 1445]. loss [777, 2296]. Lots [1870]. Lotus [1663, 1141]. Low [2522, 477]. Low-energy [2522]. low-level [477]. lower [2489]. LRU [877, 1003]. LRU-K [877]. LSDIS [1909]. Lu [2144]. luck [2230]. lunatic [2294]. Lyngbaek [707]. LyriC [1106].

M [1943]. MA [2860]. MAC [2037]. Machiavelli [506]. Machine [2805, 1047, 2583, 2609, 2608, 2701, 1008, 2639, 2809, 2350, 2627, 131, 478, 1839, 134, 274, 130, 776, 109, 1924, 419]. machine-learning [1839]. Machines [2710, 292, 513]. Machiraju [2122]. macro [635]. macro-databases [635]. made [366]. Madman [130]. Magic [591, 997, 376, 1287]. Magic-sets [997]. magnitude [1163]. mail [272, 1901]. Main [1670, 1810, 956, 1562, 1748, 2535, 784, 354, 252, 1808, 345, 1232, 2267]. Main-Memory [1562, 1810, 1232, 2267]. Maine [703]. Mainstream [1668]. maintainability [461]. Maintaining [1382, 863, 1289]. Maintenance [1405, 1404, 1702, 1431, 1752, 1378, 1403, 1719, 275, 1872, 1859, 1290, 1131, 1132, 1822, 1288, 1211, 1130]. Make [178].

Maker [2274]. Making [2710, 281, 590, 1600, 2768, 2767, 2568, 1748, 2747, 2641, 2824, 2236, 1949, 361]. Manage [2689, 2357]. Management [2654, 2593, 2238, 2370, 1328, 62, 2862, 2863, 2866, 2867, 2868, 2870, 2871, 1322, 2435, 1423, 2278, 2843, 2645, 1061, 2216, 1047, 2869, 2354, 1465, 2554, 2788, 1534, 968, 1438, 1429, 2323, 2865, 2790, 2672, 837, 1437, 945, 1588, 1492, 2722, 2108, 1548, 1554, 2839, 896, 1476, 2635, 2685, 2713, 2757, 1648, 2557, 1522, 2511, 1445, 1550, 2380, 2467, 2517, 2544, 1661, 2545, 2840, 2466, 2494, 2553, 2546, 2558, 1359, 1698, 2856, 2809, 1448, 2634, 2154, 1052, 959, 221, 1543, 1436, 2262, 2671, 2279, 2562, 1321, 1356, 2237, 1120, 525, 2390, 666, 331, 991, 2010]. management [101, 350, 1977, 469, 1781, 1099, 2326, 336, 112, 588, 1280, 2189, 1469, 2285, 624, 2276, 2482, 546, 667, 306, 2492, 2172, 161, 797, 1965, 1327, 778, 257, 1910, 456, 1187, 203, 2425, 1517, 361, 558, 2183, 621, 1360, 434, 1686, 971, 1996, 225, 2184, 2071, 2212, 347, 2127, 2093, 1954, 1311, 2403, 737, 2072, 549, 752, 2486, 100, 692, 568, 2145, 1867, 2391, 443, 608, 2369, 1873, 345, 2046, 652, 390, 1844, 99, 1283, 521, 1456, 1794, 1847, 98, 805, 1300, 329, 1842, 226, 2393, 210, 453, 2022, 756, 2446, 61, 1369, 1918, 49, 1333, 482, 109, 1203]. management [2115, 2217, 189, 300, 1348, 116, 744, 493, 924, 1967, 2020, 819, 958, 115, 2497, 639, 2407, 192, 1987, 198, 2100, 2414, 916, 1832, 838]. management-making [361]. Manager [1670, 1324, 1051, 1756, 454, 249, 351, 1186, 1254, 908, 2335]. managers [1199]. Managing [1212, 2075, 322, 2034, 1533, 472, 2088, 1007, 1614, 643, 1666, 791, 2325, 691, 2077, 503, 1795, 932, 2423, 2408]. managment [1983]. Manifesto [1322, 1321, 616, 1096, 2138, 727]. manipulating [1276]. manipulation [720, 90, 460, 608, 68, 8, 574]. Many [2505, 846, 2252]. many-book [2252]. map [2000]. Mapped [1023]. Mapping [1645, 1193, 383, 2506, 2225, 2008, 2110, 1860]. mappings [2359, 1859, 265, 1837]. MapReduce [2626, 2615, 2352, 2421, 2612]. MapReduce-Based [2612]. maps [657, 84]. Marcus [1943]. marginal [683]. Margy [2027]. Mario [2123]. Mariposa [1348]. Market [1651, 1590, 1392, 1391, 1593, 1766, 1919, 1963]. market-place [1919]. Marketing [1591]. marketplace [1587]. Marketplaces [2800, 2462, 1766]. Markets [2801, 1589]. marriage [2428]. Maryland [1362]. Mash [2274]. mashups [2347]. Mason [1461]. Massachusetts [748]. masses [1263]. massive [1820, 1828, 1163]. Massively [1041, 2714, 2594, 2571, 1308, 1307, 1254]. Master [2677]. Master-level [2677]. Mastering [2549]. match [551, 449]. Matching [2778, 1025, 593, 2661, 2635, 2777, 2112, 2634, 2662, 1432, 2114, 2231, 2111, 2395, 104, 1185, 2319]. Matchmaking [1605]. Materialization [1610, 1739, 410, 670]. Materialized [1861, 2682, 1822, 1288, 1468, 328, 1824, 1121, 1291, 1132]. Mathematical [1082, 68]. MATLANG [2670, 2669]. Matrix [2670, 2752, 2669]. matters [2156]. max [1364, 2226]. Max-Planck [2226]. May [2228, 2869, 34]. McHugh [1943]. MDM [1423]. Meaning [1082, 560, 723, 404]. Meaningful [1994, 1372, 2427]. means [721]. measure [2248]. measurement [2232]. measurements [2284]. measures [1877, 1796]. Measuring [2622]. mechanical [2093]. mechanism [2008, 363, 401, 291]. Mechanisms [1601, 1137, 159, 900]. MED [1962, 1682]. Medal [2129]. Media [1542, 1084, 1078, 1033, 2676, 1086, 511, 996, 1255, 793]. Medians [1518]. Mediated [1683, 1132]. Mediating [168]. Mediation [1557, 1181]. Mediator [1422, 1362, 1363, 1261, 1260]. Mediators [1556, 1497, 1654]. Medical [2482, 1072, 2457, 1364, 915, 455]. Medium [1671]. Medium-dimensionality [1671]. MEDUSA [1047, 2508]. meet [1444, 2853]. Meeting [34, 2650, 2861, 2860, 743]. Meets [2136, 2287, 33, 1569, 1545, 2698, 2713, 2757, 2372, 2786, 2699, 2191, 2582, 2785, 1455]. mega [2478]. mega-modeling [2478]. Melbourne [2361]. member [936]. Members [24, 1593]. memoir [2028]. memorial [2296]. memoriam [2179]. memories [106]. Memory [1670, 956, 2827, 1522, 2744, 1518, 2544, 1007, 1562, 1748, 2743, 2828, 2535, 2765, 2756, 2128, 784, 354, 1810, 112, 252, 406, 1808, 2448, 2420, 345, 381, 1104, 859, 966, 2543, 1232, 2267]. Memory-Efficient [2765]. memory-resident [381]. memoryless [679]. Men [1660]. Mendelzon [2179]. mentor [2301, 1436]. menu [604]. menu-guided [604]. Meral [2414]. merge [1114, 1272]. merge/purge [1114]. Merging [941, 940, 227]. Message [4, 11, 2170, 112, 1951, 337, 2120]. messages [2170]. messy [2165]. Meta [436, 1086, 205, 1082, 1387, 347, 1358]. Meta-data [1086]. Meta-database [205, 1082]. Meta-functions [436]. meta-information [1358]. meta-searching [1387]. Metadata [1083, 1079, 1080, 1085, 1084, 1081, 2108, 1078, 1466, 1771, 2100]. metadatabase [726]. metanoia [1840]. metasearch [1812]. Method [1651, 601, 993, 1715, 598, 360, 1857, 246, 1559, 238, 803, 781, 664, 1265, 805, 861, 855, 1257]. methodological [430]. methodologies [293]. Methodology [268, 1967, 2061, 654, 655, 491]. Methods [850, 1677, 1620, 1665, 2558, 2550, 806, 264, 287, 409, 2439, 1953, 1938, 1286, 1364, 824, 531, 821, 664, 2431, 773, 376, 694, 706]. Metric [1637, 1400]. metrics [103]. METU [1054, 1200, 1309]. Mexico [1959]. MGCS’07 [2289]. Mice [1660]. Michael [2043, 2530, 2683]. Michaels [101]. Michel [2045]. Micheline [1942]. Micro [2197, 2546]. Micro-Distributed [2546]. Micro-views [2197]. microcomputer [1102]. Microprocessors [107]. Microsoft [1861, 1734, 1530, 1553, 1579, 2857]. Microsoft.com [1548]. MICS [2237]. Mid [2803, 1491, 2812, 2846]. Mid-Career [2803, 2812, 2846]. Mid-Query [1491]. Middle [2564, 1805, 2384, 2172, 1325]. middle-sized [2384]. middle-ware [1805]. Middleware [1596, 1726, 1225, 2499, 1807, 1928]. Migration [2795, 2617, 874, 909]. Migration-Free [2795]. migration/cohabitation [909]. Miklau [2243]. Milano [2276]. Millions [2789]. Milo [2389]. Min [1364]. Min-max [1364]. Mind [1645]. miner [1871]. Minimal [637, 488, 615]. Minimization [1838, 2199]. minimum [1111]. Mining [868, 1055, 1490, 1745, 1489, 1591, 1664, 2146, 1660, 1430, 1393, 1480, 1708, 1630, 2289, 1804, 1472, 1899, 1549, 2517, 1539, 1483, 1943, 1715, 1744, 1709, 1511, 1710, 1484, 1248, 1746, 1482, 2520, 1563, 1901, 1897, 1895, 1942, 1117, 1249, 1306, 1301, 1871, 1891, 2403, 2198, 1845, 1118, 1836, 2266, 1896, 2157, 2058, 1926, 2193, 2275, 2335, 2423, 1924]. mining-based [1896]. MINOS [351]. Minute [1460, 406]. minutes [2385]. Miro [899]. mirrored [879]. mirroring [1265]. mirrors [878]. Miscellaneous [28]. Missing [367, 370, 697, 698]. mission [1207, 1226]. MITRA [2499]. MITRE [1203]. Mittman [101]. Mixed [1086]. Mixed-media [1086]. ML.NET [2789]. MLPerf [2805]. MLR [781]. MLS [1627, 1442, 1997]. MM [1906]. MMDB [684]. MMDBS [551]. MobiDE [2139]. Mobile [2139, 988, 2082, 1208, 989, 837, 2558, 1679, 1212, 1213, 1986, 2001, 2039, 1945, 1209, 1198, 1211, 1778]. Mobility [2746, 1659]. MOCHA [1726]. mod [1699]. mod-perl [1699]. Model [1540, 1423, 833, 1735, 569, 1441, 1093, 2672, 1082, 2557, 954, 1477, 2342, 2786, 2223, 2642, 2785, 2671, 2248, 1189, 1863, 505, 323, 575, 2476, 1820, 2461, 507, 817, 2076, 236, 485, 448, 303, 778, 163, 1193, 1241, 466, 1108, 425, 489, 2077, 679, 166, 673, 509, 435, 648, 790, 308, 708, 314, 481, 172, 296, 486, 2201, 383, 542, 1970, 2349, 382, 68, 2452, 815, 635, 560, 278, 2051, 717, 1771, 698, 72]. model-based [1820]. Modeling [321, 487, 2333, 1829, 2163, 1048, 938, 916, 2062, 286, 589, 2478, 1193, 996, 452, 563, 358, 1916, 849, 176, 2349, 501, 411, 2252, 195, 2027]. modelled [145]. Modelling [171, 2000, 1356, 377, 151, 1076, 349, 164, 174, 179, 181, 184, 1966, 191, 193]. Models [290, 2858, 1463, 2800, 1657, 1879, 2106, 1442, 1600, 2720, 622, 149, 150, 1781, 2273, 1859, 161, 2383, 1835, 2328, 875, 526, 204, 978, 1062, 479, 715, 2350, 188, 1694, 2241]. Moderator [1417]. Modern [2238, 2727, 2595, 2850, 700, 2728, 2425]. modes [404]. modification [1259]. modifications [69]. Modified [298]. Modify [1066]. Modular [540, 930, 640]. Modularity [900]. module [1245]. Modules [1206, 870]. Moffat [2088]. Mohan [2120]. molecular [2080]. Molina [1959]. MoMM [2263]. Mon [2283]. Monitoring [1863, 1318, 1551, 1833, 2461, 2283]. monitors [887, 1875]. Montage [1038]. Moodview [952]. Morgan [2044, 2088, 2045]. mosaic [1181]. mosaic-based [1181]. MOSAICO [1048]. Moshe [2205]. Mostly [1043]. Move [1994, 2747]. Moving [1735, 1780, 1736, 2234, 2134]. MP [1217]. MPEG [1937]. MPEG-7 [1937]. MPI [2721]. MQL [696]. MQSeries [1246]. MTW’06 [2241]. Much [1317, 1519, 1928, 2408]. MUD [2279]. Muddle [1596]. Multi [1516, 941, 1005, 1475, 1281, 1399, 1425, 1747, 1742, 2551, 2063, 812, 1191, 1729, 688, 1495, 1737, 1746, 2569, 1679, 2571, 940, 1127, 1192, 1919, 759, 1818, 1188, 708, 658, 269, 1785, 1830, 2201, 781, 821, 1822, 433, 455, 523, 793, 1909, 116, 643, 1987, 1113, 2275]. multi-agent [1919]. multi-attribute [759, 821]. multi-computer [116]. multi-continent [2275]. Multi-Database [1425, 708]. Multi-Dimensional [1475, 1399, 1742, 2063, 1281, 1747, 658, 1785, 433]. Multi-disk [688]. multi-domain [455]. multi-join [2571, 1113]. multi-key [269]. Multi-Layer [2551]. multi-level [781, 643]. multi-linear [523]. multi-media [793]. Multi-Objective [2569]. multi-organizational [1909]. multi-paradigm [1192, 1987]. multi-parametric [1818]. multi-policy [2201]. multi-query [1822]. Multi-Resolution [1679, 1830]. Multi-Similarity [1516]. Multi-Stage [1737]. Multi-Step [1005, 1495]. Multi-system [941, 940]. Multi-table [1191]. Multi-Terabyte [1746]. multi-user [1127]. Multi-vendor [812]. Multiattribute [344]. multicast [1538]. Multicore [2750]. Multidatabase [445, 618, 944, 1050, 1613, 912, 994, 588, 923, 1128, 713, 905, 1049, 1195, 823, 1289, 917, 721]. multidatabases [722, 794]. Multidimensional [1439, 1561, 1620, 1326, 1504, 1634, 1383, 1814, 307, 1808, 1269, 2349, 299]. multidisciplinary [2393]. multifunctional [111]. Multigranularity [1094]. Multikey [305, 262]. multilevel [583, 648, 582, 403, 613]. multilevel-secure [582]. Multimedia [1628, 1370, 1099, 1080, 306, 2039, 1033, 1308, 995, 1134, 1787, 1079, 1256, 1307, 351, 1117, 1165, 1973, 892, 1937, 1906, 1254, 1880, 2255, 1987, 352, 1563]. MultiMediaMiner [1563]. multimodel [713]. multipaging [492]. Multiple [1423, 1404, 1004, 1410, 1618, 1094, 184, 1505, 935, 1974, 1859, 258, 1306, 294, 599, 750, 674, 84]. Multiple-Data [1423]. multiple-level [1306]. Multiplex [2116]. multiprocessing [123]. multiprocessor [292, 512, 104]. MULTISAFE [202]. multitransputer [928]. Multiversion [601, 540, 653, 799, 531]. multiversioning [777]. MultiView [1312]. Multiway [2835, 2836, 255, 725]. musical [413]. must [495]. Mutually [1523]. MVS [1672]. My [2293, 2748]. MYRIAD [1059]. myself [2476]. myth [2348].

n [1913, 827]. N-queens [827]. n-tier

[1913]. NAGA [2329]. Nagano [1546]. Nail [862, 675]. name [2439]. Names [607, 724]. Namur [2050]. Nanyang [2645]. NAOS [1172]. NASA [1368, 749, 1204, 730, 822]. Natassa [2689]. National [1930, 135, 1220, 2473, 559]. Native [600, 1989, 1873]. Natural [2637, 2735, 2636, 2565, 2564, 2736, 2755, 922, 1957]. NAUDA [922]. Naughton [2085]. Navigation [1558, 888, 1868]. Navigational [2259, 2149]. NCCR [2237]. Near [2733, 2734]. Nearest [1570, 1401, 1725, 1109, 1495]. nearing [2192, 2480]. Nearly [1709]. NEC [2416]. necessary [402]. Need [1497, 702, 2100]. needle [1804]. Needles [2527]. needs [2176]. negation [446]. Negotiation [1399]. Neighbor [1401, 1725, 1495, 2733, 2734, 1109]. Neighbour [1570]. Neptune [334]. Nested [677, 528, 2150, 2104, 527, 374, 463, 392, 701, 2161, 933, 384]. nesting [246, 369]. NESTREAM [2104]. NET [1322]. NetDB [2372]. NetDB’05 [2191]. Network [836, 2775, 1629, 1586, 2776, 991, 2035, 273, 264, 1344, 1343, 2001, 165, 2018, 245, 1842, 1537, 2051, 1952, 910]. network-structured [165]. networked [1368, 371]. Networking [2504, 2372, 2191, 2268]. Networks [1390, 2278, 2708, 2727, 1063, 2770, 2728, 561, 4, 2033, 2039, 2470, 2056, 2037, 2024, 453, 2040, 2038, 2052, 1978, 1952, 2070]. Neuhold [2441]. Neumann [2129]. Neural [2708]. News [985, 24, 28, 2676]. NewSQL [2585]. Next [1328, 1705, 2549, 2557, 2355, 1525, 1529, 95, 2646, 1327, 1685, 971, 78, 1089]. Next-Generation [1328, 1327]. NF [384]. NF2 [356]. NGITS’99 [1685]. NiagaraCQ [1740]. NII [1074, 986]. nimble [1846]. Ninth [2152]. NISR [2174]. No [2775, 2133, 2776, 495]. node [2623]. NoDoSE [1506]. nomadic [886]. Non [377, 2756, 653, 1158, 206, 2163, 110, 502, 663, 107]. non-BCNF [206]. Non-deterministic [377, 663]. non-expert [2163]. non-interfering [653]. non-numeric [107]. non-relational [1158]. non-software [110]. non-uniform [502]. Non-Volatile [2756]. Nondeterministic [2697, 2696]. nonissue [2294]. nonstop [785, 476, 1042, 1217]. normal [722, 384, 366]. Normalization [2242, 1194]. Northwestern [620]. NoSQL [2396]. Notation [2506, 1473, 2165]. notational [180]. Note [31, 562, 201, 488, 207, 555, 615, 725]. Notebooks [2789]. Notes [28, 203, 19, 35, 1331, 743, 1246, 1663, 1141]. Nothing [1320, 1047, 1317, 1259, 512, 860, 2205]. Notices [39]. Novel [1016]. Novelty [2123]. NSF [622, 1616, 621, 945, 1945, 973, 1222, 1352, 620, 1346, 1345, 2650, 822, 965, 2069]. NSF-INRIA [622]. NT [1143, 1672]. NTHU [1073]. NTU [2217]. NTUA [1219, 1220]. Null [2317, 438]. Nulls [2316, 2616, 500, 2272]. Number [1012, 1011, 276]. numeric [1310, 107]. numerous [1204]. nurtured [2409]. nutshell [1985, 1998]. NVRAM [2355]. NVRAMOS [2355].

O [2805, 755, 735, 1667, 2388, 1597, 1199, 1266, 744]. O-O [1667]. OASIS [1613].

Object [516, 1328, 623, 1479, 1058, 1324, 1061, 1677, 1381, 1413, 1036, 1435, 1465, 1021, 1054, 1091, 1597, 1139, 1486, 1318, 1051, 569, 414, 1514, 553, 1574, 1037, 1668, 1034, 1713, 1421, 1396, 1477, 1035, 1024, 1048, 957, 838, 1039, 938, 1777, 1478, 429, 2506, 1242, 1023, 858, 857, 755, 399, 669, 1278, 1303, 408, 589, 436, 1171, 534, 1097, 349, 645, 798, 351, 808, 2208, 746, 1282, 765, 1327, 1192, 1276, 456, 711, 409, 1107, 1193, 1365, 930, 1199, 1108, 434, 766, 918, 165, 1186, 799, 515, 459, 790, 921, 398, 602, 670, 809, 967, 1159, 789, 486]. object [664, 856, 741, 1254, 353, 756, 1305, 1366, 479, 1312, 532, 1150, 2207, 908, 744, 924, 1279, 640, 644, 775, 898, 458, 639, 1936, 2294, 2284, 352, 916, 612, 505]. Object-Based [1421, 669]. Object-Language [1597]. object-orientation [746]. Object-Oriented [1036, 1435, 1054, 1091, 1486, 1574, 1037, 1034, 1048, 1039, 623, 1139, 414, 553, 429, 858, 857, 362, 755, 399, 1303, 589, 436, 1171, 534, 645, 808, 765, 1192, 1276, 456, 1365, 930, 434, 766, 165, 515, 459, 921, 398, 809, 967, 789, 486, 664, 856, 741, 353, 1366, 479, 1312, 532, 744, 924, 1279, 640, 898, 458, 639, 352, 612]. Object-Relational [1328, 1677, 1381, 1514, 1668, 2506, 1413, 2208, 1327, 1159, 2294]. object/task [349]. Objective [2569]. objectives [430, 210]. Objectivity [763]. Objectivity/DB [763]. Objects [668, 1463, 1164, 1735, 1084, 692, 1159, 954, 1736, 666, 321, 791, 507, 1106, 448, 645, 667, 654, 655, 733, 753, 2234, 673, 599, 563, 659, 514, 533, 1277, 371, 786, 674, 643, 660]. ObjectStore [810, 744]. Observability [2722]. observable [762]. Observations [967, 2103]. observing [749]. ocean [2314]. Oct. [1945]. October [2371]. ODBASE [1994]. ODBMS [764]. Ode [2295, 505, 570, 606]. OdeView [570, 606]. ODMG [1036, 967, 1064]. ODMG-93 [1036, 1064, 967]. odyssey [2007]. Off [1592, 2265, 424]. Off-Line [1592]. office [286, 415]. Officers [22, 1447, 1598, 1674, 1695, 1207, 1226, 1330, 1353]. offs [470, 1088]. OFL [1108]. OGDI [1440]. Oklahoma [1684]. OLAP [2291, 2239, 1415, 1367, 1326, 1946, 1376, 1827, 1700, 1935, 1694]. OLB [1597]. old [2156]. OLE [1263, 2157]. Oliver [2043]. OLTP [1709, 1379]. Olympic [1546]. OminiSearch [1857]. On-Line [1558, 1480, 1024, 1713, 1711, 1403, 1258, 1259, 1361, 801]. On-the-Fly [2541]. One [1155, 1518, 1075, 2307, 2013, 2252, 304]. one-startup [2252]. Onion [1741]. OnLine [770, 2724, 1591, 1162, 1641, 1876, 1384, 2672, 1630, 2801, 2607, 1638, 2606, 2622, 1757, 2671, 1135, 1801, 2470, 1266, 1894, 2235, 936]. Only [777, 607, 1544, 2716, 653, 321, 2095, 773, 579, 800, 320]. ONR [846]. Ontario [2859]. ontological [2412, 1919]. Ontologies [1977, 2342, 2117, 1994, 2258, 2414]. ontology [2074, 2225, 1917, 2395, 2118, 2065, 1966, 1967]. ontology-based [1917, 2118]. ontology-enhanced [2065]. ONTOS [744]. OODB [876, 930, 714, 1194, 774, 1115, 1090]. OODBMS [1174, 1020, 727, 1184]. OODBMSs [952]. OODBs [872, 939]. Ooi [2614]. Open [1386, 876, 1061, 1043, 930, 1361, 1319, 18, 2599, 2691, 811, 1532, 1605, 2082, 1278, 2430, 169, 2232, 933, 640, 1771, 2332, 896]. open-source [2232]. OpenDS4All [2793]. operability [915]. operand [129]. Operating [29, 2355, 257, 313, 928]. operation [758]. Operations [1711, 2632, 2669, 2670, 405]. Operator [2617, 1701, 1971]. Operators [1700, 1767, 1935, 91]. Opportunistic [2502]. Opportunities [1222, 1369, 58, 2543, 2582, 822, 2147, 2375, 1899, 1980, 619, 1238, 2264]. optical [285, 470, 530]. OPTICS [1622]. Optimal [2778, 266, 1621, 2835, 1491, 449, 2836, 2777, 1495, 2841, 2842, 369, 2073, 854, 51]. optimality [1123]. Optimisation [1701]. Optimistic [1632, 1105, 652, 219]. Optimistically [2738, 2737]. Optimization [1417, 2611, 1741, 1001, 1643, 374, 2099, 1319, 1491, 2839, 1631, 789, 2768, 2767, 2705, 2568, 656, 2840, 1729, 2704, 318, 1509, 431, 536, 2569, 2716, 1636, 2550, 1505, 1261, 1260, 1293, 1138, 1416, 1821, 1819, 808, 527, 386, 1848, 757, 1849, 1252, 875, 373, 661, 1815, 816, 1195, 1269, 432, 823, 1822, 2452, 760, 341, 1287, 387, 2258, 1253, 1807, 1933, 650, 725, 1251, 1940, 2267]. optimization-based [1821]. Optimizations [1483, 397]. Optimized [2827, 2606, 2828, 1249, 1310]. Optimizer [2806, 1750, 876, 385, 330, 792]. Optimizers [1487, 1284, 2327]. Optimizing [2429, 1256, 1015, 2603, 1824, 279, 1018, 1808, 2602, 2420, 1190, 556, 874, 450, 597, 767]. Options [946]. OR-DBMS [1434]. Oracle [1671, 1540, 1030, 1142, 1850, 1033, 1224]. Oracle8 [1534, 1541]. Oracle8i [1758]. ORCHESTRA [2318]. Order [1066, 1163, 2836, 1638, 1828, 875, 514, 369, 722, 1253, 1933, 1251, 1978]. Order-of-magnitude [1163]. Ordering [1622]. Orders [2835]. ordinary [834]. ORES [1052]. Organization [1503, 1010, 1889, 934, 177, 283, 1377, 2224]. organizational [1909]. Organizations [2142]. Organizing [587, 2017, 2471]. orientation [746]. Oriented [1479, 1036, 1435, 1054, 1091, 833, 1486, 2047, 569, 1574, 1037, 1034, 1048, 957, 1039, 858, 857, 362, 623, 755, 399, 2481, 1303, 2273, 941, 940, 1280, 408, 589, 436, 1171, 534, 645, 808, 765, 1192, 1276, 456, 2430, 1139, 409, 1365, 930, 434, 766, 918, 165, 649, 414, 515, 553, 459, 1340, 921, 398, 809, 967, 1560, 789, 486, 1923, 664, 856, 208, 741, 1999, 353, 2025, 815, 1366, 479, 1312, 532, 2207, 908, 744, 924, 1279, 429, 640, 898, 458, 639, 548, 592, 352, 612]. original [42]. origins [2212]. ORION [1199]. Orthogonal [1848, 1075]. orthogonally [1347]. orthographic [1689]. OSAM [924]. Oscar [2179]. Other [2712, 1460, 2008, 1152, 2071, 1518]. o’Ticks [1870]. Our [2818, 2385]. outer [1129, 740]. Outerjoin [1571]. Outerjoins [1019, 595]. outlandish [2230]. Outlier [1799]. Outliers [1716, 1744]. Outline [21, 68, 696]. Output [1558, 110, 576, 571]. outs [1296, 1297]. Outsourcing [1764]. overheads [2057]. Overview [1235, 1324, 249, 1683, 274, 1574, 972, 2266, 1160, 1161, 1367, 1950, 2158, 1166, 139, 734, 2452, 2360, 744, 1896]. own [2677]. Özsoyoglu [2414]. Özsu [2440].

P [2261, 2126, 98, 2017]. P-Grid [2017]. P2P [2017, 2024, 2025]. Package [2605]. Packages [2681, 1906]. packed [284]. packets [721]. Page [1020, 1572, 1003, 1869, 1889, 1999, 2330, 877]. page-level [2330]. paging [122]. Pair [1724]. PAIRSE [2481]. PAIS [2339]. panacea [689]. pane [2133, 2775, 2776]. Panel [350, 251, 340, 354, 336, 2754, 324, 1071, 2190, 2288, 293, 317, 1070, 1981, 1145, 306, 1164, 1152, 1873, 1875, 1525, 768, 1874, 300, 1524]. Panorama [2520]. Paper [2398, 2547, 2588, 2320, 191, 991, 2468, 2337, 2316, 1245, 225, 42]. Papers [2845, 13, 27, 1582, 2855, 1703, 2588, 2802, 2791, 2811, 1573, 1595, 1680, 1615, 97, 2351, 1790, 1892, 1929, 1961, 1992, 2124, 2086, 2101, 2140, 2155, 2177, 2194, 1474, 1959, 2294, 2284]. Papyrus [931]. Paradigm [580, 589, 1192, 444, 2111, 1987]. Paradise [1183]. Parallel [1464, 1041, 1160, 1370, 1047, 1029, 1475, 770, 631, 2572, 1013, 2631, 581, 1146, 1393, 1319, 1004, 2714, 2630, 1514, 782, 2594, 1727, 2421, 1549, 2573, 955, 1484, 904, 2559, 1759, 1113, 580, 2508, 1259, 2571, 1308, 1161, 469, 105, 1307, 757, 1281, 2249, 882, 1254, 928, 407, 512, 859, 1112, 725, 556].

Parallelism [2575, 1087, 513, 1217, 2574, 2750, 785, 577, 129, 758, 1163, 132, 651, 229]. Parameswaran [2534]. Parameter [600]. Parameterized [1979]. Parametric [834, 2569, 1818]. Pardes [833]. Paris [2642]. parity [1265]. PaRMa’04 [2154]. Part [1597, 1693, 1994, 1953, 1938]. Partial [2835, 2836, 1810, 449, 614, 786, 554]. Partially [853]. Participant [2558]. particle [469, 2030]. Partition [1021, 437, 1272]. Partitioned [1396, 480, 227]. Partitioning [2727, 289, 2728, 2571, 543]. Parts [2692, 2693]. Party [2833, 118]. Pass [1518]. passing [631, 337]. Passion [2591, 2319]. Passive [2795]. Past [1665, 2281, 114, 189]. Pasta [608]. Pasta-3 [608]. Patent [1567]. Patenting [2552]. Path [1000, 2041, 1285]. paths [2486]. Pattern [551, 2154, 998, 1185, 1838, 104]. Patterns [1489, 2603, 1643, 1708, 2602]. PATTY [2471]. Pavlo [2535]. PBIR [1866]. PC [1139, 1798]. PDA [2537]. PDES [691]. PDQ [2810]. peculiarities [2100]. Pedagogic [1601]. Peer [2016, 2108, 2145, 2005, 2281, 2398, 2019, 2022, 2023, 2020]. peer-based [2022]. Peer-to-Peer [2108, 2016, 2145, 2005, 2281, 2019, 2023]. Pegasus [894]. penalty [1265]. perception [1866]. perception-based [1866]. perfect [297]. Perform [1602]. Performance [331, 1390, 583, 259, 1670, 470, 772, 1028, 1380, 1262, 956, 866, 269, 601, 1373, 1562, 796, 1202, 532, 596, 686, 1010, 1023, 1067, 290, 777, 791, 268, 336, 477, 1803, 513, 1816, 530, 2208, 478, 2307, 287, 2345, 307, 759, 410, 471, 799, 1763, 769, 1304, 381, 2399, 2232, 238, 330, 925, 51, 512, 1911, 775, 1770, 1119, 1870, 2156, 2294]. Performing [2694, 1984, 2695]. perl [1699]. PersDB [2370]. Persistence [921, 325]. Persistent [1022, 1051, 2744, 1206, 2743, 1347, 798, 1186, 687, 1277, 643]. Personal [2537, 2560, 2028, 1210, 1241, 1889].

Personalization [1755]. Personalized [2370, 1551]. Perspective [2694, 2823, 2700, 2729, 2821, 2604, 2659, 2725, 2841, 2572, 2727, 2825, 1536, 2692, 2698, 2576, 2775, 2735, 2779, 2673, 2574, 2578, 2608, 2636, 2630, 2665, 2667, 2835, 2706, 2771, 2801, 2602, 2661, 2696, 2839, 2635, 2827, 2570, 2831, 2777, 2767, 2663, 1663, 2568, 2564, 2606, 2600, 2739, 2781, 2632, 2773, 2702, 2743, 2833, 2704, 2669, 2741, 2837, 2566, 2628, 2634, 2731, 2816, 2829, 2785, 2783, 2671, 2733, 2737, 2769, 857, 546, 1919, 372, 181, 2357, 183, 324, 1201, 2819]. Perspectives [2638, 2261, 968, 2126, 1459, 2543, 2204, 2427, 368]. pessimistic [538]. Petabyte [1662]. Peter [2044, 707, 2353, 2649]. Ph.D. [2091, 2240]. pharmaceutical [917]. phase [1804]. PhD [1926]. Phil [2100]. Philippe [2045]. Phoenix [1948]. phylogeny [2353]. Physical [946, 2271, 2378, 2357, 283]. physics [469, 119]. PI [2650]. Piazza [2020]. pictorial [1935, 284]. Picture [1419]. pieces [2433]. Piers [2076]. pigs [2156]. Pipelined [1004, 903, 854, 407]. Pipelines [2789]. pipelining [2161]. Pixel [1340]. Pixel-oriented [1340]. PL [37]. PL/1 [37]. Place [1590, 1919]. Placement [1017, 1727, 441, 1135]. plan [2465]. Planck [2226]. Planning [2810, 673]. Plans [1001, 1491, 1750, 1823, 535]. Platform [1530, 2817, 1531, 2413]. Platforms [1043]. please [215]. PLSI [2456]. plumbing [2353]. POB [1362]. PODS [2213, 2418, 2718]. point [1612, 239]. PointCast [1537]. Pointer [596, 83]. Pointer-Based [596]. Pointers [85]. Points [1622, 598, 184, 2247]. Policies [1404, 1021, 1410, 772, 237]. Policy [2857, 1135, 2201]. policymakers [2219]. Politecnico [2276]. Poly’19 [2719]. polymorphic [506, 1196]. Polynomial [574, 462]. Polystore [2539]. poorly [1940]. populated [1258]. population [395]. Port [1539]. portability [8]. Portable [1515]. Portals [1660]. Portland [2437]. Position [191, 225]. positional [1935]. Positions [1736]. possible [375]. POSTGRES [355, 544, 611]. Power [2708, 1723, 2523, 2669, 2038, 885, 2670, 515, 296, 2005, 572, 2156, 2193]. Powered [1567, 2534]. powerful [2348]. powerset [463]. pp [2088]. Practical [873, 1732, 225, 1017, 567, 2768, 2767, 1692, 1824, 590, 338, 557, 1924]. Practicality [1523, 1123]. Practice [2760, 2652, 2356, 1576, 2548, 2657, 2658, 1681, 2589, 1775, 451, 1444, 1903]. Practices [1584, 2227]. Practitioner [702, 63]. Pragmatics [300]. Pre [2834, 2833]. Pre- [2834, 2833]. precision [1826, 2348]. PREDATOR [1434, 1467]. Predicate [593, 874, 1017, 2196, 262]. predicate-window [2196]. Predicates [1018, 1129, 2429, 874, 1275, 2257, 2280]. predicting [1304]. Predictive [1013]. preemptible [853]. PREFER [1818]. Preferences [1733]. Prefetching [343, 1816, 873]. Prefix [2694, 2695]. Preliminary [1582, 998, 296]. Preparation [2715]. PresDB’07 [2270]. Presence [1643, 541, 1274, 1289]. Present [26, 2547, 2281, 189]. Presentation [1628, 148, 607, 1564, 1316, 351]. presented [42]. Preservation [2270, 1888]. Preserving [1745, 2481, 1803, 2058]. Press [985, 2123, 204]. pressure [2267]. prevent [1900]. Prevention [943]. previews [1368]. price [1217]. PRIMA [753]. Primary [2567, 2008]. Primitive [2532, 516]. Principles [2859, 2593, 1611, 29, 2745, 1702, 335, 1413, 2206, 894, 2089, 2401, 2292]. printed [322]. priority [657, 113]. Prism [280]. Privacy [1745, 2852, 2290, 2558, 2339, 2481, 2197, 2058, 2423]. Privacy-Preserving [1745, 2481]. Private [2659, 2841, 2842, 2660]. prize [2187]. Proactive [2857, 1538]. Probabilistic [2613, 2740, 2334, 2739, 2383, 1835, 2333, 2267]. Probabilities [2621]. Probe [1802]. Problem [999, 2566, 10, 1110, 1294, 1114, 439, 794, 565, 1333, 490, 1927, 725]. Problems [94, 219, 2562, 225, 2024, 573, 1524, 518, 702, 2085, 2100, 2415]. Procedures [594, 1349, 802, 471, 656, 396]. Proceedings [2859, 2869, 2865, 2861, 2860, 57]. ’Process [197, 440, 1588, 322, 489, 2320, 815, 1346, 1345, 137]. process-oriented [815]. processes [1863, 2283, 2383, 1916, 1909]. processibility [504]. Processing [1730, 1041, 1060, 1005, 1320, 2613, 1545, 1013, 2746, 2618, 581, 1409, 1411, 471, 166, 2640, 508, 2664, 1514, 1548, 2594, 1618, 1644, 2422, 2663, 2538, 1094, 600, 1501, 955, 2589, 1026, 1737, 1045, 2612, 2623, 2755, 2642, 580, 2687, 2508, 1881, 326, 217, 2147, 112, 871, 519, 379, 1989, 631, 1262, 128, 2099, 1142, 1985, 1998, 577, 754, 875, 537, 1876, 378, 1143, 348, 232, 2421, 345, 1923, 2305, 2499, 353, 810, 2452, 928, 788, 796, 482, 2309, 220, 2367, 2350, 2251, 859, 906, 801, 2185, 1177, 904, 1957, 267, 475, 1952, 230, 107, 2651]. processor [903, 82, 854, 1235, 132]. Processors [1012, 2508, 1270, 105, 119, 1011]. product [691, 2408]. Production [551, 2639, 550, 762, 483, 545, 484, 548, 592, 813]. products [1070, 1413, 1152, 902, 1300, 1251, 2100]. professor [2243, 2230, 2268, 2144]. Profile [2370]. Profiles [1452, 1912, 2205]. profiling [2490]. Program [38, 2655, 8, 41, 311, 69, 2386, 1185, 2409, 20]. programmer [54]. Programming [2203, 2849, 1419, 83, 389, 70, 71, 154, 676, 144, 325, 1469, 589, 160, 1693, 2387, 165, 1294, 463, 282, 499, 72, 921, 827, 2448, 585, 767, 78, 506, 49, 50, 143, 185, 678, 2353]. programming/object [362]. programs [342, 339, 1352, 184, 573, 650, 475]. Progress [2568, 986, 2036, 1949, 67]. Progressive [1830]. Project [1055, 1427, 1047, 1419, 2677, 927, 2021, 1948, 1858, 1890, 2413, 420, 111, 2482, 1990, 2077, 726, 2327, 1795, 1786, 1312, 2020, 1314, 33, 2036, 2717, 1896]. Projected [1623, 1714, 2456]. projection [1196]. projections [266]. projects [2285, 2441]. Prolog [498, 312, 427, 319, 279]. Prolog-INGRES [427]. proof [2205]. Propagation [1626, 1571, 671]. properties [75, 738, 2175]. proposal [1362, 1387, 967, 2320, 1363, 32]. proposals [135]. proposed [1276, 2321]. Prospector [1308, 1307]. protect [2197]. Protein [1567]. Proteus [365]. protocol [799, 780, 2037, 652]. Protocols [1626, 771, 583, 444, 2059, 826, 1104, 786, 1232, 250]. Prototype [1422, 1057, 1430, 1059, 1431, 1613, 1563, 1865, 923, 929, 1862, 356, 1178, 1871, 1182, 925]. Prototypes [1704]. Prototyping [1048, 456]. proud [1940]. Provenance [2228, 2646, 2292, 2711, 2656, 2621, 2443, 2164, 2356]. prover [464]. Providing [1555, 1554, 1033, 1268, 787, 1198]. Provision [1766]. Proxy [1827]. Proxy-server [1827]. Pruning [1483]. PSA [1539]. PTool [1186, 1051]. Public [2505, 2131]. publication [2211, 2221]. Publications [2512, 2402, 2186]. publish [1538, 1806]. publish/subscribe [1538, 1806]. publishable [2319]. published [2067]. Publishing [1564, 66, 2267]. Pull [1385]. Purdue [2151]. purge [1114]. Purple [2326]. purpose [71, 2413, 2249, 291]. Push [1385, 1536]. Pushing [2524]. putting [424]. puzzles [2401]. PVLDB [2586]. Pyramid [1494]. Pyramid-Technique [1494].

Qatar [2454]. QBI [1175]. QBIC [1173]. QDB [2458]. QDB10 [2426]. QUAD [305]. QUAD-trees [305]. Qualitative [2123, 783]. Quality [1787, 2430, 2548, 2068, 2661, 2426, 2625, 2213, 2662, 2035, 1988, 1803, 2458, 2054, 2469, 1926, 2409]. Quality-aware [2430]. quantification [695, 2209]. quantified [246]. quantifier [519]. quantile [1801]. Quantiles [2782, 1518, 2781]. quantitative [1776, 1248]. Quark [1542]. quarks [2138]. QUASAR [2054]. Quasi [1439]. Quasi-Cubes [1439]. queens [827]. Queensland [948, 830, 2484]. QUEL [271]. Queries [1749, 2611, 1637, 2652, 2832, 2605, 1635, 1741, 1499, 1555, 1015, 1724, 1504, 581, 1389, 1571, 1500, 2616, 1747, 1376, 1401, 1018, 2839, 1066, 1394, 1725, 1631, 2565, 2517, 1007, 2840, 2617, 1371, 1485, 2531, 1509, 1473, 2829, 2830, 1751, 2749, 1717, 1505, 1823, 377, 1838, 2094, 1893, 2443, 1129, 519, 2041, 2429, 2352, 1138, 1256, 1821, 2477, 1285, 808, 2141, 736, 243, 504, 1805, 342, 374, 1281, 1252, 1824, 1779, 378, 471, 874, 508, 1818, 597, 246, 789, 1830, 330, 248, 333, 738, 433, 656, 1268, 1109, 2367, 502, 1864, 2258, 431, 536, 800, 1870]. queries [1113, 1832, 556]. Query [1640, 1261, 1260, 1417, 2799, 1730, 2708, 2638, 2810, 607, 2729, 1706, 1793, 1060, 1320, 1056, 1645, 1416, 1002, 1567, 1740, 1819, 2613, 1016, 1720, 1001, 605, 2841, 2686, 2730, 2637, 2206, 1013, 1368, 1683, 2852, 2842, 1486, 1438, 1643, 757, 2752, 1510, 1319, 1286, 2664, 348, 373, 1642, 2636, 2667, 2806, 1492, 1491, 232, 2334, 2567, 345, 1421, 1269, 2657, 2831, 2742, 1644, 2768, 2767, 2658, 2663, 1646, 2668, 2538, 2705, 2568, 1373, 600, 810, 1501, 1655, 955, 595, 1729, 1855, 2175, 2704, 2677, 1026, 2647, 2612, 2741, 2566, 2569, 1482, 1493, 1750, 2680, 2642, 2716, 2434]. Query [1636, 969, 2550, 516, 1173, 326, 561, 2283, 2218, 1125, 876, 1293, 485, 604, 1986, 517, 527, 919, 400, 1127, 2383, 1192, 1128, 1107, 261, 1262, 386, 1108, 2387, 1849, 1252, 535, 577, 875, 537, 802, 1876, 672, 661, 1123, 2327, 279, 1195, 824, 432, 823, 1175, 750, 1822, 2048, 353, 2452, 760, 1935, 482, 814, 2350, 2066, 318, 341, 387, 1807, 1933, 1069, 801, 725, 1337, 320, 337, 1242, 792, 267, 1940, 2267, 1834, 1952, 230, 304, 438]. query-based [2283]. Querying [2374, 2450, 2475, 1974, 1056, 2603, 1658, 1732, 2735, 1629, 2844, 809, 2602, 1785, 1395, 584, 2736, 375, 1106, 2104, 736, 2077, 2333, 2073, 1420, 1935, 1883, 2282, 1987, 1936, 2134, 2322]. Quest [1055]. Question [2597]. questions [2256]. queues [578, 113, 2120]. quick [1850]. Quickly [1009]. QuickStore [1023, 1119]. quiescing [804]. quotes [1870].

R [101, 1944, 871, 1380, 782, 1527, 1111, 284, 250, 2156]. R&D [819]. R* [2120, 330]. R-trees [871, 782, 1111, 284]. R. [1331]. R/3 [1380, 1527]. ’R2T [2841, 2842]. Rabinovich [2043]. Raghu [2219]. RAID [1169, 442]. RAID5 [879, 1265]. rainbow [2294]. Rajaraman [2491]. Ralph [2027]. Ramakrishnan [2219]. Random [1519, 1639, 1266, 2607, 1638, 603]. Randomized [597]. randomness [1978]. Range [2666, 1637, 1747, 1376, 2673, 2665, 246, 2674, 497, 1275, 267]. ranked [1818, 2322]. Ranking [2277, 1971, 1999]. Rapid [1753, 1048, 1251, 456, 1139]. Rapidly [1065]. rare [1804]. RasDaMan [1561]. rate [2037]. rates [1282, 2202, 2211]. ratio [1135]. rational [817, 795]. ray [2344]. Rdb [793, 1540, 1142]. Rdb/VMS [793]. RDBMS [1154, 2700, 1043, 2701, 1089]. RDF [2498, 2374, 2019, 2367]. RDF- [2019]. Re [1784, 1491, 2656, 2397]. Re-computation [2656]. Re-designing [1784]. Re-Optimization [1491]. reach [2219, 1174]. Read [2825, 1043, 1544, 2095, 773, 2826]. Read-Mostly [1043]. Read-Only [1544, 2095, 773]. Real [2778, 1235, 1154, 1239, 1323, 2485, 2555, 1465, 1545, 1380, 1409, 956, 1747, 1230, 1229, 867, 1762, 1231, 2744, 2777, 1007, 2528, 2743, 421, 1124, 771, 1236, 1227, 2236, 797, 215, 826, 1868, 418, 1233, 1228, 1444, 422, 1854, 417, 1234, 416, 2185, 825, 1214, 1232, 792, 1870, 2275, 2415, 1778, 419]. Real-Time [1235, 1239, 1323, 2555, 1409, 956, 1007, 1230, 1229, 867, 1231, 421, 1124, 771, 1236, 1227, 826, 1868, 418, 1233, 1228, 1444, 422, 1854, 417, 1234, 416, 2185, 825, 1214, 1232, 1778, 419]. Real-World [1465, 797]. Realizing [790]. Really [2585, 423]. realtime [343]. reasoners [869]. Reasoning [2810, 2041, 1627, 2074, 2412, 159, 448, 586, 662, 514, 171]. Rebuild [1757, 879]. rebuilds [1265]. receives [2129]. recipient [2445]. reclamation [408]. Reclustering [1024]. Recognition [2644, 2723]. Recommendation [2596]. Reconciliation [1602]. Reconciling [1839]. Reconfigurable [2538]. Reconsidering [2063]. reconstructed [742, 884]. Reconstruction [2666, 2665]. Record [1540, 1580, 1656, 1009, 1196]. Record-Boundary [1656]. Records [1546, 1811, 773, 346]. recoverable [578, 1198]. Recovery [292, 2694, 1554, 1375, 1521, 1624, 993, 1104, 2695, 511, 779, 2542, 392, 780, 905, 865, 1850, 381, 803, 781, 329, 1903, 1119, 2432]. rectangles [598, 1111]. recurrence [636]. recurrent [339]. Recursion [2833, 885, 400, 401, 524, 338]. recursions [473]. Recursive [2700, 1702, 2701, 326, 527, 339, 378, 789, 338, 474, 402]. Red [1043]. RedBD [2135]. redefinitions [1121]. Reducing [1128]. Reduction [1496, 551, 1809]. Redundancy [529, 2343, 378]. Redundant [1406, 442]. Reed [1728]. Reference [323, 425]. referential [510]. Refinement [1558, 1837]. Reflections [2604]. reformulated [495]. Reformulation [2638, 2206]. refresh [327]. regarding [1959]. Regional [26, 25]. Registration [38]. regression [258]. regular [2438, 2041]. Reinventing [2552, 2294, 2483]. Rejection [2815]. Related [57, 2562]. relation [677, 483].

Relational [1328, 1677, 1320, 1381, 517, 1326, 1702, 2556, 282, 1057, 1514, 1442, 1037, 2629, 1668, 2565, 690, 2633, 248, 997, 2022, 2367, 2856, 1511, 951, 2621, 2628, 2506, 2591, 498, 71, 101, 858, 1158, 2476, 131, 2346, 669, 207, 2044, 222, 510, 1413, 133, 264, 302, 1528, 65, 367, 370, 2208, 1327, 243, 555, 1365, 342, 466, 564, 463, 1306, 701, 508, 672, 244, 509, 648, 279, 742, 921, 205, 1159, 500, 460, 922, 173, 1873, 437, 722, 383, 542, 412, 73, 1102, 902, 455, 1295, 63, 1939, 1366, 1089, 384, 502, 396, 1883, 289, 1248, 278, 310]. relational [494, 884, 548, 592, 267, 2294, 438, 698, 1832, 99, 551, 132]. relational-database [1089]. Relations [1315, 1865, 2094, 703, 561, 44, 56, 2343, 527, 356, 1332, 1331, 1164, 228, 2471, 1111, 528, 613]. Relationship [2062, 1441, 817, 1193, 542, 1939, 651]. Relationships [150, 525, 2301, 724]. Relative [2782, 2781, 1700]. Relativism [172]. Relaxed [2234, 1045, 994]. relevance [2012]. relevant [1152, 591]. Reliability [317]. Reliable [588]. Religious [1315, 1332, 1331]. reload [684]. Remarks [839, 12, 1689, 840]. Remembering [707]. Reminiscences [2845, 2855, 1703, 2802, 2791, 2811, 1790, 1892, 1929, 1961, 1992, 2124, 2086, 2101, 2140, 2155, 2177, 2194, 1474, 1573, 1595, 1680, 1615]. remote [788]. Remotely [241]. Remotely-sensed [241]. Remy [1196]. Rensselaer [726]. reorderable [595]. Reordering [1252]. reorderings [1129]. Reorganization [1318, 1713, 835, 58, 1258]. Repair [2698, 2699]. Repeatability [2411, 2362, 2286]. replacement [877]. Replica [2056, 682]. Replicated [1626, 1738, 907, 493, 582, 443, 2499, 786, 538]. Replication [1523, 1030, 1031, 1032, 989, 1224, 1149, 1264, 2369, 532, 1150, 2043]. Reply [197, 1332]. Report [2519, 2690, 1994, 2626, 2615, 2238, 2813, 2190, 2797, 2848, 2136, 2287, 62, 136, 2278, 2139, 2379, 2475, 1594, 1675, 2655, 1323, 2216, 1793, 2228, 1981, 34, 2261, 1774, 2485, 2555, 2354, 1465, 2554, 2292, 2270, 2537, 1457, 1676, 2229, 1791, 2493, 29, 2751, 944, 1685, 2106, 2241, 2323, 2068, 1930, 9, 2215, 1946, 2152, 2459, 2127, 2289, 2403, 2277, 2108, 2405, 2685, 2713, 2757, 2510, 2561, 2153, 1034, 1993, 2254, 2290, 2355, 2046, 2107, 2137, 1764, 2511, 2512, 2599, 2126, 2426, 2449, 2380, 2410, 2467, 2092, 2656, 2048, 1466, 1609, 2070, 2851, 838, 2684, 2291]. Report [2466, 2494, 2528, 2553, 1444, 2592, 2213, 2677, 620, 2214, 1215, 2191, 1345, 2495, 60, 2154, 2173, 2125, 1071, 1765, 2262, 2339, 2279, 2340, 622, 2315, 2322, 1455, 199, 1239, 2174, 222, 1362, 2363, 1931, 2356, 2338, 2036, 621, 1686, 2047, 545, 553, 2049, 1945, 2391, 296, 67, 1456, 1962, 1209, 2263, 118, 973, 1980, 2245, 2386, 2239, 2227, 2255, 2240, 2069, 2203]. Reports [25, 1465, 1346, 2203, 96]. Repositories [1479, 1256, 2400]. Repository [1298, 1601, 1305]. Representation [365, 1569, 2774, 2773, 2154, 375, 1455, 677, 5, 549, 6, 717, 139]. representations [85, 713, 732]. representing [691, 432, 542, 2134]. Reproducibility [2798, 2764]. Reproducible [2599]. reputation [1970]. requests [578]. Requirements [2788, 1154, 112, 81, 797, 2430, 888, 2119, 2031, 121, 242, 2185]. Research [2519, 2654, 2593, 2679, 2858, 2864, 982, 2435, 1594, 2454, 2645, 1578, 2796, 1457, 1965, 1462, 1065, 2712, 1325, 2751, 968, 1152, 2001, 630, 1684, 625, 2132, 2509, 2050, 1775, 845, 1461, 2711, 2198, 2717, 2424, 1446, 2657, 1579, 2469, 948, 2817, 641, 1202, 1983, 1467, 77, 1909, 2650, 115, 2520, 1071, 198, 846, 987, 2586, 2624, 2648, 2689, 2562, 2315, 140, 142, 144, 843, 964, 2016, 2368, 2082, 496, 1227, 1070, 216, 2473, 2189, 1974, 2276, 2464, 1995, 1073, 2492, 2172, 1221, 203, 1294, 2269, 2071, 704, 64, 94, 2182, 2457, 2135, 372]. research [2072, 1100, 752, 633, 2145, 2375, 963, 1300, 181, 728, 747, 2393, 1209, 981, 1599, 830, 1243, 748, 2091, 1203, 829, 2217, 617, 1167, 1980, 1238, 628, 632, 2361, 910, 702, 820, 947, 844, 825, 2171, 1244, 2226, 2264, 1926, 1964, 2100, 2244, 2275, 2252, 2284, 2335, 2348, 2409, 2408, 2441, 831, 2484, 2690, 1650]. Researcher [2762, 2753, 2812, 2747, 2689]. researchers [1958, 1874, 2142, 2252]. resemblance [718, 719]. resident [381]. residues [816]. resistant [1273]. Resolution [2849, 2688, 1679, 520, 1830, 479]. Resolving [464, 717]. Resource [1756, 2573, 2857, 1467, 2181, 1127, 361, 1281, 710, 1231]. Respect [2550]. Response [1189, 840, 1753, 1064, 1331, 1128]. restarts [312]. Restricted [2638]. restrictions [515]. result [2377, 463, 1123, 709]. Results [2637, 2636, 998, 1351, 1350, 671, 482]. Resumption [1712]. Rete [649]. Retention [1539, 772]. Rethinking [2345, 1940]. RETINA [1868]. Retrieval [530, 1085, 1492, 1421, 1650, 1014, 1240, 305, 41, 1792, 751, 1866, 264, 1965, 134, 1817, 2127, 1954, 665, 2486, 449, 269, 2358, 776, 1934, 125, 277, 1176, 855, 860, 84, 310, 2255, 2255]. retrieval-intensive [855]. Retrieving [935]. retrospective [731]. reusable [732]. reuse [2406]. Reusing [1485, 2162]. Reuter [2483]. Revealing [2706, 2707]. Reverse [1725, 2395]. Review [101, 2027, 1977, 2044, 2818, 1697, 100, 2123, 204, 99, 98, 2088, 57, 2043, 66, 2045, 2122, 1902, 1922, 1941, 1955, 1976, 2003, 2042, 2121, 2337, 2146, 316, 821, 1854]. reviewer [2337]. Reviewing [2679, 2398, 2321, 2211, 2222, 2256, 2221]. Reviews [97, 214, 2320, 48, 52]. revised [2004]. revisited [2062, 1280, 519, 637, 374, 533, 500, 2490]. Revisiting [1411, 2801, 2839, 2840, 2419]. revitalized [2385]. revolution [2305]. rewrite [760]. Rewriting [1655]. Ricardo [2275]. Richard [2678]. Richer [1477]. Rick [2552, 2560, 2597]. RIDE [1071]. RIDE-ADS [1071]. RIDL* [541]. Rigaux [2045]. Right [2747, 523, 1524, 2230]. right- [523]. rights [368]. Ripple [1641]. RISC [1270, 1008]. RMP [1590]. Road [2545]. Roadmap [1593]. Robust [2726, 2725, 2052, 2550, 598, 1821]. ROLAP [1503]. Role [913, 2850, 957, 2625, 498, 217, 1921, 1874]. Role-Based [957]. roles [674]. Roll [1719]. Romania [937]. Ron [2619]. ROSE [397]. Ross [2027, 2641]. Round [2798, 2571]. routines [1693]. Row [2682]. Rows [2748]. RS [1728]. Rubinson [2316]. Rufus [934]. Rule [1382, 546, 1284, 1487, 1013, 342, 1460, 593, 761, 347, 1630, 549, 1755, 1511, 1482, 580, 857, 851, 495, 589, 349, 386, 649, 406, 545, 1170, 2198, 1804, 760, 1776, 2169, 813, 1834, 402]. Rule-Based [1487, 342, 589, 349, 1284, 386]. Rule-Processing [580]. Rulebase [1322, 1321]. Rules [547, 1392, 1391, 1435, 1487, 970, 1460, 1393, 1753, 1472, 1408, 1407, 1483, 1754, 1484, 594, 850, 868, 762, 1249, 1310, 520, 432, 523, 1118, 1900, 1836, 1248, 2006, 544, 548, 592, 651]. Run [1522, 2825, 2826, 2294]. running [587]. Runtime [2839, 2840]. Russia [2310]. RWTH [1100]. Ryan [2445].

S [101, 1197, 2313]. S/V [2313]. S3 [1433]. sabbaticals [2193, 2433]. SACMAT’04 [2106]. Safe [1277, 2825, 2826]. safety [447, 1845]. Sagas [393]. sailboats [2294]. sailor [2295]. salesperson [1852, 1853]. SAMOS [1178]. Samples [1749]. Sampling [1519, 1639, 1510, 2672, 1638, 1715, 1750, 2671, 2733, 2734, 1273, 802, 1829, 824, 567, 603, 1834]. Sampling-Based [1510, 824]. San [2804, 2313]. Sandwiches [1621]. Sangam [1859]. SAP [1380, 2425, 1527]. Satellite [1083]. satisfying [276]. saving [129]. savings [2542]. Saying [1388]. SBA [70]. SBBD [2137, 1609]. Scalability [1042, 2525, 2219]. Scalable [2605, 2396, 1740, 1051, 1393, 2459, 2807, 1728, 2633, 1398, 2814, 1535, 1010, 1869, 1133, 1824, 2495]. Scalar [1514, 1270]. Scale [2729, 2572, 2730, 2722, 1548, 2355, 1602, 2331, 2039, 2111, 1692]. Scaleability [2308]. scaleless [657]. Scaling [2609, 2775, 2608, 2744, 2743, 2776, 860]. Scan [2532]. Scans [1003, 2420]. Scheduling [421, 2168, 1281, 1266, 867, 1231, 914]. Schema [1083, 939, 1507, 828, 835, 2359, 399, 2114, 2009, 2400, 2110, 2231, 2111, 1860, 1772, 2265, 2019, 2235, 814, 190, 2319, 2134, 1837]. schema-based [2019]. Schemas [2063, 1839, 245, 2344, 2112]. schematic [647]. Schematically [1498]. Scheme [1635, 1592, 1701, 787, 1249, 206, 514, 2358, 299, 538, 262, 91, 1778]. Schemes [1423, 575, 461, 450, 234, 315]. Schemr [2400]. Scholl [2045]. school [2405, 2125]. Science [2526, 2751, 1221, 1446, 2807, 2599, 2691, 2759, 2850, 2814, 2789, 2560, 2641, 2683, 2793, 2069, 1360, 2071, 1692, 2164, 2465, 2120, 2392, 2390]. Sciences [2215, 2541, 2081, 1891, 2072, 910]. Scientific [2390, 2208, 2229, 797, 2184, 2711, 1682, 998, 1981, 2406, 621, 2167, 737, 752, 2159, 1456, 2161, 2162, 430, 815, 2210, 1185, 2168, 2166, 2165]. Scientist [2624]. Scientists [937, 2850, 1329]. SciFlow [2229]. Scope [842]. scoring [2233]. Scrambling [1493]. SE [814]. seamless [834, 2117]. Search [1370, 1433, 1547, 1570, 2406, 2377, 2835, 1652, 1082, 1375, 2510, 2561, 1012, 2590, 2507, 2836, 1650, 1707, 2857, 2592, 2313, 1495, 1428, 2765, 2733, 2233, 784, 2734, 657, 2076, 881, 2376, 1011, 1612, 2399, 774, 2444, 1999, 2330, 2297, 1943, 284, 1176, 2503]. searches [1791]. Searching [1496, 1614, 105, 1857, 120, 1387, 256, 2414]. SEAS [2536]. Seattle [2690, 2069]. Second [875, 2684, 2213, 1159, 2495, 2238, 2713, 2561, 2371, 1466, 2494, 1765]. Second-order [875]. secondary [106]. Secretary [114, 96]. Section [1580, 972, 1355, 2159, 1599]. Secure [1409, 1515, 648, 582, 1234, 2051, 2223]. Securing [1858]. Security [629, 403, 957, 2035, 1904, 634, 464, 745, 1874, 1279, 202]. Seeded [1006]. Seek [1560]. Segment [658, 783, 117]. segmentation [1811]. Selected [633]. selecting [200]. Selection [1635, 1021, 2557, 1701, 1861, 2181, 2082, 2431, 1822, 855]. Selections [1636]. Selective [2024]. selectives [369]. Selectivity [1619, 1637, 1002, 1723, 1835, 1520, 1274, 567, 433, 1275, 2257, 2280]. Self [1633, 1727, 1024, 1726, 1535, 2017, 254, 1282, 483, 1888, 2335]. self-adapting [254]. Self-Adaptive [1024, 1282]. self-controlling [483]. Self-Extensible [1726]. self-instantiating [1888]. self-organizing [2017]. Self-Tuning [1727, 1633, 1535, 2335]. self-validating [1888]. Seller [2801]. Sellis [2648]. Semantic [1581, 2611, 1386, 1606, 1125, 1878, 1917, 944, 718, 1601, 1607, 719, 1608, 1603, 563, 1082, 2685, 2713, 2757, 816, 122, 1600, 2149, 2118, 1602, 709, 712, 1972, 1820, 2170, 2109, 713, 309, 1954, 435, 710, 2391, 1182, 204, 383, 2471, 382, 479, 2115, 387, 280, 1909, 2266, 278, 717, 721, 1971, 2450, 1968, 1974, 1969, 1975, 2451, 1973, 1980, 2252]. Semantically [2097]. Semantics [2504, 399, 443, 2507, 1602, 1994, 2094, 583, 2381, 534, 1882, 714, 2196, 1973, 1687, 742, 1231, 437, 2032, 856, 1525, 2115, 717, 2117, 548, 813]. semantics-based [583]. SEMCOG [1421]. Semi [1506, 1449, 1282, 2079, 1860]. Semi-automatic [1282, 1860]. Semi-Automatically [1506]. Semi-Structured [1506, 1449, 2079]. Semijoin [1571]. Seminar [2723, 2068, 2651, 2322, 2404, 2405]. Semint [1182]. Semiring [2443].

Semiring-annotated [2443]. Semirings [2834, 2833]. Semistructured [1450, 1606, 1653, 1451, 1445, 1452, 1454, 1507, 1655, 2044, 1789, 1905, 1856, 1453, 1305, 1448]. Senior [2762, 2753, 2747]. Sense [2357, 642]. sensed [241]. Sensing [2558, 2054]. Sensor [2278, 2030, 2070, 2035, 2033, 2034, 2039, 2053, 2056, 2031, 2037, 2032, 2040, 2038, 2051, 2052, 916, 1952]. SENTINEL [1435]. separable [473]. Separation [1397, 957, 2148, 1777]. Sequence [2545, 1026, 2076]. Sequences [1394, 1701, 1898, 1871]. sequencing [46]. Sequential [2631, 802, 2630, 113, 346]. SEQUOIA [1083, 847]. Serge [2044, 2284]. serializability [481]. Series [968, 1025, 2578, 2545, 2684, 2579, 1371, 1187, 339, 1877, 1691, 1809, 1870, 2248]. Server [1861, 1020, 1553, 1032, 2775, 1541, 1033, 1521, 993, 1039, 2776, 1374, 1140, 1308, 680, 1307, 1210, 779, 1262, 1827, 932, 1089, 700, 681, 811, 1759]. Servers [1699, 1535, 1134, 1763, 1875, 1255]. Service [2806, 1764, 1605, 2204, 2481, 1787, 1988, 1914, 2430, 1970, 2025, 2282]. service-oriented [2481, 2430, 2025]. Services [842, 1608, 1035, 2122, 2502, 1766, 1213, 1783, 1968, 1774, 1879, 2097, 1210, 2131, 457, 2158, 1894, 1359, 2210, 936, 2252]. Session [1029, 1650, 293, 306, 1164, 1152, 1873, 1875, 1874, 324, 300]. sessions [42, 1650]. Set [649, 1631, 592, 699, 65, 1817, 82, 872, 2247, 548]. Set-oriented [649, 592, 548]. Sets [1725, 1744, 2650, 375, 868, 997]. setting [1826]. seventh [2125, 2528]. Seymour [2128]. SG [222]. SGML [1090]. Shaffer [1944]. Shape [2578, 2579]. shapes [665, 2247]. Shapley [2843]. SHARE [33]. Shared [1320, 1047, 1317, 955, 1259, 1104, 512, 859, 860, 2385]. Shared-Nothing [1047, 1317, 512, 860]. Sharing [1020, 2847, 2318, 1672, 2113, 1920, 1277, 915, 2022, 723]. Shasha [2401]. Sheldon [204]. shelf [2265].

Sherman [2827]. Shore [1184]. Shoring [1022]. short [2158, 1098]. shorts [2401]. Should [2701, 2819, 2820, 145, 1149, 1940, 2219, 2348]. shouldn’t [2284]. show [2244]. Shrinking [1649]. Shrunk [1659]. Shutdown [1247]. Shyness [2624]. Sibley [98]. SIG [2563]. SIGACT [2859]. SIGACT-SIGMOD [2859]. SIGBDP [36, 42]. SIGFIDET [13, 20, 38, 14, 42, 22, 34, 36, 21]. SIGIR [1792]. SIGMETRICS [1464]. SIGMOD [1580, 2213, 2397, 2445, 2815, 2859, 2860, 991, 2679, 2190, 2862, 2863, 2866, 2867, 2868, 2870, 2871, 1207, 1226, 1330, 1353, 1447, 1598, 1674, 1695, 2655, 2411, 2869, 1696, 2849, 2130, 2865, 2387, 1697, 2547, 2364, 200, 2448, 1894, 2211, 2718, 2362, 2286, 2011, 2221, 936, 2243, 2268, 2861]. SIGMOD/PODS [2213, 2718]. SIGMOD’10 [2386]. Signal [2666, 2665]. Signature [287, 875, 872]. Silberschatz [1926]. Silico [2711]. Silos [2759]. silver [1977]. SIM [807, 435]. similar [1817, 665]. Similarity [1516, 1651, 1370, 1433, 1499, 1555, 1496, 1652, 1650, 1707, 1501, 1371, 2468, 1828, 112, 2076, 1877, 1231, 2489, 2257, 2280, 2503, 2248]. Similarity-Based [1371]. Simon [1944]. Simple [607, 2265, 795, 266, 405, 573]. simplify [69]. Simplifying [273, 2163]. SIMS [935]. simulated [373]. Simulation [2064, 2541, 1304]. Simultaneous [1383, 1505]. Singapore [1539, 2473, 559]. singing [2156]. Single [2775, 2222, 2623, 901, 2776, 2571, 576]. Single- [2222]. single-input [576]. Single-node [2623]. single-output [576]. single-round [2571]. SIRIO [836]. Sister [1696]. Site [1438, 1429, 1646, 1544, 1517, 200]. Sites [1611, 2368, 1841]. Situation [2549]. six [2148, 1772]. Sixth [2290, 2511]. Size [1397, 1572, 1379, 1810, 1155, 1273, 802, 671, 1123, 824, 201, 502, 855]. sized [2384]. sizes [231]. SJCC [34]. Skeena [2823]. Sketches [2709]. Sketching [2735, 2736]. Skew [2572, 2487, 1273]. skew-resistant [1273]. Sky [1746, 2311]. Skyline [2477, 2382]. SLCAs [2507]. Sleepers [988]. sliced [1800]. sliding [2133]. sliding-window [2133]. Sloan [1746, 2311]. SM3 [259]. Small [1978, 2061, 2384, 730]. Smalltalk [281]. smart [2237, 1852, 1853, 1560, 1940]. smart-cache [1560]. SmartCIS [2378]. SMP [1043]. Snakes [1621]. snapshot [2131, 2095, 327]. Snodgrass [2560]. Snowball [1865]. Social [2368, 2676, 2470, 2491, 2796]. Societies [1696]. Society [2290, 118, 2339]. soft [1124]. Software [2338, 1325, 2152, 2543, 1919, 88, 2047, 414, 2403, 921, 110, 928, 188, 2409]. Solomon [1728]. Solution [1548, 1859, 1869, 1110, 1824, 1264, 2321, 2448]. solutions [2024, 794, 626, 2157]. Solving [518, 10]. Some [164, 86, 75, 1689, 636, 48, 52, 729, 2747, 998, 287, 103, 64, 368, 2351]. sometimes [1940]. SONAR [1310]. Sons [2027]. Sort [1008, 1270, 576]. sorter [903]. Sorting [1390, 1522]. Sortledton [2821, 2822]. sound [438]. Sourav [2749]. Source [1602, 2181, 2232]. Sources [1556, 1449, 1606, 1645, 1726, 935, 1455, 2081, 1974, 1138, 1839, 2102]. Sourcing [2521]. South [17]. Southern [2163]. Southerner [2472]. SOX [2326]. Space [1801, 1638, 1604, 650, 530, 1135, 888, 2234, 450, 1785, 1283, 492, 1288, 725, 910]. Space-efficient [1801]. space-filling [1785]. Spaces [1714, 1400, 1452, 600, 661, 2357]. Spain [2002, 2135]. Span [2641]. Spanners [2697, 2696]. SPARC [55, 62, 222]. sparely [1258]. sparely-populated [1258]. SPARQL [2374]. Sparse [1647, 1634]. SPARTAN [1820]. Spatial [1619, 1734, 657, 1060, 1005, 1767, 1724, 1723, 1500, 1430, 1502, 1652, 1397, 1867, 1006, 1271, 1618, 353, 600, 1398, 2091, 694, 2612, 2045, 871, 888, 409, 834, 625, 256, 529, 1272, 284]. spatial-merge [1272]. Spatio [2361, 2091]. Spatio-temporal [2361, 2091]. Spatiotemporal [841, 1473, 2358, 1768, 2134]. Spatscheck [2043]. speaking [1852, 1853]. Speaks [2712, 2524, 2525, 2534, 2535, 2552, 2560, 2591, 2586, 2597, 2598, 2614, 2624, 2619, 2653, 2641, 2649, 2642, 2648, 2683, 2678, 2716, 2689, 2765, 2756, 2764, 2749, 2750, 2815, 1926, 1940, 1959, 1991, 2120, 2085, 2100, 2156, 2193, 2176, 2205, 2243, 2230, 2219, 2244, 2268, 2260, 2267, 2275, 2252, 2319, 2294, 2284, 2335, 2348, 2353, 2397, 2385, 2389, 2409, 2401, 2415, 2423, 2408, 2414, 2433, 2453, 2432, 2441, 2440, 2483, 2491, 2463, 2472]. Special [1580, 1895, 1355, 553, 2109, 888, 2301, 2159, 1599]. specialized [469, 125]. Specific [2650]. Specification [1046, 976, 155, 1988, 162, 766, 1916, 359, 465, 187, 280, 242]. specifications [2204, 517, 714, 186, 1035]. specifying [586, 875]. spectrum [2031]. Speech [1085]. Speed [1711, 2770]. spent [1940]. SPIFFI [1133]. Spirit [1669]. sponsored [620]. sponsors [1330, 1353]. Spotfire [1338]. Springer [1977, 2122]. Springer-Verlag [1977, 2122]. spruce [1873]. SQL [1861, 2316, 476, 2476, 561, 1103, 477, 1669, 1388, 2396, 785, 1553, 2604, 740, 1989, 1164, 1702, 1349, 1597, 1693, 1617, 1705, 1949, 2105, 2067, 1042, 1217, 695, 374, 1252, 562, 2317, 2616, 1066, 2447, 960, 1406, 961, 2199, 2681, 1166, 1906, 1794, 1962, 2643, 2564, 812, 1761, 2632, 1682, 1532, 1206, 814, 2272, 2677, 1564, 2150, 1218, 494, 2348, 1717, 2436, 2209, 556]. SQL-99 [1669]. SQL-based [561]. SQL-on-Hadoop [2677]. SQL/CLI [1218]. SQL/MED [1962, 1682]. SQL/MM [1906]. SQL/MP [1217]. SQL/OLB [1597]. SQL/SE [814]. SQL/XML [1949, 2105]. SQL3 [1617, 1034, 1088]. SQLEM [1761]. SQLite [2678]. SQLJ [1528, 1597, 1693]. SQoUT [2327]. SR [1401]. SR-tree [1401].

Srivastava [2415]. SRS [2788]. SSD [2448]. SSDs [2420]. Stable [30, 503, 865]. Stage [1737]. Staggered [995]. stand [215]. Standard [1036, 1880, 1362, 1349, 1166, 455, 1363, 2209, 2643]. Standardization [1705, 1075, 323]. Standardized [2795]. Standardizing [2552]. Standards [212, 1576, 1617, 1705, 984, 2026, 2560, 983, 692, 689, 1771, 1242]. STanford [1235, 2016, 1387, 633, 1201]. STAR [2796]. starburst [537, 610, 690, 760, 813]. Start [1850]. STARTS [1387]. startup [1926, 2252, 2284, 2409]. startups [1959, 2267, 2491]. State [1300, 2617, 1244, 2058, 2503, 2430, 1882, 129, 5, 6]. State-of-the-art [2058, 2503]. Statement [36, 1207, 1226]. statements [311]. Static [2251, 1279, 506, 538]. Statistical [2055, 259, 1122, 240, 1957]. Statistics [1510, 1638]. Status [136, 1034, 634, 1962, 1209, 628, 819]. Stefano [2252]. Step [1005, 1495, 590, 78, 925, 1075]. Steps [1705, 95]. Stepwise [359]. Steven [1944]. STHoles [1814]. Still [2568, 2156]. stochastic [502, 644]. stock [467]. Stonebraker [2129, 2530]. Storage [2805, 1670, 1324, 1313, 408, 1460, 1503, 1562, 791, 1278, 1840, 120, 303, 1241, 1169, 2246, 1266, 460, 1141, 1934, 796, 847, 1911, 2176]. Store [2771, 1396, 1023, 2772, 891, 643, 1058]. Stored [1206, 2008, 2343, 1349, 697, 1653]. stores [2396]. StorHouse [1840]. Storing [1653]. Story [2706, 2707, 2398, 2138]. strange [1164]. Strategies [988, 1320, 1621, 326, 759, 410, 789, 1911]. Strategy [1485, 357, 661, 725, 230, 2224]. Stratos [2598]. Stream [2618, 2640, 2786, 2651, 2785, 2687, 2099, 1996, 2198, 2496, 2185]. Streaming [2726, 2725, 2782, 2781, 2039]. streams [2064, 1124, 2094, 1893, 2374, 2104, 2146, 1797, 2196, 2405, 2133, 2055, 2224, 2475].

Street [1412]. STREETWALKER [45]. strength [2114]. strengthen [768]. strikes [690, 2487]. String [120, 104, 2503]. Strings [2738, 1742, 2737]. STRIP [1235, 1382]. striping [995]. Strong [2524, 2614]. Structural [1432, 507, 718, 719, 2098, 2175]. Structure [1622, 2784, 2821, 2822, 714, 1401, 1728, 1454, 2688, 2783, 1240, 41, 534, 2079, 586, 65, 256, 881, 524, 1830, 135, 1453]. Structure-Agnostic [2688]. Structured [1506, 1450, 1606, 2103, 1372, 1016, 2603, 1084, 2602, 2517, 50, 2017, 1449, 1303, 1811, 2331, 1292, 2079, 165, 270, 786]. Structures [1276, 1735, 2289, 1498, 2214, 1529, 791, 1810, 81, 2008, 10, 80, 783, 90, 269, 1829, 2073, 542, 180, 395, 1366, 77, 1337]. structuring [78]. STRUDEL [1429, 1517]. student [2397]. Students [2524, 2014, 2301, 2085, 2423]. Studies [1414, 2115, 1412]. Study [62, 1028, 956, 2002, 2388, 2208, 2057, 1368, 1217, 783, 1954, 401, 1612, 849, 1111, 467, 1119, 55]. studying [290]. style [1218]. Sub [1491]. Sub-Optimal [1491]. Subcommittee [25, 15, 16, 17]. Subcommittees [26]. Subgraph [2732, 2731]. sublanguage [71]. subqueries [1848]. subscribe [1538, 1806]. Subsequence [1025]. Subspace [1490]. subsystem [253]. success [2230]. Succinct [2673, 2674]. Suciu [2044, 2624]. Sudipto [2525]. Suitability [715, 2208]. Suite [1844]. Summaries [2758, 1801]. Summary [152, 55, 2838, 621, 1510, 2754, 1752, 1378, 748, 2719, 2837, 1717, 361, 108, 2391, 439, 283]. Summary-based [2838, 2837]. summer [2125]. super [1270, 927]. supercomputing [1940]. SuperSQL [1564]. supplement [824]. Supply [1920]. Support [1041, 2526, 941, 1415, 1042, 1399, 2355, 1509, 857, 1154, 1529, 1313, 851, 940, 2008, 1859, 356, 2167, 414, 602, 1205, 854, 263, 110, 412, 455, 1782, 928, 1786, 1268, 793, 674, 1907, 260, 660]. Supporting [992, 1912, 1404, 415, 970, 733, 1394, 2691, 2632, 953, 131, 400, 456, 753, 1291, 1780, 314, 524, 1692, 786, 814, 338, 396, 1150, 924, 278, 721, 1832]. Surajit [1417, 2335]. Sure [1066]. Survey [2800, 2515, 2746, 1577, 2715, 2711, 2533, 2551, 2590, 961, 2676, 2639, 2558, 2809, 2853, 2612, 959, 2584, 2550, 2687, 2273, 217, 2225, 364, 2439, 2470, 100, 2376, 1947, 960, 2421, 2489, 978, 2431, 2496, 1062, 2118, 2313, 2367, 2462, 2164, 958, 2282, 1694, 2407, 1746, 2311]. surveys [2442, 102]. Survive [1593, 1089]. Susan [2712]. swarms [2039, 2053]. SWEET’12 [2459]. SWEET’13 [2495]. SWH [2685, 2713, 2757]. Swiss [2237]. Sybase [1032, 811, 1224]. Symmetric [1030, 351]. Symposium [2859, 1464, 2203, 2106, 2153, 1609, 2137, 57]. Synchronization [137, 540, 291, 197]. synchronized [1105]. Synchronizing [1718]. Synopses [1640, 2541, 1813]. Syntax [977, 742, 548]. synthesis [117]. synthetic [1009, 2249]. SYSEO [2482]. System [1382, 1556, 2760, 1328, 1322, 1561, 992, 1060, 1567, 1740, 836, 1418, 1462, 1200, 2539, 1438, 1429, 1310, 1722, 1426, 1085, 956, 1500, 1424, 1057, 1430, 1642, 1762, 1574, 609, 1037, 1082, 1648, 1050, 1421, 2633, 1445, 1550, 1048, 997, 1094, 1562, 943, 1709, 1726, 57, 2856, 1614, 1052, 2559, 1563, 1090, 2508, 1321, 2017, 1124, 1865, 331, 2014, 153, 886, 1173, 1820, 1920, 155, 755, 259, 469, 2381, 2481, 1298, 1278, 2326, 1216, 268, 336, 477, 941, 588, 940, 895, 616, 1898, 332, 1862, 160, 1171, 111, 667, 798, 1172, 919, 281, 1327, 2307, 334, 862, 257, 1309]. system [749, 456, 1187, 1919, 1517, 1133, 1360, 457, 434, 918, 753, 309, 577, 2212, 1181, 1871, 545, 397, 701, 499, 169, 272, 1818, 1311, 1170, 435, 279, 472, 780, 752, 1188, 742, 1180, 398, 1527, 552, 568, 237, 1358, 2324, 1049, 1868, 1899, 381, 1182, 824, 204, 99, 177, 521, 73, 98, 1141, 1984, 353, 810, 218, 313, 1183, 2452, 388, 646, 675, 2025, 928, 1786, 1179, 1333, 926, 864, 389, 1359, 768, 917, 1907, 1883, 387, 280, 1176, 1854, 1348, 2187, 544, 924, 640, 2497, 900, 898, 66, 1987, 123, 916, 304, 2165, 2679, 1542, 605, 2318]. System [221, 2156, 70]. Systems [2859, 2654, 1994, 2626, 2615, 2813, 2336, 1581, 1481, 2435, 1433, 1239, 1675, 1061, 1793, 1677, 1320, 1029, 1326, 1457, 1676, 29, 2513, 1534, 1683, 968, 1533, 1409, 2790, 1411, 593, 2536, 2667, 1461, 2595, 2706, 2108, 1554, 2839, 2635, 2594, 2557, 1727, 2355, 2620, 1704, 2721, 1521, 2455, 1600, 963, 2668, 2410, 1609, 2587, 2151, 2840, 2546, 955, 1072, 2707, 1511, 2213, 550, 2217, 1602, 594, 2634, 959, 2596, 1071, 2525, 2535, 2689, 2756, 1374, 994, 1356, 312, 1261, 1260, 440, 2571, 858, 350, 1236, 2113, 55, 623, 286, 703, 340, 1099, 995, 1227, 317, 1313]. systems [354, 547, 1890, 112, 923, 408, 1931, 624, 1413, 273, 1986, 1819, 2208, 252, 631, 1368, 53, 539, 203, 1806, 2001, 88, 2345, 47, 713, 971, 564, 649, 1146, 102, 1355, 1354, 94, 1775, 380, 2102, 1266, 826, 1100, 708, 316, 314, 1945, 1937, 608, 739, 1993, 1873, 1195, 1825, 345, 2232, 978, 781, 915, 1104, 403, 1289, 626, 2048, 1525, 909, 1842, 2019, 2393, 497, 2022, 889, 928, 788, 682, 49, 50, 209, 828, 479, 324, 138, 1184, 620, 197, 219, 1903, 396, 484, 1346, 1907, 2160, 280, 1345, 1980, 189, 619, 1443, 417, 315, 116, 1126, 744]. systems [493, 1234, 190, 725, 910, 2173, 819, 958, 860, 115, 1214, 193, 721, 548, 592, 198, 2085, 2284, 2166, 1778, 890, 1832, 2069, 137, 62, 1323, 1616, 705, 1065, 944, 945, 1685, 1765]. SystemT [2324]. sytem [2100].

T [1943]. T2 [1475]. TAB [2271]. Table [2798, 2774, 2773, 2079, 679, 2479, 1191, 1341, 30]. table-level [679]. Tableau [2505]. Tables [2738, 1752, 2744, 1378, 2743, 1717, 2737, 1820, 2008, 356, 804, 283, 1939, 1248]. tailoring [1302]. Take [1329]. taken [42]. talking [2353]. Tamer [2440]. Taming [2572]. Tan [2524]. Tandem [2306]. Tapes [891]. Task [32, 29, 67, 2558, 914, 222, 349]. taxis [382]. Taxonomy [2788, 2853, 2479, 1, 301, 2166]. TC [60].

TC-2 [60]. teach [893]. Teaching [2794, 102, 2244]. Team [2053]. teams [2408]. tech [2335]. Technical [2769, 1220, 2694, 2823, 2700, 2729, 2821, 2604, 2172, 2659, 2725, 2841, 1331, 2572, 1325, 2727, 2825, 2692, 2698, 2576, 2775, 2735, 2779, 2673, 2574, 2578, 2608, 2636, 2630, 2665, 2667, 2835, 2706, 2771, 2602, 2661, 2696, 2839, 2635, 2827, 2570, 2831, 2777, 2767, 2663, 2568, 2564, 2606, 2600, 2739, 2781, 2632, 2773, 2702, 2743, 2833, 2704, 2669, 2741, 2837, 2566, 2628, 2634, 2731, 2829, 1040, 2785, 2783, 2671, 2733, 2737, 1245, 60, 2492, 2819]. Technique [1494, 1741, 695, 657, 1193, 307, 82, 665, 1934, 283, 51, 1883]. Techniques [2788, 1638, 1644, 1743, 600, 1090, 1316, 1259, 151, 1942, 1293, 1138, 511, 2337, 252, 288, 258, 1577, 1199, 1301, 1286, 658, 443, 2489, 1269, 1889, 2089, 1253, 536, 2173, 775, 1924]. Technische [2189]. Technological [2645]. Technologies [2106, 2685, 1414, 1707, 1890, 1774, 1685, 2092, 902, 2005]. Technology [1558, 2796, 1030, 1536, 2459, 1668, 1398, 1243, 1478, 2641, 1783, 1303, 1367, 285, 213, 1169, 1458, 2031, 54, 590, 1312, 1167, 1913, 2495, 2465, 2440, 1151, 1221, 830]. telecommunication [905, 453]. telecommunications [991]. telegraph [1864, 1907]. Telescope [2312]. Template [1424]. Template-Based [1424]. Temporal [703, 1216, 975, 962, 2447, 1407, 938, 1215, 1126, 628, 972, 1052, 1237, 870, 2209, 331, 1125, 974, 466, 834, 2083, 2451, 818, 790, 1197, 295, 263, 2444, 973, 1228, 2091, 411, 1807, 1069, 638, 2361, 2169, 2000]. Temporally [2672, 2671]. Ten [1460]. Tenacious [2313]. tension [464]. Tenth [2291]. tenure [2100]. Teorey [566]. Terabyte [1746]. Teradata [1041, 1254]. Term [2795, 2544, 2677, 776]. terminal [111]. termination [762]. terminological [288]. terminology [455]. Terms [1075, 2074]. ternary [1939]. TerraServer [1734, 2310]. tertiary [891, 1266, 1934].

TES [1774]. testbed [1897, 1171, 237, 482]. testing [1192, 761, 447, 462]. Tests [2681]. Text [1506, 124, 1657, 2745, 1085, 999, 978, 2517, 1014, 1090, 1865, 1240, 1811, 2429, 1138, 568, 460, 125, 2266, 1812]. Text-Mining [2517]. Textual [1499, 1555, 105]. texture [308]. tf [2257, 2280]. tf-idf} [2257, 2280]. Thanks [2298]. Thau [2268]. Their [2800, 2669, 2670, 1859, 1000, 738, 121, 1333]. them [2471]. thematic [2000]. theorem [464]. theorems [839, 840]. Theoretical [2504, 2642]. Theoretician [2619]. theorist [1979]. theorists [2808]. Theory [2652, 2356, 2548, 1779, 2657, 2658, 2589, 2242, 224, 1958, 2487, 63, 768, 1903, 1886, 267, 1940]. there [1660, 2309, 239, 250]. These [1523]. They’re [1044]. thing [1926]. Things [2537, 1149, 2552, 2483]. Think [2631, 2630]. Thinksheet [1302]. Third [2859, 616, 887, 366, 1096, 727, 1166, 2615, 2370, 2287, 1481, 2261, 2757, 2553, 2592]. Third-generation [616, 727, 1166]. third-generation/OODBMS [727]. Thor [1277]. Thoughts [372]. threat [1895]. three [791, 2316, 1102, 180, 2116, 2272, 744, 2385]. three-level [180]. three-valued [2316, 2272]. Threshold [2832]. throughput [380, 2448]. Thumb [1460]. tier [1913]. tight [1963]. Time [1235, 2010, 286, 1239, 1323, 2485, 2555, 2583, 968, 1025, 1409, 956, 1877, 2578, 2806, 1394, 2684, 1007, 2579, 2528, 1371, 1701, 421, 1124, 771, 1236, 251, 1227, 217, 1187, 1128, 996, 406, 1230, 1229, 461, 867, 508, 826, 1691, 1809, 1197, 1231, 2358, 1868, 462, 418, 1233, 412, 218, 1228, 1444, 1288, 422, 1854, 417, 301, 1234, 416, 2185, 825, 1214, 1232, 574, 1870, 304, 1778, 2248, 419]. time-based [996]. Time-Series [1025]. timeliness [1234]. Timely [949]. Timer [1687]. Timer-driven [1687]. Timestamp [315]. timing [420]. Timos [2648]. Timothy [2088]. Tips [2180]. Toby [566]. TODS} [2195, 2143, 2220, 2245, 2547, 2529, 2580, 2588, 2610, 2029, 2178, 2245]. Together [2720]. token [106]. tokens [333]. tolerance [786]. tolerant [1948, 1134, 2499, 1255, 1864, 1150]. Too [1731, 1928, 2748, 891, 2348]. Tool [1506, 1423, 1861, 541, 875, 2077, 1860, 459, 1559, 1304, 1302]. Toolbox [1046, 2412]. Toolkit [1399, 1551, 2027]. Tools [2715, 2158, 293, 2271, 2337, 10, 1947, 187, 2649, 1924, 1464]. toolset [491]. Top [2675, 240]. Top- [2675]. Top-down [240]. Topological [1111]. Toronto [2123]. Total [2514, 266]. tough [2231]. Tova [2389]. TP [887, 1875, 706]. TPC [1351, 1350, 2388, 1380, 849, 1163, 2232, 1782]. TPC-C [2388, 849, 1163, 2232]. TPC-D [1351, 1350, 1380]. TPC-E [2388]. TPCC [2232]. TPCC-UVa [2232]. Track [2679]. Tractable [2611]. Trade [985, 1088, 470]. Trade-offs [1088]. tradeoffs [680, 1262]. trades [1870]. Trading [1288, 406, 467]. Traditional [2612, 1117]. Traffic [1760, 2514, 1544, 1868]. train [2252]. Trait [667]. Trajectory [2619]. Transaction [312, 941, 2727, 434, 1409, 1548, 1094, 1045, 2728, 476, 588, 940, 586, 379, 778, 2095, 1142, 754, 392, 867, 1143, 1954, 314, 896, 443, 652, 1923, 2305, 2499, 412, 1062, 788, 465, 796, 2309, 560, 906, 404]. Transaction-oriented [941, 940]. Transactional [1538, 1756, 2771, 1903, 2825, 2826]. Transactions [992, 2823, 1681, 2819, 2820, 994, 2824, 421, 653, 798, 587, 392, 1233, 773, 933, 2309, 795, 315, 416, 2483]. transfer [2335]. transfers [231, 106]. Transformation [2748, 835, 342, 767, 1153]. transformation-based [767]. Transformations [1487, 2404]. Transforming [2342]. transforms [2255]. Transient [1509, 773]. Transition [1040, 623, 522]. Transitive [1028, 525, 401, 613, 571]. Translation [562, 2564, 1181, 2199]. Transmission [1679]. Transparent [2748, 1521]. transportation [2392]. TRANSSOFT [928]. travel [2097]. Traversal [338, 524]. Treasurer [199, 1951, 114, 96]. Tree [2603, 1632, 2602, 1012, 2828, 1838, 598, 1570, 1128, 1110, 2202, 2234, 1011, 1830, 2489, 2073, 1572, 1889, 1190, 686, 2150, 2827, 1401]. Tree-Structured [2603, 2602]. Trees [1095, 1375, 1006, 1748, 1751, 2765, 255, 305, 871, 2008, 256, 661, 782, 428, 1808, 1612, 687, 774, 1111, 595, 284, 688, 262, 1258, 255, 2246]. Trends [2758, 110, 2653, 2031, 2100]. Trento [2446]. tribute [2296, 2144, 2028]. trie [1116]. Trigger [1406, 363]. triggers [806, 522, 587, 1687]. TriGS [1574]. Triton [701]. Trobe [831]. troubleshooting [2089]. true [2085]. Truncation [2841, 2842]. Trust [2847, 2241, 1969, 2385]. Trusted [959, 1912, 819, 958]. Truth [1751, 697]. try [2294]. Tsichritzis [100]. TSIMMIS [1181, 1424, 1557]. TSQL2 [1069, 976, 1068]. TU [2654]. TUM [2518]. tumble [1817]. Tuning [1553, 1727, 2755, 1633, 925, 1412, 2089, 1535, 900, 2335, 2401]. Tuple [562]. tuples [891, 276]. Turbo [1710]. Turbo-charging [1710]. Turing [2530, 2302]. TURL [2774, 2773]. Turmoil [1204]. Tutorial [140, 142, 144, 1169, 1068, 42, 1526, 1529, 1145, 1413, 1528, 1876, 1527, 1769]. tweets [2422]. twente [2127, 964]. twig [2429]. Twin [450]. Two [1259, 2547, 2188, 839, 302, 840, 1249, 1804, 2463]. two-career [2463]. two-dimensional [1249]. two-phase [1804]. TWOMDE [2342]. Type [1722, 158, 159, 506, 76, 271, 91]. typechecking [1927]. typed [2165]. Types [146, 1434, 195, 622, 415, 426, 86, 80, 167, 72, 92, 173, 69, 178, 182, 89, 188, 193, 194, 1936].

UbiData [2119]. Ubiquitous [1535, 2119, 2366]. UC [2804]. UDB [1533]. UDK [1358]. UIS [1359]. Ulisse [111].

Ulm [1462]. UML [2227]. UML’05 [2227]. Unbundled [2771]. Unbundling [1470, 1471]. Uncertain [2279, 2244]. uncertainty [2034, 662]. Undergraduate [29, 2014]. underlying [2401]. Understanding [44, 56, 2774, 2376, 2565, 2032, 1707, 2773, 2296, 2330, 76, 1837]. Unified [2854, 1410, 1037, 2720, 883, 2134]. Uniform [2703, 2702, 163, 502]. unifying [526]. Uninterruptible [2617]. unions [266]. UniSQL [1327, 1328, 1037, 1064]. UniSQL/X [1037]. units [2319]. Unity [2123]. Univ [830]. Universal [1530, 2821, 2822, 1541, 695, 1960, 238, 451, 193, 2209]. Universität [2189]. University [1220, 2473, 2492, 1593, 1325, 1221, 2123, 196, 559, 1202, 620, 1219, 2483, 964, 2500, 703, 2645, 216, 1362, 1578, 2796, 2172, 1965, 1462, 1684, 2050, 1461, 1446, 2455, 948, 2151, 2446, 748, 1983, 1101, 2361, 820, 2171, 1244, 1964, 831, 2484]. UNIX [1041, 1672, 1089]. unknowns [1900]. Unleashing [2523]. Unnesting [1486]. unorthodox [119]. Unpacking [1602]. unstructured [1293, 2325]. Untrusted [1592]. upcoming [2130]. Update [1626, 962, 1649, 1551, 1136, 1124, 445, 1149, 2202, 1265, 482, 1237, 310]. Updates [2610, 1014, 575, 2379, 1276, 294, 881, 804, 1377]. Updating [669, 1831, 328]. Upon [1329]. Upsizing [1140]. Urbana [1964]. Urbana-Champaign [1964]. usability [2197, 223]. Usage [1591, 961, 1602, 1543, 960]. USASI [3]. USD [752]. Use [1158, 41, 1610, 2701, 177, 277, 153, 131, 1897, 2438, 2427, 158, 428, 69, 84, 2649]. Used [2505, 1199]. useful [816]. User [136, 952, 1677, 569, 1560, 67, 1678, 606, 350, 735, 1912, 1127, 1192, 425, 372, 737, 1102, 208, 1999, 310]. User-Centered [1678]. User-Defined [1677]. user-friendly [606]. User-oriented [1560]. Users [1541, 168, 739, 2163, 2353, 2649]. Using [653, 771, 858, 2074, 777, 2682, 1087, 1508, 1465, 1002, 1499, 1555, 1504, 970, 1702, 1187, 1723, 1084, 2576, 1006, 1711, 1498, 939, 1501, 1682, 1776, 1900, 859, 1509, 1634, 2332, 1636, 1717, 1105, 1823, 578, 1783, 1303, 871, 1898, 273, 2604, 285, 288, 1693, 344, 258, 1249, 697, 1835, 1824, 102, 1779, 347, 1291, 780, 1829, 789, 1785, 1728, 934, 464, 1822, 805, 911, 1761, 1655, 2814, 1935, 1777, 284, 1883, 289, 1357, 2755, 262, 267, 914, 698]. USL [551]. USSR [2738, 2737]. UT [1995]. utilities [739]. utility [1513]. utilization [530]. utilizing [232, 91]. UVa [2232].

V [2586, 2122, 932, 2313]. vacations [2230]. Valid [2507, 1701, 1197]. validating [1888]. validation [330]. validity [1953, 1938]. Value [2843, 2771, 1707, 2772, 1289, 2284, 2353]. Value-Based [1707]. valued [2316, 564, 2272, 698]. values [697, 2317, 520, 1826, 2251, 438]. vanishing [2252]. Vardi [2205]. variable [428, 1631, 1689]. variable-length [428]. variables [83, 89]. Variance [2595]. Variant [1373, 1075]. vectorization [2382]. vectorwise [2417]. vehicle [888]. vendor [812]. vendors [1089]. verifiable [88]. Verification [2513, 852, 295]. veritable [2212]. Verlag [1977, 2122]. VERSANT [1150]. Version [666, 358, 1075, 540, 1098, 727]. Versioning [939, 1887, 773]. versions [2432]. versus [530, 65, 642, 3, 2222, 1966, 1926, 2408, 2441]. Vertical [543, 1710]. Very [2341, 2371, 1759, 105, 547, 2061, 2428, 1806, 103, 380, 804, 1940, 1257]. via [2734, 2388, 873, 2609, 2672, 2608, 1804, 2607, 936, 2671, 2733]. Video [2763, 607, 1603, 1081, 1743, 1614, 1212, 1988, 1135, 1133, 1825, 932]. video-on-demand [1133, 1825]. VideoAnywhere [1614]. Vienna [820]. View [1405, 1404, 1648, 1739, 1403, 1719, 1211, 1130, 1861, 360, 1290, 356, 386, 410, 167, 2049, 119, 2431, 1822, 63, 184, 1288, 1312, 239, 901]. viewpoint [1363, 143]. Views [2682, 1702, 1610, 1503, 226, 1678, 1468, 594, 1509, 668, 1690, 1823, 669, 328, 2197, 302, 1824, 1131, 863, 1121, 1779, 168, 172, 1132, 275]. VII [29]. Viktor [2750]. Virtual [1590, 47, 1458, 2142, 1291, 520, 2096, 2357, 859, 1854]. VisDB [1180]. Vishnu [456]. Visibility [92]. Vision [2136, 2287, 2759, 2538, 2720, 2468, 1781, 2465]. Visual [605, 2735, 1421, 1395, 2736, 765, 1420, 883, 945]. visualisation [1339]. Visualization [1418, 2596, 1313, 1334, 919, 1117, 1249, 1342, 1336, 1299]. Visualizations [2505, 736, 1340]. Visualizing [736, 1180, 2422, 933]. Vitter [2472]. VLDB [2798, 2254, 2279, 2288, 2766, 1774, 1142, 215, 2754, 2807]. VLDB’99 [1582]. VLDL [2341]. VLSI [321, 667, 908]. VMS [793]. Vocabulary [1645]. VODAK [933]. Voisard [2045]. Volatile [2756]. Volcano [577]. voting [391]. VQBD [1856]. vs [2388, 718, 719, 661, 1699, 1771]. VSkyline [2382].

walk [2478]. Walks [2607]. walkthrough [2357, 1854]. Wall [1412]. WALRUS [1650]. Wander [2607]. Wandering [2606]. Wang [1943]. want [2649]. WAP [1783]. ware [1805]. Warehouse [1734, 1760, 1043, 1621, 1431, 1649, 1712, 1378, 1403, 2027, 1872, 1153, 1771]. Warehouses [1405, 1610, 1648, 1468, 1686, 2004, 2046, 2349, 1773]. Warehousing [1415, 1031, 2291, 2239, 1840, 1367, 1296, 1944, 1946, 1297, 1771, 1130]. warfare [1904]. Wars [1315, 1332, 1331]. Washington [2869, 2500]. water [2035]. Waterloo [2859]. Wave [1402]. Wave-Indices [1402]. Wavelet [1520]. Wavelet-Based [1520]. Wavelets [1634]. Way [1751, 2115]. Wayang [2854]. WDAS [2214]. weak [575]. wears [2401]. Weaving [2342]. Web [1568, 2044, 2216, 2354, 2394, 2241, 2323, 2068, 2685, 2713, 2757, 2510, 2561, 2107, 2449, 2592, 2043, 2122, 2276, 2103, 2084, 2158, 1847, 2288, 2204, 1971, 1450, 2450, 1783, 1890, 1591, 1968, 1857, 1974, 2331, 1841, 1565, 1611, 1738, 1545, 1555, 1566, 1869, 1658, 1587, 2097, 1656, 2274, 2131, 1438, 1429, 1517, 1969, 1577, 1245, 1660, 1975, 1732, 2083, 2451, 1973, 1885, 1342, 1802, 2049, 1763, 2533, 2376, 1414, 1358, 1699, 1739, 1560, 1993, 1947, 2344, 2479, 2444, 1477, 1970, 1550, 2162, 2092, 1875, 2517, 1999, 2471, 1544, 1295, 1833, 1943, 2023, 1682, 1782, 1980, 1666, 1428, 2007]. Web [2275, 2252, 1151, 1765, 1679]. Web-available [2344]. Web-Base [1550]. Web-based [1358, 1993, 1765]. Web-scale [2331]. Web-Site [1438, 1517]. WebDB [2394]. WebDB’98 [1568]. Webpage [2330]. Websphere [1844]. weight [1186]. Weikum [2348]. went [2294]. were [250]. Wesley [2043, 66]. WG5.8 [2340]. Whang [2267]. what-if [1513]. Where [22, 1478, 2853, 224]. whether [2294]. while [2197]. WHIPS [1431]. white [1245]. Who [22, 827]. Whole [2692, 2693, 2205]. Whom [22]. Whose [2806]. Wide [2083, 1342, 1295, 1943, 1428, 1436, 1151, 1791, 901, 1577]. wide-area [1791]. WIDM [2216, 2354, 2323]. Widom [2230]. Wiki [2413]. Wikipedia [2332]. wild [1847]. wildcards [1274]. Wiley [2027]. Wilkes [2176]. Wilkinson [2681]. Will [215, 1478, 1874, 1925]. win [2302]. Window [1649, 2196, 2133]. Windows [233, 1143]. winds [832]. winning [2448]. Wins [2530]. Wireless [2139, 1210, 2037, 1945, 2038]. Wisconsin [843, 216]. wisdom [2300]. Wish [2846, 250]. wishes [2397]. within [928]. Without [1633, 1499, 777, 1708, 83, 481, 804, 1525, 1984, 2617, 835, 2524]. Witten [2088]. Wizard [1553]. Won [1064]. Wonderland [2787]. won’t [1145, 1525]. WOODSS [2162]. Work [2504, 2710, 2712, 1755, 953, 2815, 55, 2036, 361, 1155, 905, 821, 1525, 1940, 2433]. work-in-progress [2036]. workability [2411, 2362]. Workaholics [988]. Workbench [1436]. Workflow [1675, 2229, 1587, 2459, 1554, 1476, 1167, 1436, 1456, 1300, 1346, 2160, 1345, 2495, 2166]. Workflow-based [1587]. Workflows [941, 940, 2406, 2167, 2159, 2163, 2161, 2162, 2210, 2168, 2165]. working [1362, 1932, 1524]. workload [1814, 2825, 2826]. workload-aware [1814]. Workloads [2856, 1379]. Workshop [2626, 2615, 2238, 2813, 2370, 2136, 2287, 13, 20, 38, 152, 2864, 1463, 1568, 2278, 2139, 2475, 1239, 1675, 1323, 2216, 1569, 2228, 1616, 2261, 2485, 2555, 2354, 1465, 2554, 2270, 2537, 1676, 2229, 2394, 2323, 2215, 2459, 2289, 2277, 945, 1763, 2685, 2713, 2757, 2510, 2561, 2391, 2254, 2290, 2107, 2372, 2511, 2341, 2371, 2512, 2599, 2426, 2449, 1456, 2410, 2467, 2656, 2070, 2851, 838, 2684, 2342, 2291, 2466, 2494, 2528, 2553, 2592, 2214, 1215, 2191, 2154, 2239, 2255, 1765, 2262, 2339, 2279, 2340, 622, 1455, 703, 2379, 1099, 2174, 2363, 2428, 1792, 1774, 1791]. workshop [2493, 2458, 1685, 2241, 621, 1686, 2047, 1946, 2127, 316, 1945, 1993, 2046, 1764, 2126, 2380, 2092, 973, 620, 1346, 1345, 2495, 429, 2173, 2227, 2240, 2069, 2292, 2356, 14, 1457, 944, 553, 2108, 2355, 2719]. Workshops [2213]. Workstations [1390, 371]. World [2778, 1465, 1380, 2777, 2816, 1154, 797, 233, 1111, 2176, 1577, 2083, 1342, 1295, 1943, 1428, 1151]. World-Wide [1577]. Worlds [2780, 2779, 2547, 375, 1978]. WorldWide [2312]. WORM [530]. worst [346]. worst-case [346]. wouldn’t [1926]. WPES [2290]. Wrapper [1449, 2079, 2173]. Wrappers [1424]. Wrapping [1720, 1885]. Write [2827, 579, 2828, 653, 780, 805, 2401, 2463]. write-ahead [780, 805]. Write-only [579, 653]. Write-Optimized [2827, 2828]. writes [786, 2156]. writing [2309]. written [1064]. WS [2204]. WS-DAI [2204]. WSQ [1732]. WSQ/DSQ [1732]. Wu [2815]. Württemberg [1359]. WWW [1358, 1768]. WWW-UDK [1358]. WWW’2004 [2107].

X [1858, 1037, 2344]. X-ray [2344]. X3 [62, 55, 222, 9, 3]. XDB [2607]. XEWA [1791]. XEWA-00 [1791]. XIME [2261, 2126]. XIME-P [2261, 2126]. XLDB [2460]. Xmas [1562]. XMill [1721]. XML [2322, 1855, 2044, 1881, 1690, 2233, 2242, 2379, 1858, 1783, 1706, 2147, 2515, 1789, 2009, 1792, 2429, 1887, 1720, 1791, 1989, 1949, 2105, 1805, 1722, 1885, 2167, 2059, 2127, 2333, 2145, 2344, 1873, 2590, 1772, 2098, 2489, 1704, 1721, 2399, 2507, 2199, 2265, 1586, 2092, 2351, 1958, 1833, 1682, 2112, 1884, 2251, 1883, 1886, 1927, 2157, 1831, 1911, 2125, 2282, 2223, 2007]. XML-to-SQL [2199]. XPath [1985, 1998, 2149, 2066, 2259]. XPRS [758]. XPS [1162]. XQJ [2087, 2365]. XQuery [2261, 2126, 2250, 1982, 2087, 2192, 2480, 1882, 2182, 2365, 1936]. XSB [1053, 1027]. XTRACT [1722].

YAGO [2329]. YAGO-NAGA [2329]. Yahoo [2269, 2275]. Yannis [2244]. Yates [2275]. Year [2807, 1040, 2401]. Years [1460, 2418]. Yogyakarta [2263]. York [16]. Young [2267]. yourself [2090, 2085].

Zaniolo [2591]. Zhang [2765]. ZOO [1437]. Zurich [2336, 982, 981].

References

Olle:1969:TDD

[1] T. William Olle. A taxonomy of data definition languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 1(1):24–28, August 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mooers:1969:DDL

[2] Calvin N. Mooers. Data descriptive languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 1(1):31–36, August 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mooers:1969:UXA

[3] Calvin N. Mooers. For USASI X3 ad hoc committee on data descriptive language: ‘external versus internal formats’. SIGMOD Record (ACM Special Interest Group on Management of Data), 1(1):37, August 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bobrow:1969:MFC

[4] Daniel Bobrow and William R. Sutherland. Message forms in computer networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 1(1):38–42, August 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hatfield:1969:FSR

[5] Don Hatfield. Finite state representation of interactive languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 1(1):43, August 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

See:1969:FSR

[6] Richard See. Finite state representation of interactive languages. SIG-MOD Record (ACM Special Interest Group on Management of Data), 1 (1):44–45, August 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McGee:1969:DDD

[7] W. C. McGee. Data description for data independence. SIGMOD Record (ACM Special Interest Group on Management of Data), 1(2):3–10, December 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rodriguez:1969:PPD

[8] J. E. Rodriguez. Program portability, data manipulation languages, and data description languages. SIG-MOD Record (ACM Special Interest Group on Management of Data), 1 (2):11–13, December 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gosden:1969:RXD

[9] J. A. Gosden. Report to X3 on data definition languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 1(2):14–25, December 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DImperio:1969:IST

[10] M. E. D’Imperio. Information structures: tools in problem solving. SIG-MOD Record (ACM Special Interest Group on Management of Data), 1 (2):25–51, December 1969. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wright:1971:CM

[11] Kendall R. Wright. Chairman’s message. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(1):1, April 1971.

OConnell:1971:ER

[12] Daniel F. O’Connell. Editor’s remarks. SIGMOD Record (ACM Special Interest Group on Management of Data), 3 (1):2–3, April 1971.

Anonymous:1971:ASW

[13] Anonymous. 1971 ACM—SIGFIDET Workshop on Data Description, Access and Control: Call for papers. SIG-MOD Record (ACM Special Interest Group on Management of Data), 3(1): 4–5, April 1971.

Codd:1971:IAS

[14] E. F. Codd. Impressions of the 1970 ACM-SIGFIDET Workshop. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(1):6, April 1971.

Hatfield:1971:NES

[15] Don Hatfield. New England subcommittee. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(1):6, April 1971.

Manelski:1971:NYS

[16] Denis Manelski. New York subcommittee. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(1):7, April 1971.

Raichelson:1971:SES

[17] Gene Raichelson. South Eastern subcommittee. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(1):7, April 1971.

Karpinski:1971:OLD

[18] R. H. Karpinski. Open letter from Dick Karpinski. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(1):8, April 1971.

OConnell:1971:EN

[19] Daniel F. O’Connell. Editor’s notes. SIGMOD Record (ACM Special Interest Group on Management of Data), 3 (2):3–5, October 1971.

Anonymous:1971:SWP

[20] Anonymous. 1971 SIGFIDET Workshop Program. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(2):6–8, October 1971.

Wright:1971:SO

[21] Kendall R. Wright. A SIGFIDET outline. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(2):9, October 1971.

Anonymous:1971:SOW

[22] Anonymous. SIGFIDET officers — who, with whom or what, where. SIG-MOD Record (ACM Special Interest Group on Management of Data), 3(2): 10, October 1971.

Tonik:1971:FF

[23] A. B. Tonik and E. F. Codd. FDT Forum. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(2):11–13, October 1971.

Anonymous:1971:MN

[24] Anonymous. Members in the news. SIGMOD Record (ACM Special Interest Group on Management of Data), 3 (2):13, October 1971.

Brown:1971:RSR

[25] Peter S. Brown and Gene Raichelson. Regional subcommittee reports. SIG-MOD Record (ACM Special Interest Group on Management of Data), 3(2): 13–14, October 1971.

Anonymous:1971:PFR

[26] Anonymous. On present and future regional subcommittees. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(2):15, October 1971.

Anonymous:1971:CP

[27] Anonymous. Calls for papers. SIG-MOD Record (ACM Special Interest Group on Management of Data), 3(2): 16, October 1971.

Anonymous:1971:MNN

[28] Anonymous. Miscellaneous news and notes. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(2):17, October 1971.

Denning:1971:ACT

[29] Peter J. Denning. Announcement CO-SINE Task Force VII report on an undergraduate course on operating systems principles. SIGMOD Record (ACM Special Interest Group on Management of Data), 3(2):18, October 1971.

Sieglaff:1972:SAT

[30] Harold P. Sieglaff. Stable or archival table. SIGMOD Record (ACM Special Interest Group on Management of Data), 4(1):11–12, April 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1972:EN

[31] Anonymous. Editor’s note. SIGMOD Record (ACM Special Interest Group on Management of Data), 4(2):3, July 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1972:BCC

[32] Anonymous. BCS Conference on CO-DASYL Data Base Task Group Proposal. SIGMOD Record (ACM Special Interest Group on Management of Data), 4(2):4–5, July 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1972:SDB

[33] Anonymous. SHARE Data Base Project’s Data Independence Committee meets. SIGMOD Record (ACM Special Interest Group on Management of Data), 4(2):5–, July 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Browne:1972:RSM

[34] Peter S. Browne. Report on SIG-FIDET meeting, SJCC, May 17, 1972. SIGMOD Record (ACM Special Interest Group on Management of Data), 4 (2):6–7, July 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

OConnell:1972:EN

[35] Daniel O’Connell. Editor’s notes. SIGMOD Record (ACM Special Interest Group on Management of Data), 4(3):3, November 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Canning:1972:SIS

[36] Richard H. Canning. Statement of interests of SIGFIDET and SIGBDP. SIGMOD Record (ACM Special Interest Group on Management of Data), 4(3):4, November 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Karpinski:1972:WHP

[37] Richard H. Karpinski. What’s happening with PL/1 data base. SIG-MOD Record (ACM Special Interest Group on Management of Data), 4(3): 5, November 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1972:FPA

[38] Anonymous. Final program for 1972 ACM-SIGFIDET Workshop on Data Description, Access, and Control – registration information. SIGMOD Record (ACM Special Interest Group on Management of Data), 4(3):6–11, November 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1972:AN

[39] Anonymous. Availability notices. SIG-MOD Record (ACM Special Interest Group on Management of Data), 4 (3):12, November 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1972:C

[40] Anonymous. Corrigenda. SIGMOD Record (ACM Special Interest Group on Management of Data), 4(3):12, November 1972. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buschman:1973:UIF

[41] A. Camille Buschman. Use of an inverted file structure for interactive retrieval of computer program abstracts. SIGMOD Record (ACM Special Interest Group on Management of Data), 5(1):3–12, June 1973. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tellier:1973:TWD

[42] Harrison Tellier. Tutorial: what is a data base? (abstracts taken from an original paper presented during the SIGBDP, SIGFIDET tutorial sessions, ACM 72). SIGMOD Record (ACM Special Interest Group on Management of Data), 5(1):13–43, June 1973. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lyon:1973:WD

[43] John K. Lyon. What is a database. SIGMOD Record (ACM Special Interest Group on Management of Data), 5(1):44–62, June 1973. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Codd:1973:UR

[44] E. F. Codd. Understanding relations. SIGMOD Record (ACM Special Interest Group on Management of Data), 5(1):63–64, June 1973. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sieglaff:1973:SF

[45] Harold P. Sieglaff. The STREET-WALKER FILE. SIGMOD Record (ACM Special Interest Group on Management of Data), 5(2):8–13, November 1973. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Little:1973:AES

[46] John L. Little. ASCII and EBCDIC sequencing. SIGMOD Record (ACM Special Interest Group on Management of Data), 5(2):14–16, November 1973. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Folinus:1974:VID

[47] Jeffery J. Folinus, Stuart E. Madnick, and Howard B. Schutzman. Virtual information in data-base systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 6(2):1–15, March 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sibley:1974:SRC

[48] E. H. Sibley. Some reviews of current literature. SIGMOD Record (ACM Special Interest Group on Management of Data), 6(2):6–7, 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rabin:1974:CIB

[49] Jonas Rabin and Cheng W. Cheng. Common interface between programming languages and data base management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 6(2):8, 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rabin:1974:SPI

[50] Jonas Rabin and Cheng W. Cheng. Structured programming and integrated data-base systems design. SIG-MOD Record (ACM Special Interest Group on Management of Data), 6(2): 8–9, 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roach:1974:ATD

[51] Vincent Roach. An automated technique for designing optimal performance IMS data bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 6(2):16–23, March 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Swenson:1974:SRC

[52] J. R. Swenson. Some reviews of current literature. SIGMOD Record (ACM Special Interest Group on Management of Data), 6(2):26–28, March 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dummer:1974:DSA

[53] D. C. Dummer and M. W. O’Reilly. Data and systems administration. SIG-MOD Record (ACM Special Interest Group on Management of Data), 6 (2):29–30, March 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Honig:1974:BDB

[54] William L. Honig. Bringing data base technology to the programmer. SIG-MOD Record (ACM Special Interest Group on Management of Data), 6(3): 2–15, July 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bachman:1974:SCW

[55] Charles W. Bachman. Summary of current work ANSI/X3/SPARC/study group: database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 6(3):16–39, July 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Codd:1974:UR

[56] E. F. Codd. Understanding relations. SIGMOD Record (ACM Special Interest Group on Management of Data), 6(3):40–42, July 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rothnie:1974:RPF

[57] James B. Rothnie. Review: Proceedings of the Fourth Annual Computer Related Information System Symposium. SIGMOD Record (ACM Special Interest Group on Management of Data), 6(3):43–44, July 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shneiderman:1974:ODB

[58] Ben Shneiderman. Opportunities for data base reorganization. SIGMOD Record (ACM Special Interest Group on Management of Data), 6(4):1–8, December 1974. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Minsky:1974:ALD

[59] Naftaly Minsky. Another look at databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 6(4):9–17, December 1974.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Taylor:1975:RIT

[60] Robert W. Taylor. Report on IFIP TC-2 conference: ’A technical in-depth evaluation of the DDL’. SIGMOD Record (ACM Special Interest Group on Management of Data), 7(1):5–9, January 1975. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Purdy:1975:GCD

[61] J. Gerry Purdy. A graduate course in database management. SIGMOD Record (ACM Special Interest Group on Management of Data), 7(1):10–15, January 1975. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1975:IRA

[62] Anonymous. Interim report: ANSI/ X3/SPARC Study Group on Data Base Management Systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 7(2):1–140, April 1975. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Olle:1975:PVR

[63] T. William Olle. A practitioner’s view of relational data base theory. SIG-MOD Record (ACM Special Interest Group on Management of Data), 7(3– 4):29–43, September 1975. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hardgrave:1975:DRS

[64] W. T. Hardgrave and E. H. Sibley. Database research: some comments on future directions. SIG-MOD Record (ACM Special Interest Group on Management of Data), 7(3– 4):44–48, September 1975. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Clymer:1975:CDD

[65] John Clymer. Comments on the debate: data-structure-set versus relational. SIGMOD Record (ACM Special Interest Group on Management of Data), 7(3–4):49–52, September 1975. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Weldon:1975:RID

[66] Jay-Louise Weldon. Review of ’An introduction to database systems’ by C. J. Date. Addison-Wesley Publishing Co. 1975. SIGMOD Record (ACM Special Interest Group on Management of Data), 7(3–4):53–54, September 1975. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lefkovits:1976:PRA

[67] Henry C. Lefkovits. A progress report on the activities of the CODA-SYL End User Facility Task Group. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(1):1–35, January 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nordstrom:1976:OMM

[68] Bengt Nordström. An outline of a mathematical model for the definition and manipulation of data. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):1–11, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Linden:1976:UAD

[69] Theodore A. Linden. The use of abstract data types to simplify program modifications. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):12–23, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zloof:1976:SBA

[70] Moshe M. Zloof and S. Peter de Jong. The System for Business Automation (SBA): Programming language. SIG-MOD Record (ACM Special Interest Group on Management of Data), 8(2): 24, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Allman:1976:ERD

[71] Eric Allman, Michael Stonebraker, and Gerald Held. Embedding a relational data sublanguage in a general purpose programming language. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):25–35, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Johnson:1976:ADT

[72] Robert T. Johnson and James B. Morris. Abstract data types in the Model programming language. SIG-MOD Record (ACM Special Interest Group on Management of Data), 8 (2):36–46, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McLeod:1976:HLD

[73] Dennis J. McLeod. High level domain definition in a relational date base system. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):47–57, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hammer:1976:DAD

[74] Michael Hammer. Data abstractions for data bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):58–59, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Horning:1976:SDP

[75] J. J. Horning. Some desirable properties of data abstraction facilities. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):60–62, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:1976:TFU

[76] Douglas T. Ross. Toward foundations for the understanding of type. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):63–65, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shaw:1976:RDA

[77] Mary Shaw. Research directions in abstract data structures. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):66–68, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mitchell:1976:NSD

[78] Jim Mitchell and Ben Wegbreit. A next step in data structuring for programming languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):69–70, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Smith:1976:DBA

[79] John M. Smith and Diane C. P. Smith. Data base abstraction. SIG-MOD Record (ACM Special Interest Group on Management of Data), 8(2): 71, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guttag:1976:ADT

[80] John Guttag. Abstract data types and the development of data structures. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):72, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bubenko:1976:IRD

[81] J. A. Bubenko, Jr., S. Berild, E. Lindencrona-Ohlin, and S. Nachmens. From information requirements to DBTG-data structures. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):73–85, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hardgrave:1976:TIS

[82] W. T. Hardgrave. A technique for implementing a set processor. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):86–94, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kieburtz:1976:PPV

[83] Richard B. Kieburtz. Programming without pointer variables. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):95–107, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vallarino:1976:UBM

[84] Oscar Vallarino. On the use of bit maps for multiple key retrieval. SIG-MOD Record (ACM Special Interest Group on Management of Data), 8 (2):108–114, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berry:1976:CDR

[85] D. M. Berry, Z. Erlich, and C. J. Lucena. Correctness of data representations (extended abstract): Pointers in high level languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):115–119, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gries:1976:SID

[86] David Gries and Narain Gehani. Some ideas on data types in high level languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):120, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Senko:1976:DIB

[87] Michael E. Senko. DIAM II: The binary infological level and its database language — FORAL. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):121–140, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Flon:1976:TCV

[88] L. Flon and A. N. Habermann. Towards the construction of verifiable software systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):141–148, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Parnas:1976:ATD

[89] D. L. Parnas, John E. Shore, and David Weiss. Abstract types defined as classes of variables. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):149–154, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Housel:1976:HLD

[90] Barron C. Housel and Nan C. Shu. A high-level data manipulation language for hierarchical data structures. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):155–169, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wells:1976:DTE

[91] Mark B. Wells and Fred L. Cornwell. A data type encapsulation scheme utilizing base language operators. SIG-MOD Record (ACM Special Interest Group on Management of Data), 8 (2):170–178, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Koster:1976:VT

[92] Cornelis H. A. Koster. Visibility and types. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(2):179–190, March 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lacroix:1976:GJ

[93] M. Lacroix and A. Pirotte. Generalized joins. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(3):14–15, September 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hardgrave:1976:PBA

[94] W. T. Hardgrave and E. H. Sibley. Problems of basic and applied research in database systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 8 (3):16–22, September 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berg:1976:DBD

[95] John L. Berg. Data base directions: the next steps. SIGMOD Record (ACM Special Interest Group on Management of Data), 8(4):.3–.4, November 1976. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Joyce:1977:STR

[96] John J. Joyce. Secretary Treasurer’s reports. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(1):3–14, June 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bontempo:1977:RBP

[97] Charles J. Bontempo. Reviews of books and papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(1):29–30, June 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Minker:1977:RED

[98] Jack Minker. Review of ’Evaluation of data-base management systems’ by J. P. Fry and E. H. Sibley. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(1):30–34, June 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manola:1977:RRD

[99] Frank Manola. Review of ’Relational data-base management systems’ by Donald D. Chamberlain. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(1):34–37, June 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kent:1977:RHD

[100] W. Kent. Review of ’Hierarchical data-base management: a survey’ by D. C. Tsichritzis and F. H. Lochovsky. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(1):38–40, June 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Altshuler:1977:RCR

[101] Gene Altshuler. Review of ’A comparison of relational and CODASYL approaches to data-base management’ by A. S. Michaels, B. Mittman, and C. R. Carlson. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(1):40–42, June 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gudes:1977:TDB

[102] Ehud Gudes. Teaching data base systems using date and computing surveys. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(1):47–49, June 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gerritsen:1977:SMD

[103] Rob Gerritsen, Howard Morgan, and Michael Zisman. On some metrics for databases or what is a very large database? SIGMOD Record (ACM Special Interest Group on Management of Data), 9(1):50–74, June 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vineberg:1977:ICS

[104] Maniel Vineberg. Implementation of character string pattern matching on a multiprocessor. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(2):1–7, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bird:1977:APP

[105] R. M. Bird, J. C. Tu, and R. M. Worthy. Associative/parallel processors for searching very large textual data bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(2):8–9, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lipovski:1977:IFT

[106] G. J. Lipovski. On imaginary fields, token transfers and floating codes in intelligent secondary memories. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(2):17–22, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zaky:1977:MNN

[107] S. G. Zaky. Microprocessors for nonnumeric processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(2):23–30, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hsiao:1977:ADC

[108] David K. Hsiao and Krishnamurthi Kannan. The architecture of a database computer — a summary. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(2):31–33, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenthal:1977:DMM

[109] Robert S. Rosenthal. The data management machine, a classification. SIG-MOD Record (ACM Special Interest Group on Management of Data), 9(2): 35–39, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McDonell:1977:TNS

[110] Ken J. McDonell. Trends in nonsoftware support for input-output functions. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(2):40–47, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cerretti:1977:UIP

[111] R. Cerretti, D. Jasilli, and D. R. Matteucci. Ulisse: An Italian project for a multifunctional terminal system. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(2):48–50, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bray:1977:DMR

[112] Olin H. Bray. Data management requirements: The similarity of memory management, database systems, and message processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(2):68–76, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Landson:1977:CSA

[113] Barry M. Landson and Robert G. Sargent. A comparison of sequential and associate computing of priority queues. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(2):77–78, May 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Joyce:1977:PST

[114] John D. Joyce. From the past Secretary Treasurer. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(3):5–10, December 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tsichritzis:1977:RDD

[115] D. Tsichritzis. Research directions in data base management systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 9 (3):26–41, December 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sinutko:1977:DMS

[116] M. Sinutko. Database management systems on a multi-computer computer. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(3):42–62, December 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Odell:1977:CSS

[117] James J. Odell. Customized segment synthesis. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(3):63–68, December 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Party:1977:BCS

[118] The Data Dictionary Systems Working Party. The British Computer Society Data Dictionary Systems Working Party report. SIGMOD Record (ACM Special Interest Group on Management of Data), 9(4):2–24, December 1977. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Korfhage:1978:DPU

[119] R. R. Korfhage, W. H. E. Day, L. L. Beck, and W. F. Appelbe. Data physics: an unorthodox view of data and its implications in data processors. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):1–7, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Copeland:1978:SSS

[120] George P. Copeland. String storage and searching for data base applications: implementation on the INDY backend kernel. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):8–17, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Otis:1978:ERD

[121] Allen J. Otis and George P. Copeland. Editing requirements for data base applications and their implementation on the INDY backend kernel. SIG-MOD Record (ACM Special Interest Group on Management of Data), 10 (1):18–29, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lipovski:1978:SPI

[122] G. Jack Lipovski. Semantic paging on intelligent discs. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):30–34, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Williams:1978:MSD

[123] Rhon Williams. A multiprocessing system for the direct execution of LISP.

SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):35–41, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bird:1978:TFI

[124] R. M. Bird, J. B. Newsbaum, and J. L. Trefftzs. Text file inversion: an evaluation. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):42–50, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roberts:1978:SCA

[125] David C. Roberts. A specialized computer architecture for text retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):51–59, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stucki:1978:CCA

[126] M. J. Stucki, J. R. Cox, G. C. Roman, and P. N. Turcu. Coordinating concurrent access in a distributed database architecture. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):60–64, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gouda:1978:HCC

[127] Mohamed G. Gouda. A hierarchical controller for concurrent accessing of distributed databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 10 (1):65–70, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gavish:1978:EAD

[128] Bezalel Gavish and Harvey Koch. An extensible architecture for data flow processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):71–76, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Harvill:1978:FPO

[129] J. B. Harvill. Functional parallelism in an operand state saving computer. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):77–84, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hutchison:1978:MM

[130] J. S. Hutchison and W. G. Roman. Madman machine. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):85–90, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Banerjee:1978:UDM

[131] Jayanta Banerjee and David K. Hsiao. The use of a database machine for supporting relational databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 10 (1):91–98, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sadowski:1978:EPR

[132] Paul J. Sadowski and S. A. Schuster. Exploiting parallelism in a Relational Associative Processor. SIG-MOD Record (ACM Special Interest Group on Management of Data), 10 (1):99–109, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:1978:BRD

[133] Hsu Chang. Bubbles for relational database. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):110–116, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

ElMasri:1978:MIR

[134] A. El Masri, J. Rohmer, and D. Tusera. A machine for information retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 10(1):117–120, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Matteucci:1978:DSA

[135] Dante R. Matteucci. A distributed structure for the automization of the catalog of the national cultural heritage: experiences and proposals. SIG-MOD Record (ACM Special Interest Group on Management of Data), 10 (1):121–133, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1978:SRA

[136] Anonymous. Status report on the activities of the CODASYL End User Facilities Committee (EUFC). SIG-MOD Record (ACM Special Interest Group on Management of Data), 10 (2–3):1–26, August 1978. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:1981:EPS

[137] Philip A. Bernstein, Marco A. Casanova, and Nathan Goodman. Errors in ’Process Synchronization in Database Systems’. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(1):9–29, January 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rowe:1981:AFD

[138] Lawrence A. Rowe and Michael Stonebraker. Architecture of future data base systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(1):30–44, January 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mylopoulos:1981:OKR

[139] John Mylopoulos. An overview of Knowledge Representation. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):5–12, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:TAI

[140] Anonymous. Tutorial on artificial intelligence research. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):13–18, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McLeod:1981:AD

[141] Dennis McLeod and John Miles Smith. Abstraction in databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):19– 25, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:TDR

[142] Anonymous. Tutorial on database research. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):26–28, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rowe:1981:DAP

[143] Lawrence A. Rowe. Data abstraction from a programming language viewpoint. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):29–35, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:TPL

[144] Anonymous. Tutorial on programming language research. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):36–39, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:WSM

[145] Anonymous. What should be modelled? SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):40–42, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:T

[146] Anonymous. Types. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):43–52, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:B

[147] Anonymous. Behaviour. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):53– 61, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:P

[148] Anonymous. Presentation. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):62–71, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:CM

[149] Anonymous. Consistency of models. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):72–76, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:RBA

[150] Anonymous. Relationships between and among models. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):77–82, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:AMT

[151] Anonymous. Application of modelling techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):83–87, February 1981.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1981:WS

[152] Anonymous. Workshop summary. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):88–90, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Albano:1981:IIS

[153] Antonio Albano and Renzo Orsini. An interactive integrated system to design and use data bases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):91–93, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Amble:1981:IDB

[154] Tore Amble. Integration of data base design in programming languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):94, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Balzer:1981:DSS

[155] Robert M. Balzer. Dynamic system specification. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):95–97, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Borgida:1981:DAE

[156] Alexander T. Borgida and Sol Greenspan. Data and activities: Exploiting hierarchies of classes. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):98–100, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brodie:1981:DAD

[157] Michael L. Brodie. Data abstraction for designing database-intensive applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):101–103, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buneman:1981:UDT

[158] Peter Buneman and Ira Winston. The use of data type information in an interactive database environment. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):104–106, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carbonell:1981:DRI

[159] Jaime G. Carbonell. Default reasoning and inheritance mechanisms on type hierarchies. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):107–109, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cattell:1981:IDS

[160] R. G. G. Cattell. Integrating a database system and programming / information environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):110– 111, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Codd:1981:DMD

[161] E. F. Codd. Data models in database management. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):112–114, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cristian:1981:SBC

[162] Flaviu Cristian. On the specification of behavioural constraints. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):115–117, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deutsch:1981:CUM

[163] L. Peter Deutsch. Constraints: a uniform model for data and control. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):118–120, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Feather:1981:SCC

[164] Martin S. Feather. Some contrasts and considerations of an approach to modelling. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):121–123, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goldstein:1981:INS

[165] Ira Goldstein. Integrating a networkstructured database into an objectoriented programming language. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):124–125, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hardgrave:1981:PDM

[166] W. Terry Hardgrave and Donald R. Deutsch. Processing data model abstractions. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):126–127, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hayes:1981:LVT

[167] Patrick J. Hayes and Gary G. Hendrix. A logical view of types. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):128–130, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hendrix:1981:MVD

[168] Gary G. Hendrix. Mediating the views of databases and database users. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):131–132, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hitchcock:1981:DDO

[169] Peter Hitchcock. Data dictionaries in open system communication. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):133–134, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Katz:1981:HDH

[170] Randy H. Katz. Heterogeneous databases and high level abstraction. SIGMOD Record (ACM Special Interest Group on Management of Data), 11 (2):135–137, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

King:1981:MCR

[171] Jonathan J. King. Modelling concepts for reasoning about access to knowledge. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):138–140, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kreps:1981:RVC

[172] Peter Kreps. Relativism and views in a conceptual data base model. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):141–143, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lacroix:1981:ATD

[173] Michel Lacroix and Alain Pirotte. Associating types with domains of relational data bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):144– 146, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Leavenworth:1981:DAA

[174] B. Leavenworth. A data abstraction approach to database modelling. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):147–149, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Levesque:1981:IKB

[175] Hector J. Levesque. Incompleteness in knowledge bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):150– 152, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lundberg:1981:AIM

[176] Bengt Lundberg and Janis A. Bubenko, Jr. Axiomatic information modeling. SIGMOD Record (ACM Special Interest Group on Management of Data), 11 (2):153–154, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mark:1981:UDO

[177] William Mark. Use of database organization in the Consul system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):155–157, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mayr:1981:MMD

[178] Heinrich C. Mayr. Make more of data types. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):158–160, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McLeod:1981:CDM

[179] Dennis McLeod. On conceptual database modelling. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):161–163, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Meyer:1981:TLA

[180] Bertrand Meyer. A three-level approach to the description of data structures, and notational framework. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):164–166, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mylopoulos:1981:PRC

[181] John Mylopoulos. A perspective for research on conceptual modelling. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):167–170, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Paolini:1981:ADT

[182] Paolo Paolini. Abstract data types and data bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):171–173, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Reiter:1981:DBL

[183] Raymond Reiter. Data bases: a logical perspective. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):174–176, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rich:1981:MPV

[184] Charles Rich. Multiple points of view in modelling programs. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):177–179, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rowe:1981:IDD

[185] Lawrence A. Rowe. Issues in the design of database programming languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):180–182, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

dosSantos:1981:TCA

[186] C. S. dos Santos, A. L. Furtado, J. M. V. de Castilho, and S. E. R. de Carvalho. Towards constructive axiomatic specifications. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):183– 185, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schmidt:1981:DAT

[187] Joachim W. Schmidt. Data abstraction tools: Design, specification and application. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):186–188, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shaw:1981:ADT

[188] Mary Shaw. Abstraction, data types, and models for software. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):189– 191, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sibley:1981:DMS

[189] E. H. Sibley. Database management systems past and present. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):192, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sowa:1981:CSK

[190] John F. Sowa. A conceptual schema for knowledge-based systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):193–195, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thatcher:1981:DAD

[191] J. W. Thatcher. Data abstraction, data bases and conceptual modelling (position paper). SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):196–197, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wasserman:1981:EDA

[192] Anthony I. Wasserman. The extension of data abstraction to database management. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):198–200, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Weber:1981:DTU

[193] Herbert Weber. Are data types universal modelling concepts for data base systems? SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):201–202, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wedekind:1981:CAD

[194] Hartmut H. Wedekind. Constructive Abstract Data types (CAD). SIG-MOD Record (ACM Special Interest Group on Management of Data), 11 (2):203–206, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zilles:1981:TAM

[195] Stephen N. Zilles. Types, algebras and modeling. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(2):207–209, February 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Knowles:1981:DAA

[196] Stephen Knowles. Database activity at Aberdeen University Computing Centre. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(3):6, April 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schlageter:1981:REP

[197] G. Schlageter. Reply to ’Errors in ’Process synchronization in database systems”. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(3):7–9, April 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wiederhold:1981:RKB

[198] Gio Wiederhold, S. Jerrold Kaplan, and Daniel Sagalowicz. Research in knowledge base management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(3):26–54, April 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berg:1981:TR

[199] John Berg. Treasurer’s report. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11(4): 7, July 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Katz:1981:CSS

[200] Randy H. Katz. On the criteria for selecting the site of the annual SIG-MOD conference. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(4):8–11, July 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenthal:1981:NES

[201] Arnon S. Rosenthal. Note on the expected size of a join. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(4):19–25, July 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Trueblood:1981:MDS

[202] Robert P. Trueblood and H. Rex Hartson. MULTISAFE: a data security architecture. SIGMOD Record (ACM Special Interest Group on Management of Data), 11(4):26–31, July 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Elmasri:1981:NDA

[203] Ramez Elmasri, Cory Devor, and Said Rahimi. Notes on DDTS: an apparatus for experimental research in distributed database management systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 11(4): 32–49, July 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lochovsky:1981:RDM

[204] Fred H. Lochovsky. Review of ’Data models: a semantic approach for database systems’ by Sheldon A. Borkin. The MIT Press 1980. SIG-MOD Record (ACM Special Interest Group on Management of Data), 12 (1):16–17, October 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kilov:1981:MDA

[205] Kn. I. Kilov and I. A. Popova. Metadatabase architecture for relational DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(1):18–25, October 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Furtado:1981:HDI

[206] A. L. Furtado. Horizontal decomposition to improve a non-BCNF scheme. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(1):26–32, October 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Beeri:1981:NDR

[207] C. Beeri and M. Y. Vardi. A note on decompositions of relational databases.

SIGMOD Record (ACM Special Interest Group on Management of Data), 12(1):33–37, October 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Miller:1981:BEU

[208] Peter B. Miller, Sergey Tetelbaum, and Kincade N. Webb. BUSINESS: an enduser oriented application development language. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(1):38–69, October 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roark:1981:ECS

[209] Mayford L. Roark. Evolution in computer systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(2):4–8, Summer–Fall 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nolan:1981:EMO

[210] Richard L. Nolan. Establishing management objectives. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(2):9–19, Summer–Fall 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Burrows:1981:ACE

[211] James H. Burrows. Actual conversion experiences. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(2):20–33, Summer–Fall 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bryce:1981:S

[212] Milt Bryce. Standards. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(2):34– 38, Summer–Fall 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fry:1981:CTA

[213] James P. Fry. Conversion technology, an assessment. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(2):39–61, Summer–Fall 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ives:1981:R

[214] Blake Ives. Reviews. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(2):62– 64, Summer–Fall 1981. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hsiao:1982:WRV

[215] David K. Hsiao. Will the real VLDB conference please stand up? SIGMOD Record (ACM Special Interest Group on Management of Data), 12(3):8–14, April 1982. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Boral:1982:DRA

[216] Haran Boral, David J. DeWitt, Randy H. Katz, and Anthony Klug. Database research activities at the University of Wisconsin. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(3):19–26, April 1982. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bolour:1982:RTI

[217] A. Bolour, T. L. Anderson, L. J. Dekeyser, and H. K. T. Wong. The role of time in information processing: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(3):27–50, April 1982. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Overmyer:1982:ITE

[218] Ricky Overmyer and Michael Stonebraker. Implementation of a time expert in a data base system. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(3):51–60, April 1982. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schlageter:1982:POC

[219] G. Schlageter. Problems of optimistic concurrency control in distributed database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(3):62–66, April 1982. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rolland:1982:DPD

[220] Colette Rolland and Christian Richard. Distributed processing of data dynamics. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(3):67–85, April 1982. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Walker:1982:ADD

[221] Henry M. Walker. Administering a distributed Data Base Management System. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(3):86–99, April 1982. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brodie:1982:FRA

[222] Michael L. Brodie and Joachim W. Schmidt. Final report of the ANSI/X3/ SPARC DBS-SG relational database task group. SIGMOD Record (ACM Special Interest Group on Management of Data), 12(4):1–62, July 1982. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Date:1983:DU

[223] C. J. Date. Database usability. SIG-MOD Record (ACM Special Interest Group on Management of Data), 13(4): 1, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:1983:DTW

[224] Philip A. Bernstein. Database theory: where has it been? where is it going? SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):2, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gray:1983:PPD

[225] Jim Gray. Practical problems in data management: a position paper. SIG-MOD Record (ACM Special Interest Group on Management of Data), 13(4): 3, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Neuhold:1983:VDM

[226] Erich J. Neuhold. Views on data management. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):4–5, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wright:1983:MPD

[227] David D. Wright. On merging partitioned databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):6–14, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Maier:1983:FR

[228] David Maier and Jeffrey D. Ullman. Fragments of relations. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):15–22, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wong:1983:DDP

[229] E. Wong and R. H. Katz. Distributing a database for parallelism. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):23–29, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yu:1983:DQP

[230] C. T. Yu and C. C. Chang. On the design of a query processing strategy in a distributed database environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):30–39, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christodoulakis:1983:EBT

[231] Stavros Christodoulakis. Estimating block transfers and join sizes. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):40–54, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kambayashi:1983:QPU

[232] Yahiko Kambayashi and Masatoshi Yoshikawa. Query processing utilizing dependencies and horizontal decomposition. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):55–67, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Maier:1983:WW

[233] David Maier, David Rozenshtein, and David S. Warren. Windows on the world. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):68–78, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Laver:1983:FDC

[234] Kent Laver, Alberto O. Mendelzon, and Marc H. Graham. Functional dependencies on cyclic database schemes. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):79–91, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Honeyman:1983:NCI

[235] Peter Honeyman and Edward Sciore. A new characterization of independence.

SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):92–96, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1983:AMD

[236] Michael J. Carey. An abstract model of database concurrency control algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):97–107, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kohler:1983:ECL

[237] Walter H. Kohler, Kenneth C. Wilner, and John A. Stankovic. An experimental comparison of locking policies in a testbed database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):108–119, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:1983:HPU

[238] David B. Lomet. A high performance, universal, key associative access method. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):120–133, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1983:DAT

[239] Michael Stonebraker, Daniel Sagalowicz, and Robert Wilensky. DBMS and AI: is there any common point of view? SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):134, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rowe:1983:TSE

[240] Neil C. Rowe. Top-down statistical estimation on a database. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):135–145, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lohman:1983:RSG

[241] Guy M. Lohman, Joseph C. Stoltzfus, Anita N. Benson, Michael D. Martin, and Alfonso F. Cardenas. Remotelysensed geophysical databases: experience and implications for generalized DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):146–160, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shu:1983:FAR

[242] Nan C. Shu, Harry K. T. Wong, and Vincent Y. Lum. Forms approach to requirements specification for database design. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):161–172, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DAtri:1983:AQR

[243] Alessandro D’Atri, Marina Moscarini, and Nicolas Spyratos. Answering queries in relational databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 13 (4):173–177, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Imielinski:1983:IID

[244] Tomasz Imielinski and Witold Lipski, Jr. Incomplete information and dependencies in relational databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):178–184, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuck:1983:DGC

[245] Sharon M. Kuck and Yehoshua Sagiv. Designing globally consistent network schemas. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):185–195, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jarke:1983:RNF

[246] Matthias Jarke and Jürgen Koch. Range nesting: a fast method to evaluate quantified queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):196– 206, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zaniolo:1983:DLG

[247] Carlo Zaniolo. The database language GEM. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):207–218, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Missikoff:1983:RQD

[248] M. Missikoff and M. Scholl. Relational queries in a domain based DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):219–227, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chan:1983:OAC

[249] Arvola Chan, Umeshwar Dayal, Stephen Fox, Nathan Goodman, Daniel R. Ries, and Dale Skeen. Overview of an Ada compatible distributed database manager. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):228– 237, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wilms:1983:WWT

[250] Paul F. Wilms, Bruce G. Lindsay, and Patricia G. Selinger. “I wish I were over there”: distributed execution protocols for data definition in R. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):238–242, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ariav:1983:PTD

[251] Gad Ariav, Jim Clifford, and Matthias Jarke. Panel on time and databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 13(4):243–245, May 1983. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DeWitt:1984:ITM

[252] David J. DeWitt, Randy H. Katz, Frank Olken, Leonard D. Shapiro, Michael R. Stonebraker, and David Wood. Implementation techniques for main memory database systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14(2): 1–8, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Simon:1984:DIE

[253] Eric Simon and Patrick Valduriez. Design and implementation of an extendible integrity subsystem. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):9–17, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Boral:1984:TSA

[254] Haran Boral and Israel Gold. Towards a self-adapting centralized concurrency control algorithm. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):18–32, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Arnow:1984:ECB

[255] David M. Arnow and Aaron M. Tenenbaum. Empirical comparison of Btrees, compact B-trees and multiway trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):33–46, 1984. CO-DEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guttman:1984:TDI

[256] Antonin Guttman. R-trees: a dynamic index structure for spatial searching. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):47–57, 1984. CODEN SRECD8.

ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Diel:1984:DMF

[257] Hans Diel, Gerald Kreissig, Norbert Lenz, Michael Scheible, and Bernd Schoener. Data management facilities of an operating system kernel. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14(2): 58–69, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fedorowicz:1984:DEU

[258] Jane Fedorowicz. Database evaluation using multiple regression techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):70–76, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Baru:1984:PES

[259] C. K. Baru and S. Y. W. Su. Performance evaluation of the statistical aggregation by categorization in the SM3 system. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):77–89, 1984. CO-DEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Spooner:1984:DSI

[260] David L. Spooner. Database support for interactive computer graphics. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14(2): 90–99, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fogg:1984:LLD

[261] Dennis Fogg. Lessons from a ‘living in a database’ graphical query interface. SIGMOD Record (ACM Special Interest Group on Management of Data), 14 (2):100–106, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Valduriez:1984:MHS

[262] Patrick Valduriez and Yann Viemont. Multikey hashing scheme using predicate trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):107–114, 1984. CO-DEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lum:1984:DDS

[263] V. Lum, P. Dadam, R. Erbe, J. Guenauer, P. Pistor, G. Walch, H. Werner, and J. Woodfill. Designing DBMS support for the temporal dimension. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14 (2):115–130, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:1984:CRN

[264] Huei huang Chen and Sharon Mc-Cure Kuck. Combining relational and network retrieval methods. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):131– 142, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keller:1984:CIM

[265] Arthur M. Keller and Jeffrey D. Ullman. On complementary and independent mappings on databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):143– 148, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chan:1984:OCT

[266] Edward P. F. Chan. Optimal computation of total projections with unions of simple chase join expressions. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14 (2):149–163, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Willard:1984:EPR

[267] Dan E. Willard. Efficient processing of relational calculus expressions using range query theory. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):164–175, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Boral:1984:MDS

[268] Haran Boral and David J. DeWitt. Methodology for database system performance evaluation. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):176–185, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kriegel:1984:PCI

[269] Hans-Peter Kriegel. Performance comparison of index structures for multikey retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):186–196, 1984. CO-DEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Motro:1984:BLS

[270] Amihai Motro. Browsing in a loosely structured database. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):197–207, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1984:QDT

[271] Michael Stonebraker, Erika Anderson, Eric Hanson, and Brad Rubinstein. QUEL as a data type. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):208–214, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hogg:1984:AMS

[272] John Hogg and Stelios Gamvroulas. Active mail system. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):215–222, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:1984:SDD

[273] Jo-Mei Chang. Simplifying distributed database systems design by using a broadcast network. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):223–233, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fishman:1984:OJD

[274] Daniel H. Fishman, Ming-Yee Lai, and W. Kevin Wilkinson. Overview of the Jasmin database machine. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):234– 239, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shmueli:1984:MV

[275] Oded Shmueli and Alon Itai. Maintenance of views. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):240–255, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Piatetsky-Shapiro:1984:AEN

[276] Gregory Piatetsky-Shapiro and Charles Connell. Accurate estimation of the number of tuples satisfying a condition. SIGMOD Record (ACM Special Interest Group on Management of Data), 14 (2):256–276, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Salton:1984:UEB

[277] Gerard Salton. Use of extended Boolean logic in information retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 14 (2):277–285, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tsur:1984:IGS

[278] Shalom Tsur and Carlo Zaniolo. Implementation of GEM — supporting a semantic data model on a relational back-end. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):286–295, 1984. CO-DEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jarke:1984:OPF

[279] Matthias Jarke, Jim Clifford, and Yannis Vassiliou. Optimizing Prolog frontend to a relational query system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14 (2):296–306, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shepherd:1984:PKB

[280] Allan Shepherd and Larry Kerschberg. Prism: a knowledge based system for semantic integrity specification and enforcement in database systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14 (2):307–315, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Copeland:1984:MSD

[281] George Copeland and David Maier. Making Smalltalk a database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 14 (2):316–325, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hall:1984:RAL

[282] Patrick A. V. Hall. Relational algebras, logic, and functional programming. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(2):326–333, 1984. CO-DEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ozsoyoglu:1985:LPO

[283] Gultekin Ozsoyoglu, Z. Meral Ozsoyoglu, and Francisco Mata. A language and a physical organization technique for summary tables. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):3–16, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roussopoulos:1985:DSS

[284] Nick Roussopoulos and Daniel Leifker. Direct spatial search on pictorial databases using packed R-trees. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14(4): 17–31, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christodoulakis:1985:IAD

[285] S. Christodoulakis. Issues in the architecture of a document archiver using optical disk technology. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):34–50, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barbic:1985:TMO

[286] F. Barbic and B. Pernici. Time modeling in office information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):51–62, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Faloutsos:1985:SFD

[287] Chris Faloutsos. Signature files: design and performance comparison of some signature extraction methods. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):63–82, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eick:1985:ATK

[288] Christoph F. Eick and Peter C. Lockemann. Acquisition of terminological knowledge using database design techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):84–94, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shin:1985:PRD

[289] D. G. Shin and K. B. Irani. Partitioning a relational database horizontally using a knowledge-based approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):95–105, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1985:MSC

[290] Rakesh Agrawal, Michael J. Carey, and Miron Livny. Models for studying concurrency control performance: alternatives and implications. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14 (4):108–121, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Robinson:1985:FGP

[291] John T. Robinson. A fast generalpurpose hardware synchronization mechanism. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):122–130, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1985:RAM

[292] Rakesh Agrawal and David J. De-Witt. Recovery architectures for multiprocessor database machines. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14 (4):131–145, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Batini:1985:DDM

[293] Carlo Batini, Stefano Ceri, Al Hershey, George Gardarin, and David Reiner. Database design: methodologies, tools, and environments (panel session). SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):148–150, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hsu:1985:ICM

[294] Arding Hsu and Tomasz Imielinski. Integrity checking for multiple updates. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):152–168, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kung:1985:VDT

[295] C. H. Kung. On verification of database temporal constraints. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):169–179, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuper:1985:EPL

[296] Gabriel M. Kuper and Moshe Y. Vardi. On the expressive power of the logical data model: preliminary report. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):180–187, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Larson:1985:EPH

[297] Per-Ake Larson and M. V. Ramakrishna. External perfect hashing. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14 (4):190–200, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kawagoe:1985:MDH

[298] Kyoji Kawagoe. Modified dynamic hashing. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):201–213, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Otoo:1985:MDH

[299] Ekow J. Otoo. A multidimensional digital hashing scheme for files with composite keys. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):214–229, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sibley:1985:PDM

[300] Edgar H. Sibley, Matthias Jarke, Cecil S. McMinn, John Murray, Randall Rustin, and Ken Sloan. Pragmatics of database management (panel session). SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):232–234, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:1985:TTD

[301] Richard Snodgrass and Ilsoo Ahn. A taxonomy of time databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14 (4):236–246, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Clifford:1985:AHR

[302] James Clifford and Abdullah Uz Tansel. On an algebra for historical relational databases: two views. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):247–265, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Copeland:1985:DSM

[303] George P. Copeland and Setrag N. Khoshafian. A decomposition storage model. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):268–279, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yu:1985:AIS

[304] C. T. Yu and C. H. Chen. Adaptive information system design: one query at a time. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):280–290, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Beckley:1985:MRK

[305] D. A. Beckley, M. W. Evens, and V. K. Raman. Multikey retrieval from K-d trees and QUAD-trees. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14 (4):291–301, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christodoulakis:1985:MDM

[306] Starvos Christodoulakis, D. Badal, A. Cardenas, P. Mantey, F. Tompa, and D. Tsichritzis. Multimedia database management (panel session). SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):304–305, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fushimi:1985:APE

[307] Shinya Fushimi, Masaru Kitsuregawa, Masaya Nakayama, Hidehiko Tanaka, and Tohru Moto-oka. Algorithm and performance evaluation of adaptive multidimensional clustering technique. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):308–318, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kamel:1985:MDD

[308] Nabil Kamel and Roger King. A model of data distribution based on texture analysis. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):319–325, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goldman:1985:IIS

[309] Kenneth J. Goldman, Sally A. Goldman, Paris C. Kanellakis, and Stanley B. Zdonik. ISIS: interface for a semantic information system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 14 (4):328–342, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vossen:1985:HLU

[310] Gottfried Vossen and Volkert Brosda. A high-level user interface for update and retrieval in relational databases— language aspects. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):343–353, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Demo:1985:ACD

[311] G. Barbara Demo and Sukhamay Kundu. Analysis of the context dependency of CODASYL find-statements with application to a database program conversion. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):354–361, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Acharya:1985:TRP

[312] Shridhar Acharya and Gael Buckley. Transaction restarts in Prolog database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):364–373, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Page:1985:GDD

[313] Thomas W. Page, Jr., Matthew J. Weinstein, and Gerald J. Popek. Genesis: a distributed database operating system. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):374–387, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Klahold:1985:TMS

[314] P. Klahold, G. Schlageter, R. Unland, and W. Wilkes. A transaction model supporting complex applications in integrated information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):388–401, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sinha:1985:TBC

[315] Mukul K. Sinha, P. D. Nandikar, and S. L. Mehndiratta. Timestamp based certification schemes for transactions in distributed database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):402–411, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kerschberg:1985:EDS

[316] Larry Kerschberg, Michael Brodie, Charles Kellogg, D. Stott Parker, Gio Wiederhold, and Carlo Zaniolo. Expert database systems (workshop review). SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):414–417, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bhargava:1985:RDD

[317] Bharat Bhargava. Reliability in distributed database systems (panel discussion). SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):420–422, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sellis:1985:OED

[318] Timos K. Sellis and Leonard Shapiro. Optimization of extended database query languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):424–436, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gray:1985:EPA

[319] P. M. D. Gray. Efficient Prolog access to CODAYSL and FDM databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):437–443, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ullman:1985:ILQ

[320] Jeffrey D. Ullman. Implementation of logical query languages for databases (abstract only). SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):444, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Batory:1985:MCV

[321] D. S. Batory and W. Kim. Modeling concepts for VLSI CAD objects (abstract only). SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):446, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blain:1985:MPC

[322] Tomas Blain, Michael Dohler, Ralph Michaelis, and Emran Qureshi. Managing the printed circuit board design process. SIGMOD Record (ACM Special Interest Group on Management of Data), 14(4):447–456, May 1985. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1986:RMD

[323] Anonymous. Reference model for DBMS standardization. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(1):19–58, March 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roussopoulos:1986:EIS

[324] Nick Roussopoulos. Engineering information systems (panel session): builders and designers perspective. SIGMOD Record (ACM Special Interest Group on Management of Data), 15 (2):1–3, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buneman:1986:IPD

[325] Peter Buneman and Malcolm Atkinson. Inheritance and persistence in database programming languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):4–15, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bancilhon:1986:AIR

[326] François Bancilhon and Raghu Ramakrishnan. An amateur’s introduction to recursive query processing strategies. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):16–52, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lindsay:1986:SDR

[327] Bruce Lindsay, Laura Haas, C. Mohan, Hamid Pirahesh, and Paul Wilms. A snapshot differential refresh algorithm. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):53–60, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blakeley:1986:EUM

[328] Jose A. Blakeley, Per-Ake Larson, and Frank Wm Tompa. Efficiently updating materialized views. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):61–71, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Moss:1986:ARM

[329] J. Eliot B. Moss, Nancy D. Griffeth, and Marc H. Graham. Abstraction in recovery management. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):72–83, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mackert:1986:ROV

[330] Lothar F. Mackert and Guy M. Lohman. R* optimizer validation and performance evaluation for local queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):84–95, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ahn:1986:PET

[331] Ilsoo Ahn and Richard Snodgrass. Performance evaluation of a temporal database management system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 15(2):96–107, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1986:LBL

[332] Michael J. Carey and Hongjun Lu. Load balancing in a locally distributed DB system. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):108–119, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Motro:1986:CQT

[333] Amihai Motro. Constructing queries from tokens. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):120–131, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Delisle:1986:NHS

[334] Norman Delisle and Mayer Schwartz. Neptune: a hypertext system for CAD applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):132–143, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Frasson:1986:PIB

[335] C. Frasson and M. Er-radi. Principles of an icons-based language. SIG-MOD Record (ACM Special Interest Group on Management of Data), 15 (2):144–152, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Boral:1986:PDS

[336] Hanan Boral. Panel: Database system performance management. SIG-MOD Record (ACM Special Interest Group on Management of Data), 15 (2):153–154, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

VanGelder:1986:MPF

[337] Allen Van Gelder. A message passing framework for logical query evaluation. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):155–165, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenthal:1986:TRP

[338] Arnon Rosenthal, Sandra Heiler, Umeshwar Dayal, and Frank Manola. Traversal recursion: a practical approach to supporting recursive applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):166–176, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gardarin:1986:EDR

[339] Georges Gardarin and Christophe de Maindreville. Evaluation of database recursive logic programs as recurrent function series. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):177– 186, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Batory:1986:PED

[340] D. S. Batory and M. Mannino. Panel: Extensible database systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 15 (2):187–190, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sellis:1986:GQO

[341] Timos K. Sellis. Global query optimization. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):191–205, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Freytag:1986:RBT

[342] Johann Christoph Freytag and Nathan Goodman. Rule-based transformation of relational queries into iterative programs. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):206–214, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wedekind:1986:PRD

[343] H. Wedekind and George Zoerntlein. Prefetching in realtime database applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):215–226, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Faloutsos:1986:MHU

[344] Christos Faloutsos. Multiattribute hashing using Gray codes. SIG-MOD Record (ACM Special Interest Group on Management of Data), 15 (2):227–238, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lehman:1986:QPM

[345] Tobin J. Lehman and Michael J. Carey. Query processing in main memory database management systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 15 (2):239–250, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Willard:1986:GWC

[346] Dan E. Willard. Good worst-case algorithms for inserting and deleting records in dense sequential files. SIG-MOD Record (ACM Special Interest Group on Management of Data), 15 (2):251–260, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Harandi:1986:RBM

[347] Mehdi T. Harandi, Thierry Schang, and Seth Cohen. Rule base management using meta knowledge. SIG-MOD Record (ACM Special Interest Group on Management of Data), 15 (2):261–267, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Imielinski:1986:QPD

[348] Tomasz Imieliński. Query processing in deductive databases with incomplete information. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):268–280, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:1986:RBO

[349] Qiming Chen. A rule-based object/task modelling approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):281–292, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anderson:1986:PUI

[350] T. Lougenia Anderson and Gad Ariav. Panel: user interfaces and database management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):293– 294, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christodoulakis:1986:MOP

[351] S. Christodoulakis, F. Ho, and M. Theodoridou. The multimedia object presentation manager of MINOS: a symmetric approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):295–310, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Woelk:1986:OOA

[352] Darrell Woelk, Won Kim, and Willis Luther. An object-oriented approach to multimedia databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):311– 325, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Orenstein:1986:SQP

[353] Jack A. Orenstein. Spatial query processing in an object-oriented database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):326–336, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bitton:1986:PEL

[354] Dina Bitton. Panel: The effect of large main memory on database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):337–339, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1986:DP

[355] Michael Stonebraker and Lawrence A. Rowe. The design of POSTGRES. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):340–355, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dadam:1986:DPS

[356] P. Dadam, K. Kuespert, F. Andersen, H. Blanken, and R. Erbe. A DBMS prototype to support extended NF2 relations: an integrated view on flat tables and hierarchies. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):356–367, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bocca:1986:ESE

[357] Jorge Bocca. On the evaluation strategy of EDUCE. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):368–378, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Katz:1986:VMC

[358] Randy H. Katz, Ellis Chang, and Rajiv Bhateja. Version modeling concepts for computer-aided design databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):379–386, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lipeck:1986:SSD

[359] Udo W. Lipeck. Stepwise specification of dynamic database behaviour. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):387–397, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Biskup:1986:FVI

[360] Joachim Biskup and Bernhard Convent. A formal view integration method. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(2):398–407, June 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fong:1986:DBD

[361] Elizabeth N. Fong and Alan H. Goldfine. Data base directions information resource management-making it work, executive summary. SIG-MOD Record (ACM Special Interest Group on Management of Data), 15 (3):3–10, September 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bancilhon:1986:LPO

[362] François Bancilhon. A logic-programming/ object-oriented cocktail. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(3):11–21, September 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dittrich:1986:ETM

[363] Klaus R. Dittrich, Angelika M. Kotz, and Jutta A. Mülle. An event/trigger mechanism to enforce complex consistency constraints in design databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 15 (3):22–36, September 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Elmagarmid:1986:SDD

[364] Ahmed K. Elmagarmid. A survey of distributed deadlock detection algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(3):37–45, September 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anderson:1986:PBR

[365] T. Lougenia Anderson, Earl F. Ecklund, and David Maier. The Proteus bibliography: Representation and interactive display in databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 15 (3):46–55, September 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Salzberg:1986:TNF

[366] Betty Salzberg. Third normal form made easy. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(4):2–18, December 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Codd:1986:MIA

[367] E. F. Codd. Missing information (applicable and inapplicable) in relational databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 15(4):53, December 1986. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Minker:1987:HRS

[368] Jack Minker. Human rights: some perspectives. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(1):18–27, March 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kumar:1987:EJS

[369] Akhil Kumar and Michael Stonebraker. The effect of join selectives on optimal nesting order. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(1):28–41, March 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Codd:1987:MCM

[370] E. F. Codd. More commentary on missing information in relational databases (applicable and inapplicable information). SIGMOD Record (ACM Special Interest Group on Management of Data), 16(1):42–50, March 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Neff:1987:DBC

[371] R. K. Neff. Data bases, compound objects, and networked workstations: Beyond distributed computing (abstract). SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):1, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ingenthron:1987:TDR

[372] Kurt Ingenthron. Thoughts on database research: a user perspective. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):2, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1987:QOS

[373] Yannis E. Ioannidis and Eugene Wong. Query optimization by simulated annealing. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):9–22, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ganski:1987:ONS

[374] Richard A. Ganski and Harry K. T. Wong. Optimization of nested SQL queries revisited. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):23–33, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:1987:RQS

[375] Serge Abiteboul, Paris Kanellakis, and Gosta Grahne. On the representation and querying of sets of possible worlds. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):34–48, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sacca:1987:MCM

[376] Domenico Sacca and Carlo Zaniolo. Magic counting methods. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):49–59, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aly:1987:NDM

[377] Hussien Aly and Z. Meral Ozsoyoglu. Non-deterministic modelling of logical queries in deductive databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):60–72, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:1987:HRP

[378] Jiawei Han and Lawrence J. Henschen. Handling redundancy in the processing of recursive database queries. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):73–81, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Daniels:1987:DLT

[379] Dean S. Daniels, Alfred Z. Spector, and Dean S. Thompson. Distributed logging for transaction processing. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):82–96, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Herman:1987:DAV

[380] Gary Herman, K. C. Lee, and Abel Weinrib. The datacycle architecture for very high throughput database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):97–103, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lehman:1987:RAH

[381] Tobin J. Lehman and Michael J. Carey. A recovery algorithm for a high-performance memory-resident database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):104–117, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nixon:1987:ICS

[382] Brian Nixon, Lawrence Chung, John Mylopoulos, David Lauzon, Alex Borgida, and M. Stanley. Implementation of a compiler for a semantic data model: Experiences with taxis. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):118–131, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lyngbaek:1987:MSD

[383] Peter Lyngbaek and Victor Vianu. Mapping a semantic database model to the relational model. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):132–142, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roth:1987:DRD

[384] Mark A. Roth and Henry F. Korth. The design of ¬1NF relational databases into nested normal form. SIGMOD Record (ACM Special Interest Group on Management of Data), 16 (3):143–159, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Graefe:1987:EOG

[385] Goetz Graefe and David J. DeWitt. The EXODUS optimizer generator. SIGMOD Record (ACM Special Interest Group on Management of Data), 16 (3):160–172, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Freytag:1987:RBV

[386] Johann Christoph Freytag. A rulebased view of query optimization. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):173–180, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shenoy:1987:SSQ

[387] Sreekumar T. Shenoy and Z. Meral Ozsoyoglu. A system for semantic query optimization. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):181–195, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Paul:1987:AID

[388] H. B. Paul, H. J. Schek, and M. H. Scholl. Architecture and implementation of the Darmstadt database kernel system. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):196–207, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Richardson:1987:PCD

[389] Joel E. Richardson and Michael J. Carey. Programming constructs for database system implementation in EXODUS. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):208–219, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lindsay:1987:DME

[390] Bruce Lindsay, John McPherson, and Hamid Pirahesh. A data management extension architecture. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):220–226, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jajodia:1987:DV

[391] Sushil Jajodia and David Mutchler. Dynamic voting. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):227–238, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Haerder:1987:CTR

[392] Theo Haerder and Kurt Rothermel. Concepts for transaction recovery in nested transactions. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):239–248, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Garcia-Molina:1987:S

[393] Hector Garcia-Molina and Kenneth Salem. Sagas. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):249–259, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Freeston:1987:BFN

[394] Michael Freeston. The BANG file: a new kind of grid file. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):260–269, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nelson:1987:PAH

[395] Randal C. Nelson and Hanan Samet. A population analysis for hierarchical data structures. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):270–277, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sellis:1987:ESP

[396] Timos K. Sellis. Efficiently supporting procedures in relational database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):278–291, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hardwick:1987:WRF

[397] Martin Hardwick. Why ROSE is fast: Five optimizations in the design of an experimental database system for CAD/CAM applications. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):292–298, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kemper:1987:OOS

[398] Alfons Kemper, Peter C. Lockemann, and Mechtild Wallrath. An objectoriented system for engineering applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):299–310, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Banerjee:1987:SIS

[399] Jay Banerjee, Won Kim, Hyoung-Joo Kim, and Henry F. Korth. Semantics and implementation of schema evolution in object-oriented databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):311–322, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cruz:1987:GQL

[400] Isabel F. Cruz, Alberto O. Mendelzon, and Peter T. Wood. A graphical query language supporting recursion. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):323–330, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:1987:STC

[401] H. V. Jagadish, Rakesh Agrawal, and Linda Ness. A study of transitive closure as a recursion mechanism. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):331–344, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhang:1987:NCD

[402] Weining Zhang and C. T. Yu. A necessary condition for a doubly recursive rule to be equivalent to a linear recursive rule. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):345–356, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Morgenstern:1987:SIM

[403] Matthew Morgenstern. Security and inference in multilevel database and knowledge-base systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3): 357–373, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stemple:1987:MMF

[404] David Stemple, Subhasish Mazumdar, and Tim Sheard. On the modes and meaning of feedback to transaction designers. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):374–386, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rubenstein:1987:BSD

[405] W. B. Rubenstein, M. S. Kubicar, and R. G. G. Cattell. Benchmarking simple database operations. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):387–394, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gray:1987:MRT

[406] Jim Gray and Franco Putzolu. The 5 minute rule for trading memory for disc accesses and the 10 byte rule for trading memory for CPU time. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):395–398, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Richardson:1987:DEP

[407] James P. Richardson, Hongjun Lu, and Krishna Mikkilineni. Design and evaluation of parallel pipelined join algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):399–409, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Butler:1987:SRO

[408] Margaret H. Butler. Storage reclamation in object oriented database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):410–425, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Faloutsos:1987:AOO

[409] Christos Faloutsos, Timos Sellis, and Nick Roussopoulos. Analysis of object oriented spatial access methods. SIGMOD Record (ACM Special Interest Group on Management of Data), 16 (3):426–439, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hanson:1987:PAV

[410] Eric N. Hanson. A performance analysis of view materialization strategies. SIGMOD Record (ACM Special Interest Group on Management of Data), 16 (3):440–453, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Segev:1987:LMT

[411] Arie Segev and Arie Shoshani. Logical modeling of temporal data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):454–466, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McKenzie:1987:ERA

[412] Edwin McKenzie and Richard Snodgrass. Extending the relational algebra to support transaction time. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):467–478, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rubenstein:1987:DDM

[413] W. Bradley Rubenstein. A database design for musical information. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):479–490, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hudson:1987:OOD

[414] Scott E. Hudson and Roger King. Object-oriented database support for software environments. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3): 491–503, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Croft:1987:SOD

[415] W. B. Croft and D. W. Stemple. Supporting office document architectures with constrained types. SIG-MOD Record (ACM Special Interest Group on Management of Data), 16 (3):504–509, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stankovic:1988:RTT

[416] John A. Stankovic and Wei Zhao. On real-time transactions. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(1):4–18, March 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Singhal:1988:IAD

[417] Mukesh Singhal. Issues and approaches to design of real-time database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(1):19–33, March 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lin:1988:EAD

[418] Kwei-Jay Lin and Ming-Ju Lin. Enhancing availability in distributed realtime databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(1):34–43, March 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

vonBultzingsloewen:1988:KDD

[419] Gunter von Bultzingsloewen and Klaus R. Dittrich. Kardamom — a dataflow database machine for realtime applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(1):44–50, March 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1988:HPC

[420] M. J. Carey, M. Livny, and R. Jauhari. The hiPAC project: combining active databases and timing constraints. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(1):51–70, March 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abbott:1988:SRT

[421] Robert Abbott and Hector Garcia-Molina. Scheduling real-time transactions. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(1):71–81, March 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sha:1988:CCD

[422] Lui Sha, Ragunathan Rajkumar, and John P. Lehooczky. Concurrency control for distributed real-time databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(1):82–98, March 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jiang:1988:DDR

[423] Bin Jiang. Deadlock detection is really cheap. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(2):2–13, June 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shaw:1988:ADD

[424] Christopher J. Shaw. Application and database design — putting it off. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(2):14–22, June 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gersting:1988:RMD

[425] John Gersting, Kate Kinsley, Nancy McDonald, John North, Mark Sastry, and Edward Stull. Reference model for DBMS user facility. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(2):23–52, June 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Date:1988:DDT

[426] C. J. Date. Defining data types in a database language. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(2):53–76, June 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ghosh:1988:IPI

[427] S. Ghosh, C. C. Lin, and T. Sellis. Implementation of a Prolog-INGRES interface. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(2):77–88, June 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keller:1988:CUB

[428] Arthur M. Keller and Gio Wiederhold. Concurrent use of B-trees with variable-length entries. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(2):89–90, June 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thatte:1988:OOD

[429] Satish M. Thatte. Object-oriented database workshop: implementation aspects. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(2):95–107, June 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Murtagh:1988:ALD

[430] F. Murtagh. Astronomy from large databases: scientific objectives and methodological approaches. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(2):108– 111, June 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Swami:1988:OLJ

[431] A. Swami and A. Gupta. Optimization of large join queries. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):8–17, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lohman:1988:GLF

[432] G. M. Lohman. Grammar-like functional rules for representing query optimization alternatives. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):18–27, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Muralikrishna:1988:EDH

[433] M. Muralikrishna and D. J. De-Witt. Equi-depth histograms for estimating selectivity factors for multidimensional queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):28–36, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Garza:1988:TMO

[434] J. F. Garza and Won Kim. Transaction management in an objectoriented database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):37–45, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagannathan:1988:SDS

[435] D. Jagannathan, R. L. Guck, B. L. Fritchman, J. P. Thompson, and D. M. Tolbert. SIM: a database system based on the semantic data model. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):46–55, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Caruso:1988:MFC

[436] M. Caruso and E. Sciore. Metafunctions and contexts in an objectoriented database language. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):56–65, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Laurent:1988:PSI

[437] D. Laurent and N. Spyratos. Partition semantics for incomplete information in relational databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):66–73, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yuan:1988:SCQ

[438] Li Yan Yuan and Ding-An Chiang. A sound and complete query evaluation algorithm for relational databases with null values. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):74–81, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Malvestuto:1988:DPS

[439] F. M. Malvestuto. The derivation problem for summary data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):82–89, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Alexander:1988:PDC

[440] W. Alexander and G. Copeland. Process and dataflow control in distributed data-intensive systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):90–98, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Copeland:1988:DPB

[441] G. Copeland, W. Alexander, E. Boughter, and T. Keller. Data placement in bubba. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):99–108, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Patterson:1988:CRA

[442] D. A. Patterson, G. Gibson, and R. H. Katz. A case for redundant arrays of inexpensive disks (RAID). SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):109–116, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kumar:1988:SBT

[443] A. Kumar and M. Stonebraker. Semantics based transaction management techniques for replicated data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):117–125, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

ElAbbadi:1988:GPC

[444] A. El Abbadi and S. Toueg. The group paradigm for concurrency control protocols. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):126–134, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Breitbart:1988:MUI

[445] Y. Breitbart and A. Silberschatz. Multidatabase update issues. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):135–142, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:1988:DFD

[446] S. Abiteboul and R. Hull. Data functions, datalog and negation. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):143–153, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Krishnamurthy:1988:FTS

[447] R. Krishnamurthy, R. Ramakrishnan, and O. Shmueli. A framework for testing safety and effective computability of extended datalog. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):154–163, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:1988:IMR

[448] Qiming Chen and G. Gardarin. An implementation model for reasoning with complex objects. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):164–172, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kim:1988:OFD

[449] M. H. Kim and S. Pramanik. Optimal file distribution for partial match retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):173–182, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hutflesz:1988:TGF

[450] A. Hutflesz, H.-W. Six, and P. Widmayer. Twin grid files: space optimizing access schemes. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):183–190, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramakrishna:1988:HPA

[451] M. V. Ramakrishna. Hashing in practice, analysis of hashing and universal hashing. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):191–199, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1988:DMD

[452] E. Y. Ioannidis and M. Livny. Data modeling in DELAB. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):200, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ono:1988:DMT

[453] K. Ono, M. Aoyama, and H. Fujimoto. Data management of telecommunication networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):201, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Alho:1988:DDM

[454] K. Alho, H. Peltonen, M. Mantyla, and R. Sulonen. A design data manager. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):202, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Naeymi-Rad:1988:RDD

[455] F. Naeymi-Rad, L. Carmony, D. Trace, C. Georgakis, M. H. Weil, and M. Evens. A relational database design in support of standard medical terminology in multi-domain knowledge bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):203, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Drew:1988:VOO

[456] P. A. Drew, M. S. Ganti, D. J. Moore, R. J. Nassif, S. Podar, and D. H. Taenzer. Vishnu: an objectoriented database management system supporting rapid prototyping. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):204, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fromm:1988:DSG

[457] H. Fromm. Database services in the German Bildschirmtext system (Btx). SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):205, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vines:1988:GOO

[458] D. Vines and T. King. Gaia: an object-oriented framework for an Ada environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):206, September 1988.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kachhwaha:1988:LOO

[459] P. Kachhwaha and R. Hogan. LCE: an object-oriented database application development tool. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):207, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Krieger:1988:ISM

[460] H. Krieger, D. Schoenbauer, and B. Strodthoff. Integrated storage and manipulation of data and text based on a relational database. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):208, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hernandez:1988:CCT

[461] H. J. Hernandez and E. P. F. Chan. A characterization of constant-time-maintainability for BCNF database schemes. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):209–217, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Leuchner:1988:PTA

[462] J. Leuchner, L. Miller, and G. Slutzki. A polynomial time algorithm for testing implications of a join dependency and embodied functional dependencies. SIGMOD Record (ACM Special Interest Group on Management of Data), 17 (3):218–224, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gyssens:1988:PAR

[463] M. Gyssens and D. Van Gucht. The powerset algebra as a result of adding programming constructs to the nested relational algebra. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):225–232, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mazumdar:1988:RTB

[464] S. Mazumdar, D. Stemple, and T. Sheard. Resolving the tension between integrity and security using a theorem prover. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):233–242, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Qian:1988:TLD

[465] Xiaolei Qian and R. Waldinger. A transaction logic for database specification. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):243–250, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gadia:1988:GMR

[466] S. K. Gadia and Chuen-Sing Yeung. A generalized model for a relational temporal database. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):251–259, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Peinl:1988:HCS

[467] P. Peinl, A. Reuter, and H. Sammer. High contention in a stock trading database: a case study. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):260–268, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Haynie:1988:DLD

[468] M. N. Haynie. A DMBS for large design automation databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):269–276, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bell:1988:SDM

[469] J. L. Bell. A specialized data management system for parallel execution of particle physics codes. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):277–285, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christodoulakis:1988:PAF

[470] S. Christodoulakis and D. A. Ford. Performance analysis and fundamental performance trade offs for CLV optical disks. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):286–294, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hanson:1988:PQA

[471] E. N. Hanson. Processing queries against database procedures: a performance analysis. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):295–302, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jarke:1988:MKA

[472] M. Jarke and T. Rose. Managing knowledge about information system evolution. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):303–311, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Naughton:1988:CSR

[473] J. F. Naughton. Compiling separable recursions. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):312–319, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Youn:1988:CRF

[474] Cheong Youn and J. L. Henschen. Classification of recursive formulas in deductive databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):320–328, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wolfson:1988:DPL

[475] O. Wolfson and A. Silberschatz. Distributed processing of logic programs. SIGMOD Record (ACM Special Interest Group on Management of Data), 17 (3):329–336, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1988:BNS

[476] Anonymous. A benchmark of Non-Stop SQL on the debit credit transaction. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):337–341, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Borr:1988:HPS

[477] A. J. Borr and F. Putzolu. High performance SQL through low-level system integration. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):342–349, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DeWitt:1988:PAG

[478] D. J. DeWitt, S. Ghandeharizadeh, and D. Schneider. A performance analysis of the gamma database machine. SIGMOD Record (ACM Special Interest Group on Management of Data), 17 (3):350–360, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roesler:1988:DRS

[479] M. Roesler and W. A. Burkhard. Deadlock resolution and semantic lock models in object-oriented distributed systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):361–370, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramarao:1988:CPD

[480] K. V. S. Ramarao. Commitment in a partitioned distributed database. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):371–378, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Korth:1988:FMC

[481] H. F. Korth and G. D. Speegle. Formal model of correctness without serializability. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):379–386, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramnarayan:1988:DKB

[482] R. Ramnarayan and Hongjun Lu. A data/knowledge base management testbed and experimental results on data/knowledge base query and update processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):387–395, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Delcambre:1988:SCI

[483] L. M. L. Delcambre and J. N. Etheredge. A self-controlling interpreter for the relation production language. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(3):396–403, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sellis:1988:ILP

[484] T. Sellis, Chih-Chen Lin, and L. Raschid. Implementation large production systems in a DBMS environment: concepts and algorithms. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):404–412, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1988:DMQ

[485] M. J. Carey, D. J. DeWitt, and S. L. Vandenberg. A data model and query language for EXODUS. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):413–423, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lecluse:1988:OOD

[486] C. Lecluse, P. Richard, and F. Velez. O_{2}, an object-oriented data model. SIGMOD Record (ACM Special Interest Group on Management of Data), 17 (3):424–433, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Borgida:1988:MCH

[487] A. Borgida. Modeling class hierarchies with contradictions. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (3):434–443, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Atkins:1988:NMC

[488] John Atkins. A note on minimal covers. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(4):16–21, December 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guyot:1988:PMD

[489] Jacques Guyot. A process model for data bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(4):22–30, December 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schiel:1988:PID

[490] Ulrich Schiel. The problem of identification (databases). SIGMOD Record (ACM Special Interest Group on Management of Data), 17(4):31–36, December 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McLaughlin:1988:IMT

[491] M. E. McLaughlin, K. Bates Hill, D. D. Brown, M. A. Rogers, A. M. Howell, and P. P. Hatch. An integrated methodology and toolset for database design. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(4):37–55, December 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Merrett:1988:DME

[492] T. H. Merrett and Louis Hamel. Dynamic multipaging in even less space. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(4):56–61, December 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Son:1988:RDM

[493] Sang Hyuk Son. Replicated data management in distributed database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(4):62–69, December 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vossen:1988:EDL

[494] Gottfried Vossen and Jim Yacabucci. An extension of the database language SQL to capture more relational concepts. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(4):70–78, December 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buff:1988:WCR

[495] H. W. Buff. Why Codd’s rule no. 6 must be reformulated. SIG-MOD Record (ACM Special Interest Group on Management of Data), 17 (4):79–80, December 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:1989:FDD

[496] P. A. Bernstein, U. Dayal, D. J. De-Witt, D. Gawlick, J. Gray, M. Jarke, B. G. Lindsay, P. C. Lockemann, D. Maier, E. J. Neuhold, A. Reuter, L. A. Rowe, H. J. Schek, J. W. Schmidt, M. Schrefl, and M. Stonebraker. Future directions in DBMS research. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(1):17–26, March 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

OHare:1989:ICR

[497] Anthony B. O’Hare and Amit P. Sheth. The interpreted-compiled range of AI/ DB systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(1):32–42, March 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abdel-Wahab:1989:RLG

[498] H. Abdel-Wahab and B. Jamison. The role of linkage-graphs in Prolog relational databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(1):43–53, March 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Heytens:1989:GED

[499] Michael L. Heytens and Rishiyur S. Nikhil. GESTALT: an expressive database programming system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 18 (1):54–67, March 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kocharekar:1989:NRD

[500] Raju Kocharekar. Nulls in relational databases: revisited. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(1):68–73, March 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pletch:1989:CMC

[501] Andrew Pletch. Conceptual modeling in the classroom. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(1):74–80, March 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Salza:1989:ESQ

[502] Silvio Salza and Mario Terranova. Evaluating the size of queries on relational databases with non-uniform distribution and stochastic dependence. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):8–14, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kolodner:1989:AGC

[503] Elliot Kolodner, Barbara Liskov, and William Weihl. Atomic garbage collection: managing a stable heap. SIG-MOD Record (ACM Special Interest Group on Management of Data), 18(2): 15–25, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dong:1989:DPD

[504] Guozhu Dong. On distributed processibility of datalog queries by decomposing databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):26–35, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1989:OOD

[505] R. Agrawal and N. H. Gehani. ODE (Object Database and Environment): the language and the data model. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):36–45, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ohori:1989:DPM

[506] Atsushi Ohori, Peter Buneman, and Val Breazu-Tannen. Database programming in Machiavelli — a polymorphic language with static type inference. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):46–57, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Borgida:1989:CSD

[507] Alexander Borgida, Ronald J. Brachman, Deborah L. McGuinness, and Lori Alperin Resnick. CLASSIC: a structural data model for objects. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):58–67, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hou:1989:PAR

[508] Wen-Chi Hou, Gultekin Özsoyoğlu, and B. K. Taneja. Processing aggregate relational queries with hard time constraints. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):68–77, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:1989:IHR

[509] H. V. Jagadish. Incorporating hierarchy in a relational model of data. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):78–87, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cammarata:1989:ERD

[510] Stephanie Cammarata, Prasadram Ramachandra, and Darrell Shane. Extending a relational database with deferred referential integrity checking and intelligent joins. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):88–97, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Copeland:1989:CHA

[511] George Copeland and Tom Keller. A comparison of high-availability media recovery techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):98–109, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schneider:1989:PEF

[512] Donovan A. Schneider and David J. DeWitt. A performance evaluation of four parallel join algorithms in a shared-nothing multiprocessor environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):110–121, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1989:PCC

[513] Michael J. Carey and Miron Livny. Parallelism and concurrency control performance in distributed database machines. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):122–133, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kifer:1989:FLH

[514] Michael Kifer and Georg Lausen. Flogic: a higher-order language for reasoning about objects, inheritance, and scheme. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):134–146, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hull:1989:AOO

[515] Richard Hull and Jianwen Su. On accessing object-oriented databases: expressive power, complexity, and restrictions. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):147–158, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:1989:OIQ

[516] Serge Abiteboul and Paris C. Kanellakis. Object identity as a query language primitive. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):159–173, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chomicki:1989:RSI

[517] Jan Chomicki and Tomasz Imieliński. Relational specifications of infinite query answers. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):174–183, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sun:1989:SIP

[518] Xian-He Sun, Nabil Kamel, and Lionel M. Ni. Solving implication problems in database applications. SIG-MOD Record (ACM Special Interest Group on Management of Data), 18 (2):185–192, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bry:1989:TEE

[519] François Bry. Towards an efficient evaluation of general queries: quantifier and disjunction processing revisited. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):193–204, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1989:CRR

[520] Yannis E. Ioannidis and Timos K. Sellis. Conflict resolution of rules assigning values to virtual attributes. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):205–214, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McCarthy:1989:AAD

[521] Dennis R. McCarthy and Umeshwar Dayal. The architecture of an active database management system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 18 (2):215–224, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cohen:1989:CCD

[522] D. Cohen. Compiling complex database transition triggers. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):225– 234, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Naughton:1989:EER

[523] J. F. Naughton, R. Ramakrishnan, Y. Sagiv, and J. D. Ullman. Efficient evaluation of right-, left-, and multilinear rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):235–242, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Larson:1989:FSS

[524] P.-A. Larson and V. Deshpande. A file structure supporting traversal recursion. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):243–252, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1989:EMT

[525] R. Agrawal, A. Borgida, and H. V. Jagadish. Efficient management of transitive relationships in large data and knowledge bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):253–262, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gyssens:1989:GBA

[526] M. Gyssens, J. Paredaens, and D. van Gucht. A grammar-based approach towards unifying hierarchical data models. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):263–272, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Colby:1989:RAQ

[527] Latha S. Colby. A recursive algebra and query optimization for nested relations. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):273–283, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tansel:1989:NHR

[528] A. U. Tansel and L. Garnett. Nested historical relations. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):284–293, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Orenstein:1989:RSD

[529] J. A. Orenstein. Redundancy in spatial databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):294–305, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christodoulakis:1989:RPV

[530] Stavros Christodoulakis and Daniel Alexander Ford. Retrieval performance versus disc space utilization on WORM optical discs. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):306–314, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:1989:AMM

[531] David Lomet and Betty Salzberg. Access methods for multiversion data. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):315–324, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shekita:1989:PET

[532] Eugene J. Shekita and Michael J. Carey. Performance enhancement through replication in an objectoriented DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):325–336, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kim:1989:COR

[533] Won Kim, Elisa Bertino, and Jorge F. Garza. Composite objects revisited. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):337–347, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:1989:EIS

[534] E. E. Chang and R. H. Katz. Exploiting inheritance and structure semantics for effective clustering and buffering in an object-oriented DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):348–357, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Graefe:1989:DQE

[535] G. Graefe and K. Ward. Dynamic query evaluation plans. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):358– 366, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Swami:1989:OLJ

[536] A. Swami. Optimization of large join queries: combining heuristics and combinatorial techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):367–376, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Haas:1989:EQP

[537] L. M. Haas, J. C. Freytag, G. M. Lohman, and H. Pirahesh. Extensible query processing in starburst.

SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):377–388, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tang:1989:SPS

[538] Jian Tang and N. Natarajan. A static pessimistic scheme for handling replicated databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):389–398, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ellis:1989:CCG

[539] C. A. Ellis and S. J. Gibbs. Concurrency control in groupware systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):399–407, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1989:MSM

[540] D. Agrawal and S. Sengupta. Modular synchronization in multiversion databases: version control and concurrency control. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):408–417, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DeTroyer:1989:RTC

[541] O. De Troyer. RIDL*: a tool for the computer-assisted engineering of large databases in the presence of integrity constraints. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):418–429, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Markowitz:1989:CRE

[542] Victor M. Markowitz and Arie Shoshani. On the correctness of representing extended entity-relationship structures in the relational model. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(2):430–439, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Navathe:1989:VPD

[543] Shamkant B. Navathe and Mingyoung Ra. Vertical partitioning for database design: a graphical algorithm. SIG-MOD Record (ACM Special Interest Group on Management of Data), 18 (2):440–450, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1989:CPR

[544] M. Stonebraker, M. Hearst, and S. Potamianos. A commentary on the POSTGRES rules system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 18 (3):5–11, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hanson:1989:IRD

[545] Eric N. Hanson. An initial report on the design of Ariel DBMS with an integrated production rule system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 18 (3):12–19, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chakravarthy:1989:RME

[546] S. Chakravarthy. Rule management and evaluation: an active DBMS perspective. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(3):20–28, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bocca:1989:RIV

[547] J. B. Bocca and J. C. Freytag. Rules for implementing very large knowledge base systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(3):29–35, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Widom:1989:SSS

[548] J. Widom and S. J. Finkelstein. A syntax and semantics for set-oriented production rules in relational database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(3):36–45, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jarke:1989:RRM

[549] M. Jarke and M. Jeusfeld. Rule representation and management in ConceptBase. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(3):46–51, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sellis:1989:DIP

[550] T. Sellis, Chih-Chen Lin, and L. Raschid. Data intensive production systems: the DIPS approach. SIG-MOD Record (ACM Special Interest Group on Management of Data), 18 (3):52–58, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Delcambre:1989:PMR

[551] L. M. L. Delcambre, J. Waramahaputi, and J. N. Etheredge. Pattern match reduction for Relational Production Language in the USL MMDBS. SIGMOD Record (ACM Special Interest Group on Management of Data), 18 (3):59–67, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kiernan:1989:DIE

[552] J. Kiernan, C. de Maindreville, and E. Simon. The design and implementation of an extendible deductive database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(3):68–77, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Joseph:1989:SRO

[553] J. Joseph, S. Thatte, C. Thompson, and D. Wells. Special report: the 1988 Object-oriented Database Workshop. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(3):78–101, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1989:CPI

[554] M. Stonebraker. The case for partial indexes (DBMS). SIGMOD Record (ACM Special Interest Group on Management of Data), 18(4):4–11, December 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Date:1989:NRC

[555] C. J. Date. A note on the relational calculus. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(4):12–16, December 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

vonBultzingsloewen:1989:OSQ

[556] G. von Bultzingsloewen. Optimizing SQL queries for parallel execution. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(4):17–22, December 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Teorey:1989:DDD

[557] T. J. Teorey. Distributed database design: a practical approach and example education. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(4):23–39, December 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fong:1989:IMD

[558] E. N. Fong and A. H. Goldfine. Information management directions: the integration challenge. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(4):40–43, December 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Narasimhalu:1989:DGI

[559] D. Narasimhalu. The database group at ISS, National University of Singapore. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(4):44–49, December 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rusinkiewicz:1990:ETM

[560] M. E. Rusinkiewicz, A. K. Elmagarmid, Y. Leu, and W. Litwin. Extending the transaction model to capture more meaning. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(1):3–7, March 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Basu:1990:SBQ

[561] Amit Basu and Rafiul Ahad. An SQL-based query language for networks of relations. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(1):8–17, March 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gogolla:1990:NTS

[562] Martin Gogolla. A note on the translation of SQL to tuple calculus. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(1):18–22, March 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jonsson:1990:SMT

[563] Dan Jonsson. Semantic modeling through identification and characterization of objects. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(1):23–28, March 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gessert:1990:FVL

[564] G. H. Gessert. Four valued logic for relational database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(1):29–35, March 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Missaoui:1990:IPI

[565] Rokia Missaoui and Robert Godin. The implication problem for inclusion dependencies: a graph approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(1):36–40, March 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Teorey:1990:LTJ

[566] Toby J. Teorey. Letter by Toby J. Teorey. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(1):43–53, March 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lipton:1990:PSE

[567] Richard J. Lipton, Jeffrey F. Naughton, and Donovan A. Schneider. Practical selectivity estimation through adaptive sampling. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):1–11, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

King:1990:BAT

[568] Roger King and Ali Morfeq. Bayan. an Arabic text database management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):12–23, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gyssens:1990:GOO

[569] Marc Gyssens, Jan Paredaens, and Dirk Van Gucht. A graph-oriented object model for database end-user interfaces. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):24–33, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1990:OGI

[570] R. Agrawal, N. H. Gehani, and J. Srinivasan. OdeView. the graphical interface to Ode. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):34–43, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ullman:1990:IOC

[571] Jeffrey D. Ullman and Mihalis Yannakakis. The input/output complexity of transitive closure. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):44–53, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sheng:1990:IEE

[572] Yeh-Heng Sheng. IDLOG. extending the expressive power of deductive database languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):54–63, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Saraiya:1990:HPS

[573] Yatin P. Saraiya. Hard problems for simple logic programs. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):64–73, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:1990:PTD

[574] Ke Wang. Polynomial time designs toward both BCNF and efficient data manipulation. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):74–83, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Atzeni:1990:EUI

[575] Paolo Atzeni and Riccardo Torlone. Efficient updates to independent schemes in the weak instance model. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):84–93, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Salzberg:1990:FDS

[576] Betty Salzberg, Alex Tsukerman, Jim Gray, Michael Stuewart, Susan Uren, and Bonnie Vaughan. FastSort: a distributed single-input single-output external sort. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):94–101, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Graefe:1990:EPV

[577] Goetz Graefe. Encapsulation of parallelism in the Volcano query processing system. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):102–111, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:1990:IRR

[578] Philip A. Bernstein, Meichun Hsu, and Bruce Mann. Implementing recoverable requests using queues. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (2):112–122, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Solworth:1990:WOD

[579] Jon A. Solworth and Cyril U. Orji. Write-only disk caches. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):123– 132, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wolfson:1990:NPP

[580] Ouri Wolfson and Aya Ozeri. A new paradigm for parallel and distributed rule-processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):133–142, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ganguly:1990:FPP

[581] Sumit Ganguly, Avi Silberschatz, and Shalom Tsur. A framework for the parallel processing of Datalog queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):143–152, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kogan:1990:CCM

[582] Boris Kogan and Sushil Jajodia. Concurrency control in multilevel-secure databases based on replicated architecture. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):153–162, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Badrinath:1990:PES

[583] B. R. Badrinath and Krithi Ramamritham. Performance evaluation of semantics-based multilevel concurrency control protocols. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):163–172, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Motro:1990:QDK

[584] Amihai Motro and Qiuhui Yuan. Querying database knowledge. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (2):173–183, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Laenens:1990:ELP

[585] Els Laenens, Domenico Sacca, and Dirk Vermeir. Extending logic programming. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):184–193, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chrysanthis:1990:AFS

[586] Panayiotis K. Chrysanthis and Krithi Ramamritham. ACTA. A framework for specifying and reasoning about transaction structure and behavior. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):194–203, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dayal:1990:OLR

[587] Umeshwar Dayal, Meichun Hsu, and Rivka Ladin. Organizing long-running activities with triggers and transactions. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):204–214, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Breitbart:1990:RTM

[588] Yuri Breitbart, Avi Silberschatz, and Glenn R. Thompson. Reliable transaction management in a multidatabase system. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):215–224, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cacace:1990:IOO

[589] F. Cacace, S. Ceri, S. Crespi-Reghizzi, L. Tanca, and R. Zicari. Integrating object-oriented data modeling with a rule-based programming paradigm. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):225–236, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kiernan:1990:MDD

[590] G. Kiernan, C. de Maindreville, and E. Simon. Making deductive database a practical technology. A step forward.

SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):237–246, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mumick:1990:MR

[591] Inderpal Singh Mumick, Sheldon J. Finkelstein, Hamid Pirahesh, and Raghu Ramakrishnan. Magic is relevant. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):247–258, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Widom:1990:SOP

[592] Jennifer Widom and Sheldon J. Finkelstein. Set-oriented production rules in relational database systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (2):259–270, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hanson:1990:PMA

[593] Eric N. Hanson, Moez Chaabouni, Chang-Ho Kim, and Yu-Wang Wang. A predicate matching algorithm for database rule systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):271– 280, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1990:RPC

[594] Michael Stonebraker, Anant Jhingran, Jeffrey Goh, and Spyros Potamianos. On rules, procedures, caching and views in data base systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (2):281–290, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenthal:1990:QGI

[595] Arnon Rosenthal and Cesar Galindo-Legaria. Query graphs, implementing trees, and freely-reorderable outerjoins. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):291–299, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shekita:1990:PEP

[596] Eugene J. Shekita and Michael J. Carey. A performance evaluation of pointer-based joins. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):300–311, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1990:RAO

[597] Yannis E. Ioannidis and Younkyung Cha Kang. Randomized algorithms for optimizing large join queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):312– 321, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Beckmann:1990:TER

[598] Norbert Beckmann, Hans-Peter Kriegel, Ralf Schneider, and Bernhard Seeger. The R*-tree: an efficient and robust access method for points and rectangles. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):322–331, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:1990:LCO

[599] H. V. Jagadish. Linear clustering of objects with multiple attributes. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):332–342, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Orenstein:1990:CSQ

[600] Jack Orenstein. A comparison of spatial query processing techniques for native and parameter spaces. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (2):343–352, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:1990:PMA

[601] David Lomet and Betty Salzberg. The performance of a multiversion access method. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):353–363, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kemper:1990:ASO

[602] Alfons Kemper and Guido Moerkotte. Access support in object bases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (2):364–374, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Olken:1990:RSH

[603] Frank Olken, Doron Rotem, and Ping Xu. Random sampling from hash files.

SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):375–386, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cha:1990:KCM

[604] Sang K. Cha and Gio Wiederhold. Kaleidoscope. A cooperative menuguided query interface. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):387, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Consens:1990:SGV

[605] Mariano P. Consens and Alberto O. Mendelzon. The G+/GraphLog Visual Query System. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):388, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1990:OUF

[606] R. Agrawal, N. H. Gehani, and J. Srinivasan. OdeView. A user-friendly graphical interface to Ode. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):389, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blum:1990:ISQ

[607] Bruce I. Blum and Ralph D. Semmel. The INA: a simple query language with only attribute names: a video presentation. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):390, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuntz:1990:PGD

[608] Michel Kuntz. Pasta-3. A graphical direct manipulation interface for knowledge base management systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19(2): 391, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kent:1990:IDS

[609] Bill Kent, Peter Lyngbaek, Samir Mathur, and Kevin Wilkinson. The Iris database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):392, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lapis:1990:SB

[610] George Lapis, Guy M. Lohman, and Hamid Pirahesh. A Starburst is born. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):393–??, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1990:PD

[611] Michael Stonebraker. Postgres DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):394–??, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zicari:1990:IIO

[612] Roberto Zicari. Incomplete information in object-oriented databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (3):5–16, September 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thuraisingham:1990:CTC

[613] Bhavani Thuraisingham. Computing transitive closures of multilevel relations. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(3):17–23, September 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lei:1990:DDA

[614] Li Lei, G.-H. Moll, and J. Kouloumdjian. A deductive database architecture based on partial evaluation. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (3):24–29, September 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nummenmaa:1990:YAN

[615] Jyrki Nummenmaa and Peter Thanisch. Yet another note on minimal covers. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(3):30, September 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

CADF:1990:TGD

[616] The Committee for Advanced DBMS Function. Third-generation database system manifesto. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(3):31–44, September 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sheth:1990:DRB

[617] Amit P. Sheth. Database research at Bellcore. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(3):45–52, September 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Breitbart:1990:MI

[618] Y. Breitbart. Multidatabase interoperability. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(3):53–60, September 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Silberschatz:1990:DSA

[619] Avi Silberschatz, Michael Stonebraker, and Jeffrey D. Ullman. Database systems: achievements and opportunities. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(4):6–22, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Scheuermann:1990:RWH

[620] P. Scheuermann, C. Yu, A. Elmagarmid, F. Manola, A. Rosenthal, H. García-Molina, D. McLeod, and M. Templeton. Report on the workshop on heterogeneous database systems held at Northwestern University Evanston, Illinois, December 11– 13, 1989 sponsored by NSF. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (4):23–31, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

French:1990:SFR

[621] James C. French, Anita K. Jones, and John L. Pfaltz. Summary of the final report of the NSF workshop on scientific database management. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (4):32–40, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:1990:NHD

[622] Serge Abiteboul, Peter Buneman, Claude Delobel, Richard Hull, Paris Kanellakis, and Victor Vianu. New hope on data models and types: report of an NSF-INRIA workshop. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (4):41–48, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bancilhon:1990:OOD

[623] François Bancilhon and Won Kim. Object-oriented database systems: in transition. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(4):49–53, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1990:EDM

[624] Michael Carey and Laura Haas. Extensible database management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(4):54–60, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guenther:1990:RIS

[625] O. Guenther and A. Buchmann. Research issues in spatial databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (4):61–68, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Motro:1990:AID

[626] Amihai Motro. Accommodating imprecision in database systems: issues and solutions. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(4):69–74, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ullman:1990:DDA

[627] Jeffrey D. Ullman and Carlo Zaniolo. Deductive databases: achievements and future directions. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (4):75–82, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:1990:TDS

[628] Richard Snodgrass. Temporal databases: status and research directions. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (4):83–89, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lunt:1990:DS

[629] Teresa F. Lunt and Eduardo B. Fernandez. Database security. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (4):90–97, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Garcia-Molina:1990:RDD

[630] Héctor García-Molina and Bruce Lindsay. Research directions for distributed databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(4):98–103, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DeWitt:1990:PDS

[631] David J. DeWitt and Jim Gray. Parallel database systems: the future of database processing or a passing fad? SIGMOD Record (ACM Special Interest Group on Management of Data), 19 (4):104–112, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1990:DBR

[632] Michael Stonebraker. Data base research at Berkeley. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(4):113–118, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keller:1990:SDR

[633] Arthur Keller, Peter Rathmann, Jeff Ullman, and Gio Wiederhold. Selected database research at Stanford. SIG-MOD Record (ACM Special Interest Group on Management of Data), 19 (4):119–122, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jajodia:1990:DSC

[634] Sushil Jajodia and Ravi Sandhu. Database security: current status and key issues. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(4):123–126, December 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rafanelli:1991:FMM

[635] M. Rafanelli and F. L. Ricci. A functional model for macro-databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(1):3–8, March 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Robinson:1991:SFA

[636] John T. Robinson. Some further analysis of the essential blocking recurrence. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(1):9–11, March 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Diederich:1991:MCR

[637] Jim Diederich. Minimal covers revisited: correct and efficient algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(1):12–13, March 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Soo:1991:BTD

[638] Michael D. Soo. Bibliography on temporal databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(1):14–23, March 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vossen:1991:BOO

[639] Gottfried Vossen. Bibliography on object-oriented database management. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(1):24–46, March 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thatte:1991:MOO

[640] Satish M. Thatte. A modular and open object-oriented database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(1):47–52, March 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Richard:1991:RA

[641] Philippe Richard. Research at Altaïr. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(1):53–59, March 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lenat:1991:KAC

[642] Douglas B. Lenat. Keynote address: computers versus common sense. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20(2): 1, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1991:MPO

[643] Michael Stonebraker. Managing persistent objects in a multi-level store. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):2–11, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tsangaris:1991:SAC

[644] Manolis M. Tsangaris and Jeffrey F. Naughton. A stochastic approach for clustering in object bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):12–21, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cheng:1991:ECC

[645] Jia-Bing R. Cheng and A. R. Hurson. Effective clustering of complex objects in object-oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):22–31, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Perrizo:1991:HHD

[646] William Perrizo, Joseph Rajkumar, and Prabhu Ram. HYDRO: a heterogeneous distributed database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):32–39, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Krishnamurthy:1991:LFI

[647] Ravi Krishnamurthy, Witold Litwin, and William Kent. Language features for interoperability of databases with schematic discrepancies. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):40–49, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jajodia:1991:TMS

[648] Sushil Jajodia and Ravi Sandhu. Toward a multilevel secure relational data model. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):50–59, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gordin:1991:SOC

[649] Douglas N. Gordin and Alexander J. Pasik. Set-oriented constructs: from Rete rule bases to database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):60–67, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sudarshan:1991:SOB

[650] S. Sudarshan, Divesh Srivastava, Raghu Ramakrishnan, and Jeffrey F. Naughton. Space optimization in the bottom-up evaluation of logic programs. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):68–77, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wolfson:1991:IER

[651] Ouri Wolfson, Hasanat M. Dewan, Salvatore J. Stolfo, and Yechiam Yemini. Incremental evaluation of rules and its relationship to parallelism. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):78–87, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Levy:1991:OCP

[652] Eliezer Levy, Henry F. Korth, and Abraham Silberschatz. An optimistic commit protocol for distributed transaction management. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):88–97, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1991:UMD

[653] D. Agrawal and V. Krishnaswamy. Using multiversion data for noninterfering execution of write-only transactions. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):98–107, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chrysanthis:1991:ECO

[654] Panos K. Chrysanthis, S. Raghuram, and Krithi Ramamritham. Extracting concurrency from objects: a methodology. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):108–117, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Crysanthis:1991:ECO

[655] P. K. Crysanthis, S. Raghuram, and K. Ramamritham. Extracting concurrency from objects: a methodology. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):108–117, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Neugebauer:1991:OED

[656] Leonore Neugebauer. Optimization and evaluation of database queries including embedded interpolation procedures. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):118–127, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Becker:1991:SPS

[657] Bruno Becker, Hans-Werner Six, and Peter Widmayer. Spatial priority search: an access technique for scaleless maps. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):128–137, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kolovson:1991:SID

[658] Curtis P. Kolovson and Michael Stonebraker. Segment indexes: dynamic indexing techniques for multidimensional interval data. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):138– 147, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keller:1991:EAC

[659] Tom Keller, Goetz Graefe, and David Maier. Efficient assembly of complex objects. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):148–157, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vandenberg:1991:ASC

[660] Scott L. Vandenberg and David J. De-Witt. Algebraic support for complex objects with arrays, identity, and inheritance. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):158–167, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1991:LDV

[661] Yannis E. Ioannidis and Younkyung Cha Kang. Left-deep vs. bushy trees: an analysis of strategy spaces and its implications for query optimization. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):168–177, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guntzer:1991:NDU

[662] U. Güntzer, W. Kießling, and H. Thöne. New directions for uncertainty reasoning in deductive databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):178–187, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sheng:1991:NDD

[663] Yeh-Heng Sheng. A non-deterministic deductive database language. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (2):188–197, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lou:1991:LOO

[664] Yanjun Lou and Z. Meral Özsoyoğlu. LLO: an object-oriented deductive language with methods and method inheritance. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):198–207, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:1991:RTS

[665] H. V. Jagadish. A retrieval technique for similar shapes. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):208–217, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ahmed:1991:VMC

[666] Rafi Ahmed and Shamkant B. Navathe. Version management of composite objects in CAD databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (2):218–227, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chiueh:1991:TAM

[667] Tzi cker Chiueh and Randy Katz. Trait: an attribute management system for VLSI design objects. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (2):228–237, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:1991:OV

[668] Serge Abiteboul and Anthony Bonner. Objects and views. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):238–247, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barsalou:1991:URD

[669] Thierry Barsalou, Niki Siambela, Arthur M. Keller, and Gio Wiederhold. Updating relational databases through object-based views. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):248–257, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kemper:1991:FMO

[670] Alfons Kemper, Christoph Kilger, and Guido Moerkotte. Function materialization in object bases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (2):258–267, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1991:PES

[671] Yannis E. Ioannidis and Stavros Christodoulakis. On the propagation of errors in the size of join results. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):268–277, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hou:1991:ECC

[672] Wen-Chi Hou, Gultekin Özsoyoğlu, and Erdogan Dogdu. Error-constrained COUNT query evaluation in relational databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):278–287, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Imielinski:1991:IOD

[673] Tomasz Imieliński, Shamim Naqvi, and Kumar Vadaparty. Incomplete objects — a data model for design and planning applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):288–297, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Richardson:1991:AEO

[674] Joel Richardson and Peter Schwarz. Aspects: extending objects to support multiple, independent roles. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (2):298–307, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Phipps:1991:GND

[675] Geoffrey Phipps, Marcia A. Derr, and Kenneth A. Ross. Glue-Nail: a deductive database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):308–317, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Annevelink:1991:DPL

[676] Jurgen Annevelink. Database programming languages: a functional approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):318–327, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:1991:NRB

[677] Qiming Chen and Yahiko Kambayashi. Nested relation based database knowledge representation. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):328–337, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shyy:1991:KHL

[678] Yuh-Ming Shyy and Stanley Y. W. Su. K: a high-level knowledge base programming language for advanced database applications. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (2):338–347, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hansen:1991:EMI

[679] S. C. Hansen and E. A. Unger. An extended memoryless inference control model: accounting for dependence in table-level controls. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):348–356, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1991:DCT

[680] Michael J. Carey, Michael J. Franklin, Miron Livny, and Eugene J. Shekita. Data caching tradeoffs in client-server DBMS architectures. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):357–366, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:1991:CCC

[681] Yongdong Wang and Lawrence A. Rowe. Cache consistency and concurrency control in a client/server DBMS architecture. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):367–376, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pu:1991:RCD

[682] Calton Pu and Avraham Leff. Replica control in distributed systems: an asynchronous approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):377– 386, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ng:1991:FBA

[683] Raymond Ng, Christos Faloutsos, and Timos Sellis. Flexible buffer allocation based on marginal gains. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (2):387–396, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gruenwald:1991:MRA

[684] Le Gruenwald and Margaret H. Eich. MMDB reload algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):397– 405, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Weikum:1991:DFA

[685] Gerhard Weikum, Peter Zabback, and Peter Scheuermann. Dynamic file allocation in disk arrays. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):406–415, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Srinivasan:1991:PBT

[686] V. Srinivasan and Michael J. Carey. Performance of B-tree concurrency control algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):416–425, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lanka:1991:FPB

[687] Sitaram Lanka and Eric Mays. Fully persistent B^{+}-trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):426–435, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Seeger:1991:MDB

[688] Bernhard Seeger and Per-Åke Larson. Multi-disk B-trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):436–445, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thompson:1991:SPH

[689] Glenn Thompson, Hector García-Molina, James Gray, Avi Silverschatz, and Michael Stonebraker. Are standards the panacea for heterogeneous distributed DBMSs? SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):446, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lohman:1991:SIE

[690] Guy M. Lohman, George Lapis, Tobin Lehman, Rakesh Agrawal, Roberta Cochrane, John McPherson, C. Mohan, Hamid Pirahesh, and Jennifer Widom. Starburst II: the extender strikes back! (relational database). SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):447, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Furlani:1991:PRM

[691] C. Furlani, J. Parks, K. C. Morris, M. Loomis, J. Fowler, and Y. Yang.

PDES: representing and managing product data. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):448, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kent:1991:OIM

[692] B. Kent, E. Fong, D. Deutsch, D. Lenkov, A. Otis, and J. Slitz. Objects in information management standards. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):449, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wade:1991:CFD

[693] D. Wade, C. Hsu, J. Schwartz, R. Mazdra, and S. Banks. CAD frameworks and databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(2):450, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Salzberg:1991:SDA

[694] Betty Salzberg and David B. Lomet. Spatial database access methods. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (3):6–15, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fratarcangeli:1991:TUQ

[695] Claudio Fratarcangeli. Technique for universal quantification in SQL. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (3):16–24, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Streeter:1991:OM

[696] Victor J. Streeter. An outline of MQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 20 (3):25–29, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gessert:1991:HMD

[697] G. H. Gessert. Handling missing data by using stored truth values. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (3):30–42, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yue:1991:MGM

[698] Kwok bun Yue. A more general model for handling missing information in relational databases using a 3-valued logic. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(3):43–49, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buff:1991:CIS

[699] H. W. Buff. A complete identity set for Codd algebras. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (3):50–51, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roussopoulos:1991:MCS

[700] Nick Roussopoulos and Alexis Delis. Modern client-server DBMS architectures. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(3):52–61, September 1991.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Harvey:1991:DTN

[701] Tina M. Harvey, Craig W. Schnepf, and Mark A. Roth. The design of the Triton nested relational database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(3):62–72, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thomas:1991:PPN

[702] Gomer Thomas, K. Kawagoe, R. Krishnamurthy, T. Imieliński, D. Reiner, and A. Wolski. Practitioner problems in need of database research. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (3):73–84, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barrera:1991:TRG

[703] Renato Barrera, Andrew Frank, and Khaled Al-Taha. Temporal relations in geographic information systems: a workshop at the University of Maine. SIGMOD Record (ACM Special Interest Group on Management of Data), 20 (3):85–91, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Haas:1991:DRI

[704] Laura M. Haas and Patricia G. Selinger. Database research at the IBM Almaden Research Center. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (3):92–98, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Copler:1991:ICD

[705] Judith Copler. The Indiana Center for Database Systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (3):99–105, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thomasian:1991:CCC

[706] Alexander Thomasian. Centralized concurrency control methods for highend TP. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(3):106–115, September 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1991:RPL

[707] Anonymous. Remembering Peter Lyngbaek. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):4, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kent:1991:BIM

[708] William Kent. The breakdown of the information model in multi-database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):10–15, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ventrone:1991:SHR

[709] V. Ventrone and S. Heiler. Semantic heterogeneity as a result of domain evolution. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):16–20, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kalinichenko:1991:AIR

[710] Leonid Kalinichenko. An anatomy of the information resource semantic abstraction. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):21–24, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eliassen:1991:IOI

[711] Frank Eliassen and Randi Karlsen. Interoperability and object identity. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (4):25–29, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Worboys:1991:SHD

[712] M. F. Worboys and S. M. Deen. Semantic heterogeneity in distributed geographic databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):30–34, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gangopadhyay:1991:SEH

[713] Dipayan Gangopadhyay and T. Barsalou. On the semantic equivalence of heterogeneous representations in multimodel multidatabase systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (4):35–39, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Geller:1991:SSO

[714] J. Geller, Y. Perl, and E. J. Neuhold. Structure and semantics in OODB class specifications. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):40–43, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Saltor:1991:SDM

[715] F. Saltor, M. Castellanos, and M. García-Solaco. Suitability of data models as canonical models for federated databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):44–48, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Spaccapietra:1991:CCA

[716] Stefano Spaccapietra and Christine Parent. Conflicts and correspondence assertions in interoperable databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):49–54, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Urban:1991:RSH

[717] Susan D. Urban and Jian Wu. Resolving semantic heterogeneity through the explicit representation of data model semantics. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):55–58, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fankhauser:1991:SVS

[718] Peter Fankhauser, Martin Kracker, and Erich J. Neuhold. Semantic vs. structural resemblance of classes. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (4):59–63, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Frankhauser:1991:SVS

[719] P. Frankhauser, M. Kracker, and E. J. Neuhold. Semantic vs. structural resemblance of classes. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):59–63, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chatterjee:1991:DMH

[720] Abhirup Chatterjee and Arie Segev. Data manipulation in heterogeneous databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):64–68, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Weishar:1991:DKP

[721] Doyle Weishar and Larry Kerschberg. Data/knowledge packets as a means of supporting semantic heterogeneity in multidatabase systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (4):69–73, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Litwin:1991:FON

[722] Witold Litwin, M. Ketabchi, and Ravi Krishnamurthy. First order normal form for relational databases and multidatabases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):74–76, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Siegel:1991:CIS

[723] Michael Siegel and Stuart E. Madnick. Context interchange: sharing the meaning of data. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):77–78, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yu:1991:DRA

[724] Clement Yu, Biao Jia, Wei Sun, and Son Dao. Determining relationships among names in heterogeneous databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):79–80, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tan:1991:NSS

[725] Kian-Lee Tan and Hongjun Lu. A note on the strategy space of multiway join query optimization problem in parallel systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):81–82, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hsu:1991:MPR

[726] Cheng Hsu. The metadatabase project at Rensselaer. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):83–90, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manola:1991:TGO

[727] Frank Manola. The third-generation/ OODBMS manifesto, commercial version. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):91, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Neimat:1991:DRH

[728] Marie-Anne Neimat and Ming-Chien Shan. Database research at HP labs. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):92–97, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tsur:1991:SRD

[729] Shalom Tsur. Some recent developments in deductive databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 20 (4):98–101, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:1991:FSU

[730] Marianne Winslett. Funding for small US businesses and from DARPA and NASA. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(4):102–104, December 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rowe:1992:RDA

[731] Lawrence A. Rowe. A retrospective on database application development frameworks. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):5–10, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

King:1992:BRD

[732] Roger King and Michael Novak. Building reusable data representations with FaceKit. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):11–17, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Flynn:1992:SDG

[733] Belinda B. Flynn and David Maier. Supporting display generation for complex database objects. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):18–24, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Paredaens:1992:OG

[734] Jan Paredaens, Jan Van den Bussche, Marc Andries, Marc Gemis, Marc Gyssens, Inge Thyssens, Dirk Van Gucht, Vijay Sarathy, and Lawrence Saxton. An overview of GOOD. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):25–31, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Borras:1992:BUI

[735] P. Borras, J. C. Mamou, D. Plateau, B. Poyet, and D. Tallot. Building user interfaces for database applications: the O_{2} experience. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):32–38, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Consens:1992:VQQ

[736] Mariano P. Consens, Isabel F. Cruz, and Alberto O. Mendelzon. Visualizing queries and querying visualizations. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):39–46, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1992:GUI

[737] Yannis E. Ioannidis, Miron Livny, and Eben M. Haber. Graphical user interfaces for the management of scientific experiments and data. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):47–53, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Motro:1992:AAT

[738] Amihai Motro. Annotating answers with their properties (database queries). SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):54–57, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuntz:1992:GGG

[739] Michel Kuntz. The gist of GIUKU: graphical interactive intelligent utilities for knowledgeable users of data base systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):58–64, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

David:1992:ACO

[740] Michael M. David. Advanced capabilities of the outer join (SQL).

SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):65–70, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mishra:1992:FCO

[741] Priti Mishra and Margaret Eich. Functional completeness in object-oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):71–83, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jonsson:1992:FSS

[742] Dan Jonsson. Formal syntax and semantics of a reconstructed relational database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):84–89, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wilson:1992:EMN

[743] Peter R. Wilson. EUG’91 meeting notes. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):90–92, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Soloviev:1992:OTC

[744] Valery Soloviev. An overview of three commercial object-oriented database management systems: ONTOS, ObjectStore, and O_{2}. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):93–104, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pernul:1992:BDS

[745] G. Pernul and G. Luef. Bibliography on database security. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (1):105–122, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Conrad:1992:ABO

[746] Stefan Conrad and Martin Gogolla. An annotated bibliography on objectorientation and deduction. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):123– 132, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Neuhold:1992:DRI

[747] Erich Neuhold and Volker Turau. Database research at IPSI. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (1):133–138, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramamritham:1992:SDR

[748] K. Ramamritham, E. Moss, J. A. Stankovic, D. Stemple, B. Croft, and D. Towsley. Summary of database research activities at the University of Massachusetts, Amherst. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (1):139–143, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dozier:1992:KAA

[749] Jeff Dozier. Keynote address: access to data in NASA’s Earth observing system. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):1, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Merz:1992:DQF

[750] Ulla Merz and Roger King. DI-RECT: a query facility for multiple databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):2, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Celentano:1992:CDB

[751] A. Celentano, M. G. Fugini, and S. Pozzi. Conceptual document browsing and retrieval in Kabiria. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):3, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Johnson:1992:UDM

[752] Rowland R. Johnson, Mandy Goldner, Mitch Lee, Keith McKay, Robert Shectman, and John Woodruff. USD — a database management system for scientific research. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):4, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gesmann:1992:PDS

[753] Michael Gesmann, Andreas Grasnickel, Theo Härder, Christoph Hübel, Wolfgang Käfer, Bernhard Mitschang, and Harald Schöning. PRIMA — a database system supporting dynamically defined composite objects. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21(2): 5, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gray:1992:DTP

[754] Jim Gray. Database and transaction processing benchmarks. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):6, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bancilhon:1992:OOD

[755] François Bancilhon. The O_{2} objectoriented database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):7, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Orenstein:1992:AOD

[756] Jack Orenstein. Architectures for object data management. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):8, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ganguly:1992:QOP

[757] Sumit Ganguly, Waqar Hasan, and Ravi Krishnamurthy. Query optimization for parallel execution. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):9–18, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hong:1992:EIO

[758] Wei Hong. Exploiting inter-operation parallelism in XPRS. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):19–28, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ghandeharizadeh:1992:PAA

[759] Shahram Ghandeharizadeh, David J. DeWitt, and Waheed Qureshi. A performance analysis of alternative multiattribute declustering strategies. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21(2): 29–38, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pirahesh:1992:ERB

[760] Hamid Pirahesh, Joseph M. Hellerstein, and Waqar Hasan. Extensible/rule based query rewrite optimization in Starburst. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):39–48, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hanson:1992:RCT

[761] Eric N. Hanson. Rule condition testing and action execution in Ariel. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):49–58, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aiken:1992:BDP

[762] Alexander Aiken, Jennifer Widom, and Joseph M. Hellerstein. Behavior of database production rules: termination, confluence, and observable determinism. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):59–68, June 1992.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wade:1992:FDO

[763] Andrew E. Wade. Full distribution in Objectivity/DB. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):69, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barry:1992:IDO

[764] Douglas K. Barry. ITASCA distributed ODBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):70, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cruz:1992:DVL

[765] Isabel F. Cruz. DOODLE: a visual language for object-oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):71–80, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gehani:1992:ESA

[766] N. H. Gehani, H. V. Jagadish, and O. Shmueli. Event specification in an active object-oriented database. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21(2): 81–90, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lieuwen:1992:TBA

[767] Daniel F. Lieuwen and David J. De-Witt. A transformation-based approach to optimizing loops in database programming languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):91– 100, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenthal:1992:WCW

[768] Arnon Rosenthal. What can we do to strengthen the connection between theory and system builders (panel). SIG-MOD Record (ACM Special Interest Group on Management of Data), 21(2): 101, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kasi:1992:HPA

[769] Jay Kasi. High performance and availability through data distribution. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21(2): 102, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Davison:1992:PIB

[770] Wayne Davison. Parallel index building in Informix OnLine 6.0. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):103, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1992:UDC

[771] D. Agrawal, A. El Abbadi, and R. Jeffers. Using delayed commitment in locking protocols for real-time databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):104–113, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dan:1992:PAC

[772] Asit Dan and Philip S. Yu. Performance analysis of coherency control policies through lock retention. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):114–123, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mohan:1992:EFM

[773] C. Mohan, Hamid Pirahesh, and Raymond Lorie. Efficient and flexible methods for transient versioning of records to avoid locking by read-only transactions. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):124–133, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Low:1992:HTD

[774] Chee Chin Low, Beng Chin Ooi, and Hongjun Lu. H-trees: a dynamic associative search index for OODB. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):134–143, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tsangaris:1992:POC

[775] Manolis M. Tsangaris and Jeffrey F. Naughton. On the performance of object clustering techniques. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (2):144–153, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ley:1992:TRA

[776] Michael Ley. The term retrieval abstract machine. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):154–163, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bamford:1992:UMI

[777] Roger Bamford. Using multiversioning to improve performance without loss of consistency (abstract only). SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):164, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Descollonges:1992:CMT

[778] Marc Descollonges. A concurrency model for transaction management. SIGMOD Record (ACM Special Interest Group on Management of Data), 21 (2):164, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Franklin:1992:CRC

[779] Michael J. Franklin, Michael J. Zwilling, C. K. Tan, Michael J. Carey, and David J. DeWitt. Crash recovery in client-server EXODUS. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (2):165–174, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jhingran:1992:ARD

[780] Anant Jhingran and Pratap Khedkar. Analysis of recovery in a database system using a write-ahead log protocol.

SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):175–184, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:1992:MRM

[781] David B. Lomet. MLR: a recovery method for multi-level systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):185–194, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kamel:1992:PRT

[782] Ibrahim Kamel and Christos Faloutsos. Parallel R-trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):195–204, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hoel:1992:QCS

[783] Erik G. Hoel and Hanan Samet. A qualitative comparison study of data structures for large line segment databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):205–214, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Analyti:1992:FSM

[784] Anastasia Analyti and Sakti Pramanik. Fast search in main memory databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):215–224, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Celis:1992:DPA

[785] Pedro Celis. Distribution, parallelism, and availability in nonstop SQL. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21(2): 225, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rabinovich:1992:IFT

[786] Michael Rabinovich and Edward D. Lazowska. Improving fault tolerance and supporting partial writes in structured coterie protocols for replicated objects. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):226–235, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bhide:1992:ESP

[787] Anupam Bhide, Ambuj Goyal, Hui-I Hsiao, and Anant Jhingran. An efficient scheme for providing high availability. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):236–245, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Polyzois:1992:ERB

[788] Christos A. Polyzois and Hector Garcia-Molina. Evaluation of remote backup algorithms for transaction processing systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):246–255, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lanzelotte:1992:OOO

[789] Rosana S. G. Lanzelotte, Patrick Valduriez, and Mohamed Zaït. Optimization of object-oriented recursive queries using cost-controlled strategies. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):256–265, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kafer:1992:RTC

[790] Wolfgang Käfer and Harald Schöning. Realizing a temporal complex-object data model. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):266–275, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Biliris:1992:PTD

[791] Alexandros Biliris. The performance of three database storage structures for managing large objects. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (2):276–285, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:1992:ERL

[792] Yun Wang. Experience from a real life query optimizer. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):286, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rengarajan:1992:RVS

[793] T. K. Rengarajan. Rdb/VMS support for multi-media databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):287, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mehrotra:1992:CCP

[794] Sharad Mehrotra, Rajeev Rastogi, Yuri Breitbart, Henry F. Korth, and Avi Silberschatz. The concurrency control problem in multidatabases: characteristics and solutions. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):288–297, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shasha:1992:SRG

[795] Dennis Shasha, Eric Simon, and Patrick Valduriez. Simple rational guidance for chopping up transactions. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):298–307, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rahm:1992:PEE

[796] Erhard Rahm. Performance evaluation of extended storage architectures for transaction processing. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (2):308–317, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cowley:1992:SDM

[797] Paula J. Cowley. Scientific data management: real-world issues and requirements. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):318, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chou:1992:DIP

[798] Hong-Tai Chou. The design and implementation of persistent transactions in an object database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):319, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Harris:1992:HPM

[799] Craig Harris, Madhu Reddy, and Carl Woolf. A high performance multiversion concurrency control protocol for object databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):320, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Terry:1992:CQA

[800] Douglas Terry, David Goldberg, David Nichols, and Brian Oki. Continuous queries over append-only databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):321–330, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Srinivasan:1992:CBL

[801] V. Srinivasan and Michael J. Carey. Compensation-based on-line query processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):331–340, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Haas:1992:SSP

[802] Peter J. Haas and Arun N. Swami. Sequential sampling procedures for query size estimation. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):341–350, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:1992:AMC

[803] David Lomet and Betty Salzberg. Access method concurrency with recovery. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):351–360, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mohan:1992:ACI

[804] C. Mohan and Inderpal Narang. Algorithms for creating indexes for very large tables without quiescing updates. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):361–370, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mohan:1992:AIE

[805] C. Mohan and Frank Levine. ARIES/ IM: an efficient and high concurrency index management method using writeahead logging. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):371–380, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carter:1992:EIM

[806] Fred Carter. Extending INGRES with methods and triggers. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):381, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bigelow:1992:IGC

[807] Richard Bigelow. Implementation of general constraints in SIM. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):382, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cluet:1992:GFO

[808] Sophie Cluet and Claude Delobel. A general framework for the optimization of object-oriented queries. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (2):383–392, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kifer:1992:QOO

[809] Michael Kifer, Won Kim, and Yehoshua Sagiv. Querying object-oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):393–402, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Orenstein:1992:QPO

[810] Jack Orenstein, Sam Haradhvala, Benson Margulies, and Don Sakahara. Query processing in the ObjectStore database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):403–412, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Melmon:1992:SOS

[811] Paul Melmon. The Sybase Open Server. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):413, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Newmann:1992:MVI

[812] Scott Newmann. Multi-vendor interoperability through SQL access. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21(2): 414, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Widom:1992:DSS

[813] Jennifer Widom. A denotational semantics for the Starburst production rule language. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(3):4–9, September 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roddick:1992:SSQ

[814] John F. Roddick. SQL/SE: a query language extension for databases supporting schema evolution. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(3):10–16, September 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pratt:1992:POS

[815] J. Michael Pratt and Maxine Cohen. A process-oriented scientific database model. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(3):17–25, September 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kowarschick:1992:SOW

[816] Wolfgang L. J. Kowarschick. Semantic optimization: what are disjunctive residues useful for? SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (3):26–32, September 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buff:1992:RMC

[817] H. W. Buff. The rational model contra entity relationship? SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (3):33–34, September 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jensen:1992:GTD

[818] C. S. Jensen, J. Clifford, S. K. Gadia, A. Segev, and R. T. Snodgrass. A glossary of temporal database concepts. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(3):35–43, September 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thuraisingham:1992:CSR

[819] Bhavani Thuraisingham. Current status of R&D in trusted database management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(3):44–50, September 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tjoa:1992:DRA

[820] A. M. Tjoa and G. Vinek. Database research activities at the University of Vienna. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(3):51–55, September 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:1992:RRW

[821] David Lomet. A review of recent work on multi-attribute access methods. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(3):56–63, September 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:1992:OUN

[822] Marianne Winslett. Opportunities in the US from NSF, DARPA, and NASA. SIGMOD Record (ACM Special Interest Group on Management of Data), 21 (3):64–67, September 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lu:1992:GMQ

[823] Hongjun Lu, Beng-Chin Ooi, and Cheng-Hian Goh. On global multidatabase query optimization. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (4):6–11, December 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ling:1992:SSB

[824] Yibei Ling and Wei Sun. A supplement to sampling-based methods for query size estimation in a database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(4):12–15, December 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ulusoy:1992:CRR

[825] Özgür Ulusoy. Current research on real-time databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(4):16–21, December 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hung:1992:LPC

[826] S. L. Hung and K. Y. Lam. Locking protocols for concurrency control in real-time database systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 21 (4):22–27, December 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kiessling:1992:CBL

[827] Werner Kießling. A complex benchmark for logic programming and deductive databases, or who can beat the N-queens? SIGMOD Record (ACM Special Interest Group on Management of Data), 21(4):28–34, December 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roddick:1992:SED

[828] John F. Roddick. Schema evolution in database systems: an annotated bibliography. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(4):35–40, December 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sacks-Davis:1992:DRC

[829] Ron Sacks-Davis and Kotagiri Ramamohanarao. Database research at CITRI. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(4):41–47, December 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Papazoglou:1992:DRQ

[830] M. P. Papazoglou, M. McLoughlin, E. Lindsay, and S. Willie. Database research at the Queensland Univ. of Technology. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(4):48–52, December 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zeleznikow:1992:DRT

[831] John Zeleznikow. Database research at La Trobe University. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(4):53–56, December 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:1992:WC

[832] Marianne Winslett. The winds of change? SIGMOD Record (ACM Special Interest Group on Management of Data), 21(4):57–61, December 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Etzion:1993:PDD

[833] Opher Etzion. Pardes — a data driven oriented active database model. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):7–14, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gadia:1993:PDS

[834] Shashi K. Gadia. Parametric databases: seamless integration of spatial, temporal, belief and ordinary data. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):15–20, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tresch:1993:STD

[835] Markus Tresch and Marc H. Scholl. Schema transformation without database reorganization. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):21–27, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Costilla:1993:SDI

[836] C. R. Costilla, M. J. Bas, and J. Villamor. SIRIO: a distributed information system over a heterogeneous computer network. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):28–33, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Imielinski:1993:DMM

[837] Tomasz Imieĺinski and B. R. Badrinath. Data management for mobile computing. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):34–39, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ozsu:1993:WRI

[838] M. Tamer Özsu, Umeshwar Dayal, and Patrick Valduriez. Workshop report: International Workshop on Distributed Object Management. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):40–54, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buff:1993:RTN

[839] H. W. Buff. Remarks on two new theorems of Date and Fagin. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):55–56, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). See response [840].

Date:1993:RRT

[840] C. J. Date and Ronald Fagin. Response to “Remarks on two new theorems of Date and Fagin”. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):57–58, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). See [839].

Al-Taha:1993:BSD

[841] Khaled K. Al-Taha, Richard T. Snodgrass, and Michael D. Soo. Bibliography on spatiotemporal databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):59–67, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barbara:1993:ESD

[842] Daniel Barbará. Extending the scope of database services. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):68–73, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1993:DRW

[843] Anonymous. Database research at Wisconsin. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):74–81, March 1993.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

UWisconsin:1993:DRW

[844] Univ. of Wisconsin. Database research at Wisconsin. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):74–81, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:1993:DRA

[845] H. V. Jagadish. Database research at AT&T Bell Laboratories. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):82–88, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:1993:DRF

[846] Marianne Winslett. Database research funding: Change at ONR, and many funding announcements elsewhere. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(1):89–95, March 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1993:SSB

[847] Michael Stonebraker, Jim Frew, Kenn Gardels, and Jeff Meredith. The SE-QUOIA 2000 storage benchmark. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22(2): 2–11, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1993:B

[848] Michael J. Carey, David J. DeWitt, and Jeffrey F. Naughton. The 007 Benchmark. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):12–21, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Leutenegger:1993:MST

[849] Scott T. Leutenegger and Daniel Dias. A modeling study of the TPC-C benchmark. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):22–31, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:1993:MR

[850] Serge Abiteboul, Georg Lausen, Heinz Uphoff, and Emmanuel Waller. Methods and rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):32–41, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brant:1993:ISR

[851] David A. Brant and Daniel P. Miranker. Index support for rule activation. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):42–48, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gupta:1993:LVG

[852] Ashish Gupta and Jennifer Widom. Local verification of global integrity constraints in distributed databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22(2):

49–58, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pang:1993:PPH

[853] HweeHwa H. Pang, Michael J. Carey, and Miron Livny. Partially preemptible hash joins. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):59–68, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lo:1993:OPA

[854] Ming-Ling Lo, Ming-Syan Chen, C. V. Ravishankar, and Philip S. Yu. On optimal processor allocation to support pipelined hash joins. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):69–78, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sun:1993:IAS

[855] Wei Sun, Yibei Ling, Naphtali Rishe, and Yi Deng. An instant and accurate size estimation method for joins and selection in a retrieval-intensive environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):79–88, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Meseguer:1993:LSO

[856] José Meseguer and Xiaolei Qian. A logical semantics for object-oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):89–98, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anwar:1993:NPR

[857] E. Anwar, L. Maugis, and S. Chakravarthy. A new perspective on rule support for object-oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):99– 108, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ananthanarayanan:1993:UCE

[858] R. Ananthanarayanan, V. Gottemukkala, W. Kaefer, T. J. Lehman, and H. Pirahesh. Using the co-existence approach to achieve combined functionality of object-oriented and relational systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):109–118, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shatdal:1993:USV

[859] Ambuj Shatdal and Jeffrey F. Naughton. Using shared virtual memory for parallel join processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):119–128, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tomasic:1993:CDS

[860] Anthony Tomasic and Héctor García-Molina. Caching and database scaling in distributed shared-nothing information retrieval systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):129– 138, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mohan:1993:EFM

[861] C. Mohan and Inderpal Narang. An efficient and flexible method for archiving a data base. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):139–146, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Derr:1993:DIG

[862] Marcia A. Derr, Shinichi Morishita, and Geoffrey Phipps. Design and implementation of the glue-nail database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):147–156, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gupta:1993:MVI

[863] Ashish Gupta, Inderpal Singh Mumick, and V. S. Subrahmanian. Maintaining views incrementally. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):157–166, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramakrishnan:1993:ICD

[864] Raghu Ramakrishnan, Divesh Srivastava, S. Sudarshan, and Praveen Seshadri. Implementation of the CORAL deductive database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):167– 176, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kolodner:1993:AIG

[865] Elliot K. Kolodner and William E. Weihl. Atomic incremental garbage collection and recovery for a large stable heap. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):177–186, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keen:1993:PEE

[866] John S. Keen and William J. Dally. Performance evaluation of ephemeral logging. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):187–196, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hong:1993:RTT

[867] D. Hong, T. Johnson, and S. Chakravarthy. Real-time transaction scheduling: a cost conscious approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):197– 206, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1993:MAR

[868] Rakesh Agrawal, Tomasz Imieliński, and Arun Swami. Mining association rules between sets of items in large databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):207–216, June 1993.

CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Borgida:1993:LDD

[869] Alex Borgida and Ronald J. Brachman. Loading data into description reasoners. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):217–226, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:1993:TMA

[870] X. Sean Wang, Sushil Jajodia, and V. S. Subrahmanian. Temporal modules: an approach toward federated temporal databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):227–236, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brinkhoff:1993:EPS

[871] Thomas Brinkhoff, Hans-Peter Kriegel, and Bernhard Seeger. Efficient processing of spatial joins using R-trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):237–246, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ishikawa:1993:ESF

[872] Yoshiharu Ishikawa, Hiroyuki Kitagawa, and Nobuo Ohbo. Evaluation of signature files as set access facilities in OODBs. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):247–256, June 1993.

CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Curewitz:1993:PPD

[873] Kenneth M. Curewitz, P. Krishnan, and Jeffrey Scott Vitter. Practical prefetching via data compression. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):257–266, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hellerstein:1993:PMO

[874] Joseph M. Hellerstein and Michael Stonebraker. Predicate migration: optimizing queries with expensive predicates. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):267–276, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guting:1993:SOS

[875] Ralf Hartmut Güting. Second-order signature: a tool for specifying data models, query processing, and optimization. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):277–286, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blakeley:1993:EBO

[876] José A. Blakeley, William J. McKenna, and Goetz Graefe. Experiences building the open OODB query optimizer. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):287–296, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

ONeil:1993:LKP

[877] Elizabeth J. O’Neil, Patrick E. O’Neil, and Gerhard Weikum. The LRU-K page replacement algorithm for database disk buffering. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):297– 306, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Orji:1993:DDM

[878] Cyril U. Orji and Jon A. Solworth. Doubly distorted mirrors. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):307– 316, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hou:1993:CRA

[879] Robert Y. Hou and Yale N. Patt. Comparing rebuild algorithms for mirrored and RAID5 disk arrays. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):317– 326, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Litwin:1993:LLH

[880] Witold Litwin, Marie-Anne A. Neimat, and Donovan A. Schneider. LH^{∗} — linear hashing for distributed files. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):327–336, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Johnson:1993:LUD

[881] Theodore Johnson and Padmashree Krishna. Lazy updates for distributed search structure. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):337–346, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:1993:ALP

[882] Jianzhong Li, Doron Rotem, and Jaideep Srivastava. Algorithms for loading parallel grid files. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):347– 356, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vadaparty:1993:TUV

[883] K. Vadaparty, Y. A. Aslandogan, and G. Özsoyoğlu. Towards a unified visual database access. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):357–366, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Watters:1993:IRR

[884] Aaron Watters. Interpreting a reconstructed relational calculus (extended abstract). SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):367–376, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Beeri:1993:PAR

[885] Catriel Beeri and Tova Milo. On the power of algebras with recursion. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):377–386, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Alonso:1993:DSI

[886] Rafael Alonso and Henry F. Korth. Database system issues in nomadic computing. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):388–392, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dayal:1993:TGT

[887] Umesh Dayal, Héctor García-Molina, Mei Hsu, Ben Kao, and Ming-Chien Shan. Third generation TP monitors: a database challenge. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):393–397, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Egenhofer:1993:WSA

[888] Max J. Egenhofer. What’s special about spatial?: database requirements for vehicle navigation in geographic space. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):398–402, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ordille:1993:DCG

[889] Joann J. Ordille and Barton P. Miller. Database challenges in global information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):403–407, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zdonik:1993:IDS

[890] Stanley B. Zdonik. Incremental database systems: databases from the ground up. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):408–412, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1993:THD

[891] Michael J. Carey, Laura M. Haas, and Miron Livny. Tapes hold data, too: challenges of tuples on tertiary store. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):413–417, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:1993:IMD

[892] H. V. Jagadish, Ed Fox, Narain Gehani, Ramesh Jain, Hanan Samet, Mike Stonebraker, and John Thompson. Issues in multimedia databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 22 (2):419, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Motro:1993:WTA

[893] Amihai Motro. What to teach about databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):420, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shan:1993:PAD

[894] Ming-Chien Shan. Pegasus architecture and design principles. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):422–425, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bukhres:1993:ISB

[895] Omran Bukhres, Jiansan Chen, and Rob Pezzoli. An InterBase system at BNR. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):426–429, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Klein:1993:ODC

[896] Johannes Klein and Francis Upton IV. Open DECdtm: constraint based transaction management. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):430– 433, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wiederhold:1993:III

[897] Gio Wiederhold. Intelligent integration of information. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):434–437, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vieille:1993:DOO

[898] Laurent Vieille. A deductive and object-oriented database system: why and how? SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):438, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1993:MD

[899] Michael Stonebraker. The Miro DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 22 (2):439, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Velez:1993:MTM

[900] Fernando Vélez. Modularity and tuning mechanisms in the O_{2} system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22(2): 440, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wade:1993:SLV

[901] Andrew E. Wade. Single logical view over enterprise-wide distributed databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):441–444, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mohan:1993:IRD

[902] C. Mohan. IBM’s relational DBMS products: features and technologies. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):445–448, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fushimi:1993:GCD

[903] Shinya Fushimi and Masaru Kitsuregawa. GREO: a commercial database processor based on a pipelined hardware sorter. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):449–452, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tseng:1993:PDP

[904] Emy Tseng and David Reiner. Parallel database processing on the KSR1 computer. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):453–455, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jin:1993:CCR

[905] W. Woody Jin, Marek Rusinkiewicz, Linda Ness, and Amit Sheth. Concurrency control and recovery of multidatabase work flows in telecommunication applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):456–459, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sherman:1993:AED

[906] Mark Sherman. Architecture of the Encina distributed transaction processing family. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):460–463, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Colton:1993:RDD

[907] Malcolm Colton. Replicated data in a distributed environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):464– 466, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Singhal:1993:DOO

[908] Anoop Singhal, Robert M. Arlein, and Chi-Yuan Lo. DDB: an object oriented design data manager for VLSI CAD. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):467–470, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nassif:1993:IAM

[909] Rodolphe Nassif and Don Mitchusson. Issues and approaches for migration/ cohabitation between legacy and new systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):471–474, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thieman:1993:IDN

[910] James R. Thieman. The International Directory Network and connected data information systems for research in the Earth and space sciences. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):475– 478, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mukhopadhyay:1993:IUA

[911] Debajyoti Mukhopadhyay. Interoperability using APPC. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):479–482, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sheth:1993:MII

[912] Amit P. Sheth and George Karabatis. Multidatabase interdependencies in industry. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):483–486, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cohen:1993:RIB

[913] David Cohen, Gary Larson, and Larry Berke. Role of interoperability in business application development. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (2):487–490, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Woelk:1993:TSU

[914] Darrell Woelk, Paul Attie, Phil Cannata, Greg Meredith, Amit Sheth, Munindar Singh, and Christine Tomlinson. Task scheduling using intertask dependencies in Carnot. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):491– 494, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mannai:1993:EIO

[915] Dhamir N. Mannai and Khaled Bugrara. Enhancing inter-operability and data sharing in medical information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):495–498, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Woyna:1993:MBS

[916] Mark A. Woyna, John H. Christiansen, Christopher W. Hield, and Kathy Lee Simunich. Modeling battlefield sensor environments with an object database management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):499–501, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schaller:1993:ICM

[917] Tony Schaller. The INtersect concept for multidatabase system integration in the pharmaceutical industry. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):502– 503, June 1993. CODEN SRECD8.

ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gemis:1993:GGO

[918] Marc Gemis, Jan Paredaens, Inge Thyssens, and Jan Van den Bussche. GOOD: a graph-oriented object database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):505–510, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Consens:1993:HHB

[919] Mariano P. Consens and Alberto O. Mendelzon. Hy+: a Hygraph-based query and visualization system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (2):511–516, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chu:1993:DIC

[920] Wesley W. Chu, Matthew A. Merzbacher, and Ladislav Berkovich. The design and implementation of CoBase. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):517–522, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keller:1993:PSB

[921] Arthur M. Keller, Richard Jensen, and Shailesh Agarwal. Persistence software: bridging object-oriented programming and relational databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):523– 528, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kupper:1993:NCN

[922] D. Küpper, M. Storbel, and D. Rösner. NAUDA: a cooperative natural language interface to relational databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):529–533, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bukhres:1993:IMP

[923] Omran A. Bukhres, Jiansan Chen, Ahmed K. Elmagarmid, Xiangning Liu, and James G. Mullen. Inter-Base: a multidatabase prototype systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):534–539, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Su:1993:OKO

[924] Stanley Y. W. Su, Herman X. Lam, Srinivasa Eddula, Javier Arroyo, Neeta Prasad, and Ronghao Zhuang. OSAM^{∗}.KBMS: an objectoriented knowledge base management system for supporting advanced applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):540–541, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Moenkeberg:1993:CPS

[925] Axel Moenkeberg, Peter Zabback, Christof Hasse, and Gerhard Weikum. The COMFORT prototype: a step towards automated database performance tuning. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):542–543, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramakrishnan:1993:CDD

[926] Raghu Ramakrishnan, William G. Roth, Praveen Seshadri, Divesh Srivastava, and S. Sudarshan. The CORAL deductive database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):544– 545 (or 541–545??), June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Andersson:1993:SP

[927] M. Andersson, A.-M. Auddino, Y. Dupont, E. Fontana, M. Gentile, and S. Spaccapietra. The “SUPER” project. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (2):546–547, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Polyachenko:1993:ICP

[928] Boris E. Polyachenko and Filipp I. Andon. Instrumental complex of parallel software system development and operating environment support for distributed processing within multitransputer systems, TRANSSOFT. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (2):548–549, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cacace:1993:LP

[929] F. Cacace, S. Ceri, S. Crespi-Reghizzi, P. Fraternali, S. Paraboschi, and L. Tanca. The LOGRES prototype. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):550–551, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ford:1993:OOM

[930] Steve Ford, José A. Blakeley, and Thomas J. Bannon. Open OODB: a modular object-oriented DBMS. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (2):552–553, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hasan:1993:PGD

[931] W. Hasan, M. Heytens, C. Kolovson, M.-A. Neimat, S. Potamianos, and D. Schneider. Papyrus GIS demonstration. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):554–555, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rakow:1993:VVS

[932] Thomas C. Rakow and Peter Muth. The V^{3} video server — managing analog and digital video clips. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):556– 557, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Muth:1993:VON

[933] Peter Muth and Thomas C. Rakow. VODAK open nested transactions — visualizing database internals. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (2):558–559, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Luniewski:1993:IOU

[934] Allen Luniewski, Peter Schwarz, Kurt Shoens, Jim Stamos, and John Thomas. Information organization using Rufus. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):560–561, June 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Arens:1993:SRI

[935] Yigal Arens and Craig Knoblock. SIMS: Retrieving and integrating information from multiple sources. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (2):562–563, June 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:1993:SGO

[936] Marianne Winslett. SIGMOD goes online: new member services via Internet. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(3):4–6, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Athanasiu:1993:HCS

[937] Irina Athanasiu. Helping computer scientists in Romania. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(3):7, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pissinou:1993:TMC

[938] Niki Pissinou, Kia Makki, and Yelena Yesha. On temporal modeling in the context of object databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (3):8–15, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Monk:1993:SEO

[939] Simon Monk and Ian Sommerville. Schema evolution in OODBs using class versioning. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(3):16–22, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Breitbart:1993:MAC

[940] Y. Breitbart, A. Deacon, H.-J. Schek, A. Sheth, and G. Weikum. Merging application-centric and data-centric approaches to support transactionoriented multi-system workflows. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (3):23–30, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Breibart:1993:MAC

[941] Y. Breibart, A. Deacon, H.-J. Schek, A. Sheth, and G. Weikum. Merging application-centric and data-centric approaches to support transactionoriented multi-system workflows. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (3):23–??, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roth:1993:DC

[942] Mark A. Roth and Scott J. Van Horn. Database compression. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (3):31–39, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Reddy:1993:DPD

[943] P. Krishna Reddy and Subhash Bhalla. Deadlock prevention in a distributed database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(3):40–46, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Drew:1993:RWS

[944] Pam Drew, Roger King, Dennis McLeod, Marek Rusinkiewicz, and Avi Silberschatz. Report of the Workshop on Semantic Heterogeneity and Interoperation in Multidatabase Systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 22 (3):47–56, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jain:1993:NWV

[945] Ramesh Jain. NSF workshop on Visual Information Management Systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(3):57–75, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Graefe:1993:OPD

[946] Goetz Graefe. Options in physical database design. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(3):76–83, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

UFlorida:1993:DRU

[947] University of Florida. Database research at the University of Florida. SIGMOD Record (ACM Special Interest Group on Management of Data), 22 (3):84–90, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Orlowska:1993:DRU

[948] Maria E. Orlowska. Database research at the University of Queensland. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (3):91–95, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:1993:TAF

[949] Marianne Winslett and Iris Sheauyin Chu. Timely access to future funding announcements. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(3):96–98, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jeffrey:1993:DCC

[950] Keith G. Jeffrey. Database conference calendar. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(3):99–101, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Scholerman:1993:RDI

[951] S. Scholerman, L. Miller, J. Tenner, S. Tomanek, and M. Zolliker. Relational database integration in the IBM AS/400. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):5–10, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Arpinar:1993:MAG

[952] İsmailcem Budak Arpinar, Asuman Doğaç, and Cem Evrendilek. Moodview: An advanced graphical user interface for OODBMSs. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):11–18, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wiil:1993:EHH

[953] Uffe Kock Wiil. Experiences with HyperBase: a hypertext database supporting collaborative work. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (4):19–25, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Levene:1993:IGB

[954] Mark Levene, Alexandra Poulovassilis, Kerima Benkerimi, Sara Schwartz, and Eran Tuv. Implementation of a graphbased data model for complex objects. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):26–31, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rahm:1993:PQP

[955] Erhard Rahm. Parallel query processing in shared disk database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):32–37, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gruenwald:1993:PSC

[956] Le Gruenwald and Sichen Liu. A performance study of concurrency control in a real-time main memory database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):38–44, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nyanchama:1993:RBS

[957] Matunda Nyanchama and Sylvia Osborn. Role-based security, object oriented databases and separation of duty. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):45–51, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thuraisingham:1993:CCT

[958] Bhavani Thuraisingham and Hai-Ping Ko. Concurrency control in trusted database management systems: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):52–59, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thurasisingham:1993:CCT

[959] Bhavani Thurasisingham and Hai-Ping Ko. Concurrency control in trusted database management systems: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):52–??, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:1993:SUS

[960] Hongjun Lee, Hock Chuan Chan, and Kwok Kee Wei. A survey on usage of SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):60–65, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lu:1993:SUS

[961] Hongjun Lu, Hock Chuan Chan, and Kwok Kee Wei. A survey on usage of SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):60–65, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kline:1993:UTD

[962] Nick Kline. An update of the temporal database bibliography. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (4):66–80, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Maier:1993:DRD

[963] David Maier, Lois Delcambre, Calton Pu, Jon Walpole, Goetz Graefe, and Len Shapiro. Database research at the Data-Intensive Systems Center. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (4):81–86, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Apers:1993:DRU

[964] P. M. G. Apers, H. Balsters, H. M. Blanken, R. A. de By, M. A. W. Houtsma, P. A. C. Verkoulen, and A. N. Wilschut. Database research at the University of Twente. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22 (4):87–91, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:1993:NHU

[965] Marianne Winslett. NSF and HPCC under attack. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):92–95, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1994:MBK

[966] Michael Stonebraker. In memory of Bob Kooi (1951–1993). SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):3, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kim:1994:OOP

[967] Won Kim. Observations on the ODMG-93 proposal for an objectoriented database language. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):4–9, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dreyer:1994:RPT

[968] Werner Dreyer, Angelika Kotz Dittrich, and Duri Schmidt. Research perspectives for time series management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):10–15, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yang:1994:HQL

[969] Li Yang. A hypertext query language for images. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):16–20, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Diaz:1994:SDD

[970] Oscar Díaz, Arturo Jaime, Norman W. Paton, and Ghassan al Qaimari. Supporting dynamic displays using active rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):21–26, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Geppert:1994:CND

[971] Andreas Geppert and Klaus R. Dittrich. Constructing the next 100 database management systems: like the handyman or like the engineer? SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):27–33, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:1994:SOT

[972] Richard T. Snodgrass. Section overview: Temporal database infrastructure. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):34–??, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pissinou:1994:TIT

[973] Niki Pissinou, Richard T. Snodgrass, Ramez Elmasri, Inderpal S. Mumick, Tamer Özsu, Barbara Pernici, Arie Segev, Babis Theodoulidis, and Umeshwar Dayal. Towards an infrastructure for temporal databases: report of an invitational ARPA/NSF workshop. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):35–51, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dyreson:1994:CGT

[974] Curtis Dyreson, Fabio Grandi, Wolfgang Käfer, Nick Kline, Nikos Lorentzos, Yannis Mitsopoulos, Angelo Montanari, Daniel Nonen, Elisa Peressi, Barbara Pernici, John F. Roddick, Nandlal L. Sarda, Maria Rita Scalas, Arie Segev, Richard T. Snodgrass, Mike D. Soo, Abdullah Tansel, Paolo Tiberio, and Gio Wiederhold. A consensus glossary of temporal database concepts. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):52–64, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jensen:1994:CGT

[975] Christian S. Jensen, J. Clifford, R. Elmasri, S. K. Gadia, P. Hayes, S. Jajodia, C. Dyreson, F. Grandi, W. Kafer, N. Kline, N. Lorentzos, Y. Mitsopoulos, A. Montanari, D. Nonen, E. Peressi, B. Pernici, J. F. Roddick, N. L. Sarda, M. R. Scalas, A. Segev, R. T. Snodgrass, M. D. Soo, A. Tansel, P. Tiberio, and G. Wiederhold. A consensus glossary of temporal database concepts. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):52–64, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:1994:TLS

[976] Richard T. Snodgrass, Ilsoo Ahn, Gadi Ariav, Don Batory, James Clifford, Curtis E. Dyreson, Ramez Elmasri, Fabio Grandi, Christian S. Jensen, Wolfgang Käfer, Nick Kline, Krishna Kulkarni, T. Y. Cliff Leung, Nikos Lorentzos, John F. Roddick, Arie Segev, Michael D. Soo, and Suryanarayana M. Sripada. TSQL2 language specification. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):65–86, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buneman:1994:CS

[977] Peter Buneman, Leonid Libkin, Dan Suciu, Val Tannen, and Limsoon Wong. Comprehension syntax. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):87–96, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Loeffen:1994:TDS

[978] Arjan Loeffen. Text databases: a survey of text models and systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):97–106, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Medeiros:1994:DG

[979] Claudia Bauzer Medeiros and Fatima Pires. Databases for GIS. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (1):107–115, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Medieros:1994:DG

[980] Claudia Bauzer Medieros and Fatima Pires. Databases for GIS. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (1):107–??, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Norrie:1994:DRG

[981] Moira C. Norrie, Stephen M. Blott, Hans-Jörg Schek, and Gerhard Weikum. The database research group at ETH Zürich. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):116–121, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1994:DRG

[982] Anonymous. The database research group at ETH Zurich. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):116–??, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gallagher:1994:IDL

[983] Leonard Gallagher. Influencing database language standards. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (1):122–127, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gallaghar:1994:DSA

[984] Leonard Gallaghar. Database standards activities. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):122–??, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Alonso:1994:TPN

[985] Rafael Alonso. Trade press news. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):128–??, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:1994:PHN

[986] Marianne Winslett. Progress on HPCC and NII. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):129–133, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:1994:DRF

[987] Marianne Winslett. Database research funding. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):129–??, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barbara:1994:SWC

[988] Daniel Barbará and Tomasz Imieliński. Sleepers and workaholics: Caching strategies in mobile environments. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):1–12, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Huang:1994:DRM

[989] Yixiu Huang, Prasad Sistla, and Ouri Wolfson. Data replication for mobile computers. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):13–24, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Imielinski:1994:EEI

[990] Tomasz Imieliński, S. Viswanathan, and B. R. Badrinath. Energy efficient indexing on air. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):25–36, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ahn:1994:SCP

[991] Ilsoo Ahn. SIGMOD challenges paper: database issues in telecommunications network management. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):37–43, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Biliris:1994:ASS

[992] A. Biliris, S. Dar, N. Gehani, H. V. Jagadish, and K. Ramamritham. AS-SET: a system for supporting extended transactions. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):44–54, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mohan:1994:ACM

[993] C. Mohan and Inderpal Narang. ARIES/CSA: a method for database recovery in client-server architectures. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):55–66, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhang:1994:ERA

[994] Aidong Zhang, Marian Nodine, Bharat Bhargava, and Omran Bukhres. Ensuring relaxed atomicity for flexible transactions in multidatabase systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):67–78, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berson:1994:SSM

[995] Steven Berson, Shahram Ghandeharizadeh, Richard Muntz, and Xiangyu Ju. Staggered striping in multimedia information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):79–90, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gibbs:1994:DMT

[996] Simon Gibbs, Christian Breiteneder, and Dennis Tsichritzis. Data modeling of time-based media. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):91–102, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mumick:1994:IMS

[997] Inderpal Singh Mumick and Hamid Pirahesh. Implementation of magicsets in a relational database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):103–114, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:1994:CPD

[998] Jason Tsong-Li Wang, Gung-Wei Chirn, Thomas G. Marr, Bruce Shapiro, Dennis Shasha, and Kaizhong Zhang. Combinatorial pattern discovery for scientific data: Some preliminary results. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):115–125, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gravano:1994:EGT

[999] Luis Gravano, Héctor García-Molina, and Anthony Tomasic. The effectiveness of GIOSS for the text database discovery problem. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):126–137, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1994:IPE

[1000] Yannis E. Ioannidis and Yezdi Lashkari. Incomplete path expressions and their disambiguation. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):138– 149, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cole:1994:ODQ

[1001] Richard L. Cole and Goetz Graefe. Optimization of dynamic query evaluation plans. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):150–160, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:1994:ASE

[1002] Chungmin Melvin Chen and Nick Roussopoulos. Adaptive selectivity estimation using query feedback. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (2):161–172, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Swami:1994:EPF

[1003] Arun Swami and K. Bernhard Schiefer.

Estimating page fetches for index scans with finite LRU buffers. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):173– 184, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hsiao:1994:PEM

[1004] Hui-I Hsiao, Ming-Syan Chen, and Philip S. Yu. On parallel execution of multiple pipelined hash joins. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):185–196, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brinkhoff:1994:MSP

[1005] Thomas Brinkhoff, Hans-Peter Kriegel, Ralf Schneider, and Bernhard Seeger. Multi-step processing of spatial joins. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):197–208, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lo:1994:SJU

[1006] Ming-Ling Lo and Chinya V. Ravishankar. Spatial joins using seeded trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):209–220, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pang:1994:MMR

[1007] HweeHwa Pang, Michael J. Carey, and Miron Livny. Managing memory for real-time queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):221–232, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nyberg:1994:ARM

[1008] Chris Nyberg, Tom Barclay, Zarka Cvetanovic, Jim Gray, and Dave Lomet. AlphaSort: a RISC machine sort. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):233–242, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gray:1994:QGB

[1009] Jim Gray, Prakash Sundaresan, Susanne Englert, Ken Baclawski, and Peter J. Weinberger. Quickly generating billion-record synthetic databases.

SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):243–252, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vingralek:1994:DFO

[1010] Radek Vingralek, Yuri Breitbart, and Gerhard Weikum. Distributed file organization with scalable cost/ performance. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):253–264, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kroll:1994:DST

[1011] Brigitte Kröll and Peter Widmayer.

Distributing a search tree among a growing number of processors. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (2):265–276, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kroell:1994:DST

[1012] B. Kroell and P. Widmayer. Distributing a search tree among a growing number of processors. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):265–??, ???? 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dewan:1994:PDL

[1013] Hasanat M. Dewan, Salvatore J.

Stolfo, Mauricio Hernández, and Jae-Jun Hwang. Predictive dynamic load balancing of parallel and distributed rule and query processing. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (2):277–288, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tomasic:1994:IUI

[1014] Anthony Tomasic, Héctor García-Molina, and Kurt Shoens. Incremental updates of inverted lists for text document retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):289–300, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Consens:1994:OQF

[1015] Mariano P. Consens and Tova Milo. Optimizing queries on files. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (2):301–312, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christophides:1994:SDN

[1016] V. Christophides, S. Abiteboul, S. Cluet, and M. Scholl. From structured documents to novel query facilities. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):313–324, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hellerstein:1994:PPP

[1017] Joseph M. Hellerstein. Practical predicate placement. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):325–335, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kemper:1994:ODQ

[1018] A. Kemper, G. Moerkotte, K. Peithner, and M. Steinbrunn. Optimizing disjunctive queries with expensive predicates. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):336–347, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Galindo-Legaria:1994:OD

[1019] César A. Galindo-Legaria. Outerjoins as disjunctions. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):348–358, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1994:FGS

[1020] Michael J. Carey, Michael J. Franklin, and Markos Zaharioudakis. Finegrained sharing in a page server OODBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):359–370, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cook:1994:PSP

[1021] Jonathan E. Cook, Alexander L. Wolf, and Benjamin G. Zorn. Partition selection policies in object database garbage collection. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):371–382, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1994:SPA

[1022] Michael J. Carey, David J. De-Witt, Michael J. Franklin, Nancy E. Hall, Mark L. McAuliffe, Jeffrey F.

Naughton, Daniel T. Schuh, Marvin H. Solomon, C. K. Tan, Odysseas G. Tsatalos, Seth J. White, and Michael J. Zwilling. Shoring up persistent applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):383–394, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

White:1994:QHP

[1023] Seth J. White and David J. De-Witt. QuickStore: a high performance mapped object store. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):395–406, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McIver:1994:SAL

[1024] William J. McIver, Jr. and Roger King. Self-adaptive, on-line reclustering of complex object data. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):407–418, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Faloutsos:1994:FSM

[1025] Christos Faloutsos, M. Ranganathan, and Yannis Manolopoulos. Fast subsequence matching in time-series databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):419–429, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Seshadri:1994:SQP

[1026] Praveen Seshadri, Miron Livny, and Raghu Ramakrishnan. Sequence query processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):430–441, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sagonas:1994:XED

[1027] Konstantinos Sagonas, Terrance Swift, and David S. Warren. XSB as an efficient deductive database engine. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (2):442–453, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dar:1994:PST

[1028] Shaul Dar and Raghu Ramakrishnan.

A performance study of transitive closure algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):454–465, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1994:SPD

[1029] Michael J. Carey. Session: Parallel database systems in the 1990’s. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23(2): 466, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Daniels:1994:OSR

[1030] Dean Daniels, Lip Boon Doo, Alan Downing, Curtis Elsbernd, Gary Hallmark, Sandeep Jain, Bob Jenkins, Peter Lim, Gordon Smith, Benny Souder, and Jim Stamos. Oracle’s symmetric replication technology and implications for application design. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):467, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dietterich:1994:DDD

[1031] Daniel J. Dietterich. DEC data distributor: for data replication and data warehousing. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):468, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gorelik:1994:SRS

[1032] Alex Gorelik, Yongdong Wang, and Mark Deppe. Sybase replication server. SIGMOD Record (ACM Special Interest Group on Management of Data), 23 (2):469, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Laursen:1994:OMS

[1033] Andrew Laursen, Jeffrey Olkin, and Mark Porter. Oracle media server: Providing consumer based interactive access to multimedia data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (2):470–477, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kulkarni:1994:OOE

[1034] Krishna G. Kulkarni. Object-oriented extensions in SQL3: a status report. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):478, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Martin:1994:CCO

[1035] Bruce E. Martin. COSS: The Common Object Services Specifications. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23(2): 479, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cattell:1994:OSO

[1036] R. G. G. Cattell. ODMG-93: a standard for object-oriented DBMSs. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23(2): 480, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kim:1994:UXU

[1037] Won Kim. UniSQL/X unified relational and object-oriented database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):481, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ubell:1994:MED

[1038] Michael Ubell. The Montage extensible DataBlade architecture. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):482, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pirahesh:1994:OOF

[1039] Hamid Pirahesh. Object-oriented features of DB2 client/server. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):483, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vaskevitch:1994:DCT

[1040] David Vaskevitch. Database in crisis and transition: a technical agenda for the year 2001. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):484–489, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ballinger:1994:ETD

[1041] Carrie Ballinger. Evolving teradata decision support for massively parallel processing with UNIX. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):490, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Englert:1994:NSS

[1042] Susanne Englert. NonStop SQL: Scalability and availability for decision support. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):491, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fernandez:1994:RBW

[1043] Phillip M. Fernandez. Red brick warehouse: a read-mostly RDBMS for open SMP platforms. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):492, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Melling:1994:EIA

[1044] Wesley P. Melling. Enterprise information architectures — they’re finally changing. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):493–504, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Singh:1994:RTP

[1045] Munindar P. Singh, Christine Tomlinson, and Darrell Woelk. Relaxed transaction processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):505, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Flokstra:1994:IDD

[1046] Jan Flokstra, Maurice van Keulen, and Jacek Skowronek. The IM-PRESS DDT: a database design toolbox based on a formal specification language. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):506, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bryan:1994:MPA

[1047] G. M. Bryan, W. E. Moore, B. J.

Curry, K. W. Lodge, and J. Geyer. The MEDUSA project: Autonomous data management in a shared-nothing parallel database machine. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):507, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Missikoff:1994:MSC

[1048] M. Missikoff and M. Toiati. MOSAICO — a system for conceptual modeling and rapid prototyping of objectoriented database application. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23(2): 508, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuhn:1994:LBM

[1049] Eva Kühn, Thomas Tschernko, and Konrad Schwarz. A language based multidatabase system. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):509, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuehn:1994:LBM

[1050] E. Kuehn, T. Tschernko, and K. Schwarz. A language based multidatabase system. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):509–??, ???? 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grossman:1994:PSP

[1051] R. L. Grossman and X. Qin. Ptool: a scalable persistent object manager. SIGMOD Record (ACM Special Interest Group on Management of Data), 23 (2):510, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Theodoulidis:1994:OTD

[1052] Babis Theodoulidis, Aziz Ait-Braham, George Andrianopoulos, Jayant Chaudhary, George Karvelis, and Simon Sou. The ORES temporal database management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):511, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sagonas:1994:XDD

[1053] Konstantinos Sagonas, Terrance Swift, and David S. Warren. XSB as a deductive database. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):512, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dogac:1994:MOO

[1054] Asuman Dogac, Budak Arpinar, Cem Evrendilek, Cetin Ozkan, Ilker Altintas, Ilker Durusoy, Mehmet Altinel, Tansel Okay, and Yuksel Saygin. METU object-oriented DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 23 (2):513, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1994:QPD

[1055] R. Agrawal, M. Carey, C. Faloutsos, S. Ghosh, M. Houtsma, T. Imieliński, B. Iyer, A. Mahboob, H. Miranda, R. Srikant, and A. Swami. Quest: a project on database mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):514, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Catarci:1994:QDG

[1056] Tiziana Catarci and Giuseppe Santucci. Query by diagram: a graphical environment for querying databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 23 (2):515, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:1994:DSP

[1057] Jiawei Han, Yongjian Fu, Yue Huang, Yandong Cai, and Nick Cercone. DBLearn: a system prototype for knowledge discovery in relational databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):516, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Biliris:1994:EEO

[1058] Alexandros Biliris and Euthimios Panagos. EOS: An Extensible Object Store. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):517, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hwang:1994:MFD

[1059] S.-Y. Hwang, E.-P. Lim, H.-R. Yang, S. Musukula, K. Mediratta, M. Ganesh, D. Clements, J. Stenoien, and J. Srivastava. The MYRIAD federated database prototype. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):518, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brinkhoff:1994:GSE

[1060] Thomas Brinkhoff, Hans-Peter Kriegel, Ralf Schneider, and Bernhard Seeger. GENESYS: a system for efficient spatial query processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):519, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blakeley:1994:OOD

[1061] José A. Blakeley. Open object database management systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23(2): 520, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mohan:1994:SCA

[1062] C. Mohan. A survey and critique of advanced transaction models. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):521, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:1994:DN

[1063] H. V. Jagadish. Databases for networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):522, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

ODMG:1994:ROC

[1064] Object Database Management Group.

Response to the ODMG-93 commentary written by Dr. Won Kim of UniSQL, Inc. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(3):3–7, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Datta:1994:RID

[1065] Anindya Datta. Research issues in databases for ARCS: Active Rapidly Changing data Systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(3):8–13, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Klein:1994:HMS

[1066] H.-J. Klein. How to modify SQL queries in order to guarantee sure answers. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(3):14–20, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yeung:1994:PEN

[1067] Chim fu Yeung, Sheung lun Hung, and Kam yiu Lam. Performance evaluation of a new distributed deadlock detection algorithm. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(3):21–26, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:1994:TT

[1068] Richard T. Snodgrass, Ilsoo Ahn, Gad Ariav, Don Batory, James Clifford, Curtis E. Dyreson, Ramez Elmasri, Fabio Grandi, Christian S. Jensen, Wolfgang Käfer, Nick Kline, Krishna Kulkarni, T. Y. Cliff Leung, Nikos Lorentzos, John F. Roddick, Arie Segev, Michael D. Soo, and Suryanarayana M. Sripada. A TSQL2 tutorial. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(3):27–33, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:1994:ATQ

[1069] Richard T. Snodgrass, Ilsoo Ahn, Gad Ariav, Don Batory, James Clifford, Curtis E. Dyreson, Ramez Elmasri, Fabio Grandi, Christian S. Jensen, Wolfgang Käfer, Nick Kline, Krishna Kulkarni, T. Y. Cliff Leung, Nikos Lorentzos, John F. Roddick, Arie Segev, Michael D. Soo, and Suryanarayana M. Sripada. Announcement — the temporal query language TSQL2 final language definition. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (3):34, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blakeley:1994:IDR

[1070] José A. Blakeley, Dan Fishman, David Lomet, and Michael Stonebraker. The impact of database research on industrial products (panel). SIGMOD Record (ACM Special Interest Group on Management of Data), 23(3):35–40, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Widom:1994:RIA

[1071] Jennifer Widom. Research issues in active database systems: Report from the closing panel at RIDE-ADS ’94. SIGMOD Record (ACM Special Interest Group on Management of Data), 23 (3):41–43, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramirez:1994:MIS

[1072] Jorge C. G. Ramirez, Lon A. Smith, and Lynn L. Peterson. Medical information systems: Characterization and challenges. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(3):44–53, September 1994.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:1994:DRN

[1073] Arbee L. P. Chen. Database research at NTHU and ITRI. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (3):54–59, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Qian:1994:JNB

[1074] Xiaolei Qian. Jumping on the NII bandwagon. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(3):61–65, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wedekind:1994:TVV

[1075] Hartmut Wedekind. Are the terms “version” and “variant” orthogonal to one another? — a critical assessment of the STEP standardization. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):3–7, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bertram:1994:DML

[1076] Martin Bertram. Data modelling in the large. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):8–12, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shin:1994:NJA

[1077] Dong Keun Shin and Arnold Charles Meltzer. A new join algorithm. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (4):13–18, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Klas:1994:MDM

[1078] W. Klas and A. Sheth. Metadata for digital media. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):19–??, ???? 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bohm:1994:MMD

[1079] Klemens Böhm and Thomas C. Rakow.

Metadata for multimedia documents. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):21–26, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bohms:1994:MMD

[1080] K. Bohms and T. C. Rakow. Metadata for multimedia documents. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (4):21–??, ???? 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jain:1994:MVD

[1081] Ramesh Jain and Arun Hampapur.

Metadata in video databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (4):27–33, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kiyoki:1994:MDS

[1082] Yasushi Kiyoki, Takashi Kitagawa, and Takanari Hayama. A meta-database system for semantic image search by a mathematical model of meaning. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (4):34–41, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anderson:1994:SMS

[1083] Jean T. Anderson and Michael Stonebraker. SEQUOIA 2000 metadata schema for satellite images. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (4):42–48, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grosky:1994:UMI

[1084] William I. Grosky, Farshad Fotouhi, Ishwar K. Sethi, and Bogdan Capatina. Using metadata for the intelligent browsing of structured media objects. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):49–56, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Glavitsch:1994:MIS

[1085] Ulrike Glavitsch, Peter Schäuble, and Martin Wechsler. Metadata for integrating speech documents in a text retrieval system. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):57–63, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:1994:MDM

[1086] Francine Chen, Marti Hearst, Julian Kupiec, Jan Pedersen, and Lynn Wilcox. Meta-data for mixed-media access. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):64–71, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barclay:1994:LDU

[1087] Tom Barclay, Robert Barnes, Jim Gray, and Prakash Sundaresan. Loading databases using dataflow parallelism. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):72–83, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mattos:1994:RDT

[1088] Nelson Mattos and Linda G. DeMichiel.

Recent design trade-offs in SQL3. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23 (4):84–90, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenblatt:1994:URN

[1089] Bill Rosenblatt. Unix RDBMS: the next generation: what are the Unix relational-database vendors doing to survive in the next generation of client/ server environments. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):91–103, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhang:1995:AOS

[1090] Jian Zhang. Application of OODB and SGML techniques in text database: An electronic dictionary system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):3–8, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dogac:1995:IAO

[1091] Asuman Dogac, Mehmet Altinel, Cetin Ozkan, and Ilker Durusoy. Implementation aspects of an object-oriented DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):9–14, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Karlapalem:1995:HAF

[1092] Kamalakar Karlapalem, Qing Li, and Chung-Dak Shum. HODFA: An architectural framework for homogenizing heterogeneous legacy databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):15–20, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hanna:1995:CLI

[1093] Magdy S. Hanna. A close look at the IFO data model. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):21–26, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). See response [1189].

Ng:1995:MLM

[1094] Shan hoi Ng and Sheung lun Hung. Multigranularity locking in multiple job classes transaction processing system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):27–32, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jannink:1995:IDB

[1095] Jan Jannink. Implementing deletion in B^{+}-trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):33–38, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Darwen:1995:TMD

[1096] Hugh Darwen and C. J. Date. The third manifesto (databases). SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):39–49, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhri:1995:ABB

[1097] Akmal B. Chaudhri. An annotated bibliography of benchmarks for object databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):50–57, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jaeger:1995:ABA

[1098] Ulrike Jaeger and Johann Christoph Freytag. An annotated bibliography on active databases (short version). SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):58–69, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berra:1995:AMC

[1099] Bruce Berra, Kingsley Nwosu, and Bhavani Thuraisingham. ACM Multimedia ’94 conference workshop on multimedia database management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):70–72, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jarke:1995:ISR

[1100] Matthias Jarke. Information systems research at RWTH Aachen. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):73–78, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schlageter:1995:DGU

[1101] Gunter Schlageter, Thomas Berkel, Eberhard Heuel, Silke Mittrach, Andreas Scherer, and Wolfgang Wilkes. The database group at University of Hagen. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):79–84, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Medsker:1995:CTU

[1102] Carl Medsker, Margaret Christensen, and Il-Yeol Song. A comparison of three user interfaces to relational microcomputer data bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):86–99, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berenson:1995:CAS

[1103] Hal Berenson, Phil Bernstein, Jim Gray, Jim Melton, Elizabeth O’Neil, and Patrick O’Neil. A critique of ANSI SQL isolation levels. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):1–10, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Molesky:1995:RPS

[1104] Lory D. Molesky and Krithi Ramamritham. Recovery protocols for shared memory database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):11–22, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adya:1995:EOC

[1105] Atul Adya, Robert Gruber, Barbara Liskov, and Umesh Maheshwari. Efficient optimistic concurrency control using loosely synchronized clocks. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24(2): 23–34, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brodsky:1995:LLQ

[1106] Alexander Brodsky and Yoram Kornatzky. The LyriC language: querying constraint objects. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):35–46, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fegaras:1995:TEC

[1107] Leonidas Fegaras and David Maier.

Towards an effective calculus for object query languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):47–58, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gardarin:1995:OFE

[1108] Georges Gardarin, Fernando Machuca, and Philippe Pucheral. OFL: a functional execution model for object query languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):59–70, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roussopoulos:1995:NNQ

[1109] Nick Roussopoulos, Stephen Kelley, and Frédéric Vincent. Nearest neighbor queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):71–79, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Freeston:1995:GSD

[1110] Michael Freeston. A general solution of the n-dimensional B-tree problem. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):80–91, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Papadias:1995:TRW

[1111] Dimitris Papadias, Timos Sellis, Yannis Theodoridis, and Max J. Egenhofer. Topological relations in the world of minimum bounding rectangles: a study with R-trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):92–103, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shatdal:1995:APA

[1112] Ambuj Shatdal and Jeffrey F. Naughton. Adaptive parallel aggregation algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):104–114, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wilschut:1995:PEM

[1113] Annita N. Wilschut, Jan Flokstra, and Peter M. G. Apers. Parallel evaluation of multi-join queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):115–126, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hernandez:1995:MPP

[1114] Mauricio A. Hernández and Salvatore J. Stolfo. The merge/purge problem for large databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (2):127–138, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramaswamy:1995:OIC

[1115] Sridhar Ramaswamy and Paris C.

Kanellakis. OODB indexing by classdivision. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):139–150, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aref:1995:HTI

[1116] Walid Aref, Daniel Barbará, and Padmavathi Vallabhaneni. The handwritten trie: indexing electronic ink. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):151–162, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Faloutsos:1995:FFA

[1117] Christos Faloutsos and King-Ip Lin. FastMap: a fast algorithm for indexing, data-mining and visualization of traditional and multimedia datasets. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):163–174, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Park:1995:EHB

[1118] Jong Soo Park, Ming-Syan Chen, and Philip S. Yu. An effective hash-based algorithm for mining association rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):175–186, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

White:1995:ICR

[1119] Seth J. White and David J. De-Witt. Implementing crash recovery in QuickStore: a performance study. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):187–198, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Acharya:1995:BDD

[1120] Swarup Acharya, Rafael Alonso, Michael Franklin, and Stanley Zdonik. Broadcast disks: data management for asymmetric communication environments. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):199–210, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gupta:1995:AMV

[1121] Ashish Gupta, Inderpal S. Mumick, and Kenneth A. Ross. Adapting materialized views after redefinitions. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):211–222, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hou:1995:EDC

[1122] Wen-Chi Hou and Zhongyang Zhang.

Enhancing database correctness: a statistical approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):223–232, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1995:BHO

[1123] Yannis E. Ioannidis and Viswanath Poosala. Balancing histogram optimality and practicality for query result size estimation. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):233–244, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adelberg:1995:AUS

[1124] B. Adelberg, H. García-Molina, and B. Kao. Applying update streams in a soft real-time database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):245–256, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bettini:1995:SAQ

[1125] Claudio Bettini, X. Sean Wang, Elisa Bertino, and Sushil Jajodia. Semantic assumptions and query evaluation in temporal databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):257–268, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sistla:1995:TCI

[1126] A. Prasad Sistla and Ouri Wolfson. Temporal conditions and integrity constraints in active database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):269–280, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Davison:1995:DRB

[1127] Diane L. Davison and Goetz Graefe. Dynamic resource brokering for multiuser query execution. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):281–292, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Du:1995:RMQ

[1128] Weimin Du, Ming-Chien Shan, and Umeshwar Dayal. Reducing multidatabase query response time by tree balancing. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):293–303, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bhargava:1995:HBR

[1129] Gautam Bhargava, Piyush Goel, and Bala Iyer. Hypergraph based reorderings of outer join queries with complex predicates. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):304–315, May 1995.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhuge:1995:VMW

[1130] Yue Zhuge, Héctor García-Molina, Joachim Hammer, and Jennifer Widom. View maintenance in a warehousing environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):316–327, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Griffin:1995:IMV

[1131] Timothy Griffin and Leonid Libkin. Incremental maintenance of views with duplicates. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):328–339, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lu:1995:EMM

[1132] James J. Lu, Guido Moerkotte, Joachim Schue, and V. S. Subrahmanian. Efficient maintenance of materialized mediated views. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):340–351, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Freedman:1995:SSV

[1133] Craig S. Freedman and David J. De-Witt. The SPIFFI scalable videoon-demand system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):352–363, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berson:1995:FTD

[1134] Steven Berson, Leana Golubchik, and Richard R. Muntz. Fault tolerant design of multimedia servers. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (2):364–375, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dan:1995:OVP

[1135] Asit Dan and Dinkar Sitaram. An online video placement policy based on bandwidth to space ratio (BSR). SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):376–385, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:1995:DIF

[1136] Serge Abiteboul, Sophie Cluet, and Tova Milo. A database interface for file update. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):386–397, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brin:1995:CDM

[1137] Sergey Brin, James Davis, and Héctor García-Molina. Copy detection mechanisms for digital documents. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (2):398–409, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:1995:JQE

[1138] Surajit Chaudhuri, Umeshwar Dayal, and Tak W. Yan. Join queries with external text sources: execution and optimization techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):410–422, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

FDTM:1995:OOR

[1139] Fox Development Team Microsoft.

Object-oriented, rapid application development in a PC database environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):423–424, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

ATM:1995:UFF

[1140] The Access Team Microsoft. Upsizing form file server to client server architectures. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):425–426, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Moore:1995:LNS

[1141] Kenneth Moore. The Lotus Notes storage system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):427–428, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gettys:1995:DOR

[1142] William L. Gettys. DIRECTV and Oracle Rdb: the challenge of VLDB transaction processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):429–430, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hope:1995:ETP

[1143] Greg Hope. Enterprise transaction processing on Windows NT. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (2):431–432, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ivinskis:1995:HAC

[1144] Kestutis Ivinskis. High availability of commercial applications. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (2):433–434, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buneman:1995:DYW

[1145] Peter Buneman and David Maier. The data that you won’t find in databases: tutorial panel on data exchange formats. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):435, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gray:1995:PDS

[1146] Jim Gray. Parallel database systems 101. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):436, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ellison:1995:KA

[1147] Larry J. Ellison. Keynote address. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):437, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Epstein:1995:KA

[1148] Robert S. Epstein. Keynote address. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):438, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goldring:1995:TEU

[1149] Rob Goldring. Things every update replication customer should know (abstract). SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):439–440, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shyy:1995:VRS

[1150] Yuh-Ming Shyy, H. Stephen Au-Yeung, and C. P. Chou. VERSANT replication: supporting fault-tolerant object databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):441–442, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Woelk:1995:CID

[1151] D. Woelk, B. Bohrer, N. Jacobs, K. Ong, C. Tomlinson, and C. Unnikrishnan. Carnot and InfoSleuth: database technology and the World Wide Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):443–444, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Edelstein:1995:RPT

[1152] Herb Edelstein. Research and products — are they relevant to each other? (panel session). SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):445, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Squire:1995:DET

[1153] Case Squire. Data extraction and transformation for the data warehouse. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):446–447, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bansal:1995:RWR

[1154] Sanju K. Bansal. Real world requirements for decision support — implications for RDBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):448, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

French:1995:OSF

[1155] Clark D. French. “one size fits all” database architectures do not work for DSS. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):449–450, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Perna:1995:LIA

[1156] Janet Perna. Leveraging the information asset. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):451–452, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Olson:1995:CYA

[1157] Michael A. Olson. Cover your assets. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):453, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Atkinson:1995:UCA

[1158] Robert Atkinson. Use of a component architecture in integrating relational and non-relational. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):454, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kleissner:1995:EOF

[1159] Charly Kleissner. Enterprise objects framework. A second generation object-relational enabler. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (2):455–459, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barn:1995:ODP

[1160] C. Barn, G. Fecteau, A. Goyal, H. i. Hsiao, A. Jhingran, S. Padmanabhan, and W. Wilson. An overview of DB2 parallel edition. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):460–462, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Baru:1995:ODP

[1161] Chaitanya Baru and Gilles Fecteau.

An overview of DB2 parallel edition. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):460–462, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gerber:1995:IOX

[1162] Bob Gerber. Informix online XPS. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):463, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Levine:1995:OMA

[1163] Charles Levine. Order-of-magnitude advantage on TPC-C through massive parallelism. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):464–465, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deutsch:1995:OSP

[1164] Donald R. Deutsch. Objects and SQL (panel session): strange relations? SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):466, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Faloutsos:1995:IMD

[1165] Christos Faloutsos. Indexing multimedia databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):467, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Melton:1995:OET

[1166] Jim Melton and Nelson Mendoca Mattos. An overview of the emerging thirdgeneration SQL standard. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):468, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sheth:1995:WAA

[1167] Amit Sheth. Workflow automation:

applications, technology and research. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):469, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Norrie:1995:IAC

[1168] Moira C. Norrie. Integration approaches for CIM. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):470, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gibson:1995:TST

[1169] Garth A. Gibson. Tutorial on storage technology: RAID and beyond. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):471, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Imielinski:1995:DIR

[1170] T. Imieliński and A. Virmani. DataMine — interactive rule discovery system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):472, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ceri:1995:ATC

[1171] Stefano Ceri, Piero Fraternali, Stefano Paraboschi, and Giuseppe Psaila. The Algres testbed of CHIMERA: an active object-oriented database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):473, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Collet:1995:NS

[1172] C. Collet and T. Coupaye. The NAOS system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):474, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ashley:1995:QIC

[1173] Jonathan Ashley, Myron Flickner, James Hafner, Denis Lee, Wayne Niblack, and Dragutin Petkovic. The query by image content (QBIC) system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):475, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buchmann:1995:RAO

[1174] A. P. Buchmann, A. Deutsch, J. Zimmermann, and M. Higa. The REACH active OODBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):476, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Massari:1995:QQI

[1175] Antonio Massari, Stefano Pavani, Lorenzo Saladini, and Panos K. Chrysanthis. QBI: query by icons. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (2):477, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shklar:1995:ISS

[1176] Leon Shklar, Amit Sheth, Vipul Kashyap, and Satish Thatte. Info-Harness: a system for search and retrieval of heterogeneous information. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):478, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thomas:1995:DIA

[1177] J. Thomas, S. Deßloch, and N. Mattos. Design and implementation of advanced knowledge processing in the KBMS KRISYS. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):479, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gatziu:1995:SAD

[1178] Stella Gatziu, Andreas Geppert, and Klaus R. Dittrich. The SAMOS active DBMS prototype. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):480, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Radeke:1995:EFD

[1179] E. Radeke, R. Böttger, B. Burkert, Y. Engel, G. Kachel, S. Kolmschlag, and D. Nolte. Efendi: federated database system of Cadlab. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):481, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keim:1995:VSV

[1180] Daniel A. Keim and Hans-Peter Kriegel. VisDB: a system for visualizing large databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):482, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hammer:1995:ITM

[1181] Joachim Hammer, Héctor García-Molina, Kelly Ireland, Yannis Papakonstantinou, Jeffrey Ullman, and Jennifer Widom. Information translation, mediation, and mosaic-based browsing in the TSIMMIS system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):483, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:1995:SSP

[1182] Wen-Syan Li and Chris Clifton. Semint: a system prototype for semantic integration in heterogeneous databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):484, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

ParadiseTeam:1995:PDS

[1183] The Paradise Team. Paradise: a database system for GIS applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):485, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

SHORETeam:1995:SCB

[1184] The SHORE Team. Shore: combining the best features of OODBMS and file systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):486, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:1995:PMP

[1185] Jason T. L. Wang, Kaizhong Zhang, and Dennis Shasha. Pattern matching and pattern discovery in scientific, program, and document databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (2):487, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grossman:1995:PLW

[1186] R. L. Grossman, D. Hanley, and X. Qin. PTool: a light weight persistent object manager. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):488, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dreyer:1995:UCT

[1187] Werner Dreyer, Angelika Kotz Dittrich, and Duri Schmidt. Using the CA-LANDA time series management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):489, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jonker:1995:EMD

[1188] Willem Jonker and Heribert Schütz.

The ECRC multi database system.

SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):490, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:1995:RCL

[1189] Serge Abiteboul and Richard Hull. Response to “A close look at the IFO data model”. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(3):4, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). See [1093].

Maelbrancke:1995:OJJ

[1190] R. Maelbrancke and H. Olivié. Optimizing Jan Jannink’s implementation of B^{+}-tree deletion. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(3):5–7, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

ONeil:1995:MTJ

[1191] Patrick O’Neil and Goetz Graefe. Multi-table joins through bitmapped join indices. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(3):8–11, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Doan:1995:DUT

[1192] Dac Khoa Doan, Norman W. Paton, and Alistair Kilgour. Design and user testing of a multi-paradigm query interface to an object-oriented database. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (3):12–17, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fong:1995:MEE

[1193] Joseph Fong. Mapping extended entity relationship model to object modeling technique. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(3):18–22, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:1995:NOD

[1194] Byung S. Lee. Normalization in OODB design. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(3):23–27, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:1995:AQO

[1195] Chiang Lee, Chia-Jung Chen, and Hongjun Lu. An aspect of query optimization in multidatabase systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (3):28–33, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wong:1995:IRF

[1196] Limsoon Wong. An introduction to Remy’s fast polymorphic record projection. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(3):34–39, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kokkotos:1995:IVT

[1197] Stavros Kokkotos, Efstathios V. Ioannidis, Themis Panayiotopoulos, and Constantine D. Spyropoulos. On the issue of valid time(s) in temporal databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(3):40–43, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pitoura:1995:FPC

[1198] Evaggelia Pitoura and Bharat Bhargava. A framework for providing consistent and recoverable agent-based access to heterogeneous mobile databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (3):44–49, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gamache:1995:ATU

[1199] André Gamache and Nadjiba Sahraoui. Addressing techniques used in database object managers O_{2} and ORION. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (3):50–55, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dogac:1995:MID

[1200] A. Dogac, C. Dengi, E. Kilic, G. Ozhan, F. Ozcan, S. Nural, C. Evrendilek, U. Halici, B. Arpinar, P. Koksal, N. Kesim, and S. Mancuhan. METU interoperable database system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (3):56–61, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yan:1995:IFD

[1201] Tak W. Yan and Héctor García-Molina. Information finding in a digital library: the Stanford perspective. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (3):62–70, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rishe:1995:FIU

[1202] Naphtali Rishe, Wei Sun, David Barton, Yi Deng, Cyril Orji, Michael Alexopoulos, Leonardo Loureiro, Carlos Ordonez, Mario Sanchez, and Artyom Shaposhnikov. Florida International University High Performance Database Research Center. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (3):71–76, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenthal:1995:DMR

[1203] Arnon Rosenthal, Len Seligman, Catherine McCollum, Barbara Blaustein, Bhavani Thuraisingham, and Edward Lafferty. Data management research at the MITRE corporation. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (3):77–82, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Qian:1995:TNN

[1204] Xiaolei Qian. Turmoil at NASA, and numerous funding announcements. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (3):83–90, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kimball:1995:WDS

[1205] Ralph Kimball and Kevin Strehlo. Why decision support fails and how to fix it. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (3):92–97, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Richey:1995:CHS

[1206] Jeff Richey. Condition handling in SQL persistent stored modules. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (3):98–103, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1995:SOC

[1207] Anonymous. SIGMOD officers, committees, awards, and mission statement. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):1–??, ???? 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dunham:1995:MCD

[1208] Margaret H. Dunham and Abdelsalam Helal. Mobile computing and databases: Anything new? SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (4):5–9, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Noble:1995:RSR

[1209] Brian D. Noble and M. Satyanarayanan. A research status report on adaptation for mobile data access.

SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):10–15, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Elmagarmid:1995:WCS

[1210] Ahmed Elmagarmid, Jin Jing, and Tetsuya Furukawa. Wireless client/ server computing for personal information services and applications. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (4):16–21, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wolfson:1995:VMM

[1211] Ouri Wolfson, Prasad Sistla, Son Dao, Kailash Narayanan, and Ramya Raj. View maintenance in mobile computing. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):22–27, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Alonso:1995:MVD

[1212] Rafael Alonso, Yuh-Lin Chang, Liviu Iftode, and V. S. Mani. Managing video data in a mobile environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):28–33, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bhargava:1995:DLS

[1213] Bharat Bhargava, Melliyal Annamalai, and Evaggelia Pitoura. Digital library services in mobile computing. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24 (4):34–39, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ulusoy:1995:ABR

[1214] Özgür Ulusoy. An annotated bibliography on real-time database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):40–45, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Segev:1995:RIW

[1215] Arie Segev, Christian S. Jensen, and Richard T. Snodgrass. Report on the 1995 International Workshop on Temporal Databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):46–52, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bohlen:1995:TDS

[1216] Michael H. Böhlen. Temporal database system implementations. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):53–60, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Englert:1995:PPC

[1217] Susanne Englert, Ray Glasstone, and Waqar Hasan. Parallelism and its price: a case study of NonStop SQL/MP. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):61–71, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Venkatrao:1995:SCN

[1218] Murali Venkatrao and Michael Pizzo.

SQL/CLI — a new binding style for SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):72–77, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

UAthens:1995:DGN

[1219] National Technical Univ. of Athens.

The database group at National Technical University of Athens (NTUA). SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):78–83, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1995:DGN

[1220] Anonymous. The database group at National Technical University of Athens (NTUA). SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):78–??, ???? 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Drew:1995:DKB

[1221] P. Drew, B. Hamidzadeh, K. Karlapalem, A. Kean, D. Lee, Q. Li, F. Lochovsky, C. D. Shum, and B. Wuthrich. Data and knowledge base research at Hong Kong University of Science and Technology. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):84–89, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Qian:1995:OAN

[1222] Xiaolei Qian. Opportunities at ARPA, NSF, and elsewhere. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):90–93, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Alonso:1995:EI

[1223] R. Alonso. Editor’s introduction. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):94–??, ???? 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stacey:1995:RDO

[1224] Doug Stacey. Replication: DB2, Oracle, or Sybase? SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):95–101, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Finkelstein:1995:NM

[1225] Rich Finkelstein. The new middleware. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):102–106, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1996:SOCa

[1226] Anonymous. SIGMOD officers, committees awards, and mission statement. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):1–??, ???? 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bestavros:1996:ART

[1227] Azer Bestavros. Advances in real-time database systems research. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):3–7, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramamritham:1996:ITR

[1228] Krithi Ramamritham, Raju Sivasankaran, John A. Stankovic, Don T. Towsley, and Ming Xiong. Integrating temporal, real-time, and active databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):8–12, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Haritsa:1996:RTI

[1229] Jayant R. Haritsa and S. Seshadri.

Real-time index concurrency control. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):13–17, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Harista:1996:RTI

[1230] J. R. Harista and S. Seshadri. Realtime index concurrency control. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (1):13–??, ???? 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuo:1996:RTD

[1231] Tei-Wei Kuo and Aloysius K. Mok.

Real-time database — similarity semantics and resource scheduling. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (1):18–22, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ulusoy:1996:EMM

[1232] Özgür Ulusoy and Alejandro Buchmann. Exploiting main-memory DBMS features to improve real-time concurrency control protocols. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25(1): 23–25, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lin:1996:EEC

[1233] Kwei-Jay Lin and Ching-Shan Peng. Enhancing external consistency in realtime transactions. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):26–28, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Son:1996:ITR

[1234] Sang H. Son, Rasikan David, and Bhavani Thuraisingham. Improving timeliness in real-time secure database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):29–33, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adelberg:1996:OSR

[1235] Brad Adelberg, Ben Kao, and Hector Garcia-Molina. Overview of the STanford Real-Time Information Processor (STRIP). SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):34–37, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Andler:1996:DTD

[1236] S. F. Andler, J. Hansson, J. Eriksson, J. Mellin, M. Berndtsson, and B. Eftring. DeeDS towards a distributed and active real-time database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):38–51, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tsotras:1996:TDB

[1237] V. J. Tsotras and A. Kumar. Temporal database bibliography update. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):41–??, ???? 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Silberschatz:1996:DRA

[1238] Avi Silberschatz, Mike Stonebraker, and Jeff Ullman. Database research: Achievements and opportunities into the 21st Century. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):52–63, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berndtsson:1996:WRF

[1239] Mikael Berndtsson and Jörgen Hansson. Workshop report: First International Workshop on Active and Real-Time Database Systems (ARTDB-95). SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):64–66, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Baeza-Yates:1996:ICS

[1240] Ricardo Baeza-Yates and Gonzalo Navarro. Integrating contents and structure in text retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):67–79, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Freeman:1996:LSM

[1241] Eric Freeman and David Gelernter. Lifestreams: a storage model for personal data. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):80–86, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wade:1996:OQS

[1242] Andrew E. Wade. Object query standards. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):87–92, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Phatak:1996:DRI

[1243] D. B. Phatak, N. L. Sarda, S. Seshadri, and S. Sudarshan. Database research at the Indian Institute of Technology, Bombay. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):93–98, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Urban:1996:DRA

[1244] Susan D. Urban, Suzanne W. Dietrich, and Forouzan Golshani. Database research at Arizona State University.

SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):99–104, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gaffney:1996:IWD

[1245] John Gaffney. Illustra’s Web DataBlade module: An Illustra technical white paper. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):105–112, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Corp:1996:MCL

[1246] Lotus Development Corp. MQSeries and CICS link for Lotus Notes. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (1):113–118, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Qian:1996:SBF

[1247] Xiaolei Qian. Shutdown, budget, and funding. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(1):119–122, March 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Srikant:1996:MQA

[1248] Ramakrishnan Srikant and Rakesh Agrawal. Mining quantitative association rules in large relational tables. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):1–12, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fukuda:1996:DMU

[1249] Takeshi Fukuda, Yasuhiko Morimoto, Shinichi Morishita, and Takeshi Toluyama. Data mining using twodimensional optimized association rules: scheme, algorithms, and visualization. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):13–23, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Selfridge:1996:IID

[1250] Peter G. Selfridge, Divesh Srivastava, and Lynn O. Wilson. IDEA: Interactive data exploration and analysis. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):24–34, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vance:1996:RBJ

[1251] Bennet Vance and David Maier. Rapid bushy join-order optimization with Cartesian products. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):35–46, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goel:1996:SQO

[1252] Piyush Goel and Bala Iyer. SQL query optimization: Reordering for a general class of queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):47–56, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Simmen:1996:FTO

[1253] David E. Simmen, Eugene J. Shekita, and Timothy Malkemus. Fundamental techniques for order optimization. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):57–67, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

OConnell:1996:TCB

[1254] William O’Connell, Ion T. Ieong, David Schrader, Alexandros Biliris, Susan Choo, Pierre Colin, Glenn Linderman, Euthimios Panagos, June Wang, and Todd Walter. A Teradata contentbased multimedia object manager for massively parallel architectures. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25(2): 68–78, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ozden:1996:FTA

[1255] Banu Özden, Rajeev Rastogi, Prashant Shenoy, and Avi Silberschatz. Faulttolerant architectures for continuous media servers. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):79–90, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:1996:OQM

[1256] Surajit Chaudhuri and Luis Gravano.

Optimizing queries over multimedia repositories. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):91–102, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhang:1996:BED

[1257] Tian Zhang, Raghu Ramakrishnan, and Miron Livny. BIRCH: An efficient data clustering method for very large databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):103–114, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zou:1996:LRS

[1258] Chendong Zou and Betty Salzberg. On-line reorganization of sparelypopulated B^{+}-trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):115–124, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Achyutuni:1996:TTL

[1259] Kiran J. Achyutuni, Edward Omiecinski, and Shamkant B. Navathe. Two techniques for on-line index modification in shared nothing parallel databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):125–136, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adali:1996:QCO

[1260] S. Adali, K. S. Candan, Y. Papakonstantinou, and V. S. Subrahmanian. Query caching and optimization in distributed mediator systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (2):137–146, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adah:1996:QCO

[1261] Sibel Adah, Kasim Selcuk Candan, Yannis Papkonstantinou, and V. S. Subrahmanian. Query caching and optimization in distributed mediator systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):137–??, ???? 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Franklin:1996:PTC

[1262] Michael J. Franklin, Björn Thór Jónsson, and Donald Kossmann. Performance tradeoffs for client-server query processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):149–160, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blakeley:1996:DAM

[1263] José A. Blakeley. Data access for the masses through OLE DB. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (2):161–172, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gray:1996:DRS

[1264] Jim Gray, Pat Helland, Patrick O’Neil, and Dennis Shasha. The dangers of replication and a solution. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (2):173–182, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mogi:1996:HMM

[1265] Kazuhiko Mogi and Masaru Kitsuregawa. Hot mirroring: a method of hiding parity update penalty and degradation during rebuilds for RAID5. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):183–194, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hillyer:1996:RSO

[1266] Bruce K. Hillyer and Avi Silberschatz. Random I/O scheduling in online tertiary storage systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):195–204, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Harinarayan:1996:IDC

[1267] Venky Harinarayan, Anand Rajaraman, and Jeffrey D. Ullman. Implementing data cubes efficiently. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (2):205–216, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rao:1996:PBS

[1268] Sudhir G. Rao, Antonio Badia, and Dirk van Gucht. Providing better support for a class of decision support queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):217–227, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Libkin:1996:QLM

[1269] Leonid Libkin, Rona Machlin, and Limsoon Wong. A query language for multidimensional arrays: Design, implementation, and optimization techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):228–239, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agarwal:1996:SSS

[1270] Ramesh C. Agarwal. A super scalar sort algorithm for RISC processors. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):240–246, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lo:1996:SHJ

[1271] Ming-Ling Lo and Chinya V. Ravishankar. Spatial hash-joins. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (2):247–258, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Patel:1996:PBS

[1272] Jignesh M. Patel and David J. De-Witt. Partition based spatial-merge join. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):259–270, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ganguly:1996:BSS

[1273] Sumit Ganguly, Phillip B. Gibbons, Yossi Matias, and Avi Silberschatz. Bifocal sampling for skew-resistant join size estimation. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):271–281, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Krishnan:1996:EAS

[1274] P. Krishnan, Jeffrey Scott Vitter, and Bala Iyer. Estimating alphanumeric selectivity in the presence of wildcards.

SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):282–293, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Poosala:1996:IHS

[1275] Viswanath Poosala, Yannis E. Ioannidis, Peter J. Haas, and Eugenge J. Shekita. Improved histograms for selectivity estimation of range predicates. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):294–305, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Doherty:1996:SMP

[1276] Michael Doherty, Richard Hull, and Mohammed Rupawalla. Structures for manipulating proposed updates in object-oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):306– 317, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Liskov:1996:SES

[1277] Barbara Liskov, Atud Adya, Miguel Castro, Mark Day, Sanjay Ghemawat, Robert Gruber, Umesh Maheshwari, Andrew J. Myers, and Liuba Shrira. Safe and efficient sharing of persistent objects in Thor. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):318–329, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blott:1996:OAO

[1278] Stephen Blott, Lukas Relly, and Hans-Jörg Schek. An open abstract-object storage system. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):330–340, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tajima:1996:SDS

[1279] Keishi Tajima. Static detection of security flaws in object-oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):341–352, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brown:1996:GOB

[1280] Kurt P. Brown, Michael J. Carey, and Miron Livny. Goal-oriented buffer management revisited. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):353–364, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Garofalakis:1996:MDR

[1281] Minos N. Garofalakis and Yannis E.

Ioannidis. Multi-dimensional resource scheduling for parallel queries. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (2):365–376, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cook:1996:SAS

[1282] Jonathan E. Cook, Artur W. Klauser, Alexander L. Wolf, and Benjamin G. Zorn. Semi-automatic, self-adaptive control of garbage collection rates in object databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):377–388, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McAuliffe:1996:TEE

[1283] Mark L. McAuliffe, Michael J. Carey, and Marvin H. Solomon. Towards effective and efficient free space management. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):389–400, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cherniack:1996:RLI

[1284] Mitch Cherniack and Stanley B. Zdonik. Rule languages and internal algebras for rule-based optimizers. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):401–412, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christophides:1996:EQG

[1285] Vassilis Christophides, Sophie Cluet, and Guido Moerkotte. Evaluating queries with generalized path expressions. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):413–422, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hellerstein:1996:QET

[1286] Joseph M. Hellerstein and Jeffrey F. Naughton. Query execution techniques for caching expensive methods. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):423–434, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Seshadri:1996:CBO

[1287] Praveen Seshadri, Joseph M. Hellerstein, Hamid Pirahesh, T. Y. Cliff Leung, Raghu Ramakrishnan, Divesh Srivastava, Peter J. Stuckey, and S. Sudarshan. Cost-based optimization for magic: Algebra and implementation. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):435–446, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:1996:MVM

[1288] Kenneth A. Ross, Divesh Srivastava, and S. Sudarshan. Materialized view maintenance and integrity constraint checking: Trading space for time. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):447–458, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Morpain:1996:MDC

[1289] Claire Morpain, Michéle Cart, Jean Ferrié, and Jean-François Pons. Maintaining database consistency in presence of value dependencies in multidatabase systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):459–468, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Colby:1996:ADV

[1290] Latha S. Colby, Timothy Griffin, Leonid Libkin, Inderpal Singh Mumick, and Howard Trickey. Algorithms for deferred view maintenance. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):469–480, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hull:1996:FSD

[1291] Richard Hull and Gang Zhou. A framework for supporting data integration using the materialized and virtual approaches. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):481–492, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chawathe:1996:CDH

[1292] Sudarshan S. Chawathe, Anand Rajaraman, Hector Garcia-Molina, and Jennifer Widom. Change detection in hierarchically structured information. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):493–504, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buneman:1996:QLO

[1293] Peter Buneman, Susan Davidson, Gerd Hillebrand, and Dan Suciu. A query language and optimization techniques for unstructured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):505– 516, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goyal:1996:GPD

[1294] Nita Goyal, Charles Hoch, Ravi Krishnamurthy, Brian Meckler, and Michael Suckow. Is GUI programming a database research problem? SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (2):517–528, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nguyen:1996:ARD

[1295] Tam Nguyen and V. Srinivasan. Accessing relational databases from the World Wide Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):529–540, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fernandez:1996:IOE

[1296] Phil Fernandez and Donovan Schneider. The ins and outs (and everything in between) of data warehousing. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):541, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schneider:1996:IOE

[1297] Donovan Schneider and Phil Fernandez. The ins and outs (and everything in between) of data warehousing. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):541–??, ???? 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:1996:RSE

[1298] Philip A. Bernstein. Repository system engineering. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):542, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keim:1996:DV

[1299] Daniel A. Keim. Databases and visualization. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):543, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mohan:1996:SAW

[1300] C. Mohan. State of the art in workflow management research and products. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):544, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:1996:DMT

[1301] Jiawei Han. Data mining techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 25 (2):545, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Piatko:1996:TTT

[1302] Peter Piatko, Roman Yangarber, Daoi Lin, and Dennis Shasha. Thinksheet: a tool for tailoring complex documents. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):546, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bohm:1996:HAS

[1303] Klemens Böhm and Karl Aberer. HyperStorM — administering structured documents using object-oriented database technology. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):547, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lefler:1996:DST

[1304] Mike Lefler, Mark Stokrp, and Craig Wong. DBSim: a simulation tool for predicting database performance. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25(2): 548, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Quass:1996:LLO

[1305] Dallan Quass, Jennifer Widom, Roy Goldman, Kevin Haas, Qingshan Luo, Jason McHugh, Svetlozar Nestorov, Anand Rajaraman, Hugo Rivero, Serge Abiteboul, Jeffrey D. Ullman, and Janet Wiener. LORE: a lightweight object REpository for semistructured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):549, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:1996:DIM

[1306] Jaiwei Han, Youngjian Fu, Wei Wang, Jenny Chiang, Osmar R. Zaïane, and Krzysztof Koperski. DBMiner: Interactive mining of multiple-level knowledge in relational databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):550, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Choo:1996:PCB

[1307] S. Choo, W. O’Connell, G. Linerman, H. Chen, K. Ganapathy, A. Biliris, E. Panagos, and D. Schrader. Prospector: a content-based multimedia server for massively parallel architectures. SIGMOD Record (ACM Special Interest Group on Management of Data), 25 (2):551, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Au:1996:PCB

[1308] Grace Au, Alexandros Biliris, Susan Choo, Kicha Ganapathy, Glenn Linderman, William O’Connell, Euthimios Panagos, and David Schrader. Prospector: a content-based multimedia server for massively parallel architectures. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):551–??, ???? 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dogac:1996:MID

[1309] Asuman Dogac, Ugur Halici, Ebru Kilic, Gokhan Ozhan, Fatma Ozcan, Sena Nural, Cevdet Dengi, Sema Mancuhan, Budak Arpinar, Pinar Koksal, and Cem Evrendilek. METU interoperable database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):552, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fukuda:1996:SSO

[1310] Takeshi Fukuda, Yasuhiko Morimoto, Shinichi Morishita, and Takeshi Tokuyama. SONAR: System for optimized numeric association rules. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (2):553, June 1996. CODEN SRECD8.

ISSN 0163-5808 (print), 1943-5835 (electronic).

Hung:1996:CAC

[1311] Patrick C. K. Hung, Helen P. Yeung, and Kamalakar Karlapalem. CapBasED-AMS: a capability based and event-driven activity management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):554, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rundensteiner:1996:MPO

[1312] Elke Rundensteiner, Harumi A. Kuno, Young-Gook Ra, Viviane Crestana-Taube, Matthew Jones, and Pedro Marron. The MultiView project: object-oriented view technology and applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):555, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Biliris:1996:BSS

[1313] Alex Biliris, Thomas Funkhouser, William O’Connell, and Euthimios Panagos. BeSS: Storage support for interactive visualization systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (2):556, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

TorkRoth:1996:GP

[1314] Mary Tork Roth, Manish Arya, Laura M. Haas, Michael J. Carey, William Cody, Ron Fagin, Peter M. Schwarz, John Thomas, and Edward L.

Wimmers. The Garlic project. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (2):557, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Camps:1996:DRR

[1315] R. Camps. Domains, relations and religious wars. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):3–9, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). See replies [1331, 1332].

Zobel:1996:GPC

[1316] Justin Zobel, Alistair Moffat, and Kotagiri Ramamohanarao. Guidelines for presentation and comparison of indexing techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):10–15, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Norman:1996:MAA

[1317] Michael G. Norman, Thomas Zurek, and Peter Thanisch. Much ado about shared-nothing. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):16–21, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gerlhof:1996:CMR

[1318] Carsten A. Gerlhof, Alfons Kemper, and Guido Moerkotte. On the cost of monitoring and reorganization of object bases for clustering. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (3):22–27, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hasan:1996:OIP

[1319] Waqar Hasan, Daniela Florescu, and Patrick Valduriez. Open issues in parallel query optimization. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (3):28–33, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brunie:1996:CSC

[1320] Lionel Brunie and Harald Kosch. Control strategies for complex relational query processing in shared nothing systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):34–39, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

ANC:1996:ADM

[1321] Act-Net Consortium. The active database management system manifesto: a rulebase of ADBMS features. SIGMOD Record (ACM Special Interest Group on Management of Data), 25 (3):40–49, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1996:ADM

[1322] Anonymous. The active database management system manifesto: a rulebase of ADBMS features the ACT-NET consortium. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):40–??, ???? 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bestavros:1996:RFI

[1323] Azer Bestavros, Kwei-Jay Lin, and Sang Son. Report on First International Workshop on Real-Time Database Systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):50–52, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Biliris:1996:BOS

[1324] Alexandros Biliris and Euthimios Panagos. The BeSS object storage manager: Architecture overview. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (3):53–58, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dogac:1996:MET

[1325] Asuman Dogac. Middle East Technical University Software Research and Development Center. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):59–63, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Colliat:1996:ORM

[1326] George Colliat. OLAP, relational, and multidimensional database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):64–69, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DAndrea:1996:UNG

[1327] Albert D’Andrea and Phil Janus. UniSQL’s next-generation objectrelational database management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):70–76, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Andrea:1996:UNG

[1328] A. D. Andrea and P. Janus. UniSQL’s next-generation object-relational databse management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):70–??, ???? 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Qian:1996:SCU

[1329] Xiaolei Qian. Scientists called upon to take actions. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):77–80, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1996:SOCb

[1330] Anonymous. SIGMOD officers, committees, awards, and institutional sponsors. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):1–??, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Date:1996:TCN

[1331] C. J. Date. Technical correspondence and notes — a response to R. Camps’ article “Domains, relations and religious wars”. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):3–5, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). See [1315].

Darwen:1996:RDR

[1332] Hugh Darwen. In reply to “Domains, relations and religious wars”. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (4):6–7, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). See [1315].

Rafanelli:1996:ADP

[1333] M. Rafanelli, A. Bezenchek, and L. Tininini. The aggregate data problem: a system for their definition and management. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):8–13, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Catarci:1996:IV

[1334] Tiziana Catarci and Isabel F. Cruz. Information visualization. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):14–15, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Catacrci:1996:GEF

[1335] T. Catacrci and I. F. Cruz. Guest Editor’s foreword. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):14–??, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1996:DIV

[1336] Yannis E. Ioannidis. Dynamic information visualization. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):16–20, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tanin:1996:IDS

[1337] Egemen Tanin, Richard Beigel, and Ben Shneiderman. Incremental data structures and algorithms for dynamic query interfaces. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):21–24, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ahlberg:1996:SIE

[1338] Christopher Ahlberg. Spotfire: an information exploration environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):25–29, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kennedy:1996:FIV

[1339] Jessie B. Kennedy, Kenneth J.

Mitchell, and Peter J. Barclay. A framework for information visualisation. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):30–34, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keim:1996:POD

[1340] Daniel A. Keim. Pixel-oriented database visualizations. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):35–39, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Santucci:1996:TTH

[1341] Giuseppe Santucci and Laura Tarantino. To table or not to table: a hypertabular answer. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):40–44, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hasan:1996:ADV

[1342] Masum Z. Hasan, Alberto O. Mendelzon, and Dimitra Vista. Applying database visualization to the World Wide Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):45–49, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cox:1996:GND

[1343] Kenneth C. Cox, Stephen G. Eick, and Taosong He. 3D geographic network displays. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):50–54, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cos:1996:GND

[1344] K. C. Cos, S. G. Eick, and T. He. 3D geographic network displays. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (4):50–??, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sheth:1996:RNW

[1345] Amit Sheth, Dimitrios Georgakopoulos, Stef M. M. Joosten, Marek Rusinkiewicz, Walt Scacchi, Jack Wileden, and Alexander L. Wolf. Report from the NSF workshop on workflow and process automation in information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):55–67, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Seth:1996:RNW

[1346] A. Seth, D. Georgakopoulos, S. M. M.

Joosten, M. Rusinkiewicz, W. Scacchi, J. Wileden, and A. Wolf. Reports from the NSF workshop on workflow and process automation in information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):55–??, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Atkinson:1996:OPJ

[1347] M. P. Atkinson, L. Daynès, M. J. Jordan, T. Printezis, and S. Spence. An orthogonally persistent Java. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (4):68–75, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sidell:1996:MDD

[1348] Jeff Sidell. The Mariposa distributed database management system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (4):76–80, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:1996:NSS

[1349] Andrew Eisenberg. New standard for stored procedures in SQL. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (4):81–88, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bhashyam:1996:TDC

[1350] Ramesh Bhashyam. TPC-D — the challenges, issues and results. SIG-MOD Record (ACM Special Interest Group on Management of Data), 25 (4):89–93, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bashyam:1996:TDC

[1351] R. Bashyam. TPC-D — the challenges, issues and results. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):89–??, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Qian:1996:NPD

[1352] Xiolai Qian. New programs at DARPA and NSF. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):94–98, December 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1997:SOCa

[1353] Anonymous. SIGMOD officers, committees, awards, and institutional sponsors. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):1–??, ???? 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gunther:1997:EIS

[1354] Oliver Günther. Environmental information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):3–4, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gunter:1997:SSE

[1355] O. Gunter. Special section on environmental information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):3–??, ???? 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abel:1997:IMS

[1356] David J. Abel, Kerry Taylor, and Dean Kun. Integrating modelling systems for environmental management information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):5–10, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tomasic:1997:IAE

[1357] Anthony Tomasic and Eric Simon. Improving access to environmental data using context information. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):11–15, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kramer:1997:WUW

[1358] Ralf Kramer, Ralf Nicholai, Arne Koschel, Claudia Rolker, Peter Lockemann, Andree Keitel, Rudolf Legat, and Konrad Tirm. WWW-UDK: a Web-based environmental metainformation system. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):16–21, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Riekert:1997:MDS

[1359] Wolf-Fritz Riekert, Roland Mayer-Föll, and Gerlinde Wiest. Management of data and services in the environmental information system (UIS) of Baden-Württemberg. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):22–26, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Frew:1997:DME

[1360] James Frew and Jeff Dozier. Data management for Earth system science. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):27–31, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gardels:1997:OGL

[1361] Kenn Gardels. Open GIS and online environmental libraries. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):32–38, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buneman:1997:MLP

[1362] Peter Buneman, Louiqa Raschid, and Jeffrey Ullman. Mediator languages — a proposal for a standard: report of an I3/POB working group held at the University of Maryland, April 12 and 13, 1996. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):39–44, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). See comments [1363].

Rosenthal:1997:CVM

[1363] Arnon Rosenthal, Erich Hughes, Scott Renner, and Len Seligman. A consumer viewpoint on “Mediator languages — a proposal for a standard”. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):45–46, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). See [1362].

Karadimitriou:1997:MMC

[1364] Kosmas Karadimitriou and John M.

Tyler. Min-max compression methods for medical image databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):47–52, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fong:1997:CRO

[1365] Joseph Fong. Converting relational to object-oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):53–58, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramanathan:1997:EOO

[1366] Shekar Ramanathan and Julia Hodges.

Extraction of object-oriented structures from existing relational databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):59–64, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:1997:ODW

[1367] Surajit Chaudhuri and Umeshwar Dayal. An overview of data warehousing and OLAP technology. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):65–74, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Doan:1997:QPN

[1368] Khoa Doan, Catherine Plaisant, Ben Shneiderman, and Tom Bruns. Query previews for networked information systems: a case study with NASA environmental data. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):75–81, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Qian:1997:OIM

[1369] Xiaolei Qian. Opportunities in information management and assurance. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(1):82–84, March 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berchtold:1997:FPS

[1370] Stefan Berchtold, Christian Böhm, Bernhard Braunmüller, Daniel A. Keim, and Hans-Peter Kriegel. Fast parallel similarity search in multimedia databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):1–12, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rafiei:1997:SBQ

[1371] Davood Rafiei and Alberto Mendelzon. Similarity-based queries for time series data. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):13–25, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chawathe:1997:MCD

[1372] Sudarshan S. Chawathe and Hector Garcia-Molina. Meaningful change detection in structured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):26–37, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

ONeil:1997:IQP

[1373] Patrick O’Neil and Dallan Quass. Improved query performance with variant indexes. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):38–49, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zaharioudakis:1997:HCC

[1374] Markos Zaharioudakis and Michael J.

Carey. Highly concurrent cache consistency for indices in client-server database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):50–61, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kornacker:1997:CRG

[1375] Marcel Kornacker, C. Mohan, and Joseph M. Hellerstein. Concurrency and recovery in generalized search trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):62–72, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ho:1997:RQO

[1376] Ching-Tien Ho, Rakesh Agrawal, Nimrod Megiddo, and Ramakrishnan Srikant. Range queries in OLAP data cubes. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):73–88, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roussopoulos:1997:COB

[1377] Nick Roussopoulos, Yannis Kotidis, and Mema Roussopoulos. Cubetree: organization of and bulk incremental updates on the data cube. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):89–99, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mumick:1997:MDC

[1378] Inderpal Singh Mumick, Dallan Quass, and Barinderpal Singh Mumick. Maintenance of data cubes and summary tables in a warehouse. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):100–111, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tsuei:1997:DBS

[1379] Thin-Fong Tsuei, Allan N. Packer, and Keng-Tai Ko. Database buffer size investigation for OLTP workloads. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):112–122, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Doppelhammer:1997:DPR

[1380] Joachen Doppelhammer, Thomas Höppler, Alfons Kemper, and Donald Kossmann. Database performance in the real world: TPC-D and SAP R/3. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):123–134, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1997:BOR

[1381] Michael J. Carey, David J. DeWitt, Jeffrey F. Naughton, Mohammad Asgarian, Paul Brown, Johannes E. Gehrke, and Dhaval N. Shah. The BUCKY object-relational benchmark. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):135–146, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adelberg:1997:SRS

[1382] Brad Adelberg, Hector Garcia-Molina, and Jennifer Widom. The STRIP rule system for efficiently maintaining derived data. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):147–158, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhao:1997:ABA

[1383] Yihong Zhao, Prasad M. Deshpande, and Jeffrey F. Naughton. An arraybased algorithm for simultaneous multidimensional aggregates. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):159– 170, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hellerstein:1997:OA

[1384] Joseph M. Hellerstein, Peter J. Haas, and Helen J. Wang. Online aggregation. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):171–182, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Acharya:1997:BPP

[1385] Swarup Acharya, Michael Franklin, and Stanley Zdonik. Balancing push and pull for data broadcast. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):183–194, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bayardo:1997:IAB

[1386] R. J. Bayardo, Jr., W. Bohrer, R. Brice, A. Cichocki, J. Fowler, A. Helal, V. Kashyap, T. Ksiezyk, G. Martin, M. Nodine, M. Rashid, M. Rusinkiewicz, R. Shea, C. Unnikrishnan, A. Unruh, and D. Woelk. InfoSleuth: Agent-based semantic integration of information in open and dynamic environments. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):195–206, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gravano:1997:SSP

[1387] Luis Gravano, Chen-Chuan K. Chang, Héctor García-Molina, and Andreas Paepcke. STARTS: Stanford proposal for Internet meta-searching. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):207–218, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1997:SEA

[1388] Michael J. Carey and Donald Kossmann. On saying “Enough Already!” in SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):219–230, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Griffin:1997:FIH

[1389] Timothy Griffin and Richard Hull. A framework for implementing hypothetical queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):231–242, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Arpaci-Dusseau:1997:HPS

[1390] Andrea C. Arpaci-Dusseau, Remzi H.

Arpaci-Dusseau, David E. Culler, Joseph M. Hellerstein, and David A. Patterson. High-performance sorting on networks of workstations. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):243– 254, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brin:1997:DIC

[1391] Sergey Brin, Rajeev Motwani, Jeffrey D. Ullman, and Shalom Tsur. Dynamic itemset counting and implication rules for market basket data.

SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):255–264, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brin:1997:BMB

[1392] Sergey Brin, Rajeev Motwani, and Craig Silverstein. Beyond market baskets: Generalizing association rules to correlations. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):265–276, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:1997:SPD

[1393] Eui-Hong Han, George Karypis, and Vipin Kumar. Scalable parallel data mining for association rules. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):277–288, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Korn:1997:ESA

[1394] Flip Korn, H. V. Jagadish, and Christos Faloutsos. Efficiently supporting ad hoc queries in large datasets of time sequences. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):289–300, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Livny:1997:DIQa

[1395] M. Livny, R. Ramakrishnan, K. Beyer, G. Chen, D. Donjerkovic, S. Lawande, J. Myllymaki, and K. Wenger. DE-Vise: Integrated querying and visual exploration of large datasets. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):301–312, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Maheshwari:1997:PGC

[1396] Umesh Maheshwari and Barbara Liskov. Partitioned garbage collection of a large object store. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):313– 323, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Koudas:1997:SSS

[1397] Nick Koudas and Kenneth C. Sevcik. Size separation spatial join. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):324–335, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Patel:1997:BSG

[1398] Jignesh Patel, JieBing Yu, Navin Kabra, Kristin Tufte, Biswadeep Nag, Josef Burger, Nancy Hall, Karthikeyan Ramasamy, Roger Lueder, Curt Ellmann, Jim Kupsch, Shelly Guo, Johan Larson, David De Witt, and Jeffrey Naughton. Building a scalable geo-spatial DBMS: Technology, implementation, and evaluation. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):336–347, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gebhardt:1997:TNS

[1399] Michael Gebhardt, Matthias Jarke, and Stephan Jacobs. A toolkit for negotiation support interfaces to multidimensional data. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):348–356, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bozkaya:1997:DBI

[1400] Tolga Bozkaya and Meral Ozsoyoglu. Distance-based indexing for highdimensional metric spaces. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):357– 368, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Katayama:1997:STI

[1401] Norio Katayama and Shin’ichi Satoh. The SR-tree: An index structure for high-dimensional nearest neighbor queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):369–380, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shivakumar:1997:WII

[1402] Narayanan Shivakumar and Héctor García-Molina. Wave-indices: Indexing evolving databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):381–392, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Quass:1997:LWV

[1403] Dallan Quass and Jennifer Widom. On-line warehouse view maintenance. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):393–404, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Colby:1997:SMV

[1404] Latha S. Colby, Akira Kawaguchi, Daniel F. Lieuwen, Inderpal Singh Mumick, and Kenneth A. Ross. Supporting multiple view maintenance policies. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):405–416, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1997:EVM

[1405] D. Agrawal, A. El Abbadi, A. Singh, and T. Yurek. Efficient view maintenance at data warehouses. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):417–427, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Llirbat:1997:ECR

[1406] François Llirbat, Françoise Fabret, and Eric Simon. Eliminating costly redundant computations from SQL trigger executions. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):428–439, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Motakis:1997:TAA

[1407] Iakovos Motakis and Carlo Zaniolo. Temporal aggregation in active database rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):440–451, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Miller:1997:ARI

[1408] R. J. Miller and Y. Yang. Association rules over interval data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):452–461, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

George:1997:STP

[1409] Binto George and Jayant Haritsa. Secure transaction processing in firm real-time database systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):462–473, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jajodia:1997:UFE

[1410] Sushil Jajodia, Pierangela Samarati, V. S. Subrahmanian, and Eliza Bertino. A unified framework for enforcing multiple access control policies. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):474–485, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gupta:1997:RCP

[1411] Ramesh Gupta, Jayant Haritsa, and Krithi Ramamritham. Revisiting commit processing in distributed database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):486–497, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shasha:1997:LWS

[1412] Dennis Shasha. Lessons from Wall Street: case studies in configuration, tuning, and distribution. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):498– 501, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1997:ORD

[1413] Michael J. Carey, Nelson M. Mattos, and Anil K. Nori. Object-relational database systems (tutorial): principles, products and challenges. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):502, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kramer:1997:DWT

[1414] Ralf Kramer. Databases on the Web: Technologies for federation architectures and case studies. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):503–506, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:1997:DWO

[1415] Surajit Chaudhuri and Umeshwar Dayal. Data warehousing and OLAP for decision support. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):507–508, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:1997:QOC

[1416] Surajit Chaudhuri. Query optimization at the crossroads. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):509, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1997:QOC

[1417] Anonymous. Query optimization at the crossroads: Moderator: Surajit Chaudhuri. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):509–??, ???? 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cruz:1997:DDV

[1418] Isabel F. Cruz, M. Averbuch, Wendy T. Lucas, Melissa Radzyminski, and Kirby Zhang. Delaunay: a database visualization system. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):510–513, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goyal:1997:PPP

[1419] Nita Goyal, Charles Hoch, Ravi Krishnamurthy, Brian Meckler, Michael Suchow, and Moshe Zloof. Picture programming project. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):514–516, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Livny:1997:DDA

[1420] M. Livny, R. Ramakrishnan, K. Beyer, G. Chen, D. Donjerkovic, S. Lawande, J. Myllymaki, and K. Wenger. DEVise (demo abstract): integrated querying and visual exploration of large datasets. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):517–520, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:1997:SOB

[1421] Wen-Syan Li, K. Selçuk Candan, Kyoji Hirata, and Yoshinori Hara. SEM-COG: An object-based image retrieval system and its visual query interface. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):521–524, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bressan:1997:CIM

[1422] S. Bressan, C. H. Goh, K. Fynn, M. Jakobisiak, K. Hussein, H. Kon, T. Lee, S. Madnick, T. Pena, J. Qu, A. Shum, and M. Siegel. The COntext INterchange mediator prototype. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):525–527, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Atzeni:1997:MMD

[1423] Paolo Atzeni and Riccardo Torlone.

MDM: a multiple-data model tool for the management of heterogeneous database schemes. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):528–531, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hammer:1997:TBW

[1424] Joachim Hammer, Héctor García-Molina, Svetlozar Nestorov, Ramana Yerneni, Marcus Breunig, and Vasilis Vassalos. Template-based wrappers in the TSIMMIS system. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):532– 535, June 1997. CODEN SRECD8.

ISSN 0163-5808 (print), 1943-5835 (electronic).

Gingras:1997:LMD

[1425] Frédéric Gingras, Laks V. S. Lakshmanan, Iyer N. Subramanian, Despina Papoulis, and Nematollaah Shiri. Languages for multi-database interoperability. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):536–538, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Genesereth:1997:III

[1426] Michael R. Genesereth, Arthur M. Keller, and Oliver M. Duschka. Infomaster: An information integration system. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):539–542, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bayardo:1997:IP

[1427] R. J. Bayardo, Jr., W. Bohrer, R. Brice, A. Cichocki, J. Fowler, A. Helal, V. Kashyap, T. Ksiezyk, G. Martin, M. Nodine, M. Rashid, M. Rusinkiewicz, R. Shea, C. Unnikrishnan, A. Unruh, and D. Woelk. The InfoSleuth project. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):543–545, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tomasic:1997:DIS

[1428] Anthony Tomasic, Rémy Amouroux, Philippe Bonnet, Olga Kapitskaia, Hubert Naacke, and Louiqa Raschid. The Distributed Information Search Component (Disco) and the World Wide Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):546–548, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fernandez:1997:SWS

[1429] Mary Fernandez, Daniela Florescu, Jaewoo Kang, Alon Levy, and Dan Suciu. STRUDEL: a Web site management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):549–552, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:1997:GSP

[1430] Jaiwei Han, Krzysztof Koperski, and Nebojsa Stefanovic. GeoMiner: a system prototype for spatial data mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):553–556, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Labio:1997:WPD

[1431] Wilburt J. Labio, Yue Zhuge, Janet L.

Wiener, Himanshu Gupta, Héctor García-Molina, and Jennifer Widom. The WHIPS prototype for data warehouse creation and maintenance. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (2):557–559, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:1997:SMD

[1432] Jason Tsong-Li Wang, Dennis Shasha, George J. S. Chang, Liam Relihan, Kaizhong Zhang, and Girish Patel. Structural matching and discovery in document databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):560–563, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berchtold:1997:SSS

[1433] Stefan Berchtold and Hans-Peter Kriegel. S3: Similarity search in CAD database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):564–567, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Seshadri:1997:PDE

[1434] Praveen Seshadri and Mark Paskin. PREDATOR: An OR-DBMS with enhanced data types. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):568–571, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chakravarthy:1997:SOO

[1435] S. Chakravarthy. SENTINEL: An object-oriented DBMS with eventbased rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):572–575, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wodtke:1997:MWE

[1436] Dirk Wodtke, Jeanine Weissenfels, Gerhard Weikum, Angelika Kotz Dittrich, and Peter Muth. The MENTOR workbench for enterprise-wide workflow management. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):576–579, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ioannidis:1997:ZDE

[1437] Yannis E. Ioannidis, Miron Livny, Anastassia Ailamaki, Anand Narayanan, and Andrew Therber. ZOO: a desktop experiment management environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):580–583, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fernandez:1997:QLW

[1438] Mary Fernandez, Daniela Florescu, Alon Levy, and Dan Suciu. A query language for a Web-site management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(3):4–11, September 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barbara:1997:QCE

[1439] Daniel Barbará and Mark Sullivan.

Quasi-cubes: Exploiting approximations in multidimensional databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 26 (3):12–17, September 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Clement:1997:OTI

[1440] Gilles Clément, Christian Larouche, Denis Gouin, Paul Morin, and Henry Kucera. OGDI: Toward interoperability among geospatial databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (3):18–23, September 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hadzilacos:1997:EER

[1441] Thanasis Hadzilacos and Nectaria Tryfona. An extended entity-relationship model for geographic applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 26 (3):24–29, September 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jukic:1997:ABM

[1442] Nenad A. Jukic and Susan V. Vrbsky. Asserting beliefs in MLS relational models. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(3):30–35, September 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Silberschatz:1997:DSB

[1443] Avi Silberschatz and Stan Zdonik. Database systems — breaking out of the box. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(3):36–50, September 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramamritham:1997:RDD

[1444] Krithi Ramamritham and Nandit Soparkar. Report on DART ’96: databases: active and real-time (concepts meet practice). SIGMOD Record (ACM Special Interest Group on Management of Data), 26(3):51–53, September 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McHugh:1997:LDM

[1445] Jason McHugh, Serge Abiteboul, Roy Goldman, Dallas Quass, and Jennifer Widom. Lore: a database management system for semistructured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 26 (3):54–66, September 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Konig-Ries:1997:RDD

[1446] Brigitta König-Ries and Peter C. Lockermann. Research in databases and data-intensive applications Computer Science Dept. and FZI, University of Karlsruhe. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(3):67–72, September 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1997:SOCb

[1447] Anonymous. SIGMOD officers, committees, and awards. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):1–??, ???? 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Suciu:1997:FMS

[1448] Dan Suciu. Foreword: Management of semistructured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):4–7, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ashish:1997:WGS

[1449] Naveen Ashish and Craig A. Knoblock.

Wrapper generation for semi-structured Internet sources. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):8–15, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Atzeni:1997:SSD

[1450] Paolo Atzeni, Giansalvatore Mecca, and Paolo Merialdo. Semistructured and structured data in the Web: Going back and forth. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):16–23, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McHugh:1997:IDF

[1451] Jason McHugh and Jennifer Widom. Integrating dynamically-fetched external information into a DBMS for semistructured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):24–31, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nado:1997:EEP

[1452] Robert A. Nado and Scott B. Huffman. Extracting entity profiles from semistructured information spaces. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):32–38, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nestorov:1997:ISS

[1453] Svetlozer Nestorov, Serge Abiteboul, and Rajeev Motwani. Inferring structure in semistructured data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (4):39–43, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nestoroc:1997:ISS

[1454] S. Nestoroc, S. Abiteboul, and R. Motwani. Inferring structure in semistructured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):39–??, ???? 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Baader:1997:IAH

[1455] Franz Baader, Manfred A. Jeusfeld, and Werner Nutt. Intelligent access to heterogeneous information sources: report on the 4th workshop on knowledge representation meets databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):44–48, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McClatchey:1997:WWM

[1456] R. McClatchey and G. Vossen. Workshop on workflow management in scientific and engineering applications — report. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):49–53, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Conrad:1997:RIF

[1457] S. Conrad, B. Eaglestone, W. Hasselbring, M. Roantree, M. Schöhoff, M. Strässler, M. Vermeer, and F. Saltor. Research issues in federated database systems: Report of EFDBS ’97 Workshop. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):54–56, December 1997.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gupta:1997:VDT

[1458] Ashish Gupta, Venky Harinarayan, and Anand Rajaraman. Virtual database technology. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):57–61, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Seligman:1997:IP

[1459] L. Seligman. Industry perspectives. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):62–??, ???? 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gray:1997:FMR

[1460] Jim Gray and Goetz Graefe. The fiveminute rule ten years later, and other computer storage rules of thumb. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (4):63–68, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jajodia:1997:ISR

[1461] Sushil Jajodia, Daniel Barbará, Alex Brodsky, Larry Kerschberg, Ami Motro, Edgar Sibley, and X. Sean Wang. Information systems research at George Mason University. SIG-MOD Record (ACM Special Interest Group on Management of Data), 26 (4):69–74, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dadam:1997:DIS

[1462] Peter Dadam and Wolfgang Klas. The Database and Information System Research Group at the University of Ulm. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):75–79, December 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1997:WFM

[1463] Anonymous. Workshop on foundations of models and languages for data and objects. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):80–??, ???? 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1997:SSP

[1464] Anonymous. SIGMETRICS Symposium on Parallel and Distributed Tools. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(4):81–??, ???? 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhri:1998:RWR

[1465] Akmal B. Chaudhri. Reports: Workshop report on experiences using object data management in the real-world. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):5–10, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Musick:1998:RSI

[1466] Ron Musick and Chris Miller. Report on the Second IEEE Metadata Conference (Metadata ’97). SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):11–15, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Seshadri:1998:PRD

[1467] Praveen Seshadri. PREDATOR: a resource for database research. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):16–20, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roussopoulos:1998:MVD

[1468] Nick Roussopoulos. Materialized views and data warehouses. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):21–26, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Burton:1998:AJP

[1469] Bradley F. Burton and Victor W. Marek. Applications of JAVA programming language to database management. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):27–34, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gatziu:1998:UAF

[1470] Stella Gatziu, Arne Koschel, Günter von Bültzingsloewen, and Hans Fritschi. Unbundling active functionality. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):35–40, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gatzui:1998:UAF

[1471] S. Gatzui, A. Koschel, G. von Bultzingsloewen, and H. Fritschi. Unbundling active functionality. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (1):35–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuok:1998:MFA

[1472] Chan Man Kuok, Ada Fu, and Man Hon Wong. Mining fuzzy association rules in databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):41–46, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tsotras:1998:ENS

[1473] Vassilis J. Tsotras, Christian S. Jensen, and Richard T. Snodgrass. An extensible notation for spatiotemporal index queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):47–53, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:1998:RIPa

[1474] Richard Snodgrass, Elisa Bertino, Mike Carey, Jim Gray, Henry F. Korth, Betty Salzberg, and Dennis Shasha. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):54–57, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:1998:TCP

[1475] Chialin Chang, Anurag Acharya, Alan Sussman, and Joel Saltz. T2: a customizable parallel database for multidimensional data. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):58–66, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Koksal:1998:WHM

[1476] Pinar Koksal, Sena Nural Arpinar, and Asuman Dogac. Workflow history management. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):67–75, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manola:1998:TRW

[1477] Frank Manola. Towards a richer Web object model. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):76–80, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenthal:1998:WWO

[1478] Arnon Rosenthal. Where will object technology drive data administration? SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):81–87, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:1998:ROO

[1479] Philip A. Bernstein. Repositories and object oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):88–96, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:1998:TLA

[1480] Jiawei Han. Towards on-line analytical mining in large databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (1):97–107, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1998:TIC

[1481] Anonymous. Third IFCIS Conference on Cooperative Information Systems (CoopIS ’98). SIGMOD Record (ACM Special Interest Group on Management of Data), 27(1):108–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tsur:1998:QFG

[1482] Dick Tsur, Jeffrey D. Ullman, Serge Abiteboul, Chris Clifton, Rajeev Motwani, Svetlozar Nestorov, and Arnon Rosenthal. Query flocks: a generalization of association-rule mining. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27(2): 1–12, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ng:1998:EMP

[1483] Raymond T. Ng, Laks V. S. Lakshmanan, Jiawei Han, and Alex Pang. Exploratory mining and pruning optimizations of constrained associations rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):13–24, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shintani:1998:PMA

[1484] Takahiko Shintani and Masaru Kitsuregawa. Parallel mining algorithms for generalized association rules with classification hierarchy. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):25–36, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rao:1998:RIN

[1485] Jun Rao and Kenneth A. Ross. Reusing invariants: a new strategy for correlated queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):37–48, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fegaras:1998:QUO

[1486] Leonidas Fegaras. Query unnesting in object-oriented databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):49–60, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cherniack:1998:CRT

[1487] Mitch Cherniack and Stan Zdonik. Changing the rules: Transformations for rule-based optimizers. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):61–72, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guha:1998:CEC

[1488] Sudipto Guha, Rajeev Rastogi, and Kyuseok Shim. CURE: An efficient clustering algorithm for large databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):73–84, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bayardo:1998:EML

[1489] Roberto J. Bayardo, Jr. Efficiently mining long patterns from databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):85–93, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:1998:ASC

[1490] Rakesh Agrawal, Johannes Gehrke, Dimitrios Gunopulos, and Prabhakar Raghavan. Automatic subspace clustering of high dimensional data for data mining applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):94–105, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kabra:1998:EMQ

[1491] Navin Kabra and David J. DeWitt.

Efficient mid-query re-optimization of sub-optimal query execution plans. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):106–117, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jonsson:1998:IQE

[1492] Björn T. Jónsson, Michael J. Franklin, and Divesh Srivastava. Interaction of query evaluation and buffer management for information retrieval. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):118–129, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Urhan:1998:CBQ

[1493] Tolga Urhan, Michael J. Franklin, and Laurent Amsaleg. Cost based query scrambling for initial delays. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):130–141, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berchtold:1998:PTT

[1494] Stefan Berchtold, Christian Böhm, and Hans-Peter Kriegal. The pyramidtechnique: Towards breaking the curse of dimensionality. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):142–153, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Seidl:1998:OMS

[1495] Thomas Seidl and Hans-Peter Kriegel. Optimal multi-step k-nearest neighbor search. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):154–165, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kanth:1998:DRS

[1496] K. V. Ravi Kanth, Divyakant Agrawal, and Ambuj Singh. Dimensionality reduction for similarity searching in dynamic databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):166–176, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cluet:1998:YMN

[1497] Sophie Cluet, Claude Delobel, Jérôme Siméon, and Katarzyna Smaga. Your mediators need data conversion! SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):177–188, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Miller:1998:USH

[1498] Renée J. Miller. Using schematically heterogeneous structures. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):189–200, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cohen:1998:IHD

[1499] William W. Cohen. Integration of heterogeneous databases without common domains using queries based on textual similarity. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):201–212, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grumbach:1998:DSC

[1500] Stéphane Grumbach, Philippe Rigaux, and Luc Segoufin. The DEDALE system for complex spatial queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):213–224, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Papadopoulos:1998:SQP

[1501] Apostolos N. Papadopoulos and Yannis Manolopoulos. Similarity query processing using disk arrays. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):225–236, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hjaltason:1998:IDJ

[1502] Gísli R. Hjaltason and Hanan Samet. Incremental distance join algorithms for spatial databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):237–248, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kotidis:1998:ASO

[1503] Yannis Kotidis and Nick Roussopoulos. An alternative storage organization for ROLAP aggregate views based on cubetrees. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):249–258, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deshpande:1998:CMQ

[1504] Prasad M. Deshpande, Karthikeyan Ramasamy, Amit Shukla, and Jeffrey F. Naughton. Caching multidimensional queries using chunks. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):259–270, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhao:1998:SOE

[1505] Yihong Zhao, Prasad M. Deshpande, Jeffrey F. Naughton, and Amit Shukla. Simultaneous optimization and evaluation of multiple dimensional queries.

SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):271–282, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adelberg:1998:NTS

[1506] Brad Adelberg. NoDoSE: a tool for semi-automatically extracting structured and semi-structured data from text documents. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):283–294, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nestorov:1998:ESS

[1507] Svetlozar Nestorov, Serge Abiteboul, and Rajeev Motwani. Extracting schema from semistructured data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):295–306, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chakrabarti:1998:EHC

[1508] Soumen Chakrabarti, Byron Dom, and Piotr Indyk. Enhanced hypertext categorization using hyperlinks. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):307–318, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Subramanian:1998:CBO

[1509] Subbu N. Subramanian and Shivakumar Venkataraman. Cost-based optimization of decision support queries using transient views. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):319–330, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gibbons:1998:NSB

[1510] Phillip B. Gibbons and Yossi Matias. New sampling-based summary statistics for improving approximate query answers. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):331–342, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sarawagi:1998:IAR

[1511] Sunita Sarawagi, Shiby Thomas, and Rakesh Agrawal. Integrating association rule mining with relational database systems: Alternatives and implications. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):343–354, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chan:1998:BID

[1512] Chee-Yong Chan and Yannis E. Ioannidis. Bitmap index design and evaluation. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):355–366, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:1998:AWI

[1513] Surajit Chaudhuri and Vivek Narasayya. AutoAdmin “what-if” index analysis utility. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):367–378, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jaedicke:1998:PPA

[1514] Michael Jaedicke and Bernhard Mitschang.

On parallel processing of aggregate and scalar functions in object-relational DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):379–389, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Godfrey:1998:SPD

[1515] Michael Godfrey, Tobias Mayr, Praveen Seshadri, and Thorsten von Eicken. Secure and portable database extensibility. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):390–401, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adali:1998:MSA

[1516] Sibel Adali, Pietro Bonatti, Maria Luisa Sapino, and V. S. Subrahmanian. A multi-similarity algebra. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):402– 413, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fernandez:1998:CBS

[1517] Mary Fernández, Daniela Florescu, Jaewoo Kang, Alon Levy, and Dan Suciu. Catching the boat with Strudel: experiences with a Web-site management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):414–425, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manku:1998:AMO

[1518] Gurmeet Singh Manku, Sridhar Rajagopalan, and Bruce G. Lindsay. Approximate medians and other quantiles in one pass and with limited memory. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):426–435, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:1998:RSH

[1519] Surajit Chaudhuri, Rajeev Motwani, and Vivek Narasayya. Random sampling for histogram construction: How much is enough? SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):436–447, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Matias:1998:WBH

[1520] Yossi Matias, Jeffrey Scott Vitter, and Min Wang. Wavelet-based histograms for selectivity estimation. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):448–459, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:1998:ETA

[1521] David Lomet and Gerhard Weikum. Efficient transparent application recovery in client-server information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):460–471, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Larson:1998:MMD

[1522] Per-Åke Larson and Goetz Graefe.

Memory management during run generation in external sorting. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):472–483, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anderson:1998:RCP

[1523] Todd Anderson, Yuri Breitbart, Henry F. Korth, and Avishai Wool. Replication, consistency, and practicality: Are these mutually exclusive? SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):484–495, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:1998:WWR

[1524] Michael Stonebraker. Are we working on the right problems? (panel). SIGMOD Record (ACM Special Interest Group on Management of Data), 27 (2):496, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mylopoulos:1998:NGD

[1525] John Mylopoulos. Next generation database systems won’t work without semantics! (panel). SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):497, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adam:1998:ECT

[1526] Nabil R. Adam and Yelena Yesha. Electronic commerce: tutorial. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):498, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kemper:1998:SRT

[1527] Alfons Kemper, Donald Kossmann, and Florian Matthes. SAP R/3 (tutorial): a database application system. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):499, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Clossman:1998:JRD

[1528] Gray Clossman, Phil Shaw, Mark Hapner, Johannes Klein, Richard Pledereder, and Brian Becker. Java and relational databases (tutorial): SQLJ. SIGMOD Record (ACM Special Interest Group on Management of Data), 27 (2):500, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berchtold:1998:HDI

[1529] Stefan Berchtold and Daniel A. Keim. High-dimensional index structures: database support for next decade’s applications (tutorial). SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):501, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blakeley:1998:MUD

[1530] José A. Blakeley and Michael J. Pizzo. Microsoft universal data access platform. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):502–503, June 1998.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

White:1998:EJP

[1531] Seth White, Rick Cattell, and Shel Finkelstein. Enterprise Java platform data access. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):504–505, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Reinwald:1998:SOH

[1532] Berthold Reinwald and Hamid Pirahesh. SQL open heterogeneous data access. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):506–507, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eaton:1998:MLS

[1533] Chris Eaton. Managing large systems with DB2 UDB. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):508–509, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Doherty:1998:DSM

[1534] C. Gregory Doherty. Database systems management and Oracle8. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):510–511, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Spiro:1998:UST

[1535] Peter Spiro. Ubiquitous, self-tuning, scalable servers. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):512–515, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Franklin:1998:DYF

[1536] Michael Franklin and Stan Zdonik. “data in your face”: Push technology in perspective. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):516–519, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramakrishnan:1998:PNA

[1537] Satish Ramakrishnan and Vibha Dayal. The PointCast network (abstract). SIGMOD Record (ACM Special Interest Group on Management of Data), 27 (2):520, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chan:1998:TPS

[1538] Arvola Chan. Transactional publish/ subscribe: the proactive multicast of database changes (abstract). SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):521, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ng:1998:DMA

[1539] KianSing Ng, Huan Liu, and Hwee-Bong Kwah. A data mining application: Customer retention at the Port of Singapore Authority (PSA). SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):522–525, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anderson:1998:ORR

[1540] Richard Anderson, Gopalan Arun, and Richard Frank. Oracle Rdb’s record caching model. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):526–527, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lahiri:1998:UOU

[1541] Tirthankar Lahiri, Ashok Joshi, Amit Jasuja, and Sumanta Chatterjee. 50,000 users on an Oracle8 universal server database. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):528–530, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aulakh:1998:AQD

[1542] Kamar Aulakh. About Quark Digital Media System. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):531–532, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Whelan:1998:FID

[1543] Daniel S. Whelan. FileNet integrated document management: Database usage and issues. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):533, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nauman:1998:DHT

[1544] John Nauman and Ray Suorsa. Developing a high traffic, read-only Web site. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):534–535, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chong:1998:RBP

[1545] James Chong and Charles Schwab. Real business processing meets the Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):536, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lassettre:1998:ORD

[1546] Edwin R. Lassettre. Olympic records for data at the 1998 Nagano Games. SIGMOD Record (ACM Special Interest Group on Management of Data), 27 (2):537, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brewer:1998:DHA

[1547] Eric A. Brewer. Delivering high availability for Inktomi search engines. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27(2): 538, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kennamer:1998:MCH

[1548] Sherri Kennamer. Microsoft.com: a high-scale data management and transaction processing solution. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):539– 540, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:1998:FPD

[1549] Bin Li and Dennis Shasha. Free parallel data mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):541–543, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mecca:1998:AWB

[1550] Giansalvatore Mecca, Paolo Atzeni, Alessandro Masci, Paolo Merialdo, and Giuseppe Sindoni. The ARANEUS Web-base management system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):544–546, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Liu:1998:CPU

[1551] Ling Liu, Calton Pu, Wei Tang, David Buttler, John Biggs, Tong Zhou, Paul Benninghoff, Wei Han, and Fenghua Yu. CQ: a personalized update monitoring toolkit. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):547–549, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Olston:1998:D

[1552] Chris Olston, Allison Woodruff, Alexander Aiken, Michael Chu, Vuk Ercegovac, Mark Lin, Mybrid Spalding, and Michael Stonebraker. DataSplash. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):550–552, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:1998:MIT

[1553] Surajit Chaudhuri and Vivek Narasayya.

Microsoft index tuning wizard for SQL Server 7.0. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):553–554, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kiepuszewski:1998:FPB

[1554] Bartek Kiepuszewski, Ralf Muhlberger, and Maria E. Orlowska. FlowBack: Providing backward recovery for workflow management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):555– 557, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cohen:1998:PDL

[1555] William W. Cohen. Providing database-like access to the Web using queries based on textual similarity. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):558–560, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ambite:1998:ASC

[1556] José Luis Ambite, Naveen Ashish, Greg Barish, Craig A. Knoblock, Steven Minton, Pragnesh J. Modi, Ion Muslea, Andrew Philpot, and Sheila Tejada. ARIADNE: a system for constructing mediators for Internet sources. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):561–563, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:1998:CBM

[1557] Chen Li, Ramana Yerneni, Vasilis Vassalos, Hector Garcia-Molina, Yannis Papakonstantinou, Jeffrey Ullman, and Murty Valiveti. Capability based mediation in TSIMMIS. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):564–566, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Avnur:1998:CCO

[1558] Ron Avnur, Joseph M. Hellerstein, Bruce Lo, Chris Olston, Bhaskaran Raman, Vijayshankar Raman, Tali Roth, and Kirk Wylie. CONTROL: Continuous output and navigation technology with refinement on-line. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):567–569, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kornacker:1998:AAM

[1559] Marcel Kornacker, Mehul Shah, and Joseph M. Hellerstein. AMDB: an access method debugging tool. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):570–571, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lacroix:1998:UOS

[1560] Zoé Lacroix, Arnaud Sahuguet, and Raman Chandrasekar. User-oriented smart-cache for the Web: What You Seek is What You Get! SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (2):572–574, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Baumann:1998:MDS

[1561] Peter Baumann, Andreas Dehmel, Paula Furtado, Roland Ritsch, and Norbert Widmann. The multidimensional database system RasDaMan. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):575–577, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Park:1998:XEM

[1562] Jang Ho Park, Yong Sik Kwon, Ki Hong Kim, Sang Ho Lee, Byoung Dae Park, and Sang Kyun Cha. Xmas: An extensible main-memory storage system for high-performance applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):578–580, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zaiane:1998:MSP

[1563] Osmar R. Zaïane, Jiawei Han, Ze-Nian Li, Sonny H. Chee, and Jenny Y. Chiang. MultiMediaMiner: a system prototype for MultiMedia data mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):581–583, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Toyama:1998:SES

[1564] Motomichi Toyama. SuperSQL: An extended SQL for database publishing and presentation. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):584–586, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ceri:1998:IWL

[1565] Stefano Ceri, Piero Fraternali, and Stefano Paraboschi. The IDEA Web lab. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):587–589, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dar:1998:DDD

[1566] Shaul Dar, Gadi Entin, Shai Geva, and Eran Palmon. DTL’s DataSpot: database exploration as easy as browsing the Web . . .. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):590–592, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:1998:PPQ

[1567] Jing Chen, Limsoon Wong, and Louxin Zhang. A protein patent query system powered by Kleisli. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):593–595, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Atzeni:1998:WIW

[1568] P. Atzeni and A. Mendelzon. WebDB’98:

International Workshop on the Web and Databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):4–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Borgida:1998:KIW

[1569] A. Borgida, V. Chaudhri, and M. Staudt. KRDB’98: The 5th International Workshop on Knowledge Representation Meets Databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27(3): 10–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cheung:1998:ENN

[1570] K. Cheung and A. Fu. Enhanced nearest neighbour search on the R-tree. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):16–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Griffin:1998:ACP

[1571] T. Griffin and B. Kumar. Algebraic change propagation for semijoin and outerjoin, queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):22–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:1998:BTP

[1572] D. Lomet. B-tree page size when caching is considered. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):28–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:1998:RIPb

[1573] R. Snodgrass, H. Garcia-Molina, T. Imielisnki, D. Maier, P. Selinger, and J. Ullman. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):33–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kappel:1998:TAO

[1574] G. Kappel and W. Retschitzegger. The TriGS active object-oriented database system: An overview. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):36–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keeton:1998:CID

[1575] K. Keeton, D. Patterson, and J. Hellerstein. A case for intelligent disks (IDISKs). SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):42–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:1998:SP

[1576] A. Eisenberg and J. Melton. Standards in practice. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):53–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Florescu:1998:DTW

[1577] D. Florescu, A. Levy, and A. Mendelzon. Database techniques for the World-Wide Web: a survey. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (3):59–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:1998:DRC

[1578] S. Chang, L. Gravano, G. Kaiser, K. Ross, and S. Stolfo. Database research at Columbia University. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (3):75–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:1998:MDR

[1579] D. Lomet, R. Barga, S. Chaudhri, and P. Larson. The Microsoft Database Research Group. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):81–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1998:ASR

[1580] Anonymous. ACM SIGMOD Record December 1998 Special Section on Electronic Commerce. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):86–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1998:SIG

[1581] Anonymous. Semantic interoperability in global information systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27 (3):87–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1998:VCP

[1582] Anonymous. VLDB’99 conference preliminary conference announcement and call for papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(3):88–??, ???? 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dogac:1998:I

[1583] A. Dogac. Introduction. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):5–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bichler:1998:CBC

[1584] M. Bichler, A. Segev, and J. Zhao.

Component-based E-commerce: Assessment of current practices. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27(4): 7–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Danish:1998:BDD

[1585] S. Danish. Building database-driven electronic catalogs. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):15–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Meltzer:1998:XEC

[1586] B. Meltzer and R. Glushko. XML and electronic commerce: Enabling the network economy. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):21–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dogac:1998:WBE

[1587] A. Dogac, I. Durusoy, S. Arpinar, N. Tatbul, P. Koskal, I. Cingil, and N. Dimililer. A workflow-based electronic marketplace on the Web. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27(4): 25–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jennings:1998:AAB

[1588] N. Jennings, T. Norman, and P. Faratin. ADEPT: An agent-based approach to business process management. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):32–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Reich:1998:CAD

[1589] A. Reich and I. Ben-Shaul. A componentized architecture for dynamic electronic markets. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):40–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Boll:1998:DIR

[1590] S. Boll, W. Klas, and B. Battaglin. Design and implementation of RMP — a virtual electronic market place. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27(4): 48–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buchner:1998:DIM

[1591] A. Buchner and M. Mulvenna. Discovering Internet marketing intelligence through online analytical Web usage mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):54–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Domingo-Ferrer:1998:AEC

[1592] J. Domingo-Ferrer and J. Herrera-Joancomarti. An anonymous electronic commerce scheme with an off-line authority and untrusted agents. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):62–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christoffel:1998:EMR

[1593] M. Christoffel, S. Pilkowski, B. Schmitt, and P. Lockemann. Electronic market: The roadmap for university libraries and members to survive the information jungle. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):68–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:1998:ARD

[1594] P. Bernstein, M. Brodie, S. Ceri, D. Dewitt, M. Franklin, H. Garcia-Molina, J. Gray, J. Held, J. Hellerstein, H. Jagadish, M. Lesk, D. Maier, J. Naughton, H. Pirahesh, M. Stonebraker, and J. Ullman. The Asilomar report on database research. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):74–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:1998:RIPc

[1595] R. Snodgrass, L. Haas, A. Mendelzohn, M. Ozsoyoglu, J. Paredaens, K. Ramamritham, N. Roussopoulos, J. Widom, and P. Yu. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):81–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ritter:1998:MM

[1596] D. Ritter. The middleware muddle. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):86–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:1998:SPN

[1597] A. Eisenberg and J. Melton. SQLJ part O, now known as SQL/OLB (objectlanguage bindings). SIGMOD Record (ACM Special Interest Group on Management of Data), 27(4):94–??, 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1999:SOCa

[1598] Anonymous. SIGMOD officers, committees, and awards. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):1–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ouksel:1999:BIR

[1599] A. Ouksel and A. Sheth. A brief introduction to the research area and the special section. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):5–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Mackay:1999:SIE

[1600] D. Mackay. Semantic integration of environmental models for application to global information systems and decision-making. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):13–??, 1999.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Forte:1999:SPI

[1601] E. Forte, F. Haenni, K. Warkentyne, E. Duval, K. Cardinaels, E. Vervaet, K. Hendrikx, M. Forte, and F. Simillion. Semantic and pedagogic interoperability mechanisms in the ARI-ADNE educational repository. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28(1): 20–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Smith:1999:USS

[1602] K. Smith and L. Obrst. Unpacking the semantics of source and usage to perform semantic reconciliation in largescale information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):26–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Hampapur:1999:SVI

[1603] A. Hampapur. Semantic video indexing: Approach and issues. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):32–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Papazoglou:1999:CIS

[1604] M. Papazoğlou and J. Hoppenbrouwers. Contextualizing the information space in federated digital libraries. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):40–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Sycara:1999:SMA

[1605] K. Sycara, M. Klusch, S. Widoff, and J. Lu. 3dynamic service matchmaking among agents in open information environments. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):47–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Bergamaschi:1999:SIS

[1606] S. Bergamaschi, S. Castano, and M. Vincini. Semantic integration of semistructured and structured data sources. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):54–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Fowler:1999:ABS

[1607] J. Fowler, B. Perry, M. Nodine, and B. Bargmeyer. Agent-based semantic interoperability in InfoSleuth. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28(1):

60–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Gal:1999:SII

[1608] A. Gal. Semantic interoperability in information services: Experiencing with CoopWARE. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):68–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Nascimento:1999:RBS

[1609] M. Nascimento and C. Medeiros. Report on the 13th Brazilian Symposium on Database Systems (SBBD ’98). SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):76–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

FariasdeSouza:1999:EMU

[1610] M. Farias de Souza and M. Costa Sampaio. Efficient materialization and use of views in data warehouses. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):78–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). Special section on semantic interoperability in global information systems.

Ceri:1999:DPD

[1611] S. Ceri, P. Fraternali, and S. Paraboschi. Design principles for dataintensive Web sites. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):84–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuan:1999:SDP

[1612] J. Kuan and P. Lewis. A study on data point search for HG-trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):90–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roantree:1999:OMP

[1613] M. Roantree, J. Murphy, and W. Hasselbring. The OASIS multidatabase prototype. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):97–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sheth:1999:VSS

[1614] A. Sheth, C. Bertram, and K. Shah. VideoAnywhere: a system for searching and managing distributed heterogeneous video assets. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):104–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snograss:1999:RIP

[1615] R. Snograss, S. Abiteboul, S. Cluet, M. Franklin, G. Lohman, D. Lomet, G. Ozsoyoglu, R. Ramakrishnan, K. Ross, T. Sellis, and P. Valduriez. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 28 (1):110–??, 1999. CODEN SRECD8.

ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1999:NWI

[1616] M. Carey and L. Seligman. NSF Workshop on Industrial/Academic Cooperation in Database Systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):115– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:1999:SSF

[1617] A. Eisenberg and J. Melton. Standards — SQL:1999, formerly known as SQL3. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(1):131–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mamoulis:1999:ISJ

[1618] Nikos Mamoulis and Dimitris Papadias. Integration of spatial join algorithms for processing multiple inputs. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):1–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Acharya:1999:SES

[1619] Swarup Acharya, Viswanath Poosala, and Sridhar Ramaswamy. Selectivity estimation in spatial databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28(2): 13–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chakrabarti:1999:ECC

[1620] Kaushik Chakrabarti and Sharad Mehrotra. Efficient concurrency control in multidimensional access methods. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):25–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:1999:SSO

[1621] H. V. Jagadish, Laks V. S. Lakshmanan, and Divesh Srivastava. Snakes and sandwiches: Optimal clustering strategies for a data warehouse. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28(2): 37–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ankerst:1999:OOP

[1622] Mihael Ankerst, Markus M. Breunig, Hans-Peter Kriegel, and Jorg Sander. OPTICS: Ordering points to identify the clustering structure. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):49–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aggarwal:1999:FAP

[1623] Charu C. Aggarwal, Cecilia Procopiuc, Joel L. Wolf, Philip S. Yu, and Jong Soo Park. Fast algorithms for projected clustering. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):61–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:1999:LLE

[1624] David Lomet and Mark Tuttle. Logical logging to extend recovery to new domains. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):73–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shanmugasundaram:1999:ECC

[1625] Jayavel Shanmugasundaram, Arvind Nithrakashyap, Rajendran Sivasankaran, and Krithi Ramamritham. Efficient concurrency control for broadcast environments. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):85–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Breitbart:1999:UPP

[1626] Yuri Breitbart, Raghavan Komondoor, Rajeev Rastogi, S. Seshadri, and Avi Silberschatz. Update propagation protocols for replicated databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28(2): 97–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jamil:1999:BRM

[1627] Hasan M. Jamil. Belief reasoning in MLS deductive databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):109– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adali:1999:MPA

[1628] Sibel Adali, Maria-Luisa Sapino, and V. S. Subrahmanian. A multimedia presentation algebra. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):121–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:1999:QND

[1629] H. V. Jagadish, Laks V. S. Lakshmanan, Tova Milo, Divesh Srivastava, and Dimitra Vista. Querying network directories. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):133–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hidber:1999:OAR

[1630] Christian Hidber. Online association rule mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):145–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lakshmanan:1999:OCF

[1631] Laks V. S. Lakshmanan, Raymond T. Ng, Jiawei Han, and Alex Pang. Optimization of constrained frequent set queries with 2-variable constraints. SIGMOD Record (ACM Special Interest Group on Management of Data), 28 (2):157–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gehrke:1999:BOD

[1632] Johannes Gehrke, Venkatesh Ganti, Raghu Ramakrishnan, and Wei-Yin Loh. BOAT — optimistic decision tree construction. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):169–??, 1999.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aboulnaga:1999:STH

[1633] Ashraf Aboulnaga and Surajit Chaudhuri. Self-tuning histograms: Building histograms without looking at data. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):181–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vitter:1999:ACM

[1634] Jeffrey Scott Vitter and Min Wang.

Approximate computation of multidimensional aggregates of sparse data using wavelets. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):193–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chan:1999:EBE

[1635] Chee-Yong Chan and Yannis E. Ioannidis. An efficient bitmap encoding scheme for selection queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):215– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wu:1999:QOS

[1636] Ming-Chuan Wu. Query optimization for selections using bitmaps. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):227– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blohsfeld:1999:CSE

[1637] Bjorn Blohsfeld, Dieter Korus, and Bernhard Seeger. A comparison of selectivity estimators for range queries on metric attributes. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):239–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manku:1999:RST

[1638] Gurmeet Singh Manku, Sridhar Rajagopalan, and Bruce G. Lindsay. Random sampling techniques for space efficient online computation of order statistics of large datasets. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):251– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:1999:RSJ

[1639] Surajit Chaudhuri, Rajeev Motwani, and Vivek Narasayya. On random sampling over joins. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):263–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Acharya:1999:JSA

[1640] Swarup Acharya, Phillip B. Gibbons, Viswanath Poosala, and Sridhar Ramaswamy. Join synopses for approximate query answering. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):275– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Haas:1999:RJO

[1641] Peter J. Haas and Joseph M. Hellerstein. Ripple joins for online aggregation. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):287–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ives:1999:AQE

[1642] Zachary G. Ives, Daniela Florescu, Marc Friedman, Alon Levy, and Daniel S. Weld. An adaptive query execution system for data integration. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):299–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Florescu:1999:QOP

[1643] Daniela Florescu, Alon Levy, Ioana Manolescu, and Dan Suciu. Query optimization in the presence of limited access patterns. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):311–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Marathe:1999:QPT

[1644] Arunprasad P. Marathe and Kenneth Salem. Query processing techniques for arrays. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):323–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:1999:MYV

[1645] Chen-Chuan K. Chang and Hector Garcia-Molina. Mind your vocabulary:

Query mapping across heterogeneous information sources. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):335–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mayr:1999:CSQ

[1646] Tobias Mayr and Praveen Seshadri. Client-site query extensions. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):347– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Beyer:1999:BCS

[1647] Kevin Beyer and Raghu Ramakrishnan. Bottom-up computation of sparse and Iceberg CUBEs. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):359–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kotidis:1999:DDV

[1648] Yannis Kotidis and Nick Roussopoulos. DynaMat: a dynamic view management system for data warehouses. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):371–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Labio:1999:SWU

[1649] Wilburt Juan Labio, Ramana Yerneni, and Hector Garcia-Molina. Shrinking the warehouse update window. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28(2): 383–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Natsev:1999:RSS

[1650] Apostol Natsev, Rajeev Rastogi, and Kyuseok Shim. Research Sessions — session: Similarity search — WAL-RUS: a similarity retrieval algorithm for image databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):395–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aggarwal:1999:NMS

[1651] Charu C. Aggarwal, Joel L. Wolf, and Philip S. Yu. A new method for similarity indexing of market basket data. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):407–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keim:1999:EGB

[1652] Daniel A. Keim. Efficient geometrybased similarity search of 3D spatial databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):419–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deutsch:1999:SSD

[1653] Alin Deutsch, Mary Fernandez, and Dan Suciu. Storing semistructured data with STORED. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):431–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yerneni:1999:CCM

[1654] Ramana Yerneni, Chen Li, Hector Garcia-Molina, and Jeffrey Ullman. Computing capabilities of mediators. SIGMOD Record (ACM Special Interest Group on Management of Data), 28 (2):443–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Papakonstantinou:1999:QRU

[1655] Yannis Papakonstantinou and Vasilis Vassalos. Query rewriting using semistructured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):455–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Embley:1999:RBD

[1656] D. W. Embley, Y. S. Jiang, and Y-K Ng. Record-boundary discovery in Web documents. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):467–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Callan:1999:ADL

[1657] Jamie Callan, Margaret Connell, and Aiqun Du. Automatic discovery of language models for text databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28(2): 479–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Davulcu:1999:LAQ

[1658] Hasan Davulcu, Juliana Freire, Michael Kifer, and I. V. Ramakrishnan. A layered architecture for querying dynamic Web content. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):491–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Seshadri:1999:HSD

[1659] Praveen Seshadri. “Honey, I shrunk the DBMS”: Footprint, mobility, and beyond. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):503–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Garofalakis:1999:CPM

[1660] Minos N. Garofalakis, Shridhar Ramaswamy, Rajeev Rastogi, and Kyuseok Shim. Of crawlers, portals, mice and men: Is there more to mining the Web? SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):504–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ozsu:1999:DMI

[1661] M. Tamer Ozsu. Data management issues in electronic commerce. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):505– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dullmann:1999:PD

[1662] Dirk Dullmann. Petabyte databases.

SIGMOD Record (ACM Special Interest Group on Management of Data), 28 (2):506–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mohan:1999:DPL

[1663] C. Mohan. A database perspective on Lotus Domino/Notes. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):507–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chakrabarti:1999:HDD

[1664] Soumen Chakrabarti. Hypertext databases and data mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):508– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hinneburg:1999:CML

[1665] Alexander Hinneburg and Daniel A. Keim. Clustering methods for large databases: From the past to the future. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):509–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Suciu:1999:MWD

[1666] Dan Suciu. Managing Web data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28(2): 510–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:1999:WHD

[1667] M. Carey, D. Chamberlin, S. Narayanan, B. Vance, D. Doole, S. Rielau, R. Swagerman, and N. Mattos. O-O, what’s happening to DB2? SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):511–??, 1999.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Krishnamurthy:1999:BOR

[1668] Vishu Krishnamurthy, Sandeepan Banerjee, and Anil Nori. Bringing object-relational technology to mainstream. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):513–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brown:1999:ISS

[1669] Paul Brown. Implementing the spirit of SQL-99. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):515–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Baulier:1999:DSM

[1670] J. Baulier, P. Bohannon, S. Gogate, C. Gupta, S. Haldar, S. Joshi, A. Khivesera, H. Korth, P. McIlroy, J. Miller, P. P. S. Narayan, M. Nemeth, R. Rastogi, S. Seshardi, A. Silberschatz, S. Sudarshan, M. Wilder, and C. Wei. DataBlitz storage manager: Main memory database performance for critical applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):519– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kanth:1999:IMD

[1671] K. V. Ravi Kanth, Siva Ravada, Jayant Sharma, and Jay Banerjee. Indexing medium-dimensionality data in oracle. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):521–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kohler:1999:EIS

[1672] Walt Kohler. EMC information sharing: Direct access to MVS data from Unix and NT. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):523–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Greer:1999:DFG

[1673] Rick Greer. Daytona and the fourthgeneration language cymbal. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(2):525– ??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1999:SOCb

[1674] Anonymous. SIGMOD officers, committees, and awards. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(3):1–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:1999:TAW

[1675] A. Bernstein, C. Dellarocas, and M. Klein. Towards adaptive workflow systems (CSCW-98 workshop report). SIGMOD Record (ACM Special Interest Group on Management of Data), 28(3):7–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Conrad:1999:EFI

[1676] S. Conrad, W. Hasselbring, U. Hohenstien, R. Kutsche, M. Roantree, G. Saake, and F. Saltor. Engineering federated information systems (EFIS ’99 workshop report). SIGMOD Record (ACM Special Interest Group on Management of Data), 28(3):9–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Boulos:1999:CEU

[1677] J. Boulos and K. Ono. Cost estimation of user-defined methods in objectrelational database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(3):22–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenthal:1999:FCV

[1678] A. Rosenthal and E. Sciore. Firstclass views: a key to user-centered computing. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(3):29–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yau:1999:MRD

[1679] S. Yau, H. Leong, D. McLeod, and A. Si. On multi-resolution document transmission in mobile Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(3):37–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:1999:RIP

[1680] R. Snodgrass, R. Agrawal, P. Bernstien, D. Chamberlin, J. Gray, D. Hadede, W. Kim, B. Lindsay, C. Mohan, R. Morrison, R. Reiter, and C. Zaniolo. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(3):43–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ram:1999:DTP

[1681] P. Ram, L. Do, and P. Drew. Distributed transactions in practice. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28(3): 49–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Paplani:1999:DSD

[1682] M. Paplani, J. Wason, A. Dunlop, and D. Nicole. A distributed scientific data archive using the Web, XML and SQL/ MED. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(3):56–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Domenig:1999:OCM

[1683] P. Domenig and K. Dittrich. An overview and classification of mediated query systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(3):63–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gruenwald:1999:DRU

[1684] L. Gruenwald, L. Brown, R. Dirckze, S. Guinepain, C. Sanchez, B. Summers, and S. Vanichayobon. Database research at the University of Oklahoma. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(3):73–??, 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Etzion:1999:RNF

[1685] Opher Etzion. Report on NGITS’99:

The fourth international workshop on next generation information technologies and systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(4):3–6, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gatziu:1999:DMD

[1686] Stella Gatziu, Manfred Jeusfeld, Martin Staudt, and Yannis Vassiliou. Design and management of data warehouses report on the DMDW’99 workshop. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(4):7–10, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hanson:1999:TDD

[1687] Eric N. Hanson and Lloyd X. Noronha.

Timer-driven database triggers and alerters: semantics and a challenge. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(4):11–16, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kempster:1999:DA

[1688] Tim Kempster, Colin Stirling, and Peter Thanisch. Diluting ACID. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28 (4):17–23, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Libkin:1999:SRV

[1689] Leonid Libkin. Some remarks on variable independence, closure, and orthographic dimension in constraint databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(4):24–28, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:1999:VX

[1690] Serge Abiteboul. On views and XML. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(4):30–38, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jacob:1999:FFT

[1691] Kaippallimalil J. Jacob and Dennis Shasha. FinTime: a financial time series benchmark. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(4):42–48, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Musick:1999:PLS

[1692] Ron Musick and Terence Critchlow. Practical lessons in supporting largescale computational science. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28 (4):49–57, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:1999:SPS

[1693] Andrew Eisenberg and Jim Melton. SQLJ part 1: SQL routines using the Java programming language. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28 (4):58–63, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vassiliadis:1999:SLM

[1694] Panos Vassiliadis and Timos Sellis. A survey of logical models for OLAP databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(4):64–69, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:2000:SOCa

[1695] Anonymous. SIGMOD officers, committees, and awards. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):1–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ceri:2000:SSS

[1696] S. Ceri, L. Lalinichenko, M. Kitsuregawa, H. Lu, M. Ozsoyovğlu, R. Snodgrass, and V. Vianu. SIGMOD sister societies. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):4–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:2000:ASD

[1697] H. Jagadish. ACM – SIGMOD digital review. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):16–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roddick:2000:ECD

[1698] J. Roddick, L. Al-Jadir, L. Bertossi, M. Dumas, F. Estrella, H. Gregersen, K. Hornsby, J. Lufter, F. Mandreoli, T. Mannisto, E. Mayol, and L. Wedemeijer. Evolution and change in data management — issues and directions. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):21–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Labrinidis:2000:GDC

[1699] A. Labrinidis and N. Roussopoulos. Generating dynamic content at database-backed Web servers: Cgi-bin vs. mod-perl. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):26–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pourabbas:2000:HRO

[1700] E. Pourabbas and M. Rafanelli. Hierarchies and relative operators in the OLAP environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):32–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vassilakis:2000:OSC

[1701] C. Vassilakis. An optimisation scheme for coalesce / valid time selection operator sequences. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):38–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dong:2000:DPI

[1702] G. Dong and J. Su. Database principles — incremental maintenance of recursive views using relational calculus / SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):44–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Faloutsos:2000:RIP

[1703] C. Faloutsos, A. Levy, P. O’Neil, E. Simon, D. Srivastava, V. Vianu, and G. Weikum. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):52–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Liefke:2000:SPE

[1704] H. Liefke and D. Suciu. Systems and prototypes — an extensible compressor for XML data. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):57–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:2000:SSS

[1705] A. Eisenberg and J. Melton. Standards — SQL standardization: The next steps. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):63–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bonifati:2000:CAF

[1706] A. Bonifati and S. Ceri. Comparative analysis of five XML query languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):68–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Paepcke:2000:BDS

[1707] A. Paepcke, H. Garcia-Molina, G. Rodriguez-Mula, and J. Cho. Beyond document similarity: Understanding valuebased search and browsing technologies. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(1):80–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:2000:MFP

[1708] Jiawei Han, Jian Pei, and Yiwen Yin. Mining frequent patterns without candidate generation. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):1–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Riedel:2000:DMO

[1709] Erik Riedel, Christos Faloutsos, Gregory R. Ganger, and David F. Nagle. Data mining on an OLTP system (nearly) for free. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):13–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shenoy:2000:TCV

[1710] Pradeep Shenoy, Jayant R. Haritsa, S. Sudarshan, Gaurav Bhalotia, Mayank Bawa, and Devavrat Shah. Turbo-charging vertical mining of large databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):22–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:2000:HSL

[1711] David B. Lomet. High speed on-line backup when using logical log operations. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):34–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Labio:2000:ERI

[1712] Wilburt Juan Labio, Janet L. Wiener, Hector Garcia-Molina, and Vlad Gorelik. Efficient resumption of interrupted warehouse loads. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):46–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lakhamraju:2000:LRO

[1713] Mohana K. Lakhamraju, Rajeev Rastogi, S. Seshadri, and S. Sudarshan. On-line reorganization in object databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):58–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aggarwal:2000:FGP

[1714] Charu C. Aggarwal and Philip S. Yu.

Finding generalized projected clusters in high dimensional spaces. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):70–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Palmer:2000:DBS

[1715] Christopher R. Palmer and Christos Faloutsos. Density biased sampling:

An improved method for data mining and clustering. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):82–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Breunig:2000:LID

[1716] Markus M. Breunig, Hans-Peter Kriegel, Raymond T. Ng, and Jorg Sander. LOF: Identifying density-based local outliers. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):93–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zaharioudakis:2000:ACS

[1717] Markos Zaharioudakis, Roberta Cochrane, George Lapis, Hamid Pirahesh, and Monica Urata. Answering complex SQL queries using automatic summary tables. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):105–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cho:2000:SDI

[1718] Junghoo Cho and Hector Garcia-Molina. Synchronizing a database to improve freshness. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):117–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Salem:2000:HRJ

[1719] Kenneth Salem, Kevin Beyer, Bruce Lindsay, and Roberta Cochrane. How to roll a join: Asynchronous incremental view maintenance. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):129–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christophides:2000:WQL

[1720] Vassilis Christophides, Sophie Cluet, and Jerome Simeon. On wrapping query languages and efficient XML integration. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):141–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Liefke:2000:XEC

[1721] Hartmut Liefke and Dan Suciu. XMill:

An efficient compressor for XML data. SIGMOD Record (ACM Special Interest Group on Management of Data), 29 (2):153–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Garofalakis:2000:XSE

[1722] Minos Garofalakis, Aristides Gionis, Rajeev Rastogi, S. Seshadri, and Kyuseok Shim. XTRACT: a system for extracting document type descriptors from XML documents. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):165– ??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Faloutsos:2000:SJS

[1723] Christos Faloutsos, Bernhard Seeger, Agma Traina, and Caetano Traina, Jr. Spatial join selectivity using power laws. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):177–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Corral:2000:CPQ

[1724] Antonio Corral, Yannis Manolopoulos, Yannis Theodoridis, and Michael Vassilakopoulos. Closest pair queries in spatial databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):189–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Korn:2000:ISB

[1725] Flip Korn and S. Muthukrishnan. Influence sets based on reverse nearest neighbor queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):201–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rodriguez-Martinez:2000:MSE

[1726] Manuel Rodriguez-Martinez and Nick Roussopoulos. MOCHA: a selfextensible database middleware system for distributed data sources. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):213– ??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:2000:TST

[1727] Mong Li Lee, Masaru Kitsuregawa, Beng Chin Ooi, Kian-Lee Tan, and Anirban Mondal. Towards self-tuning data placement in parallel database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):225–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Litwin:2000:LRH

[1728] Witold Litwin and Thomas Schwarz.

LH*/RS: a high-availability scalable distributed data structure using Reed Solomon codes. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):237–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roy:2000:EEA

[1729] Prasan Roy, S. Seshadri, S. Sudarshan, and Siddhesh Bhobe. Efficient and extensible algorithms for multi query optimization. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):249–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Avnur:2000:ECA

[1730] Ron Avnur and Joseph M. Hellerstein.

Eddies: Continuously adaptive query processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):261–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Popa:2000:CTF

[1731] Lucian Popa, Alin Deutsch, Arnaud Sahuguet, and Val Tannen. A chase too far? SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):273–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goldman:2000:WDP

[1732] Roy Goldman and Jennifer Widom.

WSQ/DSQ: a practical approach for combined querying of databases and the Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):285–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:2000:FEC

[1733] Rakesh Agrawal and Edward L. Wimmers. A framework for expressing and combining preferences. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):297– ??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barclay:2000:MTS

[1734] Tom Barclay, Jim Gray, and Don Slutz. Microsoft TerraServer: a spatial data warehouse. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):307–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Forlizzi:2000:DMD

[1735] Luca Forlizzi, Ralf Hartmut Guting, Enrico Nardelli, and Markus Schneider. A data model and data structures for moving objects databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):319– ??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Saltenis:2000:IPC

[1736] Simonas Saltenis, Christian S. Jensen, Scott T. Leutenegger, and Mario A. Lopez. Indexing the positions of continuously moving objects. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):331– ??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shin:2000:AMS

[1737] Hyoseop Shin, Bongki Moon, and Sukho Lee. Adaptive multi-stage distance join processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):343–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cho:2000:FRW

[1738] Junghoo Cho, Narayanan Shivakumar, and Hector Garcia-Molina. Finding replicated Web collections. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):355– ??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Labrinidis:2000:WVM

[1739] Alexandros Labrinidis and Nick Roussopoulos. Web view materialization. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):367–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:2000:NSC

[1740] Jianjun Chen, David J. DeWitt, Feng Tian, and Yuan Wang. NiagaraCQ: a scalable continuous query system for Internet databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):379–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:2000:OTI

[1741] Yuan-Chi Chang, Lawrence Bergman, Vittorio Castelli, Chung-Sheng Li, Ming-Ling Lo, and John R. Smith. The onion technique: Indexing for linear optimization queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):391–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:2000:EMD

[1742] H. V. Jagadish, Nick Koudas, and Divesh Srivastava. On effective multidimensional indexing for strings. SIG-MOD Record (ACM Special Interest Group on Management of Data), 29(2): 403–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Oh:2000:ECE

[1743] JungHwan Oh and Kien A. Hua. Efficient and cost-effective techniques for browsing and indexing large video databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):415–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramaswamy:2000:EAM

[1744] Sridhar Ramaswamy, Rajeev Rastogi, and Kyuseok Shim. Efficient algorithms for mining outliers from large data sets. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):427–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:2000:PPD

[1745] Rakesh Agrawal and Ramakrishnan Srikant. Privacy-preserving data mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):439–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Szalay:2000:DMM

[1746] Alexander S. Szalay, Peter Z. Kunszt, Ani Thakar, Jim Gray, Don Slutz, and Robert J. Brunner. Designing and mining multi-terabyte astronomy archives: The Sloan Digital Sky Survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):451–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gunopulos:2000:AMD

[1747] Dimitrios Gunopulos, George Kollios, Vassilis J. Tsotras, and Carlotta Domeniconi. Approximating multidimensional aggregate range queries over real attributes. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):463–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rao:2000:MBT

[1748] Jun Rao and Kenneth A. Ross. Making B^{+}-trees cache conscious in main memory. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):475–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Acharya:2000:CSA

[1749] Swarup Acharya, Phillip B. Gibbons, and Viswanath Poosala. Congressional samples for approximate answering of group-by queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):487–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Waas:2000:CES

[1750] Florian Waas and Cesar Galindo-Legaria. Counting, enumerating, and sampling of execution plans in a costbased query optimizer. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):499– ??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wattez:2000:BQT

[1751] Fanny Wattez, Sophie Cluet, Veronique Benzaken, Guy Ferran, and Christian Fiegel. Benchmarking queries over trees: Learning the hard truth the hard way. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):510–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lehner:2000:MCA

[1752] Wolfgang Lehner, Richard Sidle, Hamid Pirahesh, and Roberta Cochrane. Maintenance of cube automatic summary tables. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):512–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Huber:2000:CAD

[1753] Val Huber. Challenges in automating declarative business rules to enable rapid business response. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):514– ??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2000:EBR

[1754] Ronald G. Ross. Expressing business rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):515–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kintzer:2000:GBP

[1755] Eric Kintzer. Going beyond personalization: Rule engines at work. SIG-MOD Record (ACM Special Interest Group on Management of Data), 29(2): 517–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hsiao:2000:DTR

[1756] Hui-I Hsiao and Inderpal Narang.

DLFM: a transactional resource manager. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):518–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ponnekanti:2000:OIR

[1757] Nagavamsi Ponnekanti and Hanuma Kodavalla. Online index rebuild. SIG-MOD Record (ACM Special Interest Group on Management of Data), 29(2): 529–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Annamalai:2000:IIO

[1758] Melliyal Annamalai, Rajiv Chopra, Samuel DeFazio, and Susan Mavris. Indexing images in Oracle8i. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):539– ??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Weininger:2000:HVL

[1759] Andreas Weininger. Handling very large databases with Informix Extended Parallel Server. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):548– ??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:2000:ITW

[1760] Chung-Min Chen, Munir Cochinwala, Claudio Petrone, Marc Pucci, Sunil Samtani, Patrizia Santa, and Marco Mesiti. Internet traffic warehouse. SIG-MOD Record (ACM Special Interest Group on Management of Data), 29(2): 550–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ordonez:2000:SFC

[1761] Carlos Ordonez and Paul Cereghini. SQLEM: Fast clustering in SQL using the EM algorithm. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):559–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jhingran:2000:ARC

[1762] Anant Jhingran. Anatomy of a real E-commerce system. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(2):571–??, 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kant:2000:WPA

[1763] Krishna Kant and Prasant Mohapatra. Workshop on performance and architecture of Web servers. SIG-MOD Record (ACM Special Interest Group on Management of Data), 29 (3):12–14, September 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ludwig:2000:RIC

[1764] Heiko Ludwig and Paul Grefen. Report on ISDO ’00: the CAiSE*00 workshop on “Infrastructures for Dynamic Business-to-Business Service Outsourcing”. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(3):15–18, September 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wu:2000:RSI

[1765] Kun-Lung Wu and Philip S. Yu. Report on Second International Workshop on Advanced Issues of E-commerce and Web-based information systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 29 (3):19–23, September 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Arpinar:2000:PMS

[1766] Sena Arpinar and Asuman Dogac. Provision of market services for eCo compliant electronic marketplaces. SIG-MOD Record (ACM Special Interest Group on Management of Data), 29 (3):24–30, September 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Clementini:2000:SO

[1767] Eliseo Clementini and Paolino Di Felice. Spatial operators. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(3):31–38, September 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Theodoridis:2000:GSD

[1768] Yannis Theodoridis and Mario A. Nascimento. Generating spatiotemporal datasets on the WWW. SIG-MOD Record (ACM Special Interest Group on Management of Data), 29 (3):39–43, September 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

VandenBussche:2000:CDT

[1769] Jan Van den Bussche. Constraint databases: a tutorial introduction. SIGMOD Record (ACM Special Interest Group on Management of Data), 29 (3):44–51, September 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Westmann:2000:IPC

[1770] Till Westmann, Donald Kossmann, Sven Helmer, and Guido Moerkotte. The implementation and performance of compressed databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(3):55–67, September 2000. CODEN SRECD8.

ISSN 0163-5808 (print), 1943-5835 (electronic).

Vetterli:2000:MSD

[1771] Thomas Vetterli, Anca Vaduva, and Martin Staudt. Metadata standards for data warehousing: open information model vs. common warehouse metadata. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(3):68–75, September 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:2000:CAS

[1772] Dongwon Lee and Wesley W. Chu.

Comparative analysis of six XML schema languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(3):76–87, September 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Palpanas:2000:KDD

[1773] Themistoklis Palpanas. Knowledge discovery in data warehouses. SIG-MOD Record (ACM Special Interest Group on Management of Data), 29 (3):88–100, September 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Casati:2000:RVW

[1774] Fabio Casati, Umesh Dayal, and Ming-Chien Shan. Report on the VLDB workshop on technologies for e-services (TES). SIGMOD Record (ACM Special Interest Group on Management of Data), 29(4):11–15, December 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hasselbring:2000:RPF

[1775] W. Hasselbring, W.-J. van den Heuvel, G. J. Houben, R.-D. Kutsche, B. Rieger, M. Roantree, and K. Subieta. Research and practice in federated information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(4):16–18, December 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Possas:2000:UQI

[1776] B. Pôssas, M. Carvalho, R. Resende, and W. Meita. Using quantitative information for efficient association rule generation. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(4):19–25, December 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rashid:2000:ODE

[1777] Awais Rashid and Peter Sawyer. Object database evolution using separation of concerns. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(4):26–33, December 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yuen:2000:CIS

[1778] Joe Chun-Hung Yuen, Edward Chan, Kam-Yiu Lam, and H. W. Leung. Cache invalidation scheme for mobile computing systems with real-time data. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(4):34–39, December 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Halevy:2000:TAQ

[1779] Alon Y. Halevy. Theory of answering queries using views. SIG-MOD Record (ACM Special Interest Group on Management of Data), 29 (4):40–47, December 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jhingran:2000:MFC

[1780] Anant Jhingran. Moving up the food chain: supporting e-commerce applications on databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(4):50–54, December 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:2000:VMC

[1781] Phillip A. Bernstein, Alon Y. Halevy, and Rachel A. Pottinger. A vision for management of complex models. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(4):55–63, December 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Poess:2000:NTB

[1782] Meikel Poess and Chris Floyd. New TPC benchmarks for decision support and Web commerce. SIGMOD Record (ACM Special Interest Group on Management of Data), 29(4):64–71, December 2000. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Biancheri:2001:EDW

[1783] Chiara Biancheri, Jean-Christophe Pazzaglia, and Gavino Paddeu. EIHA?!?: deploying Web and WAP services using XML technology. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):5–12, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aggarwal:2001:RDD

[1784] Charu C. Aggarwal. Re-designing distance functions and distance-based applications for high dimensional data. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):13–18, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lawder:2001:QMD

[1785] J. K. Lawder and P. J. H. King. Querying multi-dimensional data indexed using the Hilbert space-filling curve. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):19–24, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pu:2001:IPS

[1786] Calton Pu, Karsten Schwan, and Jonathan Walpole. Infosphere project: system support for information flow applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):25–34, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bertino:2001:QSM

[1787] Elisa Bertino, Ahmed K. Elmagarmid, and Mohand-Saïd Hacid. Quality of service in multimedia digital libraries. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):35–40, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Feng:2001:TKB

[1788] Ling Feng, Marfred A. Jeusfeld, and Jeroen Hoppenbrouwers. Towards knowledge-based digital libraries. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (1):41–46, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buneman:2001:CSD

[1789] Peter Buneman, Wenfei Fan, Jérôme Simééon, and Scott Weinstein. Constraints for semistructured data and XML. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):47–54, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2001:RIPa

[1790] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):55–57, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Critchlow:2001:RXX

[1791] Terence Critchlow. Report on XEWA-00: the XML enabled wide-area searches for bioinformatics workshop. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):58–61, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carmel:2001:XIR

[1792] David Carmel, Yoelle Maarek, and Aya Soffer. XML and information retrieval: a SIGIR 2000 workshop. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):62–65, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bosc:2001:RFI

[1793] Patrick Bosc, Amihai Motro, and Gabriella Pasi. Report on the Fourth International Conference on Flexible Query Answering Systems (FQAS 2000). SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):66–69, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Melton:2001:SME

[1794] Jim Melton, Jan-Eike Michels, Vanja Josifovski, Krishna Kulkarni, Peter Schwarz, and Kathy Zeidenstein. SQL and management of external data. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):70–77, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Miller:2001:CPM

[1795] Renée J. Miller, Mauricio A. Hernández, Laura M. Haas, Lingling Yan, C. T. Howard Ho, Ronald Fagin, and Lucian Popa. The Clio project: managing heterogeneity. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(1):78–83, March 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:2001:ECI

[1796] Jiawei Han, Jian Pei, Guozhu Dong, and Ke Wang. Efficient computation of Iceberg cubes with complex measures. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):1–12, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gehrke:2001:CCA

[1797] Johannes Gehrke, Flip Korn, and Divesh Srivastava. On computing correlated aggregates over continual data streams. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):13–24, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ng:2001:ICC

[1798] Raymond T. Ng, Alan Wagner, and Yu Yin. Iceberg-cube computation with PC clusters. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):25–36, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aggarwal:2001:ODH

[1799] Charu C. Aggarwal and Philip S.

Yu. Outlier detection for high dimensional data. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):37–46, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rinfret:2001:BSI

[1800] Denis Rinfret, Patrick O’Neil, and Elizabeth O’Neil. Bit-sliced index arithmetic. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):47–57, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Greenwald:2001:SEO

[1801] Michael Greenwald and Sanjeev Khanna. Space-efficient online computation of quantile summaries. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):58–66, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ipeirotis:2001:PCC

[1802] Panagiotis G. Ipeirotis, Luis Gravano, and Mehran Sahami. Probe, count, and classify: categorizing hidden Web databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):67–78, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Breunig:2001:DBQ

[1803] Markus M. Breunig, Hans-Peter Kriegel, Peer Kröger, and Jörg Sander. Data bubbles: quality preserving performance boosting for hierarchical clustering. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):79–90, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Joshi:2001:MNH

[1804] Mahesh V. Joshi, Ramesh C. Agarwal, and Vipin Kumar. Mining needle in a haystack: classifying rare classes via two-phase rule induction. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):91–102, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fernandez:2001:EEX

[1805] Mary Fernandez, Atsuyuki Morishima, and Dan Suciu. Efficient evaluation of XML middle-ware queries. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (2):103–114, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fabret:2001:FAI

[1806] Françoise Fabret, H. Arno Jacobsen, François Llirbat, João Pereira, Kenneth A. Ross, and Dennis Shasha. Filtering algorithms and implementation for very fast publish/subscribe systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):115–126, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Slivinskas:2001:AQO

[1807] Giedrius Slivinskas, Christian S.

Jensen, and Richard T. Snodgrass. Adaptable query optimization and evaluation in temporal middleware. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):127–138, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kim:2001:OMI

[1808] Kihong Kim, Sang K. Cha, and Keunjoo Kwon. Optimizing multidimensional index trees for main memory access. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):139–150, June 2001.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keogh:2001:LAD

[1809] Eamonn Keogh, Kaushik Chakrabarti, Michael Pazzani, and Sharad Mehrotra. Locally adaptive dimensionality reduction for indexing large time series databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):151–162, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bohannon:2001:MMI

[1810] Philip Bohannon, Peter Mcllroy, and Rajeev Rastogi. Main-memory index structures with fixed-size partial keys. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):163–174, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Borkar:2001:AST

[1811] Vinayak Borkar, Kaustubh Deshmukh, and Sunita Sarawagi. Automatic segmentation of text into structured records. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):175–186, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yu:2001:EEM

[1812] Clement Yu, Weiyi Meng, Wensheng Wu, and King-Lup Liu. Efficient and effective metasearch for text databases incorporating linkages among documents. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):187–198, June 2001.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deshpande:2001:IGD

[1813] Amol Deshpande, Minos Garofalakis, and Rajeev Rastogi. Independence is good: dependency-based histogram synopses for high-dimensional data. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):199–210, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bruno:2001:SMW

[1814] Nicolas Bruno, Surajit Chaudhuri, and Luis Gravano. STHoles: a multidimensional workload-aware histogram. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):211–222, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:2001:GOH

[1815] H. V. Jagadish, Hui Jin, Beng Chin Ooi, and Kian-Lee Tan. Global optimization of histograms. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):223– 234, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:2001:IIP

[1816] Shimin Chen, Phillip B. Gibbons, and Todd C. Mowry. Improving index performance through prefetching. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (2):235–246, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gionis:2001:ETS

[1817] Aristides Gionis, Dimitrios Gunopulos, and Nick Koudas. Efficient and tumble similar set retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):247–258, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hristidis:2001:PSE

[1818] Vagelis Hristidis, Nick Koudas, and Yannis Papakonstantinou. PREFER: a system for the efficient execution of multi-parametric ranked queries. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (2):259–270, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:2001:QOC

[1819] Zhiyuan Chen, Johannes Gehrke, and Flip Korn. Query optimization in compressed database systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (2):271–282, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Babu:2001:SMB

[1820] Shivnath Babu, Minos Garofalakis, and Rajeev Rastogi. SPARTAN: a modelbased semantic compression system for massive data tables. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):283–294, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:2001:ROB

[1821] Surajit Chaudhuri, Gautam Das, and Vivek Narasayya. A robust, optimization-based approach for approximate answering of aggregate queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):295–306, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mistry:2001:MVS

[1822] Hoshi Mistry, Prasan Roy, S. Sudarshan, and Krithi Ramamritham. Materialized view selection and maintenance using multi-query optimization. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):307–318, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Afrati:2001:GEP

[1823] Foto N. Afrati, Chen Li, and Jeffrey D.

Ullman. Generating efficient plans for queries using views. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):319–330, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goldstein:2001:OQU

[1824] Jonathan Goldstein and Per-Åke Larson. Optimizing queries using materialized views: a practical, scalable solution. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):331–342, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:2001:DBA

[1825] Sang-Ho Lee, Kyu-Young Whang, Yang-Sae Moon, and Il-Yeol Song. Dynamic buffer allocation in video-on-demand systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):343–354, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Olston:2001:APS

[1826] Chris Olston, Boon Thau Loo, and Jennifer Widom. Adaptive precision setting for cached approximate values. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):355–366, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kalnis:2001:PSA

[1827] Panos Kalnis and Dimitris Papadias. Proxy-server architectures for OLAP. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):367–378, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bohm:2001:EGO

[1828] Christian Böhm, Bernhard Braunmüller, Florian Krebs, and Hans-Peter Kriegel. Epsilon grid order: an algorithm for the similarity join on massive highdimensional data. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):379–388, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lang:2001:MHD

[1829] Christian A. Lang and Ambuj K. Singh. Modeling high-dimensional index structures using sampling. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (2):389–400, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lazaridis:2001:PAA

[1830] Iosif Lazaridis and Sharad Mehrotra. Progressive approximate aggregate queries with a multi-resolution tree structure. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):401–412, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tatarinov:2001:UX

[1831] Igor Tatarinov, Zachary G. Ives, Alon Y. Halevy, and Daniel S. Weld. Updating XML. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):413–424, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhang:2001:SCQ

[1832] Chun Zhang, Jeffrey Naughton, David DeWitt, Qiong Luo, and Guy Lohman. On supporting containment queries in relational database management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):425–436, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nguyen:2001:MXD

[1833] Benjamin Nguyen, Serge Abiteboul, Grégory Cobena, and Mihaí Preda. Monitoring XML data on the Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):437–448, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wu:2001:AGR

[1834] Yi-Leh Wu, Divyakant Agrawal, and Amr El Abbadi. Applying the golden rule of sampling for query estimation. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):449–460, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Getoor:2001:SEU

[1835] Lise Getoor, Benjamin Taskar, and Daphne Koller. Selectivity estimation using probabilistic models. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (2):461–472, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schuster:2001:CED

[1836] Assaf Schuster and Ran Wolff. Communication-efficient distributed mining of association rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):473– 484, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yan:2001:DDU

[1837] Ling Ling Yan, Renée J. Miller, Laura M. Haas, and Ronald Fagin. Data-driven understanding and refinement of schema mappings. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (2):485–496, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Amer-Yahia:2001:MTP

[1838] Sihem Amer-Yahia, SungRan Cho, Laks V. S. Lakshmanan, and Divesh Srivastava. Minimization of tree pattern queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):497–508, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Doan:2001:RSD

[1839] AnHai Doan, Pedro Domingos, and Alon Y. Halevy. Reconciling schemas of disparate data sources: a machinelearning approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):509–520, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carino:2001:SMN

[1840] Felipe Cariño Jr., Pekka Kostamaa, Art Kaufmann, and John Burgess. StorHouse metanoia — new applications for database, storage & data warehousing. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):521–531, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Candan:2001:EDC

[1841] K. Selçuk Candan, Wen-Syan Li, Qiong Luo, Wang-Pin Hsiung, and Divyakant Agrawal. Enabling dynamic content caching for database-driven Web sites. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):532–543, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Navas:2001:NDD

[1842] Julio C. Navas and Michael Wynblatt. The network is the database: data management for highly distributed systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):544–551, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:2001:CIB

[1843] Michael Stonebraker and Joseph M. Hellerstein. Content integration for e-business. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):552–560, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Maguire:2001:CMW

[1844] Thomas Maguire. Catalog management in Websphere Commerce Suite. SIGMOD Record (ACM Special Interest Group on Management of Data), 30 (2):561, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nazeri:2001:EMA

[1845] Zohreh Nazeri, Eric Bloedorn, and Paul Ostwald. Experiences in mining aviation safety data. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):562–566, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Draper:2001:NIE

[1846] Denise Draper, Alon Y. Halevy, and Daniel S. Weld. The nimble integration engine. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):567–568, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Meseck:2001:DML

[1847] Reed M. Meseck. Data management:

lasting impact on wild, wild, web. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):569–570, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Galindo-Legaria:2001:OOS

[1848] César Galindo-Legaria and Milind Joshi. Orthogonal optimization of subqueries and aggregation. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):571– 581, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Godfrey:2001:ECL

[1849] Parke Godfrey, Jarek Gryz, and Calisto Zuzarte. Exploiting constraintlike data characterizations in query optimization. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):582–592, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lahiri:2001:FSQ

[1850] Tirthankar Lahiri, Amit Ganesh, Ron Weiss, and Ashok Joshi. Fast-Start: quick fault recovery in Oracle. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (2):593–598, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deolasee:2001:DDD

[1851] Pavan Deolasee, Amol Katkar, Ankur Panchbudhe, Krithi Ramamritham, and Prashant Shenoy. Dissemination of dynamic data. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):599, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kiessling:2001:CYS

[1852] Werner Kießling, Stefan Holland, Stefan Fischer, and Thorsten Ehm. COSIMA — your smart, speaking Esalesperson. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):600, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kiesssling:2001:CYS

[1853] Werner Kiesßling, Stefan Holland, Stefan Fischer, and Thorsten Ehm. COSIMA — your smart, speaking Esalesperson. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):600, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shou:2001:RRT

[1854] L. Shou, C. H. Chionh, Z. Huang, Y. Ruan, and Kian-Lee Tan. RE-VIEW: a real-time virtual walkthrough system. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):601, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sahuguet:2001:KMT

[1855] Arnaud Sahuguet. Kweelt: more than just “Yet Another Framework to Query XML!”. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):602, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chawathe:2001:VES

[1856] Sudarshan S. Chawathe, Thomas Baby, and Jihwang Yoo. VQBD: exploring semistructured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):603, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buttler:2001:OMS

[1857] David Buttler, Ling Liu, Calton Pu, Henrique Paques, Wei Han, and Wei Tang. OminiSearch: a method for searching dynamic content on the Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):604, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bertino:2001:SXD

[1858] Elisa Bertino, Silvana Castano, and Elena Ferrari. Securing XML documents: the Author-X project demonstration. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):605, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Claypool:2001:SSS

[1859] Kajal T. Claypool, Elke A. Rundensteiner, Xin Zhang, Su Hong, Harumi Kuno, Wang chien Lee, and Gail Mitchell. Sangam — a solution to support multiple data models, their mappings and maintenance. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):606, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hernandez:2001:CSA

[1860] Mauricio A. Hernández, Renée J. Miller, and Laura M. Haas. Clio: a semi-automatic tool for schema mapping. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):607, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:2001:MVI

[1861] Sanjay Agrawal, Surajit Chaudhuri, and Vivek Narasayya. Materialized view and index selection tool for Microsoft SQL Server 2000. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):608, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Catarci:2001:PDS

[1862] Tiziana Catarci and Giuseppe Santucci. The prototype of the DARE system. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):609, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adii:2001:MBP

[1863] Asaf Adii, David Botzer, Opher Etzion, and Tali Yatzkar-Haham. Monitoring business processes through event correlation based on dependency model. SIGMOD Record (ACM Special Interest Group on Management of Data), 30 (2):610, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shah:2001:FTL

[1864] Mehul A. Shah and Sirish Chandrasekaran. Fault-tolerant, loadbalancing queries in telegraph. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30(2): 611, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agichtein:2001:SPS

[1865] Eugene Agichtein, Luis Gravano, Jeff Pavel, Viktoriya Sokolova, and Aleksandr Voskoboynik. Snowball: a prototype system for extracting relations from large text collections. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):612, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:2001:PPB

[1866] Edward Chang, Kwang-Ting Cheng, and Lihyuarn L. Chang. PBIR — perception-based image retrieval. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30(2): 613, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kriegel:2001:SDM

[1867] Hans-Peter Kriegel, Andreas Müller, Marco Pötke, and Thomas Seidl. Spatial data management for computer aided design. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):614, June 2001.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lam:2001:RRT

[1868] Kam-Yiu Lam, Edward Chan, Tei-Wei Kuo, S. W. Ng, and Dick Hung. RETINA: a real-time traffic navigation system. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):615, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Datta:2001:DCA

[1869] Anindya Datta, Kaushik Dutta, Krithi Ramamritham, Helen Thomas, and Debra VanderMeer. Dynamic content acceleration: a caching solution to enable scalable dynamic Web page generation. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):616, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Whitney:2001:LOR

[1870] Arthur Whitney and Dennis Shasha. Lots o’Ticks: real time high performance time series queries on billions of trades and quotes. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):617, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:2001:DMS

[1871] Jiawei Han, Hasan Jamil, Ying Lu, Liangyou Chen, Yaqin Liao, and Jian Pei. DNA-miner: a system prototype for mining DNA sequences. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):618, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:2001:DDW

[1872] Jun Chen, Xin Zhang, Songting Chen, Andreas Koeller, and Elke A. Rundensteiner. DyDa: data warehouse maintenance in fully concurrent environments. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):619, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Larson:2001:XDM

[1873] Per-Åke Larson, Dana Florescu, Goetz Graefe, Guido Moerkotte, Hamid Pirahesh, and Harald Schöning. XML data management (panel session): go native or spruce up relational systems? SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):620, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenthal:2001:WDR

[1874] Arnon Rosenthal, Klaus Dittrich, Jim Donahue, and Bill Maimone. Will database researchers have any role in data security? (panel session). SIG-MOD Record (ACM Special Interest Group on Management of Data), 30(2): 621, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mohan:2001:ASP

[1875] C. Mohan, Larry Cable, Matthieu Devin, Scott Dietzen, Pat Helland, and Dan Wolfson. Application servers (panel session): born-again TP monitors for the Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):622, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Haas:2001:OQP

[1876] Peter J. Haas and Joseph M. Hellerstein. Online query processing: a tutorial. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):623, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gunopulos:2001:TSS

[1877] Dimitrios Gunopulos and Gautam Das. Time series similarity measures and time series indexing. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):624, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bussler:2001:SBI

[1878] Christoph Bussler. Semantic B2B integration. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):625, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Casati:2001:MLD

[1879] Fabio Casati and Ming-Chien Shan. Models and languages for describing and discovering E-services. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):626, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Smith:2001:SMD

[1880] John R. Smith. Standard for multimedia databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(2):627, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2001:AXD

[1881] Karl Aberer. Advanced XML data processing: guest editor’s introduction. SIGMOD Record (ACM Special Interest Group on Management of Data), 30 (3):11–13, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fankhauser:2001:XFS

[1882] Peter Fankhauser. XQuery formal semantics state and challenges. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (3):14–19, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shanmugasundaram:2001:GTQ

[1883] Jayavel Shanmugasundaram, Eugene Shekita, Jerry Kiernan, Rajasekar Krishnamurthy, Efstratios Viglas, Jeffrey Naughton, and Igor Tatarinov. A general technique for querying XML documents using a relational database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(3):20–26, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schmidt:2001:WHB

[1884] Albrecht Schmidt, Florian Waas, Martin Kersten, Daniela Florescu, Michael J. Carey, Ioana Manolescu, and Ralph Busse. Why and how to benchmark XML databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (3):27–32, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Han:2001:WWD

[1885] Wei Han, David Buttler, and Calton Pu. Wrapping Web data into XML. SIGMOD Record (ACM Special Interest Group on Management of Data), 30 (3):33–38, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Suciu:2001:DTX

[1886] Dan Suciu. On database theory and XML. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(3):39–45, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chien:2001:XDV

[1887] Shu Yao Chien, Vassilis J. Tsotras, and Carlo Zaniolo. XML document versioning. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(3):46–53, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ludascher:2001:PDD

[1888] Bertram Ludäscher, Richard Marciano, and Reagan Moore. Preservation of digital data with self-validating, selfinstantiating knowledge-based archives. SIGMOD Record (ACM Special Interest Group on Management of Data), 30 (3):54–63, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lomet:2001:EEB

[1889] David Lomet. The evolution of effective B-tree: page organization and techniques: a personal account. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (3):64–69, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bouganim:2001:EPD

[1890] Luc Bouganim, Maria Claudia Cavalcanti, Françoise Fabret, Maria Luiza Campos, François Llirbat, Marta Mattoso, Rubens Melo, Ana Maria Moura, Esther Pacitti, Fabio Porto, Margareth Simoes, Eric Simon, Asterio Tanaka, and Patrick Valduriez. The Ecobase project: database and Web technologies for environmental information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(3):70–75, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Huyn:2001:DAM

[1891] Nam Huyn. Data analysis and mining in the life sciences. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(3):76–85, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2001:RIPb

[1892] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(3):101–104, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Babu:2001:CQD

[1893] Shivnath Babu and Jennifer Widom. Continuous queries over data streams. SIGMOD Record (ACM Special Interest Group on Management of Data), 30 (3):109–120, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Labrinidis:2001:CES

[1894] Alexandros Labrinidis and Alberto O. Mendelzon. Career-enhancing services at SIGMOD online. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (3):121–126, September 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barbara:2001:SID

[1895] Daniel Barbará. Special issue on data mining for intrusion detection and threat analysis. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):4, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stolfo:2001:DMB

[1896] Salvatore J. Stolfo, Wenke Lee, Philip K. Chan, Wei Fan, and Eleazar Eskin. Data mining-based intrusion detectors: an overview of the Columbia IDS Project. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):5–14, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barbara:2001:ATE

[1897] Daniel Barbará, Julia Couto, Sushil Jajodia, and Ningning Wu. ADAM: a testbed for exploring the use of data mining in intrusion detection. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (4):15–24, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cabrera:2001:DCI

[1898] João B. D. Cabrera, Lundy Lewis, and Raman K. Mehra. Detection and classification of intrusions and faults using sequences of system calls. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (4):25–34, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:2001:MSA

[1899] Wenke Lee and Wei Fan. Mining system audit data: opportunities and challenges. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):35–44, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Saygin:2001:UUP

[1900] Yücel Saygin, Vassilios S. Verykios, and Chris Clifton. Using unknowns to prevent discovery of association rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):45–54, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

deVel:2001:MMC

[1901] O. de Vel, A. Anderson, M. Corney, and G. Mohay. Mining e-mail content for author identification forensics. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):55–64, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2001:BRC

[1902] Karl Aberer. Book review column. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):65–66, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Scholl:2001:TIS

[1903] Marc H. Scholl. Transactional information systems: theory, algorithms, and the practice of concurrency control and recovery. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):67–68, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:2001:IWS

[1904] H. V. Jagadish. Information warfare and security. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):69–70, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cardelli:2001:DSD

[1905] Luca Cardelli. Describing semistructured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):80–85, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Melton:2001:SMA

[1906] Jim Melton and Andrew Eisenberg.

SQL multimedia and application packages (SQL/MM). SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):97–102, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shah:2001:JSD

[1907] Mehul A. Shah, Michael J. Franklin, Samuel Madden, and Joseph M. Hellerstein. Java support for data-intensive systems: experiences building the telegraph dataflow system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 30 (4):103–114, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cetintemel:2001:AIC

[1908] Uğur Çetintemel. On the academic interview circuit: an end-to-end discussion. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):116–122, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sheth:2001:RMO

[1909] Amit Sheth, John Miller, Krys Kochut, and Budak Arpinar. Research in multiorganizational processes and semantic information brokering at the LS-DIS lab. SIGMOD Record (ACM Special Interest Group on Management of Data), 30(4):123–128, December 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dogac:2002:DMI

[1910] Asuman Dogac. Data management issues in electronic commerce: guest editor’s introduction. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):3–4, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tian:2002:DPE

[1911] Feng Tian, David J. DeWitt, Jianjun Chen, and Chun Zhang. The design and performance evaluation of alternative XML storage strategies. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):5–10, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cingil:2002:SGU

[1912] Ibrahim Cingil. Supporting global user profiles through trusted authorities. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):11–17, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sommer:2002:THT

[1913] Rainer A. Sommer, Thomas R. Gulledge, and David Bailey. The ntier hub technology. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):18–23, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cilia:2002:AFS

[1914] M. Cilia and A. P. Buchmann. An active functionality service for e-business applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):24–30, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hummer:2002:CDE

[1915] W. Hümmer, W. Lehner, and H. Wedekind. Contracting in the days of eBusiness. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):31–36, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kim:2002:CMS

[1916] HyoungDo Kim. Conceptual modeling and specification generation for B2B business processes based on ebXML. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):37–42, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cui:2002:SBI

[1917] Zhan Cui, Dean Jones, and Paul O’Brien. Semantic B2B integration: issues in ontology-based approaches. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):43–48, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Quix:2002:BDM

[1918] Christoph Quix, Mareike Schoop, and Manfred Jeusfeld. Business data management for business-to-business electronic commerce. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):49–54, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Erdur:2002:MAS

[1919] Riza Cenk Erdur and Oğuz Dikenelli. A multi-agent system infrastructure for software component market-place: an ontological perspective. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):55–60, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ball:2002:SCI

[1920] Michael O. Ball, Meng Ma, Louiqa Raschid, and Zhengying Zhao. Supply chain infrastructures: system integration and information sharing. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):61–66, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bussler:2002:RBE

[1921] Christoph Bussler. The role of B2B engines in B2B integration architectures. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):67–72, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2002:BRCa

[1922] Karl Aberer. Book review column. SIGMOD Record (ACM Special Interest Group on Management of Data), 31 (1):73, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lewis:2002:DTP

[1923] Philip M. Lewis, Arthur Bernstein, and Michael Kifer. Databases and transaction processing: an applicationoriented approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):74–75, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Witten:2002:DMP

[1924] Ian H. Witten and Eibe Frank. Data mining: practical machine learning tools and techniques with Java implementations. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):76–77, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dertouzos:2002:WW

[1925] Michael Dertouzos. What will be.

SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):78–79, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2002:ASS

[1926] Marianne Winslett. Avi Silberschatz speaks out: on academia versus industrial labs, startup fever, database PhD quality, the future of data mining as a research area, IP issues, how he wouldn’t change a thing, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):80–88, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Suciu:2002:XTP

[1927] Dan Suciu. The XML typechecking problem. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):89–96, March 2002.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:2002:TMM

[1928] Michael Stonebraker. Too much middleware. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):97–106, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2002:RIPa

[1929] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):107– 108, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goble:2002:RBN

[1930] Carole Goble and Brian Read. Report on the 18^{th} British National Conference on Databases (BNCOD). SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (1):109–112, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Caplinskas:2002:ADI

[1931] Albertas Caplinskas, Johann Eder, and Olegas Vasilecas. Advances in databases and information systems: report of 5^{th} East European Conference ADBIS’2001. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):113–117, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2002:WLW

[1932] Richard T. Snodgrass. Why I like working in academia. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(1):118– 121, March 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Slivinskas:2002:BOQ

[1933] Giedrius Slivinskas, Christian S.

Jensen, and Richard T. Snodgrass. Bringing order to query optimization. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):5–14, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Orlandic:2002:DRT

[1934] Ratko Orlandic, Jack Lukaszuk, and Craig Swietlik. The design of a retrieval technique for high-dimensional data on tertiary storage. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):15–21, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pourabbas:2002:PQL

[1935] Elaheh Pourabbas and Maurizio Rafanelli. A pictorial query language for querying geographic databases using positional and OLAP operators. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):22–27, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wiegand:2002:IXQ

[1936] Nancy Wiegand. Investigating XQuery for querying across database object types. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):28–33, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kosch:2002:MMD

[1937] Harald Kosch. MPEG-7 and multimedia database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):34–39, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Halkidi:2002:CVM

[1938] Maria Halkidi, Yannis Batistakis, and Michalis Vazirgiannis. Cluster validity methods: part I. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):40–45, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pare:2002:TRR

[1939] Rafael Camps Paré. From ternary relationship to relational tables: a case against common beliefs. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):46–49, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2002:DDS

[1940] Marianne Winslett. David DeWitt speaks out: on rethinking the CS curriculum, why the database community should be proud, why query optimization doesn’t work, how supercomputing funding is sometimes very poorly spent, how he’s not a good coder and isn’t smart enough to do DB theory, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):50–62, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2002:BRCb

[1941] Karl Aberer. Book review column.

SIGMOD Record (ACM Special Interest Group on Management of Data), 31 (2):65, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berzal:2002:DMC

[1942] Fernando Berzal and Nicolfás Matín.

Data mining: concepts and techniques by Jiawei Han and Micheline Kamber. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):66–68, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ouksel:2002:MWW

[1943] Aris Ouksel. Mining the World Wide Web: an information search approach by George Chang, Marcus J. Healey (editor), James A. M. McHugh, Jason T. L. Wang. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):69–70, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goethals:2002:DWB

[1944] Frank G. Goethals. Data warehousing and business intelligence for ecommerce: by Alan R. Simon & Steven L. Shaffer. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):71–72, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Konig-Ries:2002:BIM

[1945] Birgitta König-Ries, Kia Makki, Sam Makki, Charles Perkins, Niki Pissinou, Peter Reiher, Peter Scheuermann, Jari Veijalainen, and Ouri Wolfson. On building an infrastructure for mobile and wireless systems: report on the NSF workshop on an infrastructure for mobile and wireless systems, Oct. 15, 2001. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):73–79, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hammer:2002:RAF

[1946] Joachim Hammer. Report on the ACM fourth international workshop on data warehousing and OLAP (DOLAP 2001). SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):80–83, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Laender:2002:BSW

[1947] Alberto H. F. Laender, Berthier A. Ribeiro-Neto, Altigran S. da Silva, and Juliana S. Teixeira. A brief survey of Web data extraction tools. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):84–93, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barga:2002:PPF

[1948] Roger Barga and David Lomet. Phoenix project: fault-tolerant applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):94–100, June 2002.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:2002:SXM

[1949] Andrew Eisenberg and Jim Melton.

SQL/XML is making good progress. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):101–108, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fagin:2002:CFI

[1950] Ronald Fagin. Combining fuzzy information: an overview. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(2):109–118, June 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hammer:2002:TM

[1951] Joachim Hammer. Treasurer’s message. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(3):5, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yao:2002:CAN

[1952] Yong Yao and Johannes Gehrke.

The Cougar approach to in-network query processing in sensor networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(3):9–18, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL http://www.cs.cornell.edu/
database/cougar/.

Halkidi:2002:CVC

[1953] Maria Halkidi, Yannis Batistakis, and Michalis Vazirgiannis. Clustering validity checking methods: part II. SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (3):19–27, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Huang:2002:SMS

[1954] Shi Ming Huang, Irene Kwan, and Chih He Li. A study on the management of semantic transaction for efficient data retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(3):28–33, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2002:BRCc

[1955] Karl Aberer. Book review column. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(3):34, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuijpers:2002:ICD

[1956] Bart Kuijpers. Introduction to constraint databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(3):35–36, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Weikum:2002:FSN

[1957] Gerhard Weikum. Foundations of statistical natural language processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 31 (3):37–38, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Neven:2002:ATX

[1958] Frank Neven. Automata theory for XML researchers. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(3):39–46, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2002:HGM

[1959] Marianne Winslett. Hector Garcia-Molina speaks out: regarding startups, how life is getting harder, delta papers, CS in Mexico, life as a department chair, and more. SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (3):47–54, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lightstone:2002:TAC

[1960] Sam S. Lightstone, Guy Lohman, and Danny Zilio. Toward autonomic computing with DB2 universal database. SIGMOD Record (ACM Special Interest Group on Management of Data), 31 (3):55–61, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2002:RIPb

[1961] Kenneth A. Ross, Theodore Johnson, and Richard T. Snodgrass. Reminiscences on influential papers. SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (3):62–63, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Melton:2002:SMS

[1962] Jim Melton, Jan Eike Michels, Vanja Josifovski, Krishna Kulkarni, and Peter Schwarz. SQL/MED: a status report. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(3):81–89, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ives:2002:IDT

[1963] Zachary G. Ives and Qiong Luo. Interviewing during a tight job market. SIGMOD Record (ACM Special Interest Group on Management of Data), 31 (3):90–96, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2002:DRU

[1964] M. Winslett, K. Chang, A. Doan, J. Han, C. Zhai, and Y. Zhou. Database research at the University of Illinois at Urbana-Champaign. SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (3):97–102, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cruz:2002:RAD

[1965] Isabel Cruz, Ashfaq Khokhar, Bing Liu, Prasad Sistla, Ouri Wolfson, and Clement Yu. Research activities in database management and information retrieval at University of Illinois at Chicago. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(3):103–108, September 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Spyns:2002:DMV

[1966] Peter Spyns, Robert Meersman, and Mustafa Jarrar. Data modelling versus ontology engineering. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):12–17, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sure:2002:MDE

[1967] York Sure, Steffen Staab, and Rudi Studer. Methodology for development and employment of ontology based knowledge management applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):18–23, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bussler:2002:CAS

[1968] Christoph Bussler, Dieter Fensel, and Alexander Maedche. A conceptual architecture for Semantic Web enabled Web services. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):24–29, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Finin:2002:ATI

[1969] Tim Finin and Anupam Joshi. Agents, trust, and information access on the Semantic Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):30–35, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Maximilien:2002:CMW

[1970] E. Michael Maximilien and Munindar P. Singh. Conceptual model of Web service reputation. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):36–41, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anyanwu:2002:ODR

[1971] Kemafor Anyanwu and Amit Sheth. The ρ operator: discovering and ranking associations on the Semantic Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):42–47, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2002:FSG

[1972] Karl Aberer, Philippe Cudré-Mauroux, and Manfred Hauswirth. A framework for semantic gossiping. SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (4):48–53, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grosky:2002:ESM

[1973] W. I. Grosky, D. V. Sreenath, and F. Fotouhi. Emergent semantics and the multimedia Semantic Web. SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (4):54–58, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buttler:2002:QMB

[1974] David Buttler, Matthew Coleman, Terence Critchlow, Renato Fileto, Wei Han, Calton Pu, Daniel Rocco, and Li Xiong. Querying multiple bioinformatics information sources: can Semantic Web research help? SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (4):59–64, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goble:2002:GAS

[1975] Carole Goble and David De Roure.

The Grid: an application of the Semantic Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):65–70, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2002:BRCd

[1976] Karl Aberer. Book review column.

SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):71, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Badia:2002:ROS

[1977] Antonio Badia. Review of “Ontologies: a silver bullet for knowledge management and electronic commerce” by Dieter Fensel. Springer-Verlag, 2002. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):72–73, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wu:2002:SWD

[1978] Jie Wu and Duncan J. Watts. Small worlds: the dynamics of networks between order and randomness. SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (4):74–75, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grohe:2002:PCD

[1979] Martin Grohe. Parameterized complexity for the database theorist. SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (4):86–96, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sheth:2002:ARD

[1980] Amit Sheth and Robert Meersman. Amicalola report: database and information systems research challenges and opportunities in Semantic Web and enterprises. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):98–106, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Boucelma:2002:REP

[1981] Omar Boucelma, Silvana Castano, Carole Goble, Vanja Josifovski, Zoé Lacroix, and Bertram Ludäscher. Report on the EDBT’02 panel on scientific data integration. SIGMOD Record (ACM Special Interest Group on Management of Data), 31(4):107–112, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:2002:ELX

[1982] Andrew Eisenberg and Jim Melton. An early look at XQuery. SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (4):113–120, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roantree:2002:RIM

[1983] Mark Roantree and Alan F. Smeaton.

Research in information managment at Dublin City University. SIG-MOD Record (ACM Special Interest Group on Management of Data), 31 (4):121–126, December 2002. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

OConnell:2003:PJD

[1984] S. J. O’Connell and N. Winterbottom.

Performing joins without decompression in a compressed database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):6–11, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gottlob:2003:XPNa

[1985] Georg Gottlob, Christoph Koch, and Reinhard Pichler. XPath processing in a nutshell. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):12–19, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:2003:GQL

[1986] Ya-Hui Chang. A graphical query language for mobile information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):20–25, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wen:2003:MPQ

[1987] Ji-Rong Wen, Qing Li, Wei-Ying Ma, and Hong-Jiang Zhang. A multiparadigm querying approach for a generic multimedia database management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):26–34, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bertino:2003:DAQ

[1988] Elisa Bertino, Ahmed K. Elmagarmid, and Mohand-Saïd Hacid. A database approach to quality of service specification in video databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):35–40, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

David:2003:ASH

[1989] Michael M. David. ANSI SQL hierarchical processing can fully integrate native XML. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):41–46, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DeMarchi:2003:AED

[1990] Fabien De Marchi, Stéphane Lopes, Jean-Marc Petit, and Farouk Toumani. Analysis of existing databases at the logical level: the DBA companion project. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):47–52, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2003:JGS

[1991] Marianne Winslett. Jim Gray speaks out. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):53–61, March 2003.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2003:RIPb

[1992] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):62–63, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lacroix:2003:RFI

[1993] Zoé Lacroix and Omar Boucelma. Report on the first international workshop on efficient Web-based information systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):64–68, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2003:RFI

[1994] Karl Aberer. Report on the first international conference on ontologies, databases and applications of semantics (ODBASE): part of the federated conference “On the Move to Meaningful Internet Systems 2002”. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):69–72, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chakravarthy:2003:DRU

[1995] Sharma Chakravarthy, Alp Aslandogan, Ramez Elmasri, Leonidas Fegaras, and JungHwan Oh. Database research at UT Arlington. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(1):73–80, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Golab:2003:IDS

[1996] Lukasz Golab and M. Tamer Özsu. Issues in data stream management. SIG-MOD Record (ACM Special Interest Group on Management of Data), 32(2): 5–14, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jukic:2003:CKL

[1997] Nenad Jukic, Svetlozar Nestorov, and Susan Vrbsky. Closing the key loophole in MLS databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):15–20, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gottlob:2003:XPNb

[1998] Georg Gottlob, Christoph Koch, and Reinhard Pichler. XPath processing in a nutshell. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):21–27, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mukhopadhyay:2003:ACB

[1999] Debajyoti Mukhopadhyay, Debasis Giri, and Sanasam Ranbir Singh. An approach to confidence based page ranking for user oriented Web search. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):28–33, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

dOnofrio:2003:MTT

[2000] Alberto d’Onofrio and Elaheh Pourabbas. Modelling temporal thematic map contents. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):34–41, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fife:2003:RID

[2001] Leslie D. Fife and Le Gruenwald. Research issues for data communication in mobile ad-hoc network database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):42–47, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Calero:2003:TDB

[2002] Coral Calero, Mario Piattini, and Francisco Ruiz. Towards a database body of knowledge: a study from Spain. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):48–53, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2003:BRCa

[2003] Karl Aberer. Book review column.

SIGMOD Record (ACM Special Interest Group on Management of Data), 32 (2):54, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hoffner:2003:FDW

[2004] Vernon Hoffner. Fundamentals of data warehouses: 2^{nd} revised and extended edition. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):55–56, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nascimento:2003:PPH

[2005] Mario A. Nascimento. Peer-to-peer: harnessing the power of disruptive technologies. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):57–58, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stirling:2003:IR

[2006] Dale A. Stirling. Information rules. SIGMOD Record (ACM Special Interest Group on Management of Data), 32 (2):59, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vianu:2003:WOC

[2007] Victor Vianu. A Web odyssey: from Codd to XML. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):68–77, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chong:2003:MMS

[2008] Eugene Inseok Chong, Jagannathan Srinivasan, Souripriya Das, Chuck Freiwald, Aravind Yalamanchi, Mahesh Jagannath, Anh-Tuan Tran, Ramkumar Krishnan, and Richard Jiang. A mapping mechanism to support bitmap index and other auxiliary structures on tables stored as primary B^{+}-trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):78–88, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Campbell:2003:XS

[2009] Charles E. Campbell, Andrew Eisenberg, and Jim Melton. XML schema.

SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):96–101, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ailamaki:2003:TMN

[2010] Anastassia Ailamaki and Johannes Gehrke. Time management for new faculty. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(2):102–106, June 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nascimento:2003:ASC

[2011] Mario A. Nascimento, Jörg Sander, and Jeffrey Pound. Analysis of SIG-MOD’s co-authorship graph. SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (3):8–10, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2003:JR

[2012] Richard Snodgrass. Journal relevance.

SIGMOD Record (ACM Special Interest Group on Management of Data), 32 (3):11–15, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Halevy:2003:LAD

[2013] Alon Y. Halevy. Learning about data integration challenges from day one. SIGMOD Record (ACM Special Interest Group on Management of Data), 32 (3):16–17, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ailamaki:2003:EUS

[2014] Anastassia Ailamaki and Joseph M. Hellerstein. Exposing undergraduate students to database system internals. SIGMOD Record (ACM Special Interest Group on Management of Data), 32 (3):18–20, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2003:GEI

[2015] Karl Aberer. Guest editor’s introduction. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(3):21–22, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bawa:2003:PPR

[2016] Mayank Bawa, Brian F. Cooper, Arturo Crespo, Neil Daswani, Prasanna Ganesan, Hector Garcia-Molina, Sepandar Kamvar, Sergio Marti, Mario Schlosser, Qi Sun, Patrick Vinograd, and Beverly Yang. Peer-to-peer research at Stanford. SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (3):23–28, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2003:PGS

[2017] Karl Aberer, Philippe Cudré-Mauroux, Anwitaman Datta, Zoran Despotovic, Manfred Hauswirth, Magdalena Punceva, and Roman Schmidt. P-Grid: a self-organizing structured P2P system. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(3):29–33, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hellerstein:2003:TND

[2018] Joseph M. Hellerstein. Toward network data independence. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(3):34–40, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nejdl:2003:DIC

[2019] Wolfgang Nejdl, Wolf Siberski, and Michael Sintek. Design issues and challenges for RDF- and schemabased peer-to-peer systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (3):41–46, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tatarinov:2003:PPD

[2020] Igor Tatarinov, Zachary Ives, Jayant Madhavan, Alon Halevy, Dan Suciu, Nilesh Dalvi, Xin (Luna) Dong, Yana Kadiyska, Gerome Miklau, and Peter Mork. The Piazza peer data management project. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(3):47–52, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Arenas:2003:HPD

[2021] Marcelo Arenas, Vasiliki Kantere, Anastasios Kementsietsidis, Iluju Kiringa, Renée J. Miller, and John Mylopoulos. The Hyperion project: from data integration to data coordination. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(3):53–58, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ooi:2003:RDS

[2022] Beng Chin Ooi, Yanfeng Shu, and Kian-Lee Tan. Relational data sharing in peer-based data management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(3):59–64, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ouksel:2003:CPP

[2023] Aris M. Ouksel. In-context peer-to-peer information filtering on the Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 32 (3):65–70, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Koubarakis:2003:SID

[2024] Manolis Koubarakis, Christos Tryfonopoulos, Stratos Idreos, and Yannis Drougas. Selective information dissemination in P2P networks: problems and solutions. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(3):71–76, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pitoura:2003:DSO

[2025] Evaggelia Pitoura, Serge Abiteboul, Dieter Pfoser, George Samaras, and Michalis Vazirgiannis. DBGlobe: a service-oriented P2P system for global computing. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(3):77–82, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Malaika:2003:SDG

[2026] Susan Malaika, Andrew Eisenberg, and Jim Melton. Standards for databases on the grid. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(3):92–100, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anisimov:2003:RDW

[2027] Alexander A. Anisimov. Review of The data warehouse toolkit: the complete guide to dimensional modeling (2nd edition) by Ralph Kimball, Margy Ross. John Wiley & Sons, Inc. 2002. SIGMOD Record (ACM Special Interest Group on Management of Data), 32 (3):101–102, September 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Date:2003:EFC

[2028] C. J. Date. Edgar F. Codd: a tribute and personal memoir. SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (4):4–13, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2003:DAT

[2029] Richard Snodgrass. Developments at ACM TODS. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):14–15, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kumar:2003:SAC

[2030] Vijay Kumar. Sensor: the atomic computing particle. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):16–21, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hellerstein:2003:SST

[2031] Joseph M. Hellerstein, Wei Hong, and Samuel R. Madden. The sensor spectrum: technology, trends, and requirements. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):22–27, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mani:2003:USS

[2032] Murali Mani. Understanding the semantics of sensor data. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):28–34, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bonnet:2003:BBS

[2033] Philippe Bonnet, Allan Beaufour, Mads Bondo Dydensborg, and Martin Leopold. Bluetooth-based sensor networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):35–40, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cheng:2003:MUS

[2034] Reynold Cheng and Sunil Prabhakar. Managing uncertainty in sensor database. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):41–46, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ailamaki:2003:ESN

[2035] Anastassia Ailamaki, Christos Faloutos, Paul S. Fischbeck, Mitchell J. Small, and Jeanne VanBriesen. An environmental sensor network to determine drinking water quality and security. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):47–52, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Demers:2003:CPW

[2036] Alan Demers, Johannes Gehrke, Rajmohan Rajaraman, Niki Trigoni, and Yong Yao. The Cougar Project: a work-in-progress report. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):53–59, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kannan:2003:ERB

[2037] Rajgopal Kannan, Ram Kalidindi, S. S.

Iyengar, and Vijay Kumar. Energy and rate based MAC protocol for wireless sensor networks. SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (4):60–65, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tan:2003:PED

[2038] Hüseyin Özgür Tan and Ibrahim Körpeoğlu. Power efficient data gathering and aggregation in wireless sensor networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):66–71, December 2003.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gerla:2003:MSL

[2039] Mario Gerla and Kaixin Xu. Multimedia streaming in large-scale sensor networks with mobile swarms. SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (4):72–76, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Palpanas:2003:DDD

[2040] Themistoklis Palpanas, Dimitris Papadopoulos, Vana Kalogeraki, and Dimitrios Gunopulos. Distributed deviation detection in sensor networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):77–82, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Calvanese:2003:RRP

[2041] D. Calvanese, G. De Giacomo, M. Lenzerini, and M. Y. Vardi. Reasoning on regular path queries. SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (4):83–92, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2003:BRCb

[2042] Karl Aberer. Book review column. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):106, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:2003:RWC

[2043] Qiang Wang and Brian D. Davison.

Review of Web caching and replication by Michael Rabinovich and Oliver Spatscheck. Addison Wesley 2002. SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (4):107–108, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berzal:2003:RDW

[2044] Fernando Berzal and Nicolás Marín.

Review of Data on the Web: from relational to semistructured data and XML by Serge Abiteboul, Peter Buneman, and Dan Suciu. Morgan Kaufmann 1999. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):109–110, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wiegand:2003:RSD

[2045] Nancy Wiegand. Review of Spatial databases with application to GIS by Philippe Rigaux, Michel Scholl, and Agnes Voisard. Morgan Kaufmann 2002. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):111–112, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lenz:2003:RIW

[2046] Hans J. Lenz, Panos Vassiliadis, Manfred Jeusfeld, and Martin Staudt. Report on the 5^{th} international workshop on the design and management of data warehouses (DMDW’03). SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (4):113–116, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Giorgini:2003:AOS

[2047] Paolo Giorgini. Agent-oriented software engineering report on the 4^{th} AOSE workshop (AOSE 2003). SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (4):117–119, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Motro:2003:RFF

[2048] Amihai Motro and Troels Andreasen. Report on FQAS 2002: fifth international conference on flexible query answering systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):120–121, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kambhampati:2003:IIW

[2049] Subbarao Kambhampati and Craig A. Knoblock. Information integration on the Web: a view from AI and databases (report on IIWeb-03). SIG-MOD Record (ACM Special Interest Group on Management of Data), 32 (4):122–123, December 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hainaut:2003:RDE

[2050] Jean-Luc Hainaut. Research in database engineering at the University of Namur. SIGMOD Record (ACM Special Interest Group on Management of Data), 32(4):124–128, December 2003.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tubaishat:2004:SHM

[2051] Malik Tubaishat, Jian Yin, Biswajit Panja, and Sanjay Madria. A secure hierarchical model for sensor network. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):7–13, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:2004:RKE

[2052] Yongge Wang. Robust key establishment in sensor networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):14–19, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gerla:2004:TCA

[2053] Mario Gerla and Yunjung Yi. Team communications among autonomous sensor swarms. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):20–25, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lazaridis:2004:QQA

[2054] Iosif Lazaridis, Qi Han, Xingbo Yu, Sharad Mehrotra, Nalini Venkatasubramanian, Dmitri V. Kalashnikov, and Weiwen Yang. QUASAR: quality aware sensing architecture. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):26–31, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Park:2004:SGB

[2055] Nam Hun Park and Won Suk Lee. Statistical grid-based clustering over data streams. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):32–37, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hara:2004:RAC

[2056] Takahiro Hara, Norishige Murakami, and Shojiro Nishio. Replica allocation for correlated data items in ad hoc sensor networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):38–43, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deshpande:2004:ISO

[2057] Amol Deshpande. An initial study of overheads of eddies. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):44–49, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Verykios:2004:SAP

[2058] Vassilios S. Verykios, Elisa Bertino, Igor Nai Fovino, Loredana Parasiliti Provenza, Yucel Saygin, and Yannis Theodoridis. State-of-the-art in privacy preserving data mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):50–57, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Helmer:2004:ELB

[2059] Sven Helmer, Carl-Christian Kanne, and Guido Moerkotte. Evaluating lock-based protocols for cooperation on XML documents. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):58–63, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Voruganti:2004:LBD

[2060] Kaladhar Voruganti, Jai Menon, and Sandeep Gopisetty. Land below a DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):64–70, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bolchini:2004:CAM

[2061] C. Bolchini, F. A. Schreiber, and L. Tanca. A context-aware methodology for very small data base design. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):71–76, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Badia:2004:ERM

[2062] Antonio Badia. Entity-relationship modeling revisited. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):77–82, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Martyn:2004:RMD

[2063] Tim Martyn. Reconsidering multidimensional schemas. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):83–88, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abdulla:2004:SDD

[2064] Ghaleb Abdulla, Terence Critchlow, and William Arrighi. Simulation data as data streams. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):89–94, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sim:2004:TOE

[2065] Kwang Mong Sim. Toward an ontology-enhanced information filtering agent. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):95–100, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schwentick:2004:XQC

[2066] Thomas Schwentick. XPath query containment. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):101–109, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:2004:SBP

[2067] Andrew Eisenberg, Jim Melton, Krishna Kulkarni, Jan-Eike Michels, and Fred Zemke. SQL:2003 has been published. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):119–126, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gertz:2004:RDS

[2068] Michael Gertz, M. Tamer Özsu, Gunter Saake, and Kai-Uwe Sattler. Report on the Dagstuhl Seminar: “Data Quality on the Web”. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):127–132, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

staff:2004:SDI

[2069] ACM SIGMOD Record staff. Science of design for information systems: report of the NSF workshop, Seattle, 2003. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):133–137, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nittel:2004:RFW

[2070] S. Nittel, A. Stefanidis, I. Cruz, M. Egenhofer, D. Goldin, A. Howard, A. Labrinidis, S. Madden, A. Voisard, and M. Worboys. Report from the First Workshop on Geo Sensor Networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(1):141–144, March 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gupta:2004:LSR

[2071] Amarnath Gupta. Life science research and data management—what can they give each other? SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):12–14, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:2004:DML

[2072] H. V. Jagadish and Frank Olken.

Database management for life sciences research. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):15–20, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Liu:2004:OAQ

[2073] Hsiao-Fei Liu, Ya-Hui Chang, and Kun-Mao Chao. An optimal algorithm for querying tree structures and its applications in bioinformatics. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):21–26, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bada:2004:URG

[2074] Michael Bada, Daniele Turi, Robin McEntire, and Robert Stevens. Using reasoning to guide annotation with gene ontology terms in GOAT. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33(2): 27–32, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aoki:2004:MAC

[2075] Kiyoko F. Aoki, Nobuhisa Ueda, Atsuko Yamaguchi, Tatsuya Akutsu, Minoru Kanehisa, and Hiroshi Mamitsuka. Managing and analyzing carbohydrate data. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):33–38, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cao:2004:PEM

[2076] Xia Cao, Shuai Cheng Li, Beng Chin Ooi, and Anthony K. H. Tung. Piers: an efficient model for similarity search in DNA sequence databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):39–44, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hammer:2004:GPD

[2077] Joachim Hammer and Markus Schneider. The GenAlg project: developing a new integrating data model, language, and tool for managing and querying genomic information. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):45–50, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Herbert:2004:BAE

[2078] Katherine G. Herbert, Narain H.

Gehani, William H. Piel, Jason T. L. Wang, and Cathy H. Wu. BIO-AJAX: an extensible framework for biological data cleaning. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):51–57, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:2004:ACW

[2079] Liangyou Chen, Hasan M. Jamil, and Nan Wang. Automatic composite wrapper generation for semi-structured biological data based on table structure identification. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):58–64, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Singh:2004:DDB

[2080] Ambuj K. Singh, B. S. Manjunath, and Robert F. Murphy. A distributed database for bio-molecular images. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):65–71, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bleiholder:2004:BCE

[2081] Jens Bleiholder, Felix Naumann, Zoé Lacroix, Louiqa Raschid, Hyma Murthy, and Maria-Esther Vidal. Bio-Fast: challenges in exploring linked life sciences sources. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):72–77, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernard:2004:MDS

[2082] Guy Bernard, Jalel Ben-othman, Luc Bouganim, Gérôme Canals, Sophie Chabridon, Bruno Defude, Jean Ferrié, Stéphane Gançarski, Rachid Guerraoui, Pascal Molli, Philippe Pucheral, Claudia Roncancio, Patricia Serrano-Alvarado, and Patrick Valduriez. Mobile databases: a selection of open issues and research directions. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):78–83, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grandi:2004:IAB

[2083] Fabio Grandi. Introducing an annotated bibliography on temporal and evolution aspects in the World Wide Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):84–86, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gottlob:2004:LBW

[2084] Georg Gottlob and Christoph Koch. Logic-based web information extraction. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):87–94, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2004:JNS

[2085] Marianne Winslett. Jeffrey Naughton speaks out on database systems as control freaks, how to choose students, how to choose problems, how to get attention, the importance of being true to yourself, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):95–102, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2004:RIPa

[2086] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):103–104, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:2004:ELX

[2087] Andrew Eisenberg and Jim Melton. An early look at XQuery API for Java (XQJ). SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):105–111, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nagaraj:2004:RMG

[2088] S. V. Nagaraj. Review of “Managing gigabytes: compressing and indexing documents and images”, by Ian H. Witten, Alistair Moffat, and Timothy C. Bell, Morgan Kauffman, 1999, 519 pp. ISBN 1-55860-570-3. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):113–114, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shasha:2004:DTP

[2089] Dennis Shasha, Philippe Bonnet, and Nancy Hartline Bercich. Database tuning principles, experiments, and troubleshooting techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):115– 116, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2004:BY

[2090] Richard Snodgrass and Merrie Brucks. Branding yourself. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):117–125, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roddick:2004:STS

[2091] John F. Roddick, Erik Hoel, Max J. Egenhofer, Dimitris Papadias, and Betty Salzberg. Spatial, temporal and spatio-temporal databases — hot issues and directions for Ph.D. research. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):126–131, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mesiti:2004:REW

[2092] Marco Mesiti, Barbara Catania, Giovanna Guerrini, and Akmal Chaudhri. Report on the EDBT’04 workshop on database technologies for handling XML information on the Web. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (2):132–134, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hislop:2004:IME

[2093] Dave Hislop, Zoé Lacroix, and Gerald Moeller. Issues in mechanical engineering design management. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (2):135–138, June 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Arasu:2004:DSC

[2094] Arvind Arasu and Jennifer Widom.

A denotational semantics for continuous queries over streams and relations. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(3):6–11, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fekete:2004:ROT

[2095] Alan Fekete, Elizabeth O’Neil, and Patrick O’Neil. A read-only transaction anomaly under snapshot isolation. SIGMOD Record (ACM Special Interest Group on Management of Data), 33 (3):12–14, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mazzoleni:2004:CCV

[2096] Pietro Mazzoleni, Elisa Bertino, Elena Ferrari, and Stefano Valtolina. CiVeDi: a customized virtual environment for database interaction. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(3):15–20, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dogac:2004:SEW

[2097] A. Dogac, Y. Kabak, G. Laleci, S. Sinir, A. Yildiz, S. Kirbas, and Y. Gurcan. Semantically enriched Web services for the travel industry. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (3):21–27, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:2004:EXI

[2098] Hanyu Li, Mong Li Lee, Wynne Hsu, and Chao Chen. An evaluation of XML indexes for structural join. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (3):28–33, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Getta:2004:ODS

[2099] Janusz R. Getta and Ehsan Vossough. Optimization of data stream processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(3):34–39, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2004:PBS

[2100] Marianne Winslett. Phil Bernstein speaks out on trends at industrial research labs: what metadata management products need, peculiarities of the tenure sytem, how to fix the problems with database research conferences, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(3):40–48, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2004:RIPb

[2101] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(3):49–50, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hernandez:2004:IBS

[2102] Thomas Hernandez and Subbarao Kambhampati. Integration of biological sources: current systems and challenges ahead. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(3):51–60, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:2004:SDW

[2103] Kevin Chen-Chuan Chang, Bin He, Chengkai Li, Mitesh Patel, and Zhen Zhang. Structured databases on the web: observations and implications. SIGMOD Record (ACM Special Interest Group on Management of Data), 33 (3):61–70, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chatziantoniou:2004:NQN

[2104] Damianos Chatziantoniou and Achilleas Anagnostopoulos. NESTREAM: querying nested streams. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(3):71–78, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:2004:ASX

[2105] Andrew Eisenberg and Jim Melton.

Advancements in SQL/XML. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (3):79–86, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ferrari:2004:RAS

[2106] Elena Ferrari. Report on the 9th ACM Symposium on Access Control Models and Technologies (SACMAT’04). SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (3):88–90, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Levene:2004:RWD

[2107] Mark Levene and Alexandra Poulovassilis. Report on the 3rd Web Dynamics Workshop, at WWW’2004. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (3):91–95, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Keenoy:2004:RWM

[2108] Kevin Keenoy, Alexandra Poulovassilis, Vassilis Christophides, George Loizou, Giorgos Kokkinidis, George Samaras, and Nicolas Spyratos. Report on the Workshop on Metadata Management in Grid and Peer-to-Peer Systems, London, December 16, 2003. SIGMOD Record (ACM Special Interest Group on Management of Data), 33 (3):96–100, September 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Doan:2004:ISI

[2109] AnHai Doan, Natalya F. Noy, and Alon Y. Halevy. Introduction to the special issue on semantic integration. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):11–13, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Embley:2004:ADI

[2110] David W. Embley, Li Xu, and Yihong Ding. Automatic direct and indirect schema mapping: experiences and lessons learned. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):14–19, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

He:2004:HPL

[2111] Bin He and Kevin Chen-Chuan Chang.

A holistic paradigm for large scale schema matching. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):20–25, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rahm:2004:MLX

[2112] Erhard Rahm, Hong-Hai Do, and Sabine Maßmann. Matching large XML schemas. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):26–31, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Andritsos:2004:KAE

[2113] Periklis Andritsos, Ariel Fuxman, Anastasios Kementsietsidis, Renée J. Miller, and Yannis Velegrakis. Kanata: adaptation and evolution in data sharing systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):32–37, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:2004:ISS

[2114] Philip A. Bernstein, Sergey Melnik, Michalis Petropoulos, and Christoph Quix. Industrial-strength schema matching. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):38–43, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rosenthal:2004:SIS

[2115] Arnon Rosenthal, Len Seligman, and Scott Renner. From semantic integration to semantics management: case studies and a way forward. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (4):44–50, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Motro:2004:MFA

[2116] Amihai Motro, Jacob Berlin, and Philipp Anokhin. Multiplex, Fusionplex and Autoplex: three generations of information integration. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (4):51–57, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Uschold:2004:OSS

[2117] Michael Uschold and Michael Gruninger. Ontologies and semantics for seamless connectivity. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):58–64, December 2004.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Noy:2004:SIS

[2118] Natalya F. Noy. Semantic integration:

a survey of ontology-based approaches. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):65–70, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Helal:2004:URA

[2119] Abdelsalam Helal and Joachim Hammer. UbiData: requirements and architecture for ubiquitous data access. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):71–76, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2004:CMS

[2120] Marianne Winslett. C. Mohan speaks out: on R*, message queues, computer science in India, how ARIES came about, life as an IBM fellow, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):77–84, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2004:BRC

[2121] Karl Aberer. Book review column.

SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):85, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wodtke:2004:RWS

[2122] Dirk Wodtke. Review of “Web Services” by G. Alonso, F. Casati, H. Kuno, and V. Machiraju, Springer-Verlag, 2004, ISBN 30540-44008-9. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):86–87, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kilov:2004:REC

[2123] Haim Kilov. Review of “Emergence and Convergence: Qualitative Novelty and the Unity of Knowledge” by Mario Bunge, University of Toronto Press, 2004, ISBN 0-8020-8860-0. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (4):88–90, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2004:RIP

[2124] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 33(4):91–92, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Torlone:2004:RSE

[2125] Riccardo Torlone and Paolo Atzeni. Report on the seventh EDBT summer school: XML & databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (4):93–96, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manolescu:2004:RFX

[2126] Ioana Manolescu and Yannis Papakonstantinou. Report on the first “XQuery Implementation, Experience and Perspectives” workshop (XIME-P). SIG-MOD Record (ACM Special Interest Group on Management of Data), 33 (4):97–99, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hiemstra:2004:RFT

[2127] Djoerd Hiemstra and Vojkan Mihajlović. Report on the first twente data management workshop on XML databases and information retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 33 (4):100–102, December 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:2005:MSG

[2128] S. Abiteboul, R. Hull, and V. Vianu.

In memory of Seymour Ginsburg 1928– 2004. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):5–12, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DeWitt:2005:SRI

[2129] David DeWitt, Michael Carey, and Joseph M. Hellerstein. Stonebraker receives IEEE John von Neumann Medal. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):13–15, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Franklin:2005:CUA

[2130] Mike Franklin. Candidates for the upcoming ACM SIGMOD elections. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):16–23, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fan:2005:SPW

[2131] Jianchun Fan and Subbarao Kambhampati. A snapshot of public Web services. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):24–32, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guinepain:2005:RIA

[2132] Sylvain Guinepain and Le Gruenwald. Research issues in automatic database clustering. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):33–38, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:2005:NPN

[2133] Jin Li, David Maier, Kristin Tufte, Vassilis Papadimos, and Peter A. Tucker. No pane, no gain: efficient evaluation of sliding-window aggregates over data streams. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):39–44, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Xie:2005:USS

[2134] Rong Xie and Ryosuke Shibasaki. A unified spatiotemporal schema for representing and querying moving features. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):45–50, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Illarramendi:2005:RDR

[2135] Arantza Illarramendi, Esperanza Marcos, and Carmen Costilla. RedBD: the database research community in Spain. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):51–56, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Amsaleg:2005:RFI

[2136] Laurent Amsaleg, Björn Pór Jónsson, and Vincent Oria. Report from the First International Workshop on Computer Vision Meets Databases (CVDB 2004). SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):57–60, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lifschitz:2005:RBS

[2137] Sérgio Lifschitz and Alberto H. F.

Laender. Report on the 19^{th} Brazilian symposium on databases (SBBD 2004). SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):61–62, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jones:2005:AMS

[2138] Cliff Jones, David Lomet, Alexander Romanovsky, Gerhard Weikum, Alan Fekete, Marie-Claude Gaudel, Henry F. Korth, Rogerio de Lemos, Eliot Moss, Ravi Rajwar, Krithi Ramamritham, Brian Randell, and Luis Rodrigues. The atomic manifesto: a story in four quarks. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):63–69, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Banerjee:2005:RMI

[2139] Sujata Banerjee, Mitch Cherniack, Panos K. Chrysanthis, Vijay Kumar, and Alexandros Labrinidis. Report on MobiDE 2003: the 3rd international ACM Workshop on Data Engineering for Wireless and Mobile Access. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):70–73, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2005:RIPa

[2140] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):74–76, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cohen:2005:CAQ

[2141] Sara Cohen. Containment of aggregate queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):77–85, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2005:DVO

[2142] Marianne Winslett. Databases in virtual organizations: a collective interview and call for researchers. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):86–89, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2005:DAT

[2143] Richard Snodgrass. Developments at ACM TODS. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(1):90, March 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:2005:TPH

[2144] Michael J. Carey and Jiawei Han. A tribute to Professor Hongjun Lu. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34(2): 5, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Koloniari:2005:PPM

[2145] Georgia Koloniari and Evaggelia Pitoura. Peer-to-peer management of XML data: issues and research challenges. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):6–17, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gaber:2005:MDS

[2146] Mohamed Medhat Gaber, Arkady Zaslavsky, and Shonali Krishnaswamy. Mining data streams: a review. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34(2): 18–26, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bordawekar:2005:APX

[2147] Rajesh R. Bordawekar and Christian A. Lang. Analytical processing of XML documents: opportunities and challenges. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):27–32, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Elmacioglu:2005:SDS

[2148] Ergin Elmacioglu and Dongwon Lee. On six degrees of separation in DBLP-DB and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):33–40, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Marx:2005:SCN

[2149] Maarten Marx and Maarten de Rijke. Semantic characterizations of navigational XPath. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):41–46, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tropashko:2005:NIT

[2150] Vadim Tropashko. Nested intervals tree encoding in SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):47–52, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ouzzani:2005:ICD

[2151] Mourad Ouzzani, Walid G. Aref, Elisa Bertino, Ann Christine Catlin, Christopher W. Clifton, Wing-Kai Hon, Ahmed K. Elmagarmid, Arif Ghafoor, Susanne E. Hambrusch, Sunil Prabhakar, Jeffrey S. Vitter, and Xiang Zhang. The Indiana Center for Database Systems at Purdue University. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):53–58, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hernandez:2005:RNC

[2152] Juan Hernández, Ernesto Pimentel, and Ambrosio Toval. Report on the Ninth conference on Software Engineering and Databases (JISBD 2004). SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):59–61, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kuijpers:2005:RIS

[2153] Bart Kuijpers and Peter Revesz. Report on the 1st International Symposium on the Applications of Constraint Databases (CDB’04). SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):62–64, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Theodoridis:2005:RIW

[2154] Yannis Theodoridis and Panos Vassiliadis. Report on the International Workshop on Pattern Representation and Management (PaRMa’04). SIG-MOD Record (ACM Special Interest Group on Management of Data), 34(2): 65–67, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2005:RIPb

[2155] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):68–70, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2005:BLS

[2156] Marianne Winslett. Bruce Lindsay speaks out: on System R, benchmarking, life as an IBM fellow, the power of DBAs in the old days, why performance still matters, Heisenbugs, why he still writes code, singing pigs, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):71–79, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tang:2005:BDM

[2157] Zhaohui Tang, Jamie Maclennan, and Peter Pyungchul Kim. Building data mining solutions with OLE DB for DM and XML for analysis. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):80–85, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hull:2005:TCW

[2158] Richard Hull and Jianwen Su. Tools for composite web services: a short overview. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(2):86–95, June 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ludascher:2005:GEI

[2159] Bertram Ludäscher and Carole Goble.

Guest editors’ introduction to the special section on scientific workflows. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):3–4, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shankar:2005:IDW

[2160] Srinath Shankar, Ameet Kini, David J.

DeWitt, and Jeffrey Naughton. Integrating databases and workflow systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):5–11, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McPhillips:2005:APN

[2161] Timothy M. McPhillips and Shawn Bowers. An approach for pipelining nested collections in scientific workflows. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):12–17, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Medeiros:2005:WWA

[2162] C. B. Medeiros, J. Perez-Alcazar, L. Digiampietri, G. Z. Pastorello, Jr., A. Santanche, R. S. Torres, E. Madeira, and E. Bacarin. WOODSS and the Web: annotating and reusing scientific workflows. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):18–23, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Maechling:2005:SCC

[2163] Philip Maechling, Hans Chalupsky, Maureen Dougherty, Ewa Deelman, Yolanda Gil, Sridhar Gullapalli, Vipin Gupta, Carl Kesselman, Jihic Kim, Gaurang Mehta, Brian Mendenhall, Thomas Russ, Gurmeet Singh, Marc Spraragen, Garrick Staples, and Karan Vahi. Simplifying construction of complex workflows for non-expert users of the Southern California Earthquake Center Community Modeling Environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):24–30, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Simmhan:2005:SDP

[2164] Yogesh L. Simmhan, Beth Plale, and Dennis Gannon. A survey of data provenance in e-science. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):31–36, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhao:2005:NSE

[2165] Yong Zhao, Jed Dobson, Ian Foster, Luc Moreau, and Michael Wilde. A notation and system for expressing and executing cleanly typed workflows on messy scientific data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34 (3):37–43, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yu:2005:TSW

[2166] Jia Yu and Rajkumar Buyya. A taxonomy of scientific workflow systems for grid computing. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):44–49, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hastings:2005:XDS

[2167] Shannon Hastings, Matheus Ribeiro, Stephen Langella, Scott Oster, Umit Catalyurek, Tony Pan, Kun Huang, Renato Ferreira, Joel Saltz, and Tahsin Kurc. XML database support for distributed execution of dataintensive scientific workflows. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34 (3):50–55, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wieczorek:2005:SSW

[2168] Marek Wieczorek, Radu Prodan, and Thomas Fahringer. Scheduling of scientific workflows in the ASKALON grid environment. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):56–62, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Verma:2005:ECB

[2169] Keshri Verma and O. P. Vyas. Efficient calendar based temporal association rule. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):63–70, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bicer:2005:AME

[2170] Veli Bicer, Gokce B. Laleci, Asuman Dogac, and Yildiray Kabak. Artemis message exchange framework: semantic interoperability of exchanged messages in the healthcare domain. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34 (3):71–76, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ulusoy:2005:DRB

[2171] Özgür Ulusoy. Database research at Bilkent University. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):77–80, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cicekli:2005:DMR

[2172] Nihan K. Cicekli, Ahmet Cosar, Asuman Dogac, Faruk Polat, Pinar Senkul, I. Hakki Toroslu, and Adnan Yazici. Data management research at the Middle East Technical University. SIGMOD Record (ACM Special Interest Group on Management of Data), 34 (3):81–84, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Thiran:2005:RWW

[2173] Ph. Thiran, T. Risch, C. Costilla, J. Henrard, Th. Kabisch, J. Petrini, W-J. van den Heuvel, and J-L. Hainaut. Report on the workshop on wrapper techniques for legacy data systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 34 (3):85–86, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bertossi:2005:EIC

[2174] Leopoldo Bertossi, Jan Chomicki, Parke Godfrey, Phokion G. Kolaitis, Alex Thomo, and Calisto Zuzarte. Exchange, integration, and consistency of data: report on the ARISE/NISR workshop. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):87–90, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Scarcello:2005:QAE

[2175] Francesco Scarcello. Query answering exploiting structural properties. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34 (3):91–99, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2005:JWS

[2176] Marianne Winslett. John Wilkes speaks out: on what the DB community needs to know about storage, how the DB and storage communities can join forces and change the world, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):100–110, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2005:RIPc

[2177] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(3):111–113, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2005:CT

[2178] Richard Snodgrass. CMM and TODS. SIGMOD Record (ACM Special Interest Group on Management of Data), 34 (3):114–117, September 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Miller:2005:MAO

[2179] Renèe J. Miller. In memoriam Alberto Oscar Mendelzon. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):7–12, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fernandez:2005:TGG

[2180] Mary Fernández. Tips on giving a good demo. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):13–14, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aksoy:2005:ISS

[2181] Demet Aksoy. Information source selection for resource constrained environments. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):15–20, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hidders:2005:LFF

[2182] Jan Hidders, Philippe Michiels, Jan Paredaens, and Roel Vercammen. LiX-Query: a formal foundation for XQuery research. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):21–26, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Franklin:2005:DDN

[2183] Michael Franklin, Alon Halevy, and David Maier. From databases to dataspaces: a new abstraction for information management. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):27–33, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gray:2005:SDM

[2184] Jim Gray, David T. Liu, Maria Nieto-Santisteban, Alex Szalay, David J. De-Witt, and Gerd Heber. Scientific data management in the coming decade. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):34–41, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:2005:RRT

[2185] Michael Stonebraker, Uğur Çetintemel, and Stan Zdonik. The 8 requirements of real-time stream processing. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34 (4):42–47, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rahm:2005:CAD

[2186] Erhard Rahm and Andreas Thor. Citation analysis of database publications. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):48–53, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sidiropoulos:2005:CBS

[2187] Antonis Sidiropoulos and Yannis Manolopoulos. A citation-based system to assist prize awarding. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34 (4):54–60, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:2005:AAC

[2188] Philip A. Bernstein, Elisa Bertino, Andreas Heuer, Christian S. Jensen, Holger Meyer, M. Tamer Özsu, and Richard T. Snodgrass. An applesto-apples comparison of two database journals. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):61–64, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buchmann:2005:DMR

[2189] A. Buchmann and M. Cilia. Data management research at Technische Universität Darmstadt. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):65–70, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Amer-Yahia:2005:RDI

[2190] Sihem Amer-Yahia, Pat Case, Thomas Rölleke, Jayavel Shanmugasundaram, and Gerhard Weikum. Report on the DB/IR panel at SIGMOD 2005. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34 (4):71–74, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shahabi:2005:RFI

[2191] Cyrus Shahabi, Ramesh Govindan, and Karl Aberer. Report on the first IEEE International Workshop on Networking Meets Databases (NetDB’05). SIG-MOD Record (ACM Special Interest Group on Management of Data), 34 (4):75–77, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:2005:XNC

[2192] Andrew Eisenberg and Jim Melton.

XQuery 1.0 is nearing completion. SIG-MOD Record (ACM Special Interest Group on Management of Data), 34 (4):78–84, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2005:CFS

[2193] Marianne Winslett. Christos Faloutsos speaks out: on power laws, fractals, the future of data mining, sabbaticals, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):85–89, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2005:RIPd

[2194] Kenneth A. Ross. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):90–91, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2005:CTE

[2195] Richard Snodgrass. Changes to the TODS editorial board. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):92, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ghanem:2006:EPW

[2196] Thanaa M. Ghanem, Walid G. Aref, and Ahmed K. Elmagarmid. Exploiting predicate-window semantics over data streams. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(1):3–8, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Byun:2006:MVH

[2197] Ji-Won Byun and Elisa Bertino. Microviews, or on how to protect privacy while enhancing data usability: concepts and challenges. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(1):9–13, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jiang:2006:RID

[2198] Nan Jiang and Le Gruenwald. Research issues in data stream association rule mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(1):14–19, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mani:2006:JMX

[2199] Murali Mani, Song Wang, Dan Dougherty, and Elke A. Rundensteiner. Join minimization in XML-to-SQL translation: an algebraic approach. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(1):20–25, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aguilar-Saborit:2006:DCF

[2200] J. Aguilar-Saborit, P. Trancoso, V. Muntes-Mulero, and J. L. Larriba-Pey. Dynamic count filters. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(1):26–32, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:2006:TDM

[2201] Zude Li and Xiaojun Ye. Towards a dynamic multi-policy dissemination control model: (DMDCON). SIGMOD Record (ACM Special Interest Group on Management of Data), 35(1):33–38, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Graefe:2006:BTI

[2202] Goetz Graefe. B-tree indexes for high update rates. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(1):39–44, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bierman:2006:ERR

[2203] Gavin Bierman and Christoph Koch.

Event reports: Report on the 10th International Symposium on Database Programming Languages: (DBPL 2005). SIGMOD Record (ACM Special Interest Group on Management of Data), 35(1):45–47, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Antonioletti:2006:IPW

[2204] Mario Antonioletti, Amy Krause, Norman W. Paton, Andrew Eisenberg, Simon Laws, Susan Malaika, Jim Melton, and Dave Pearson. Industry perspectives: The WS-DAI family of specifications for Web service data access and integration. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(1):48–55, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2006:DDP

[2205] Marianne Winslett. Distinguished database profiles: Moshe Vardi speaks out on the proof, the whole proof, and nothing but the proof. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(1):56–64, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deutsch:2006:DPQ

[2206] Alin Deutsch, Lucian Popa, and Val Tannen. Database principles: Query reformulation with constraints. SIG-MOD Record (ACM Special Interest Group on Management of Data), 35 (1):65–73, March 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sicilia:2006:EOD

[2207] Miguel-Ángel Sicilia and Elena García-Barriocanal. Extending object database interfaces with fuzziness through aspect-oriented design. SIG-MOD Record (ACM Special Interest Group on Management of Data), 35 (2):4–9, June 2006. CODEN SRECD8.

ISSN 0163-5808 (print), 1943-5835 (electronic).

Cohen:2006:SFO

[2208] Shirley Cohen, Patrick Hurley, Karl W.

Schulz, William L. Barth, and Brad Benton. Scientific formats for objectrelational database systems: a study of suitability and performance. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):10–15, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zimanyi:2006:TAT

[2209] Esteban Zimányi. Temporal aggregates and temporal universal quantification in standard SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):16–21, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tsalgatidou:2006:DSW

[2210] A. Tsalgatidou, G. Athanasopoulos, M. Pantazoglou, C. Pautasso, T. Heinis, R. Grønmo, Hjørdis Hoff, Arne-Jørgen Berre, M. Glittum, and S. Topouzidou. Developing scientific workflows from heterogeneous services. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):22–28, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Madden:2006:IDB

[2211] Samuel Madden and David DeWitt.

Impact of double-blind reviewing on SIGMOD publication rates. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):29–32, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Haigh:2006:VBF

[2212] Thomas Haigh. “A veritable bucket of facts” origins of the data base management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):33–49, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Scannapieco:2006:RFS

[2213] Monica Scannapieco and Laure Berti-Équille. Report from the First and Second International Workshops on Information Quality in Information Systems: IQIS 2004 and IQIS 2005 in conjunction with ACM SIGMOD/PODS Conferences. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):50–52, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schwarz:2006:RWD

[2214] Thomas Schwarz and Mark Manasse. Report on the 7^{th} Workshop on Distributed Data and Structures: (WDAS 2006). SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):53–55, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gupta:2006:RIW

[2215] Amarnath Gupta, Bertram Ludäscher, and Louiqa Raschid. Report on the 2nd International Workshop on Data Integration in the Life Sciences: (DILS’05). SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):56–58, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bonifati:2006:RAI

[2216] Angela Bonifati and Dongwon Lee. Report on the 7th ACM International Workshop on Web Information and Data Management: (WIDM 2005). SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):59–61, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sellis:2006:DMR

[2217] Timos Sellis and Yannis Vassiliou.

Data management research at the Knowledge and Database Systems Lab: (NTU Athens). SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):62–67, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bertossi:2006:CQA

[2218] Leopoldo Bertossi. Consistent query answering in databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):68–76, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2006:RRS

[2219] Marianne Winslett. Raghu Ramakrishnan speaks out on deductive databases, what lies beyond scalability, how he burned through $20m briskly, why we should reach out to policymakers, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):77–85, June 2006.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2006:CTE

[2220] Richard Snodgrass. Changes to the TODS editorial board. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(2):86, June 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tung:2006:IDB

[2221] Anthony K. H. Tung. Impact of double blind reviewing on SIGMOD publication: a more detail analysis. SIG-MOD Record (ACM Special Interest Group on Management of Data), 35 (3):6–7, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2006:SVD

[2222] Richard Snodgrass. Single- versus double-blind reviewing: an analysis of the literature. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(3):8–21, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vela:2006:MDD

[2223] Belén Vela, Eduardo Fernández-Medina, Esperanza Marcos, and Mario Piattini. Model driven development of secure XML databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(3):22–27, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhang:2006:ACO

[2224] Yi Zhang and Xiaoming Jin. An automatic construction and organization strategy for ensemble learning on data streams. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(3):28–33, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Choi:2006:SOM

[2225] Namyoun Choi, Il-Yeol Song, and Hyoil Han. A survey on ontology mapping. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(3):34–41, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Weikum:2006:DRG

[2226] Gerhard Weikum. The database research group at the Max-Planck Institute for Informatics. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(3):42–47, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Trujillo:2006:RFI

[2227] Juan Trujillo. A report on the first international workshop on best practices of UML: (BP-UML’05). SIG-MOD Record (ACM Special Interest Group on Management of Data), 35 (3):48–50, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bose:2006:RIP

[2228] Rajendra Bose, Ian Foster, and Luc Moreau. Report on the International Provenance and Annotation Workshop: (IPAW’06) 3–5 May 2006, Chicago. SIGMOD Record (ACM Special Interest Group on Management of Data), 35 (3):51–53, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cooper:2006:RSI

[2229] Brian F. Cooper and Roger Barga. Report on SciFlow 2006: the IEEE International Workshop on Workflow and Data Flow for Scientific Applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 35 (3):54–56, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2006:JWS

[2230] Marianne Winslett. Jennifer Widom speaks out: on luck, what constitutes success, when to get out of an area, the importance of choosing the right husband, outlandish vacations, how hard it is to be an assistant professor, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(3):57–65, September 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gal:2006:WSM

[2231] Avigdor Gal. Why is schema matching tough and what can we do about it? SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):2–5, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Llanos:2006:TUO

[2232] Diego R. Llanos. TPCC-UVa: an open-source TPC-C implementation for global performance measurement of computer systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):6–15, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Amer-Yahia:2006:XSL

[2233] Sihem Amer-Yahia and Mounia Lalmas. XML search: languages, INEX and scoring. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):16–23, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guo:2006:RSB

[2234] Shuqiao Guo, Zhiyong Huang, H. V.

Jagadish, Beng Chin Ooi, and Zhenjie Zhang. Relaxed space bounding for moving objects: a case for the buddy tree. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):24–29, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rahm:2006:OBS

[2235] Erhard Rahm and Philip A. Bernstein.

An online bibliography on schema evolution. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):30–31, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blakeley:2006:ANE

[2236] José A. Blakeley, David Campbell, S. Muralidhar, and Anil Nori. The ADO.NET entity framework: making the conceptual level real. SIG-MOD Record (ACM Special Interest Group on Management of Data), 35 (4):32–39, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aberer:2006:DMS

[2237] Karl Aberer, Gustavo Alonso, and Donald Kossmann. Data management for a smart earth: the Swiss NCCR– MICS initiative. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):40–45, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ailamaki:2006:RSI

[2238] Anastassia Ailamaki, Peter Boncz, and Stefan Manegold. Report on the Second International Workshop on Data Management on Modern Hardware (DaMoN’06). SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):46–47, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Trujillo:2006:REA

[2239] Juan Trujillo and Il-Yeol Song. A report on the Eighth ACM International Workshop on Data Warehousing and OLAP (DOLAP’05). SIG-MOD Record (ACM Special Interest Group on Management of Data), 35 (4):48–50, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yee:2006:IPD

[2240] Wai Gen Yee and Shamkant B. Navathe. ICDE 2006 Ph.D. workshop report. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):51–53, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Finin:2006:RMT

[2241] Tim Finin, Lalana Kagal, and Daniel Olmedilla. Report on the Models of Trust for the Web workshop (MTW’06). SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):54–56, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Arenas:2006:NTX

[2242] Marcelo Arenas. Normalization theory for XML. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):57–64, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2006:GMS

[2243] Marianne Winslett. Gerome Miklau speaks out on his SIGMOD distinguished dissertation award, how great it is to be a professor, and more. SIG-MOD Record (ACM Special Interest Group on Management of Data), 35 (4):65–68, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2006:YIS

[2244] Marianne Winslett. Yannis Ioannidis speaks out on database research funding in Europe, the importance of being uncertain, teaching as show business, the history of histograms, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):69–76, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2006:TRC

[2245] Richard Snodgrass. TODS report: Changes to the TODS editorial board. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):77, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Graefe:2007:ECS

[2246] Goetz Graefe. Efficient columnar storage in B-trees. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):3–6, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ross:2007:FSS

[2247] Kenneth A. Ross, David Vespe, David Hessing, and Pranay Jain. Finding shapes in a set of points. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):7–12, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zuo:2007:GHM

[2248] Xinqiang Zuo and Xiaoming Jin. General Hierarchical Model (GHM) to measure similarity of time series. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):13–18, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hoag:2007:PGP

[2249] Joseph E. Hoag and Craig W. Thompson. A parallel general-purpose synthetic data generator. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):19–24, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Braga:2007:XL

[2250] Daniele Braga, Alessandro Campi, Stefano Ceri, and Paola Spoletini. XQuery layers. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):25–30, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Segoufin:2007:SAX

[2251] Luc Segoufin. Static analysis of XML processing with data values. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):31–38, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2007:SCS

[2252] Marianne Winslett. Stefano Ceri speaks out on many-book researchers and one-startup researchers, Web modeling, the vanishing US–Europe research gap, the Semantic Web services train, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):39–46, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bahuman:2007:DII

[2253] Anil Bahuman, Chaitra Bahuman, Malati Baru, Subhasri Duttagupta, and Krithi Ramamritham. Developmental informatics at IIT Bombay. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):47–53, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:2007:RFI

[2254] Dongwon Lee and Chen Li. Report on the First International VLDB Workshop on Clean Databases (CleanDB 2006). SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):54–56, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:2007:HDT

[2255] James Z. Wang, Nozha Boujemaa, and Yixin Chen. High diversity transforms multimedia information retrieval into a cross-cutting field: report on the 8th Workshop on Multimedia Information Retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):57–59, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Snodgrass:2007:FAQ

[2256] Richard Snodgrass. Frequently-asked questions about double-blind reviewing. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(1):60–62, March 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tata:2007:ESTa

[2257] Sandeep Tata and Jignesh M. Patel. Estimating the selectivity of tf-idf based cosine similarity predicates. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(2):7–12, June 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shironoshita:2007:CEO

[2258] E. Patrick Shironoshita, Michael T.

Ryan, and Mansur R. Kabuka. Cardinality estimation for the optimization of queries on ontologies. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(2):13–18, June 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

tenCate:2007:NXC

[2259] Balder ten Cate and Maarten Marx.

Navigational XPath: calculus and algebra. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(2):19–26, June 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2007:GGS

[2260] Marianne Winslett. Georg Gottlob speaks out. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(2):27–33, June 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carey:2007:RTI

[2261] Michael Carey and Torsten Grust. Report on the Third International Workshop on XQuery Implementation, Experience and Perspectives (XIME-P 2006). SIGMOD Record (ACM Special Interest Group on Management of Data), 36(2):35–37, June 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Xiong:2007:RAW

[2262] Li Xiong and Yuni Xia. Report on ACM Workshop on Health Information and Knowledge Management (HIKM 2006). SIGMOD Record (ACM Special Interest Group on Management of Data), 36(2):39–42, June 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pardede:2007:ERI

[2263] Eric Pardede. Event report on iiWAS 2006 and MoMM 2006, Yogyakarta, December 2006. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(2):43–45, June 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

White:2007:DRO

[2264] Walker White, Christoph Koch, Nitin Gupta, Johannes Gehrke, and Alan Demers. Database research opportunities in computer games. SIG-MOD Record (ACM Special Interest Group on Management of Data), 36 (3):7–13, September 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Martens:2007:SSA

[2265] Wim Martens, Frank Neven, and Thomas Schwentick. Simple off the shelf abstractions for XML schema. SIGMOD Record (ACM Special Interest Group on Management of Data), 36 (3):15–22, September 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stavrianou:2007:OSI

[2266] Anna Stavrianou, Periklis Andritsos, and Nicolas Nicoloyannis. Overview and semantic issues of text mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 36 (3):23–34, September 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2007:KYW

[2267] Marianne Winslett. Kyu-Young Whang speaks out: on academia and startups in Korea, probabilistic counting, main-memory query optimization, how to avoid being a hostage of pressure publishing, and more. SIG-MOD Record (ACM Special Interest Group on Management of Data), 36 (3):35–40, September 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2007:BTL

[2268] Marianne Winslett. Boon Thau Loo speaks out: on his SIGMOD dissertation award, better networking through datalog, life as an assistant professor, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(3):41–45, September 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Group:2007:CSR

[2269] Community Systems Group. Community systems research at Yahoo! SIG-MOD Record (ACM Special Interest Group on Management of Data), 36 (3):47–54, September 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christophides:2007:RFI

[2270] Vassilis Christophides and Peter Buneman. Report on the First International Workshop on Database Preservation (PresDB’07). SIGMOD Record (ACM Special Interest Group on Management of Data), 36(3):55–58, September 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bruno:2007:CLT

[2271] Nicolas Bruno. A critical look at the TAB benchmark for physical design tools. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(4):7–12, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rubinson:2007:NTV

[2272] Claude Rubinson. Nulls, three-valued logic, and ambiguity in SQL: critiquing date’s critique. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(4):13–17, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bolchini:2007:DOS

[2273] Cristiana Bolchini, Carlo A. Curino, Elisa Quintarelli, Fabio A. Schreiber, and Letizia Tanca. A data-oriented survey of context models. SIG-MOD Record (ACM Special Interest Group on Management of Data), 36 (4):19–26, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ennals:2007:IMM

[2274] Rob Ennals, Eric Brewer, Minos Garofalakis, Michael Shadle, and Prashant Gandhi. Intel Mash Maker: join the Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(4):27–33, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2007:RBY

[2275] Marianne Winslett and Ricardo Baeza-Yates. Ricardo Baeza-Yates speaks out: on CS research in Latin America, his multi-continent commute for Yahoo!, how to get real data in academia, and Web mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(4):35–41, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ceri:2007:DWM

[2276] Stefano Ceri, Cristiana Bolchini, Daniele Braga, Marco Brambilla, Alessandro Campi, Sara Comai, Piero Fraternali, Pier Luca Lanzi, Marco Masseroli, Maristella Matera, Mauro Negri, Giuseppe Pelagatti, Giuseppe Pozzi, Elisa Quintarelli, Fabio A. Schreiber, and Letizia Tanca. Data and web management research at Politecnico di Milano. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(4):43–48, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ilyas:2007:RFI

[2277] Ihab F. Ilyas and Gautam Das. Report on the First International Workshop on Ranking in Databases (DBRank’07).

SIGMOD Record (ACM Special Interest Group on Management of Data), 36(4):49–51, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Balazinska:2007:RFI

[2278] Magdalena Balazinska, Amol Deshpande, Alexandros Labrinidis, Qiong Luo, Samuel Madden, and Jun Yang. Report on the Fourth International Workshop on Data Management for Sensor Networks (DMSN 2007). SIG-MOD Record (ACM Special Interest Group on Management of Data), 36 (4):53–55, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

deKeijzer:2007:RFV

[2279] Ander de Keijzer, Maurice van Keulen, and Alex Dekhtyar. Report on the First VLDB Workshop on Management of Uncertain Data (MUD). SIG-MOD Record (ACM Special Interest Group on Management of Data), 36 (4):57–58, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tata:2007:ESTb

[2280] Sandeep Tata and Jignesh M. Patel. Estimating the selectivity of tfidf based cosine similarity predicates. SIGMOD Record (ACM Special Interest Group on Management of Data), 36(4):75–80, December 2007. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bonifati:2008:DDP

[2281] Angela Bonifati, Panos K. Chrysanthis, Aris M. Ouksel, and Kai-Uwe Sattler. Distributed databases and peerto-peer databases: past and present. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):5–11, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Unay:2008:SQE

[2282] Ozan Ünay and Taflan I. Gündem.

A survey on querying encrypted XML documents for databases as a service. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):12–20, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Beeri:2008:BMQ

[2283] Catriel Beeri, Anat Eyal, Tova Milo, and Alon Pilberg. BP-Mon: querybased monitoring of BPEL business processes. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):21–24, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2008:SAS

[2284] Marianne Winslett. Serge Abiteboul speaks out: on building a research group in Europe, how he got involved in a startup, why systems papers shouldn’t have to include measurements, the value of object databases, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):25–33, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cafarella:2008:DMP

[2285] Michael Cafarella, Edward Chang, Andrew Fikes, Alon Halevy, Wilson Hsieh, Alberto Lerner, Jayant Madhavan, and S. Muthukrishnan. Data management projects at Google. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):34–38, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manolescu:2008:RES

[2286] I. Manolescu, L. Afanasiev, A. Arion, J. Dittrich, S. Manegold, N. Polyzotis, K. Schnaitter, P. Senellart, S. Zoupanos, and D. Shasha. The repeatability experiment of SIGMOD 2008. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):39–45, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Amsaleg:2008:RTI

[2287] Laurent Amsaleg, Björn pór Jónsson, and Vincent Oria. Report from the Third International Workshop on Computer Vision Meets Databases (CVDB 2007). SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):46–48, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Amer-Yahia:2008:DWP

[2288] Sihem Amer-Yahia, Volker Markl, Alon Halevy, AnHai Doan, Gustavo Alonso, Donald Kossmann, and Gerhard Weikum. Databases and Web 2.0 panel at VLDB 2007. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):49–52, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Holder:2008:RFI

[2289] Lawrence B. Holder and Xifeng Yan.

Report on the First International Workshop on Mining Graphs and Complex Structures (MGCS’07). SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):53–55, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:2008:RSA

[2290] Adam J. Lee and Ting Yu. Report on the Sixth ACM Workshop on Privacy in the Electronic Society (WPES 2007). SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):56–58, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pedersen:2008:RTA

[2291] Torben Bach Pedersen and Il-Yeol Song. Report on the Tenth ACM International Workshop on Data Warehousing and OLAP (DOLAP’07). SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):59–61, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cheney:2008:RPP

[2292] James Cheney, Peter Buneman, and Bertram Ludäscher. Report on the Principles of Provenance Workshop. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(1):62–65, March 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2008:MBF

[2293] Marianne Winslett. My biggest fan. SIGMOD Record (ACM Special Interest Group on Management of Data), 37 (2):4, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2008:JGS

[2294] Marianne Winslett. Jim Gray speaks out: on chasing the object-relational rainbow, why performance is a nonissue, bad ideas that went good, reinventing the field, sailboats, lunatic fringe papers, whether to try for a home run, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):5–15, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Carnes:2008:OS

[2295] Donna Carnes. Ode to a sailor. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):16–18, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Boss:2008:TMU

[2296] Pauline Boss. A tribute, not a memorial: understanding ambiguous loss. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):19–20, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Olson:2008:AS

[2297] Michael Olson. The amateur search.

SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):21–24, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hawthorn:2008:TUC

[2298] Paula Hawthorn. Thanks to the US Coast Guard. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):25, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Harrison:2008:JGB

[2299] Michael A. Harrison. Jim Gray at Berkeley. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):26–27, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Helland:2008:KW

[2300] Pat Helland. Knowledge and wisdom.

SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):28–29, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lazowska:2008:SRJ

[2301] Ed Lazowska. 500 special relationships: Jim as a mentor to faculty and students. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):30–32, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:2008:WDJ

[2302] Michael Stonebraker. Why did Jim Gray win the Turing Award? SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):33–34, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vaskevitch:2008:JGH

[2303] David Vaskevitch. Jim Gray: his contribution to industry. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):35, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rashid:2008:GB

[2304] Richard Rashid. A ’gap bridger’. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):36–37, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lindsay:2008:JGI

[2305] Bruce G. Lindsay. Jim Gray at IBM: the transaction processing revolution. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):38–40, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nauman:2008:JGT

[2306] John Nauman and Wendy Bartlett. Jim Gray’s Tandem contributions. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):41–44, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DeWitt:2008:JCC

[2307] David J. DeWitt and Charles Levine.

Not just correct, but correct and fast: a look at one of Jim Gray’s contributions to database system performance. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):45–49, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bell:2008:SI

[2308] C. Gordon Bell. Scaleability and immortality. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):50–53, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Reuter:2008:TLO

[2309] Andreas Reuter. Is there life outside transactions?: writing the transaction processing book. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):54–58, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barclay:2008:TRA

[2310] Tom Barclay. TerraServer and the Russia adventure . . .. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):59–60, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Szalay:2008:SDS

[2311] Alexander S. Szalay. The Sloan Digital Sky Survey and beyond. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):61–66, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wong:2008:BWT

[2312] Curtis Wong. Building the World-Wide Telescope. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):67–69, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Saade:2008:SSV

[2313] Ed Saade. Search survey for S/V Tenacious: Gulf of Farallones and approaches to San Francisco Bay. SIG-MOD Record (ACM Special Interest Group on Management of Data), 37(2): 70–77, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bellingham:2008:EOD

[2314] James G. Bellingham and Mike Godin. Exploring ocean data. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(2):78–82, June 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agrawal:2008:CRD

[2315] Rakesh Agrawal, Anastasia Ailamaki, Philip A. Bernstein, Eric A. Brewer, Michael J. Carey, Surajit Chaudhuri, AnHai Doan, Daniela Florescu, Michael J. Franklin, Hector Garcia-Molina, Johannes Gehrke, Le Gruenwald, Laura M. Haas, Alon Y. Halevy, Joseph M. Hellerstein, Yannis E. Ioannidis, Hank F. Korth, Donald Kossmann, Samuel Madden, Roger Magoulas, Beng Chin Ooi, Tim O’Reilly, Raghu Ramakrishnan, Sunita Sarawagi, Michael Stonebraker, Alexander S. Szalay, and Gerhard Weikum. The Claremont report on database research. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(3):9–19, September 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Date:2008:CCR

[2316] C. J. Date. A critique of Claude Rubinson’s paper Nulls, three-valued logic, and ambiguity in SQL: critiquing Date’s critique. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(3):20–22, September 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grant:2008:NVS

[2317] John Grant. Null values in SQL. SIG-MOD Record (ACM Special Interest Group on Management of Data), 37 (3):23–25, September 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ives:2008:OCD

[2318] Zachary G. Ives, Todd J. Green, Grigoris Karvounarakis, Nicholas E. Taylor, Val Tannen, Partha Pratim Talukdar, Marie Jacob, and Fernando Pereira. The ORCHESTRA Collaborative Data Sharing System. SIG-MOD Record (ACM Special Interest Group on Management of Data), 37 (3):26–32, September 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2008:ADS

[2319] Marianne Winslett. AnHai Doan speaks out on his ACM Dissertation Award, schema matching, following your passion, least publishable units, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(3):33–35, September 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Korth:2008:PPR

[2320] Henry F. Korth, Philip A. Bernstein, Mary Fernandez, Le Gruenwald, Phokion G. Kolaitis, Kathryn McKinley, and Tamer Ozsu. Paper and proposal reviews: is the process flawed? SIGMOD Record (ACM Special Interest Group on Management of Data), 37 (3):36–39, September 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:2008:CRC

[2321] H. V. Jagadish. The conference reviewing crisis and a proposed solution. SIGMOD Record (ACM Special Interest Group on Management of Data), 37 (3):40–45, September 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Amer-Yahia:2008:DIR

[2322] Sihem Amer-Yahia, Djoerd Hiemstra, Thomas Roelleke, Divesh Srivastava, and Gerhard Weikum. DB&IR integration: report on the Dagstuhl seminar ‘Ranked XML querying’. SIG-MOD Record (ACM Special Interest Group on Management of Data), 37 (3):46–49, September 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fundulaki:2008:RIW

[2323] Irini Fundulaki and Neoklis Polyzotis. Report on the 9th International Workshop on Web Information and Data Management (WIDM 2007). SIG-MOD Record (ACM Special Interest Group on Management of Data), 37 (3):50–52, September 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Krishnamurthy:2008:SSD

[2324] Rajasekar Krishnamurthy, Yunyao Li, Sriram Raghavan, Frederick Reiss, Shivakumar Vaithyanathan, and Huaiyu Zhu. SystemT: a system for declarative information extraction. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):7–13, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Doan:2008:IEC

[2325] AnHai Doan, Jeffrey F. Naughton, Raghu Ramakrishnan, Akanksha Baid, Xiaoyong Chai, Fei Chen, Ting Chen, Eric Chu, Pedro DeRose, Byron Gao, Chaitanya Gokhale, Jiansheng Huang, Warren Shen, and Ba-Quy Vuong. Information extraction challenges in managing unstructured data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 37 (4):14–20, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bohannon:2008:PSE

[2326] Philip Bohannon, Srujana Merugu, Cong Yu, Vipul Agarwal, Pedro DeRose, Arun Iyer, Ankur Jain, Vinay Kakade, Mridul Muralidharan, Raghu Ramakrishnan, and Warren Shen. Purple SOX extraction management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):21–27, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jain:2008:BQO

[2327] Alpa Jain, Panagiotis Ipeirotis, and Luis Gravano. Building query optimizers for information extraction: the SQoUT project. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):28–34, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gupta:2008:DAI

[2328] Rahul Gupta and Sunita Sarawagi. Domain adaptation of information extraction models. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):35–40, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kasneci:2008:YNA

[2329] Gjergji Kasneci, Maya Ramanath, Fabian Suchanek, and Gerhard Weikum. The YAGO-NAGA approach to knowledge discovery. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):41–47, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nie:2008:WUB

[2330] Zaiqing Nie, Ji-Rong Wen, and Wei-Ying Ma. Webpage understanding: beyond page-level search. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):48–54, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cafarella:2008:WSE

[2331] Michael J. Cafarella, Jayant Madhavan, and Alon Halevy. Web-scale extraction of structured data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 37 (4):55–61, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Weld:2008:UWB

[2332] Daniel S. Weld, Raphael Hoffmann, and Fei Wu. Using Wikipedia to bootstrap open information extraction. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):62–68, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kimelfeld:2008:MQP

[2333] Benny Kimelfeld and Yehoshua Sagiv.

Modeling and querying probabilistic XML data. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):69–77, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Koch:2008:QAP

[2334] Christoph Koch. On query algebras for probabilistic databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 37 (4):78–85, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2008:SCS

[2335] Marianne Winslett. Surajit Chaudhuri speaks out on how data mining led him to self-tuning databases: how he does tech transfer, life as a research manager, the fragmentation of database research, and more. SIG-MOD Record (ACM Special Interest Group on Management of Data), 37 (4):86–93, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Alonso:2008:EZS

[2336] Gustavo Alonso, Donald Kossmann, Timothy Roscoe, Nesime Tatbul, Andrew Baumann, Carsten Binnig, Peter Fischer, Oriana Riva, and Jens Teubner. The ETH Zurich Systems Group and Enterprise Computing Center. SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):94–99, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cormode:2008:HRP

[2337] Graham Cormode. How NOT to review a paper: the tools and techniques of the adversarial reviewer. SIG-MOD Record (ACM Special Interest Group on Management of Data), 37 (4):100–104, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cuesta:2008:RFE

[2338] Carlos E. Cuesta and Esperanza Marcos. A report on the first European Conference on Software Architecture (ECSA’2007). SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):105–107, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Xiong:2008:RIW

[2339] Li Xiong, Traian Marius Truta, and Farshad Fotouhi. Report on International Workshop on Privacy and Anonymity in the Information Society (PAIS 2008). SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):108–111, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

vanSinderen:2008:RIW

[2340] Marten van Sinderen, Pontus Johnson, and Lea Kutvonen. Report on the IFIP WG5.8 International Workshop on Enterprise Interoperability (IWEI 2008). SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):112–114, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manghi:2008:FWV

[2341] Paolo Manghi, Pasquale Pagano, and Pavel Zezula. First Workshop on Very Large Digital Libraries – VLDL 2008. SIGMOD Record (ACM Special Interest Group on Management of Data), 37 (4):115–117, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Parreiras:2008:FWT

[2342] Fernando Silva Parreiras, Jeff Z. Pan, Uwe Assmann, and Jakob Herinksson. First Workshop on Transforming and Weaving Ontologies in Model Driven Engineering (TWOMDE 2008).

SIGMOD Record (ACM Special Interest Group on Management of Data), 37(4):126, December 2008. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Codd:2009:DRC

[2343] E. F. Codd. Derivability, redundancy and consistency of relations stored in large data banks. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(1):17–36, March 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Laender:2009:XRW

[2344] Alberto H. F. Laender, Mirella M. Moro, Cristiano Nascimento, and Patrícia Martins. An X-ray on Webavailable XML schemas. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(1):37–42, March 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Florescu:2009:RCP

[2345] Daniela Florescu and Donald Kossmann. Rethinking cost and performance of database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(1):43–48, March 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barcelo:2009:LFR

[2346] Pablo Barceló. Logical foundations of relational data exchange. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(1):49–58, March 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DiLorenzo:2009:DIM

[2347] Giusy Di Lorenzo, Hakim Hacid, Hye young Paik, and Boualem Benatallah. Data integration in mashups. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(1):59–66, March 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2009:GWS

[2348] Marianne Winslett. Gerhard Weikum speaks out on why we should go for the grand challenges, why SQL is too powerful, the myth of precision, how to have a big research group in Germany, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(1):67–74, March 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mazon:2009:HMD

[2349] Jose-Norberto Mazón and Juan Trujillo. A hybrid model driven development framework for the multidimensional modeling of data warehouses! SIGMOD Record (ACM Special Interest Group on Management of Data), 38(2):12–17, June 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schweikardt:2009:MMQ

[2350] Nicole Schweikardt. Machine models for query processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(2):18–28, June 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Moro:2009:XSP

[2351] Mirella M. Moro, Vanessa Braganholo, Carina F. Dorneles, Denio Duarte, Renata Galante, and Ronaldo S. Mello. XML: some papers in a haystack. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(2):29–34, June 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chatziantoniou:2009:AQD

[2352] Damianos Chatziantoniou and Elias Tzortzakakis. ASSET queries: a declarative alternative to MapReduce. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(2):35–41, June 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2009:PBS

[2353] Marianne Winslett. Peter Buneman speaks out on phylogeny, the integration of databases and programming languages, curated databases, British plumbing, the value of talking to users, when to ignore the literature, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(2):42–49, June 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chan:2009:RIW

[2354] Chee-Yong Chan and Neoklis Polyzotis. Report on the 10th International Workshop on Web Information and Data Management (WIDM). SIGMOD Record (ACM Special Interest Group on Management of Data), 38(2):50–52, June 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:2009:RWO

[2355] Sang-Won Lee, Sooyong Kang, Youjip Won, and Jongmoo Choi. Report on Workshop on Operating Systems Support for Next Generation Large Scale NVRAM (NVRAMOS 2009). SIGMOD Record (ACM Special Interest Group on Management of Data), 38(2):53–56, June 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cheney:2009:WTP

[2356] James Cheney. Workshop on Theory and Practice of Provenance event report. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(2):57–60, June 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ooi:2009:SPW

[2357] Beng Chin Ooi, Kian Lee Tan, and Anthony Tung. Sense the physical, walkthrough the virtual, manage the co(existing) spaces: a database perspective. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(3):5–10, September 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lagogiannis:2009:TEI

[2358] G. Lagogiannis, N. Lorentzos, S. Sioutas, and E. Theodoridis. A time efficient indexing scheme for complex spatiotemporal retrieval. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(3):11–16, September 2009.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Arenas:2009:CIS

[2359] Marcelo Arenas, Jorge Pérez, Juan Reutter, and Cristian Riveros. Composition and inversion of schema mappings. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(3):17–28, September 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shmueli:2009:DEO

[2360] Erez Shmueli, Ronen Vaisenberg, Yuval Elovici, and Chanan Glezer. Database encryption: an overview of contemporary challenges and design considerations. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(3):29–34, September 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tanin:2009:STD

[2361] Egemen Tanin, Rui Zhang, and Lars Kulik. Spatio-temporal database research at the University of Melbourne. SIGMOD Record (ACM Special Interest Group on Management of Data), 38 (3):35–39, September 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manegold:2009:RWE

[2362] S. Manegold, I. Manolescu, L. Afanasiev, J. Feng, G. Gou, M. Hadjieleftheriou, S. Harizopoulos, P. Kalnis, K. Karanasos, D. Laurent, M. Lupu, N. Onose, C. Ré, V. Sans, P. Senellart, T. Wu, and D. Shasha. Repeatability & workability evaluation of SIGMOD 2009.

SIGMOD Record (ACM Special Interest Group on Management of Data), 38 (3):40–43, September 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cali:2009:LDR

[2363] Andrea Calì, Laks V. S. Lakshmanan, and Davide Martinenghi. Logic in databases: report on the LID 2008 workshop. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(3):44–49, September 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jonsson:2009:SBD

[2364] Björn Pór Jónsson. SIGMOD 2009 best demonstration competition. SIG-MOD Record (ACM Special Interest Group on Management of Data), 38 (3):50–51, September 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

VanCappellen:2009:XXJ

[2365] Marc Van Cappellen, Zhen Hua Liu, Jim Melton, and Maxim Orgiyan. XQJ: XQuery Java API is completed. SIG-MOD Record (ACM Special Interest Group on Management of Data), 38 (4):7–13, December 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Whang:2009:UD

[2366] Kyu-Young Whang, Il-Yeol Song, Taek-Yoon Kim, and Ki-Hoon Lee. The ubiquitous DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(4):14–22, December 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sakr:2009:RPR

[2367] Sherif Sakr and Ghazi Al-Naymat. Relational processing of RDF queries: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(4):23–28, December 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bercovitz:2009:SSR

[2368] Benjamin Bercovitz, Filip Kaliszan, Georgia Koutrika, Henry Liou, Aditya Parameswaran, Petros Venetis, Zahra Mohammadi Zadeh, and Hector Garcia-Molina. Social sites research through CourseRank. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(4):29–34, December 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lang:2009:EML

[2369] Willis Lang, Jignesh M. Patel, and Jeffrey F. Naughton. On energy management, load balancing and replication. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(4):35–42, December 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Amer-Yahia:2009:TIW

[2370] Sihem Amer-Yahia and Georgia Koutrika. Third International Workshop on ‘Personalized Access, Profile Management, and Context Awareness in Databases’ (PersDB 2009). SIG-MOD Record (ACM Special Interest Group on Management of Data), 38 (4):43–45, December 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manghi:2009:SWV

[2371] Paolo Manghi, Pasquale Pagano, and Yannis Ioannidis. Second Workshop on Very Large Digital Libraries: in Conjunction with the European Conference on Digital Libraries, Corfu, Greece, 2 October 2009. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(4):46–48, December 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Loo:2009:IWN

[2372] Boon Thau Loo and Stefan Saroiu. 5th International Workshop on Networking Meets Databases (NetDB 2009). SIGMOD Record (ACM Special Interest Group on Management of Data), 38(4):49–50, December 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hellerstein:2010:DIE

[2373] Joseph M. Hellerstein. The declarative imperative: experiences and conjectures in distributed logic. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(1):5–19, March 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barbieri:2010:QRS

[2374] Davide Francesco Barbieri, Daniele Braga, Stefano Ceri, Emanuele Della Valle, and Michael Grossniklaus. Querying RDF streams with C-SPARQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(1):20–26, March 2010.

CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kot:2010:BIR

[2375] Lucja Kot, Nitin Gupta, Sudip Roy, Johannes Gehrke, and Christoph Koch. Beyond isolation: research opportunities in declarative data-driven coordination. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(1):27–32, March 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Khare:2010:UDW

[2376] Ritu Khare, Yuan An, and Il-Yeol Song. Understanding deep Web search interfaces: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(1):33–40, March 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Drosou:2010:SRD

[2377] Marina Drosou and Evaggelia Pitoura. Search result diversification. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(1):41–47, March 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Liu:2010:SID

[2378] Mengmeng Liu, Svilen R. Mihaylov, Zhuowei Bao, Marie Jacob, Zachary G. Ives, Boon Thau Loo, and Sudipto Guha. SmartCIS: integrating digital and physical environments. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(1):48–53, March 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Benedikt:2010:REI

[2379] Michael Benedikt, Daniela Florescu, Philippa Gardner, Giovanna Guerrini, Marco Mesiti, and Emmanuel Waller. Report on the EDBT/ICDT 2010 workshop on updates in XML. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(1):54–57, March 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Meng:2010:RFI

[2380] Xiaofeng Meng, Jiaheng Lu, Jie Qiu, Ying Chen, and Haixun Wang. Report on the First International workshop on Cloud Data Management (CloudDB 2009). SIGMOD Record (ACM Special Interest Group on Management of Data), 39(1):58–60, March 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bellahsene:2010:FFD

[2381] Zohra Bellahsene, Salima Benbernou, Helene Jaudoin, François Pinet, Olivier Pivert, and Farouk Toumani. FORUM: a flexible data integration system based on data semantics. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(2):11–18, June 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cho:2010:VVE

[2382] Sung-Ryoung Cho, Jongwuk Lee, Seung-Won Hwang, Hwansoo Han, and Sang-Won Lee. VSkyline: vectorization for efficient skyline computation. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(2):19–26, June 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deutch:2010:MQL

[2383] Daniel Deutch and Tova Milo. On models and query languages for probabilistic processes. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(2):27–38, June 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grabova:2010:BIS

[2384] Oksana Grabova, Jerome Darmont, Jean-Hugues Chauchat, and Iryna Zolotaryova. Business intelligence for small and middle-sized enterprises. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(2):39–50, June 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2010:EBS

[2385] Marianne Winslett. Elisa Bertino speaks out: how she accrued 301 coauthors, revitalized a department, cut her commute to three minutes, enhanced our trust in shared data, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(2):51–57, June 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tao:2010:CRS

[2386] Yufei Tao. The Chair’s report on SIGMOD’10 demonstration program. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(2):58–60, June 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Genzmer:2010:SPC

[2387] Clément Genzmer, Volker Hudlet, Hyunjung Park, Daniel Schall, and Pierre Senellart. The SIGMOD 2010 programming contest: a distributed query engine. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(2):61–64, June 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:2010:TVT

[2388] Shimin Chen, Anastasia Ailamaki, Manos Athanassoulis, Phillip B. Gibbons, Ryan Johnson, Ippokratis Pandis, and Radu Stoica. TPC-E vs. TPC-C: characterizing the new TPC-E benchmark via an I/O comparison study. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(3):5–10, September 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2010:TMS

[2389] Marianne Winslett. Tova Milo speaks out. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(3):11–17, September 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ahmad:2010:SDM

[2390] Yanif Ahmad, Randal Burns, Michael Kazhdan, Charles Meneveau, Alex Szalay, and Andreas Terzis. Scientific data management at The Johns Hopkins Institute for Data Intensive Engineering and Science. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(3):18–23, September 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Krummenacher:2010:WSD

[2391] Reto Krummenacher, Karl Aberer, Atanas Kiryakov, and Rajaraman Kanagasabai. Workshop on semantic data management: a summary report. SIGMOD Record (ACM Special Interest Group on Management of Data), 39 (3):24–26, September 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winter:2010:TCT

[2392] Stephan Winter, Monika Sester, Ouri Wolfson, and Glenn Geers. Towards a computational transportation science. SIGMOD Record (ACM Special Interest Group on Management of Data), 39 (3):27–32, September 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nica:2010:EMR

[2393] Anisoara Nica, Fabian M. Suchanek, and Aparna S. Varde. Emerging multidisciplinary research across database management systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 39 (3):33–36, September 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dong:2010:IWW

[2394] Xin Luna Dong and Felix Naumann. 13th International Workshop on the Web and Databases: WebDB 2010. SIGMOD Record (ACM Special Interest Group on Management of Data), 39 (3):37–39, September 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Martinez-Gil:2010:ROM

[2395] Jorge Martinez-Gil and Jose F. Aldana-Montes. Reverse ontology matching. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(4):5–11, December 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cattell:2010:SSN

[2396] Rick Cattell. Scalable SQL and NoSQL data stores. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(4):12–27, December 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2010:CRS

[2397] Marianne Winslett. Christopher Ré speaks out on his ACM SIGMOD Jim Gray Dissertation Award, what he wishes he had known as a graduate student, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(4):28–30, December 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dittrich:2010:PBA

[2398] Jens Dittrich. Paper bricks: an alternative to complete-story peer reviewing. SIGMOD Record (ACM Special Interest Group on Management of Data), 39(4):31–36, December 2010. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lin:2011:IPI

[2399] Rung-Ren Lin, Ya-Hui Chang, and Kun-Mao Chao. Improving the performance of identifying contributors for XML keyword search. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(1):5–10, March 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:2011:ESR

[2400] Kuang Chen, Akshay Kannan, Jayant Madhavan, and Alon Halevy. Exploring schema repositories with Schemr. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(1):11–16, March 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2011:DSSa

[2401] Marianne Winslett. Dennis Shasha speaks out: on how puzzles helped his career, what drives him to write, how we can help biologists, the principles underlying database tuning, why he wears shorts all year, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(1):17–25, March 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aumuller:2011:AAD

[2402] David Aumüller and Erhard Rahm. Affiliation analysis of database publications. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(1):26–31, March 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hwang:2011:RDI

[2403] Seung won Hwang. Report on dataintensive software management and mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(1):32–34, March 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hu:2011:DSB

[2404] Zhenjiang Hu, Andy Schurr, Perdita Stevens, and James F. Terwilliger. Dagstuhl seminar on bidirectional transformations (BX). SIGMOD Record (ACM Special Interest Group on Management of Data), 40(1):35–39, March 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kolaitis:2011:RDA

[2405] Phokion G. Kolaitis, Maurizio Lenzerini, and Nicole Schweikardt. Report on DEIS’10: advanced school on data exchange, information, and streams (A GI-Dagstuhl Seminar). SIGMOD Record (ACM Special Interest Group on Management of Data), 40(1):40–42, March 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cohen-Boulakia:2011:SAR

[2406] Sarah Cohen-Boulakia and Ulf Leser.

Search, adapt, and reuse: the future of scientific workflows. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(2):6–16, June 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:2011:SEE

[2407] Jun Wang, Ling Feng, Wenwei Xue, and Zhanjiang Song. A survey on energy-efficient data management. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(2):17–23, June 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2011:LHS

[2408] Marianne Winslett and Vanessa Braganholo. Laura Haas speaks out on managing teams versus children, research versus product development, and much more. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(2):24–32, June 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2011:ALS

[2409] Marianne Winslett. Alberto Laender speaks out on why Google bought his startup, how to evaluate graduate program quality, how to do high-impact research in a developing country, how hyperinflation nurtured Brazil’s software industry, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(2):33–39, June 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Meng:2011:RFI

[2410] Xiaofeng Meng, Peiquan Jin, Wei Cao, and Lihua Yue. Report on the First International Workshop on Flash-based Database Systems (FlashDB 2011). SIGMOD Record (ACM Special Interest Group on Management of Data), 40(2):40–44, June 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bonnet:2011:RWE

[2411] Philippe Bonnet, Stefan Manegold, Matias Bjørling, Wei Cao, Javier Gonzalez, Joel Granados, Nancy Hall, Stratos Idreos, Milena Ivanova, Ryan Johnson, David Koop, Tim Kraska, René Müller, Dan Olteanu, Paolo Papotti, Christine Reilly, Dimitris Tsirogiannis, Cong Yu, Juliana Freire, and Dennis Shasha. Repeatability and workability evaluation of SIG-MOD 2011. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(2):45–48, June 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cali:2011:LTO

[2412] Andrea Calì, Georg Gottlob, Thomas Lukasiewicz, and Andreas Pieris. A logical toolbox for ontological reasoning. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(3):5–14, September 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buneman:2011:DWP

[2413] Peter Buneman, James Cheney, Sam Lindley, and Heiko Mueller. The database Wiki project: a generalpurpose platform for data curation and collaboration. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(3):15–20, September 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2011:MOS

[2414] Marianne Winslett and Vanessa Braganholo. Meral Özsoyoglu speaks out: on genealogical data management, searching ontologies, and more. SIG-MOD Record (ACM Special Interest Group on Management of Data), 40 (3):21–26, September 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2011:DSSb

[2415] Marianne Winslett and Vanessa Braganholo. Divesh Srivastava speaks out: on the importance of looking at real data, abstracting problems and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 40 (3):27–37, September 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Group:2011:DMR

[2416] Data Management Research Group. Data management research at NEC Labs. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(3):38–44, September 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Inkster:2011:IVI

[2417] Doug Inkster, Marcin Zukowski, and Peter Boncz. Integration of vectorwise with ingres. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(3):45–53, September 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ameloot:2011:YPF

[2418] Tom J. Ameloot, Maarten Marx, Wim Martens, Frank Neven, and Justin Van Wees. 30 Years of PODS in facts and figures. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(3):54–60, September 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Badia:2011:CAR

[2419] Antonio Badia and Daniel Lemire. A call to arms: revisiting database design. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(3):61–69, September 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:2011:OIS

[2420] Eun-Mi Lee, Sang-Won Lee, and Sangwon Park. Optimizing index scans on flash memory SSDs. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(4):5–10, December 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lee:2011:PDP

[2421] Kyong-Ha Lee, Yoon-Joon Lee, Hyunsik Choi, Yon Dohn Chung, and Bongki Moon. Parallel data processing with MapReduce: a survey. SIG-MOD Record (ACM Special Interest Group on Management of Data), 40 (4):11–20, December 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Marcus:2011:PVD

[2422] Adam Marcus, Michael S. Bernstein, Osama Badar, David R. Karger, Samuel Madden, and Robert C. Miller. Processing and visualizing the data in tweets. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(4):21–27, December 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2011:JHS

[2423] Marianne Winslett and Vanessa Braganholo. Jiawei Han speaks out: on data mining, privacy issues and managing students. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(4):28–38, December 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kersten:2011:DAR

[2424] Martin Kersten, Stefan Manegold, and Sjoerd Mullender. The Database Architectures Research Group at CWI. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(4):39–44, December 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Farber:2011:SHD

[2425] Franz Färber, Sang Kyun Cha, Jürgen Primsch, Christof Bornhövd, Stefan Sigg, and Wolfgang Lehner. SAP HANA database: data management for modern business applications. SIG-MOD Record (ACM Special Interest Group on Management of Data), 40 (4):45–51, December 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Maurino:2011:RIW

[2426] Andrea Maurino, Cinzia Cappiello, Panos Vassiliadis, and Kai-Uwe Sattler. Report on the 8th International Workshop on Quality in Databases (QDB10). SIGMOD Record (ACM Special Interest Group on Management of Data), 40(4):52–55, December 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bizer:2011:MUB

[2427] Christian Bizer, Peter Boncz, Michael L. Brodie, and Orri Erling. The meaningful use of big data: four perspectives – four challenges. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(4):56–60, December 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Candela:2011:FWV

[2428] Leonardo Candela, Paolo Manghi, and Yannis Ioannidis. Fourth workshop on very large digital libraries: on the marriage between very large digital libraries and very large data archives. SIGMOD Record (ACM Special Interest Group on Management of Data), 40(4):61–64, December 2011. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:2012:OXT

[2429] Ya-Hui Chang. Optimizing XML twig queries with full-text predicates. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(1):5–10, March 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dustdar:2012:QAS

[2430] Schahram Dustdar, Reinhard Pichler, Vadim Savenkov, and Hong-Linh Truong. Quality-aware service-oriented data integration: requirements, state of the art and open challenges. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(1):11–19, March 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mami:2012:SVS

[2431] Imene Mami and Zohra Bellahsene.

A survey of view selection methods.

SIGMOD Record (ACM Special Interest Group on Management of Data), 41(1):20–29, March 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2012:DLS

[2432] Marianne Winslett and Vanessa Braganholo. David Lomet speaks out: on database recovery, logs, versions and more. . .. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(1):30–39, March 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2012:CBS

[2433] Marianne Winslett and Vanessa Braganholo. Catriel Beeri speaks out: on his favorite pieces of work and on the importance of Sabbaticals. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(1):40–49, March 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wood:2012:QLG

[2434] Peter T. Wood. Query languages for graph databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(1):50–60, March 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:2012:IMI

[2435] Anonymous. Institute for the Management of Information Systems Athena Research Center. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(1):61–66, March 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic).  URL http://www.imis.
athena-innovation.gr.

Zemke:2012:WNS

[2436] Fred Zemke. What’s new in SQL:2011.

SIGMOD Record (ACM Special Interest Group on Management of Data), 41(1):67–73, March 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grossniklaus:2012:CFP

[2437] Michael Grossniklaus and David Maier.

The curriculum forecast for Portland: cloudy with a chance of data. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(1):74–77, March 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barcelo:2012:ARL

[2438] Pablo Barceló. Algorithms for regular languages that use algebra. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(2):5–14, June 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ferreira:2012:BSA

[2439] Anderson A. Ferreira, Marcos André Gonçalves, and Alberto H. F. Laender. A brief survey of automatic methods for author name disambiguation. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(2):15–26, June 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2012:TOS

[2440] Marianne Winslett and Vanessa Braganholo. Tamer Özsu speaks out: on journals, conferences, encyclopedias and technology. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(2):27–36, June 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2012:ENS

[2441] Marianne Winslett and Vanessa Braganholo. Erich Neuhold speaks out: on industry research versus academic research, funding projects, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(2):37–46, June 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:2012:CS

[2442] Philip A. Bernstein, Christian S. Jensen, and Kian-Lee Tan. A call for surveys. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(2):47, June 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barcelo:2012:SAD

[2443] Pablo Barceló. Semiring-annotated data: queries and provenance? SIG-MOD Record (ACM Special Interest Group on Management of Data), 41 (3):5–14, September 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manica:2012:HTI

[2444] Edimar Manica, Carina F. Dorneles, and Renata Renata Galante. Handling temporal information in Web search engines. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(3):15–23, September 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2012:RJR

[2445] Marianne Winslett and Vanessa Braganholo. Ryan Johnson: recipient of the 2012 ACM SIGMOD Jim Gray dissertation award. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(3):24–27, September 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Palpanas:2012:DDI

[2446] Themis Palpanas and Yannis Velegrakis. dbTrento: the data and information management group at the University of Trento. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(3):28–33, September 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kulkarni:2012:TFS

[2447] Krishna Kulkarni and Jan-Eike Michels.

Temporal features in SQL:2011. SIG-MOD Record (ACM Special Interest Group on Management of Data), 41 (3):34–43, September 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kissinger:2012:HTM

[2448] Thomas Kissinger, Benjamin Schlegel, Matthias Boehm, Dirk Habich, and Wolfgang Lehner. A high-throughput in-memory index, durable on flashbased SSD: insights into the winning solution of the SIGMOD programming contest 2011. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(3):44–50, September 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mazon:2012:RIW

[2449] Jose-Norberto Mazón, Irene Garrigós, Florian Daniel, and Malu Castellanos. Report of the International Workshop on Business IntelligencE and the Web: BEWEB 2011. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(3):51–53, September 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barcelo:2012:QSD

[2450] Pablo Barceló. Querying Semantic Data on the Web? SIGMOD Record (ACM Special Interest Group on Management of Data), 41(4):6–17, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grandi:2012:IAB

[2451] Fabio Grandi. Introducing an annotated bibliography on temporal and evolution aspects in the Semantic Web. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(4):18–21, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Park:2012:ODS

[2452] Hyunjung Park, Richard Pang, Aditya Parameswaran, Hector Garcia-Molina, Neoklis Polyzotis, and Jennifer Widom.

An overview of the Deco system: data model and query language; query processing and optimization. SIG-MOD Record (ACM Special Interest Group on Management of Data), 41 (4):22–27, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2012:DAS

[2453] Marianne Winslett and Vanessa Braganholo. Daniel Abadi speaks out. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(4):28–32, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Beskales:2012:DAG

[2454] George Beskales, Gautam Das, Ahmed K.

Elmagarmid, Ihab F. Ilyas, Felix Naumann, Mourad Ouzzani, Paolo Papotti, Jorge Quiane-Ruiz, and Nan Tang. The Data Analytics Group at the Qatar Computing Research Institute. SIG-MOD Record (ACM Special Interest Group on Management of Data), 41 (4):33–38, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lu:2012:DCD

[2455] Hua Lu, Torben Bach Pedersen, Simonas Saltenis, Bent Thomsen, Lone Leth Thomsen, and Kristian Torp. Daisy: the Center for Data-Intensive Systems at Aalborg University. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(4):39–44, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Aggarwal:2012:EPP

[2456] Charu C. Aggarwal. On the equivalence of PLSI and projected clustering. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(4):45–50, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hristidis:2012:CCM

[2457] Vagelis Hristidis. Challenges and communities of medical informatics research. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(4):51–54, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dong:2012:IWQ

[2458] Xin Luna Dong and Eduard Constantin Dragut. 10th international workshop on quality in databases: QDB 2012. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(4):55–59, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hidders:2012:RFW

[2459] Jan Hidders, Jacek Sroka, and Paolo Missier. Report from the First Workshop on Scalable Workflow Enactment Engines and Technology (SWEET’12). SIGMOD Record (ACM Special Interest Group on Management of Data), 41(4):60–64, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Meng:2012:XAF

[2460] Xiaofeng Meng and Fusheng Wang. XLDB Asia 2012: the First Extremely Large Databases Conference at Asia. SIGMOD Record (ACM Special Interest Group on Management of Data), 41(4):65–67, December 2012. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barcelo:2013:CDM

[2461] Pablo Barceló. The continuous distributed monitoring model. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(1):5–14, March 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Schomm:2013:MDI

[2462] Fabian Schomm, Florian Stahl, and Gottfried Vossen. Marketplaces for data: an initial survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(1):15–26, March 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2013:HKS

[2463] Marianne Winslett and Vanessa Braganholo. Hank Korth speaks out on two-career issues, why not to write a book in the beginning of your career, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(1):27–37, March 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chakrabarti:2013:DBR

[2464] Soumen Chakrabarti, Ganesh Ramakrishnan, Krithi Ramamritham, Sunita Sarawagi, and S. Sudarshan. Databased research at IIT Bombay. SIG-MOD Record (ACM Special Interest Group on Management of Data), 42 (1):38–43, March 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Stonebraker:2013:IBD

[2465] Michael Stonebraker, Sam Madden, and Pradeep Dubey. Intel “big data” science and technology center vision and execution plan. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(1):44–49, March 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pedersen:2013:RFI

[2466] Torben Bach Pedersen, Wolfgang Lehner, and Gregor Hackenbroich. Report on the First International Workshop on Energy Data Management (EnDM 2012). SIGMOD Record (ACM Special Interest Group on Management of Data), 42(1):50–52, March 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Meng:2013:RFI

[2467] Xiaofeng Meng, Fusheng Wang, and Adam Silberstein. Report on the Fourth International Workshop on Cloud Data Management (CloudDB 2012). SIGMOD Record (ACM Special Interest Group on Management of Data), 42(1):53–55, March 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barcelo:2013:TEI

[2468] Pablo Barceló. Towards efficient indexing of arbitrary similarity: vision paper. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(2):5–10, June 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Montolio:2013:REI

[2469] Sergio Lopez Montolio, David Dominguez-Sal, and Josep Lluis Larriba-Pey. Research endogamy as an indicator of conference quality. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(2):11–16, June 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guille:2013:IDO

[2470] Adrien Guille, Hakim Hacid, Cecile Favre, and Djamel A. Zighed. Information diffusion in online social networks: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(2):17–28, June 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Nakashole:2013:DSR

[2471] Ndapandula Nakashole, Gerhard Weikum, and Fabian Suchanek. Discovering semantic relations from the Web and organizing them with PATTY. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(2):29–34, June 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2013:JVS

[2472] Marianne Winslett and Vanessa Braganholo. Jeff Vitter speaks out on being a Southerner, duties of a Dean, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(2):35–45, June 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bressan:2013:DRN

[2473] Stephane Bressan, Chee Yong Chan, Wynne Hsu, Mong-Li Lee, Tok-Wang Ling, Beng Chin Ooi, Kian-Lee Tan, and Anthony K. H. Tung. Database research at the National University of Singapore. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(2):46–51, June 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cormode:2013:WDA

[2474] Graham Cormode. What does an Associate Editor actually do? SIGMOD Record (ACM Special Interest Group on Management of Data), 42(2):52–58, June 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Benedikt:2013:RFW

[2475] Michael Benedikt and Dan Olteanu. Report on the First Workshop on Innovative Querying of Streams. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(2):59–63, June 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Atzeni:2013:RMD

[2476] Paolo Atzeni, Christian S. Jensen, Giorgio Orsi, Sudha Ram, Letizia Tanca, and Riccardo Torlone. The relational model is dead, SQL is dead, and I don’t feel so good myself. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(2):64–68, June 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chomicki:2013:SQF

[2477] Jan Chomicki, Paolo Ciaccia, and Niccolo’ Meneghetti. Skyline queries, front and back. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(3):6–18, September 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ceri:2013:TMM

[2478] Stefano Ceri, Themis Palpanas, Emanuele Della Valle, Dino Pedreschi, Johann-Christoph Freytag, and Roberto Trasarti. Towards megamodeling: a walk through data analysis experiences. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(3):19–27, September 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lautert:2013:WTT

[2479] Larissa R. Lautert, Marcelo M.

Scheidt, and Carina F. Dorneles. Web table taxonomy and formalization. SIGMOD Record (ACM Special Interest Group on Management of Data), 42 (3):28–33, September 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Eisenberg:2013:XNC

[2480] Andrew Eisenberg. XQuery 3.0 is nearing completion. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(3):34–41, September 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Benslimane:2013:PPP

[2481] Djamal Benslimane, Mahmoud Barhamgi, Frederic Cuppens, Franck Morvan, Bruno Defude, and Ebrahim Nageba. PAIRSE: a privacy-preserving serviceoriented data integration system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 42 (3):42–47, September 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chabane:2013:MDM

[2482] Yahia Chabane, Laurent d’Orazio, Le Gruenwald, Baraa Mohamad, and Christophe Rey. Medical data management in the SYSEO project. SIG-MOD Record (ACM Special Interest Group on Management of Data), 42 (3):48–53, September 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2013:ARSa

[2483] Marianne Winslett and Vanessa Braganholo. Andreas Reuter speaks out on transactions, reinventing things, creating a university, and more. SIG-MOD Record (ACM Special Interest Group on Management of Data), 42 (3):54–62, September 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhou:2013:DCR

[2484] Xiaofang Zhou and Shazia Sadiq. Data centric research at the University of Queensland. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(3):63–68, September 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Castellanos:2013:RIW

[2485] Malu Castellanos, Umeshwar Dayal, and Elke Rundensteiner. Report on the 6th International Workshop on Business Intelligence for the Real Time Enterprise. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(3):69–70, September 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kacimi:2013:DIR

[2486] Mouna Kacimi, Fabian M. Suchanek, and Aparna Varde. Databases, information retrieval and knowledge management: exploring paths and crossing bridges. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(3):71–74, September 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ngo:2013:SSB

[2487] Hung Q. Ngo, Christopher Ré, and Atri Rudra. Skew strikes back: new developments in the theory of join algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(4):5–16, December 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bordawekar:2013:AA

[2488] Rajesh Bordawekar, Bob Blainey, and Chidanand Apte. Analyzing analytics. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(4):17–28, December 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:2013:STE

[2489] Fei Li, Hongzhi Wang, Jianzhong Li, and Hong Gao. A survey on tree edit distance lower bound estimation techniques for similarity join on XML data.

SIGMOD Record (ACM Special Interest Group on Management of Data), 42(4):29–39, December 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Naumann:2013:DPR

[2490] Felix Naumann. Data profiling revisited. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(4):40–49, December 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2013:ARSb

[2491] Marianne Winslett and Vanessa Braganhol. Anand Rajaraman speaks out on startups and social data. SIG-MOD Record (ACM Special Interest Group on Management of Data), 42 (4):50–60, December 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christodoulakis:2013:DMR

[2492] Stavros Christodoulakis, Minos Garofalakis, Euripides G. M. Petrakis, Antonios Deligiannakis, Vasilis Samoladas, Ekaterini Ioannou, Odysseas Papapetrou, and Stelios Sotiriadis. Data management research at the Technical University of Crete. SIGMOD Record (ACM Special Interest Group on Management of Data), 42(4):61–66, December 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Darmont:2013:RFI

[2493] Jérôme Darmont and Torben Bach Pedersen. Report on the first international workshop on cloud intelligence (Cloud-I 2012). SIGMOD Record (ACM Special Interest Group on Management of Data), 42(4):67–69, December 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pedersen:2013:RSI

[2494] Torben Bach Pedersen and Wolfgang Lehner. Report on the Second International Workshop on Energy Data Management (EnDM 2013). SIG-MOD Record (ACM Special Interest Group on Management of Data), 42 (4):70–72, December 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sroka:2013:RSW

[2495] Jacek Sroka, Jan Hidders, and Paolo Missier. Report from the second workshop on scalable workflow enactment engines and technology (SWEET’13). SIGMOD Record (ACM Special Interest Group on Management of Data), 42(4):73–77, December 2013. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McGregor:2014:GSA

[2496] Andrew McGregor. Graph stream algorithms: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(1):9–20, March 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tu:2014:SEE

[2497] Yi-Cheng Tu, Xiaorui Wang, Bo Zeng, and Zichen Xu. A system for energyefficient data management. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(1):21–26, March 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Angles:2014:LDB

[2498] Renzo Angles, Peter Boncz, Josep Larriba-Pey, Irini Fundulaki, Thomas Neumann, Orri Erling, Peter Neubauer, Norbert Martinez-Bazan, Venelin Kotsev, and Ioan Toma. The Linked Data Benchmark Council: a graph and RDF industry benchmarking effort. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(1):27–31, March 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Luiz:2014:MBF

[2499] Aldelir Fernando Luiz, Lau Cheuk Lung, and Miguel Correia. MI-TRA: Byzantine fault-tolerant middleware for transaction processing on replicated databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(1):32–38, March 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Balazinska:2014:DGU

[2500] Magdalena Balazinska, Bill Howe, and Dan Suciu. The database group at the University of Washington. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(1):39–44, March 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pawlish:2014:CEE

[2501] Michael Pawlish, Aparna S. Varde, Stefan A. Robila, and Anand Ranganathan. A call for energy efficiency in data centers. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(1):45–51, March 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anciaux:2014:ODS

[2502] Nicolas Anciaux, Luc Bouganim, Thierry Delot, Sergio Ilarri, Leila Kloul, Nathalie Mitton, and Philippe Pucheral. Opportunistic data services in least developed countries: benefits, challenges and feasibility issues. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(1):52–63, March 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wandelt:2014:SAS

[2503] Sebastian Wandelt, Dong Deng, Stefan Gerdjikov, Shashwat Mishra, Petar Mitankin, Manish Patil, Enrico Siragusa, Alexander Tiskin, Wei Wang, Jiaying Wang, and Ulf Leser. State-of-the-art in string similarity search and join. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(1):64–76, March 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ameloot:2014:DNR

[2504] Tom J. Ameloot. Declarative networking: Recent theoretical work on coordination, correctness, and declarative semantics. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(2):5–16, June 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Morton:2014:PDV

[2505] Kristi Morton, Magdalena Balazinska, Dan Grossman, Robert Kosara, and Jock Mackinlay. Public data and visualizations: How are many eyes and tableau public used for collaborative analytics? SIGMOD Record (ACM Special Interest Group on Management of Data), 43(2):17–22, June 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Torres:2014:EEN

[2506] Alexandre Torres, Renata Galante, and Marcelo Pimenta. ENORM: an essential notation for object-relational mapping. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(2):23–28, June 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lin:2014:LVS

[2507] Rung-Ren Lin, Ya-Hui Chang, and Kun-Mao Chao. Locating valid SLCAs for XML keyword search with NOT semantics. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(2):29–34, June 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhong:2014:MPG

[2508] Jianlong Zhong and Bingsheng He. Medusa: a parallel graph processing system on graphics processors. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(2):35–40, June 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Haas:2014:IRA

[2509] Laura Haas, Melissa Cefkin, Cheryl Kieliszewski, Wil Plouffe, and Mary Roth. The IBM Research Accelerated Discovery Lab. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(2):41–48, June 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Koutrika:2014:RFI

[2510] Georgia Koutrika, Laks V. S. Lakshmanan, Mirek Riedewald, and Kostas Stefanidis. Report on the First International Workshop on Exploratory Search in Databases and the Web (ExploreDB 2014). SIGMOD Record (ACM Special Interest Group on Management of Data), 43(2):49–52, June 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ma:2014:RSI

[2511] Shuai Ma, Xiaofeng Meng, and Fusheng Wang. Report on the Sixth International Workshop on Cloud Data Management (CloudDB 2014). SIG-MOD Record (ACM Special Interest Group on Management of Data), 43(2): 53–56, June 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manghi:2014:RFW

[2512] Paolo Manghi, Lukasz Bolikowski, Nikos Houssos, and Jochen Schirrwagen. Report on the First Workshop on Linking and Contextualizing Publications and Datasets. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(2):57–60, June 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deutsch:2014:AVD

[2513] Alin Deutsch, Richard Hull, and Victor Vianu. Automatic verification of database-centric systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 43 (3):5–17, September 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guo:2014:TTT

[2514] Chenjuan Guo, Christian S. Jensen, and Bin Yang. Towards total traffic awareness. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(3):18–23, September 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Braganholo:2014:SXF

[2515] Vanessa Braganholo and Marta Mattoso. A survey on XML fragmentation. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(3):24–35, September 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Santos:2014:ACD

[2516] Ricardo Jorge Santos, Jorge Bernardino, and Marco Vieira. Approaches and challenges in database intrusion detection. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(3):36–47, September 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mousavi:2014:TMS

[2517] Hamid Mousavi, Maurizio Atzori, Shi Gao, and Carlo Zaniolo. Textmining, structured queries, and knowledge management on Web document corpora. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(3):48–54, September 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kemper:2014:DGT

[2518] Alfons Kemper and Thomas Neumann.

The database group at TUM. SIG-MOD Record (ACM Special Interest Group on Management of Data), 43 (3):55–60, September 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abadi:2014:BRD

[2519] Daniel Abadi, Rakesh Agrawal, Anastasia Ailamaki, Magdalena Balazinska, Philip A. Bernstein, Michael J. Carey, Surajit Chaudhuri, Jeffrey Dean, An-Hai Doan, Michael J. Franklin, Johannes Gehrke, Laura M. Haas, Alon Y. Halevy, Joseph M. Hellerstein, Yannis E. Ioannidis, H. V. Jagadish, Donald Kossmann, Samuel Madden, Sharad Mehrotra, Tova Milo, Jeffrey F. Naughton, Raghu Ramakrishnan, Volker Markl, Christopher Olston, Beng Chin Ooi, Christopher Ré, Dan Suciu, Michael Stonebraker, Todd Walter, and Jennifer Widom. The Beckman report on database research. SIGMOD Record (ACM Special Interest Group on Management of Data), 43 (3):61–70, September 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Varde:2014:PID

[2520] Aparna S. Varde and Nikolaj Tatti.

A panorama of imminent doctoral research in data mining. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(3):71–74, September 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Amsterdamer:2014:FCD

[2521] Yael Amsterdamer and Tova Milo. Foundations of crowd data sourcing. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(4):5–14, December 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zliobaite:2014:THD

[2522] Indre Zliobaite, Jaakko Hollmen, Lauri Koskinen, and Jukka Teittinen. Towards hardware-driven design of lowenergy algorithms for data analysis. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(4):15–20, December 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lissandrini:2014:UPI

[2523] Matteo Lissandrini, Davide Mottin, Themis Palpanas, Dimitra Papadimitriou, and Yannis Velegrakis. Unleashing the power of information graphs. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(4):21–26, December 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2014:KLT

[2524] Marianne Winslett and Vanessa Braganholo. Kian-Lee Tan speaks out on how to build a strong DB group without pushing students hard. SIG-MOD Record (ACM Special Interest Group on Management of Data), 43 (4):27–32, December 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2014:SSS

[2525] Marianne Winslett and Vanessa Braganholo. Sudipto Das speaks out on scalability and elasticity of database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(4):33–36, December 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bentayeb:2014:DSD

[2526] Fadila Bentayeb, Julien Velcin, Stephane Bonnevay, and Jerome Darmont. Data science and decision support at ERIC. SIGMOD Record (ACM Special Interest Group on Management of Data), 43(4):37–42, December 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ellis:2014:EBD

[2527] Jason Ellis, Achille Fokoue, Oktie Hassanzadeh, Anastasios Kementsietsidis, Kavitha Srinivas, and Michael J. Ward. Exploring big data with Helix: Finding needles in a big haystack. SIG-MOD Record (ACM Special Interest Group on Management of Data), 43 (4):43–54, December 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pedersen:2014:RSI

[2528] Torben Bach Pedersen, Malu Castellanos, and Umesh Dayal. Report on the Seventh International Workshop on Business Intelligence for the Real Time Enterprise (BIRTE 2013). SIG-MOD Record (ACM Special Interest Group on Management of Data), 43 (4):55–58, December 2014. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jensen:2015:CTE

[2529] Christian S. Jensen. Changes to the TODS Editorial Board. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(1):5, March 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shriver:2015:MSW

[2530] Bruce Shriver and Marie Gentile. Michael Stonebraker wins the 2014 ACM Turing Award. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(1):6–9, March 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Segoufin:2015:CDE

[2531] Luc Segoufin. Constant delay enumeration for conjunctive queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(1):10–17, March 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Power:2015:IEG

[2532] Jason Power, Yinan Li, Mark D. Hill, Jignesh M. Patel, and David A. Wood. Implications of emerging 3D GPU architecture on the scan primitive. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(1):18–23, March 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kantorski:2015:AFH

[2533] Gustavo Zanini Kantorski, Viviane Pereira Moreira, and Carlos Alberto Heuser. Automatic filling of hidden Web forms: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(1):24–35, March 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2015:APSa

[2534] Marianne Winslett and Vanessa Braganholo. Aditya Parameswaran speaks out on human-powered computation. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(1):36–40, March 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2015:APSb

[2535] Marianne Winslett and Vanessa Braganholo. Andy Pavlo speaks out on main memory database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(1):41–45, March 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Idreos:2015:DDS

[2536] Stratos Idreos. DASlab: The Data Systems Laboratory at Harvard SEAS. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(1):46–51, March 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Christophides:2015:RFI

[2537] Vassilis Christophides and Themis Palpanas. Report on the First International Workshop on Personal Data Analytics in the Internet of Things (PDA@IOT 2014). SIGMOD Record (ACM Special Interest Group on Management of Data), 44(1):52–55, March 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Najafi:2015:FVF

[2538] Mohammadreza Najafi, Mohammad Sadoghi, and Hans-Arno Jacobsen. The FQP vision: Flexible query processing on a reconfigurable computing fabric. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(2):5–10, June 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Duggan:2015:BPS

[2539] Jennie Duggan, Aaron J. Elmore, Michael Stonebraker, Magda Balazinska, Bill Howe, Jeremy Kepner, Sam Madden, David Maier, Tim Mattson, and Stan Zdonik. The BigDAWG polystore system. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(2):11–16, June 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buoncristiano:2015:DCE

[2540] Marcello Buoncristiano, Giansalvatore Mecca, Elisa Quintarelli, Manuel Roveri, Donatello Santoro, and Letizia Tanca. Database challenges for exploratory computing. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(2):17–22, June 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Heinis:2015:FDS

[2541] Thomas Heinis and David A. Ham. Onthe-fly data synopses: Efficient data exploration in the simulation sciences. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(2):23–28, June 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Graefe:2015:IRD

[2542] Goetz Graefe. Instant recovery for data center savings. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(2):29–34, June 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tan:2015:MDC

[2543] Kian-Lee Tan, Qingchao Cai, Beng Chin Ooi, Weng-Fai Wong, Chang Yao, and Hao Zhang. In-memory databases: Challenges and opportunities from software and hardware perspectives. SIG-MOD Record (ACM Special Interest Group on Management of Data), 44(2): 35–40, June 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Niederee:2015:FDM

[2544] Claudia Niederee, Nattiya Kanhabua, Francesco Gallo, and Robert H. Logie. Forgetful digital memory: Towards brain-inspired long-term data and information management. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(2):41–46, June 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Palpanas:2015:DSM

[2545] Themis Palpanas. Data series management: The road to big sequence analytics. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(2):47–52, June 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pirk:2015:LCA

[2546] Holger Pirk. … like commanding an anthill: a case for micro-distributed (data) management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(2):53–58, June 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jensen:2015:BTW

[2547] Christian S. Jensen. The best of two worlds: Present your TODS paper at SIGMOD. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(2):59–60, June 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fan:2015:DQT

[2548] Wenfei Fan. Data quality: From theory to practice. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(3):7–18, September 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gawlick:2015:MSA

[2549] Dieter Gawlick, Eric S. Chan, Adel Ghoneimy, and Zhen Hua Liu. Mastering situation awareness: The next big challenge? SIGMOD Record (ACM Special Interest Group on Management of Data), 44(3):19–24, September 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yin:2015:RQO

[2550] Shaoyi Yin, Abdelkader Hameurlain, and Franck Morvan. Robust query optimization methods with respect to estimation errors: a survey. SIG-MOD Record (ACM Special Interest Group on Management of Data), 44 (3):25–36, September 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kim:2015:CDM

[2551] Jungeun Kim and Jae-Gil Lee. Community detection in multi-layer graphs: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(3):37–48, September 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2015:RCS

[2552] Marianne Winslett and Vanessa Braganholo. Rick Cattell speaks out on patenting, reinventing and standardizing things. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(3):49–56, September 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pedersen:2015:RTI

[2553] Torben Bach Pedersen. Report on the Third International Workshop on Energy Data Management (EnDM 2014). SIGMOD Record (ACM Special Interest Group on Management of Data), 44 (3):57–59, September 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:2015:RIW

[2554] Shimin Chen, Qiong Luo, and Xiaofeng Meng. Report on the International Workshop on Big Data Management on Emerging Hardware (HardBD 2015). SIGMOD Record (ACM Special Interest Group on Management of Data), 44 (3):60–61, September 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Castellanos:2015:RIW

[2555] Malu Castellanos, Umeshwar Dayal, Nesime Tatbul, Damianos Chatziantoniou, and Qiming Chen. Report on the 8th International Workshop on Business Intelligence for the Real-Time Enterprise (BIRTE’14). SIG-MOD Record (ACM Special Interest Group on Management of Data), 44 (3):62–63, September 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fagin:2015:RFI

[2556] Ronald Fagin, Benny Kimelfeld, Frederick Reiss, and Stijn Vansummeren. A relational framework for information extraction. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(4):5–16, December 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kumar:2015:MSM

[2557] Arun Kumar, Robert McCann, Jeffrey Naughton, and Jignesh M. Patel. Model selection management systems: The next frontier of advanced analytics. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(4):17–22, December 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pournajaf:2015:PPM

[2558] Layla Pournajaf, Daniel A. Garcia-Ulloa, Li Xiong, and Vaidy Sunderam. Participant privacy in mobile crowd sensing task management: a survey of methods and challenges. SIG-MOD Record (ACM Special Interest Group on Management of Data), 44 (4):23–34, December 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:2015:CPB

[2559] Hongzhi Wang, Mingda Li, Yingyi Bu, Jianzhong Li, Hong Gao, and Jiacheng Zhang. Cleanix: a parallel big data cleaning system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 44 (4):35–40, December 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2015:RSS

[2560] Marianne Winslett and Vanessa Braganholo. Rick Snodgrass speaks out on standards, personal brands and science. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(4):41–48, December 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Koutrika:2015:RSI

[2561] Georgia Koutrika, Laks V. S. Lakshmanan, Mirek Riedewald, Mohamed A. Sharaf, and Kostas Stefanidis. Report on the Second International Workshop on Exploratory Search in Databases and the Web (ExploreDB 2015). SIG-MOD Record (ACM Special Interest Group on Management of Data), 44 (4):49–52, December 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

deMelo:2015:DRP

[2562] Gerard de Melo, Mouna Kacimi, and Aparna S. Varde. Dissertation research problems in data management and related areas. SIGMOD Record (ACM Special Interest Group on Management of Data), 44(4):53–56, December 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Benevenuto:2015:HCA

[2563] Fabricio Benevenuto, Alberto H. F. Laender, and Bruno L. Alves. How connected are the ACM SIG communities? SIGMOD Record (ACM Special Interest Group on Management of Data), 44(4):57–63, December 2015. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Naughton:2016:TPN

[2564] Jeffrey F. Naughton. Technical perspective: Natural language to SQL translation by iteratively exploring a middle ground. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):5, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:2016:UNL

[2565] Fei Li and H. V. Jagadish. Understanding natural language queries over relational databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):6–13, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tan:2016:TPA

[2566] Wang-Chiew Tan. Technical perspective: Attacking the problem of consistent query answering. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):14, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Koutris:2016:CQA

[2567] Paraschos Koutris and Jef Wijsen.

Consistent query answering for primary keys. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):15–22, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Naughton:2016:TPB

[2568] Jeffrey F. Naughton. Technical perspective: Broadening and deepening query optimization yet still making progress. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):23, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Trummer:2016:MOP

[2569] Immanuel Trummer and Christoph Koch. Multi-objective parametric query optimization. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):24–31, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Libkin:2016:TPD

[2570] Leonid Libkin. Technical perspective:

Data distribution for fast joins. SIG-MOD Record (ACM Special Interest Group on Management of Data), 45(1): 32, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ameloot:2016:DPS

[2571] Tom J. Ameloot, Gaetano Geck, Bas Ketsman, Frank Neven, and Thomas Schwentick. Data partitioning for single-round multi-join evaluation in massively parallel systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):33–40, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DeWitt:2016:TPT

[2572] David J. DeWitt. Technical perspective: Taming hardware skew as parallel DBMSs scale out. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):41, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:2016:RBP

[2573] Jiexing Li, Jeffrey Naughton, and Rimma V. Nehme. Resource bricolage for parallel DBMSs on heterogeneous clusters. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):42–49, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ives:2016:TPI

[2574] Zachary G. Ives. Technical perspective: Implicit parallelism through deep language embedding. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):50, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Alexandrov:2016:IPT

[2575] Alexander Alexandrov, Asterios Katsifodimos, Georgi Krastev, and Volker Markl. Implicit parallelism through deep language embedding. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):51–58, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Halevy:2016:TPI

[2576] Alon Halevy. Technical perspective:

Incremental knowledge base construction using DeepDive. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):59, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sa:2016:DDK

[2577] Christopher De Sa, Alex Ratner, Christopher Ré, Jaeho Shin, Feiran Wang, Sen Wu, and Ce Zhang. Deep-Dive: Declarative knowledge base construction. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):60–67, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ives:2016:TPS

[2578] Zachary G. Ives. Technical perspective:

k-shape: Efficient and accurate clustering of time series. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):68, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Paparrizos:2016:SEA

[2579] John Paparrizos and Luis Gravano. kshape: Efficient and accurate clustering of time series. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):69–76, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jensen:2016:CTE

[2580] Christian S. Jensen. Changes to the TODS Editorial Board. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(1):77, March 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Olteanu:2016:FD

[2581] Dan Olteanu and Maximilian Schleich. Factorized databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(2):5–16, June 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:2016:DMD

[2582] Wei Wang, Meihui Zhang, Gang Chen, H. V. Jagadish, Beng Chin Ooi, and Kian-Lee Tan. Database meets deep learning: Challenges and opportunities. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(2):17–22, June 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dong:2016:TMI

[2583] Xin Luna Dong, Anastasios Kementsietsidis, and Wang-Chiew Tan. A time machine for information: Looking back to look forward. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(2):23–32, June 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:2016:SAD

[2584] Yihan Wang, Shaoxu Song, and Lei Chen. A survey on accessing dataspaces. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(2):33–44, June 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pavlo:2016:WRN

[2585] Andrew Pavlo and Matthew Aslett.

What’s really new with NewSQL? SIGMOD Record (ACM Special Interest Group on Management of Data), 45(2):45–55, June 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2016:HVJ

[2586] Marianne Winslett and Vanessa Braganholo. H V Jagadish speaks out on PVLDB, CoRR and data-driven research. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(2):56–62, June 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Naumann:2016:ISG

[2587] Felix Naumann and Ralf Krestel. The information systems group at HPI. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(2):63–68, June 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jensen:2016:DCT

[2588] Christian S. Jensen. The dark citations of TODS papers and what to do about it: Or: Cite the journal paper. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(2):69–70, June 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Saleh:2016:PED

[2589] Eyad Saleh, Ahmad Alsa’deh, Ahmad Kayed, and Christoph Meinel. Processing over encrypted data: Between theory and practice. SIG-MOD Record (ACM Special Interest Group on Management of Data), 45 (3):5–16, September 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Le:2016:SKS

[2590] Thuy Ngoc Le and Tok Wang Ling. Survey on keyword search over XML documents. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(3):17–28, September 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2016:CZS

[2591] Marianne Winslett and Vanessa Braganholo. Carlo Zaniolo speaks out on his passion for relational databases and logic. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(3):29–34, September 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roy:2016:RTI

[2592] Senjuti Basu Roy, Kostas Stefanidis, Georgia Koutrika, Laks V. S. Lakshmanan, and Mirek Riedewald. Report on the Third International Workshop on Exploratory Search in Databases and the Web (ExploreDB 2016). SIG-MOD Record (ACM Special Interest Group on Management of Data), 45 (3):35–38, September 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abiteboul:2016:RDP

[2593] Serge Abiteboul, Marcelo Arenas, Pablo Barceló, Meghyn Bienvenu, Diego Calvanese, Claire David, Richard Hull, Eyke Hüllermeier, Benny Kimelfeld, Leonid Libkin, Wim Martens, Tova Milo, Filip Murlak, Frank Neven, Magdalena Ortiz, Thomas Schwentick, Julia Stoyanovich, Jianwen Su, Dan Suciu, Victor Vianu, and Ke Yi. Research directions for principles of data management (abridged). SIGMOD Record (ACM Special Interest Group on Management of Data), 45(4):5–17, December 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Koutris:2016:GFA

[2594] Paraschos Koutris and Dan Suciu. A guide to formal analysis of join processing in massively parallel systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(4):18–27, December 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kamat:2016:CLV

[2595] Niranjan Kamat and Arnab Nandi. A closer look at variance implementations in modern database systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 45 (4):28–33, December 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Vartak:2016:TVR

[2596] Manasi Vartak, Silu Huang, Tarique Siddiqui, Samuel Madden, and Aditya Parameswaran. Towards visualization recommendation systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 45 (4):34–39, December 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2016:RHS

[2597] Marianne Winslett and Vanessa Braganholo. Rick Hull speaks out on asking the new question. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(4):40–45, December 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2016:SIS

[2598] Marianne Winslett and Vanessa Braganholo. Stratos Idreos speaks out on database cracking. SIGMOD Record (ACM Special Interest Group on Management of Data), 45(4):46–48, December 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Manghi:2016:RFI

[2599] Paolo Manghi, Jochen Schirrwagen, Oscar Corcho, and Amir Aryani. Report on the first international workshop on reproducible open science. SIG-MOD Record (ACM Special Interest Group on Management of Data), 45 (4):49–52, December 2016. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Olteanu:2017:TPJ

[2600] Dan Olteanu. Technical perspective:

Juggling functions inside a database. SIGMOD Record (ACM Special Interest Group on Management of Data), 46 (1):5, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Khamis:2017:JFI

[2601] Mahmoud Abo Khamis, Hung Q. Ngo, and Atri Rudra. Juggling functions inside a database. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):6–13, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kimelfeld:2017:TPO

[2602] Benny Kimelfeld. Technical perspective: Optimizing tree patterns for querying graph- and tree-structured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):14, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Czerwinski:2017:OTP

[2603] Wojciech Czerwinski, Wim Martens, Matthias Niewerth, and Pawel Parys. Optimizing tree patterns for querying graph- and tree-structured data. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):15–22, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chaudhuri:2017:TPR

[2604] Surajit Chaudhuri. Technical perspective: Reflections on extending SQL using constraints. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):23, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brucato:2017:SEE

[2605] Matteo Brucato, Azza Abouzied, and Alexandra Meliou. A scalable execution engine for package queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):24–31, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Naughton:2017:TPO

[2606] Jeffrey F. Naughton. Technical perspective: Optimized wandering for online aggregation. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):32, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:2017:WJX

[2607] Feifei Li, Bin Wu, Ke Yi, and Zhuoyue Zhao. Wander join and XDB: Online aggregation via random walks. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):33–40, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ives:2017:TPS

[2608] Zachary G. Ives. Technical perspective: Scaling machine learning via compressed linear algebra. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):41, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Elgohary:2017:SML

[2609] Ahmed Elgohary, Matthias Boehm, Peter J. Haas, Frederick R. Reiss, and Berthold Reinwald. Scaling machine learning via compressed linear algebra. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):42–49, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jensen:2017:UTE

[2610] Christian S. Jensen. Updates to the TODS Editorial Board. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(1):50, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barcelo:2017:SOT

[2611] Pablo Barceló, Andreas Pieris, and Miguel Romero. Semantic optimization in tractable classes of conjunctive queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(2):5–17, June 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Singh:2017:STM

[2612] Hari Singh and Seema Bawa. A survey of traditional and MapReduce-based spatial query processing approaches. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(2):18–29, June 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:2017:AEQ

[2613] Yang Chen, Xiaofeng Zhou, Kun Li, and Daisy Zhe Wang. Archimedes: Efficient query processing over probabilistic knowledge bases. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(2):30–35, June 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2017:BCO

[2614] Marianne Winslett and Vanessa Braganholo. Beng Chin Ooi speaks out on building a strong database group. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(2):36–42, June 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Afrati:2017:RTW

[2615] Foto N. Afrati, Jan Hidders, Christopher Ré, Jacek Sroka, and Jeffrey Ullman. Report from the Third Workshop on Algorithms and Systems for MapReduce and Beyond (BeyondMR’16). SIGMOD Record (ACM Special Interest Group on Management of Data), 46(2):43–48, June 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Guagliardo:2017:CSQ

[2616] Paolo Guagliardo and Leonid Libkin. Correctness of SQL queries on databases with nulls. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(3):5–16, September 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pham:2017:UMC

[2617] Thao N. Pham, Nikos R. Katsipoulakis, Panos K. Chrysanthis, and Alexandros Labrinidis. Uninterruptible migration of continuous queries without operator state migration. SIG-MOD Record (ACM Special Interest Group on Management of Data), 46 (3):17–22, September 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Esteves:2017:ESP

[2618] Sergi Esteves, Nico Janssens, Bart Theeten, and Luis Veiga. Empowering stream processing through edge clouds. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(3):23–28, September 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2017:RFS

[2619] Marianne Winslett and Vanessa Braganholo. Ron Fagin speaks out on his trajectory as a database theoretician. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(3):29–35, September 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lehner:2017:DDS

[2620] Wolfgang Lehner. The Dresden Database Systems Group. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(3):36–41, September 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Senellart:2017:PPR

[2621] Pierre Senellart. Provenance and probabilities in relational databases. SIG-MOD Record (ACM Special Interest Group on Management of Data), 46 (4):5–15, December 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pitoura:2017:MBO

[2622] Evaggelia Pitoura, Panayiotis Tsaparas, Giorgos Flouris, Irini Fundulaki, Panagiotis Papadakos, Serge Abiteboul, and Gerhard Weikum. On measuring bias in online information. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(4):16–21, December 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Spyropoulos:2017:DBD

[2623] Vasilis Spyropoulos and Yannis Kotidis Kotidis. Digree: Building a distributed graph processing engine out of single-node graph database installations. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(4):22–27, December 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2017:DSS

[2624] Marianne Winslett and Vanessa Braganholo. Dan Suciu speaks out on research, shyness and being a scientist. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(4):28–34, December 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sadiq:2017:DQR

[2625] Shazia Sadiq, Tamraparni Dasu, Xin Luna Dong, Juliana Freire, Ihab F. Ilyas, Sebastian Link, Miller J. Miller, Felix Naumann, Xiaofang Zhou, and Divesh Srivastava. Data quality: The role of empiricism. SIGMOD Record (ACM Special Interest Group on Management of Data), 46(4):35–43, December 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Afrati:2017:RFW

[2626] Foto N. Afrati, Jan Hidders, Paraschos Koutris, Jacek Sroka, and Jeffrey Ullman. Report from the Fourth Workshop on Algorithms and Systems for MapReduce and Beyond (BeyondMR ’17). SIGMOD Record (ACM Special Interest Group on Management of Data), 46(4):44–48, December 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tandon:2017:CKM

[2627] Niket Tandon, Aparna S. Varde, and Gerard de Melo. Commonsense knowledge in machine intelligence. SIG-MOD Record (ACM Special Interest Group on Management of Data), 46 (4):49–52, December 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tan:2018:TPR

[2628] Wang-Chiew Tan. Technical perspective: a relational framework for classifier engineering. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(1):5, March 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kimelfeld:2018:RFC

[2629] Benny Kimelfeld and Christopher Ré. A relational framework for classifier engineering. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(1):6–13, March 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ives:2018:TPT

[2630] Zachary G. Ives. Technical perspective: From think parallel to think sequential. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(1):14, March 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Fan:2018:TPT

[2631] Wenfei Fan, Yang Cao, Jingbo Xu, Wenyuan Yu, Yinghui Wu, Chao Tian, Jiaxin Jiang, and Bohan Zhang. From think parallel to think sequential. SIG-MOD Record (ACM Special Interest Group on Management of Data), 47 (1):15–22, March 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Papakonstantinou:2018:TPS

[2632] Yannis Papakonstantinou. Technical perspective: Supporting linear algebra operations in SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(1):23, March 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Luo:2018:SLA

[2633] Shangyu Luo, Zekai J. Gao, Michael Gubanov, Luis L. Perez, and Christopher Jermaine. Scalable linear algebra on a relational database system.

SIGMOD Record (ACM Special Interest Group on Management of Data), 47(1):24–31, March 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tan:2018:TPT

[2634] Wang-Chiew Tan. Technical perspective: Toward building entity matching management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(1):32, March 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Konda:2018:TPT

[2635] Pradap Konda, Sanjib Das, Paul Suganthan G. C., Philip Martinkus, Adel Ardalan, Jeffrey R. Ballard, Yash Govind, Han Li, Fatemah Panahi, Haojun Zhang, Jeff Naughton, Shishir Prasad, Ganesh Krishnan, Rohit Deep, and Vijay Raghavendra. Technical perspective: Toward building entity matching management systems. SIG-MOD Record (ACM Special Interest Group on Management of Data), 47 (1):33–40, March 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ives:2018:TPN

[2636] Zachary G. Ives. Technical perspective: Natural language explanations for query results. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(1):41, March 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Deutch:2018:NLE

[2637] Daniel Deutch, Nave Frost, and Amir Gilad. Natural language explanations for query results. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(1):42–49, March 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Benedikt:2018:LBP

[2638] Michael Benedikt. Logic-based perspectives on query reformulation over restricted interfaces. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(2):5–16, June 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Polyzotis:2018:DLC

[2639] Neoklis Polyzotis, Sudip Roy, Steven Euijong Whang, and Martin Zinkevich. Data lifecycle challenges in production machine learning: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(2):17–28, June 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hirzel:2018:SPL

[2640] Martin Hirzel, Guillaume Baudart, Angela Bonifati, Emanuele Della Valle, Sherif Sakr, and Akrivi Akrivi Vlachou. Stream processing languages in the big data era. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(2):29–40, June 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2018:KRS

[2641] Marianne Winslett and Vanessa Braganholo. Kenneth Ross speaks out on making contributions that span technology and science. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(2):41–47, June 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2018:PKS

[2642] Marianne Winslett and Vanessa Braganholo. Paris Koutris speaks out on a theoretical model for query processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(2):48–50, June 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Michels:2018:NIS

[2643] Jan Michels, Keith Hare, Krishna Kulkarni, Calisto Zuzarte, Zhen Hua Liu, Beda Hammerschmidt, and Fred Zemke. The new and improved SQL: 2016 Standard. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(2):51–60, June 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Alevizos:2018:CER

[2644] Elias Alevizos, Alexander Artikis, Nikos Katzouris, Evangelos Michelioudakis, Georgios Paliouras, and Georgios Paliouras. The complex event recognition group. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(2):61–66, June 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bhowmick:2018:DDI

[2645] Sourav S. Bhowmick and Xiaokui Xiaokui Xiao. DANTE: Data and information management research at Nanyang Technological University. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(2):67–72, June 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buneman:2018:DPW

[2646] Peter Buneman and Wang-Chiew Tan. Data provenance: What next? SIG-MOD Record (ACM Special Interest Group on Management of Data), 47 (3):5–16, September 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Shay:2018:DEA

[2647] Richard Shay, Uri Blumenthal, Vijay Gadepally, Ariel Hamlin, John Darby Mitchell, and Robert K. Cunningham. Don’t even ask: Database access control through query control. SIG-MOD Record (ACM Special Interest Group on Management of Data), 47 (3):17–22, September 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2018:TSS

[2648] Marianne Winslett and Vanessa Braganholo. Timos Sellis speaks out on research in Australia and Greece. SIG-MOD Record (ACM Special Interest Group on Management of Data), 47 (3):23–28, September 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2018:PBS

[2649] Marianne Winslett and Vanessa Braganholo. Peter Bailis speaks out on building tools users want to use. SIG-MOD Record (ACM Special Interest Group on Management of Data), 47 (3):29–31, September 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Singh:2018:NBP

[2650] Lisa Singh, Amol Deshpande, Wenchao Zhou, Arindam Banerjee, Alex Bowers, Sorelle Friedler, H. V. Jagadish, George Karypis, Zoran Obradovic, Anil Vullikanti, and Wangda Zuo. NSF BIG-DATA PI meeting — domain-specific research directions and data sets. SIG-MOD Record (ACM Special Interest Group on Management of Data), 47 (3):32–35, September 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sakr:2018:DSB

[2651] Sherif Sakr, Tilmann Rabl, Martin Hirzel, Paris Carbone, and Martin Strohbach. Dagstuhl Seminar on Big Stream Processing. SIG-MOD Record (ACM Special Interest Group on Management of Data), 47 (3):36–39, September 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bonifati:2018:GQT

[2652] Angela Bonifati and Stefania Dumbrava. Graph queries: From theory to practice. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(4):5–16, December 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2018:ACS

[2653] Marianne Winslett and Vanessa Braganholo. Andrew Chien speaks out on the impact of current trends in architecture. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(4):17–22, December 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Abedjan:2018:DMS

[2654] Ziawasch Abedjan, Sebastian Bresß, Volker Markl, Tilmann Rabl, and Juan Soto. Data management systems research at TU Berlin. SIG-MOD Record (ACM Special Interest Group on Management of Data), 47 (4):23–28, December 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bernstein:2018:SPC

[2655] Philip A. Bernstein. SIGMOD 2018 Program Committee Chair’s report. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(4):29–34, December 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Missier:2018:RFI

[2656] Paolo Missier, Tanu Malik, and Jacek Cala. Report on the First International Workshop on Incremental Recomputation: Provenance and Beyond. SIGMOD Record (ACM Special Interest Group on Management of Data), 47(4):35–38, December 2018. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Libkin:2019:RHB

[2657] Leonid Libkin. Research highlights: Bridging theory and practice with query log analysis. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):5, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Martens:2019:BTP

[2658] Wim Martens and Tina Trautner.

Bridging theory and practice with query log analysis. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):6–13, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cormode:2019:TPK

[2659] Graham Cormode. Technical perspective: ϵKTELO: A framework for defining differentially-private computations. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):14, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhang:2019:FDD

[2660] Dan Zhang, Ryan McKenna, Ios Kotsogiannis, George Bissias, Michael Hay, Ashwin Machanavajjhala, and Gerome Miklau. ϵKTELO: a framework for defining differentially-private computations. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):15–22, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kimelfeld:2019:TPE

[2661] Benny Kimelfeld and Wim Martens.

Technical perspective: Entity matching with quality and error guarantees. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):23, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tao:2019:EMQ

[2662] Yufei Tao. Entity matching with quality and error guarantees. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):24–31, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Martens:2019:TPE

[2663] Wim Martens. Technical perspective: Efficient query processing for dynamically changing datasets. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):32, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Idris:2019:EQP

[2664] Muhammad Idris, Martín Ugarte, Stijn Vansummeren, Hannes Voigt, and Wolfgang Lehner. Efficient query processing for dynamically changing datasets. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):33–40, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ives:2019:TPE

[2665] Zachary G. Ives. Technical perspective: Efficient signal reconstruction for a broad range of applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):41, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Asudeh:2019:ESR

[2666] Abolfazl Asudeh, Jees Augustine, Azade Nazi, Saravanan Thirumuruganathan, Nan Zhang, Gautam Das, and Divesh Srivastava. Efficient signal reconstruction for a broad range of applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):42–49, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Jagadish:2019:TPH

[2667] H. V. Jagadish. Technical perspective: How do humans and data systems establish a common query language? SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):50, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McCamish:2019:HDH

[2668] Ben McCamish, Vahid Ghadakchi, Arash Termehchy, Liang Huang, and Behrouz Touri. How do humans and data systems establish a common query language? SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):51–58, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Suciu:2019:TPM

[2669] Dan Suciu. Technical perspective for: MATLANG: Matrix operations and their expressive power. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):59, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brijder:2019:MMO

[2670] Robert Brijder, Floris Geerts, Jan Van den Bussche, and Timmy Weerwag. MATLANG: Matrix operations and their expressive power. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):60–67, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Yi:2019:TPO

[2671] Ke Yi. Technical perspective: Online model management via temporally biased sampling. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):68, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hentschel:2019:OMM

[2672] Brian Hentschel, Peter J. Haas, and Yuanyuan Tian. Online model management via temporally biased sampling. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):69–76, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Idreos:2019:TPS

[2673] Stratos Idreos. Technical perspective: Succinct range filters. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):77, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Zhang:2019:SRF

[2674] Huanchen Zhang, Hyeontaek Lim, Viktor Leis, David G. Andersen, Kimberly Keeton, and Andrew Pavlo. Succinct range filters. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(1):78–85, March 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rahul:2019:GDT

[2675] Saladi Rahul and Yufei Tao. A guide to designing top-k indexes. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(2):6–17, June 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Pierri:2019:FNS

[2676] Francesco Pierri and Stefano Ceri.

False news on social media: a datadriven survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(2):18–27, June 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Scherzinger:2019:BYO

[2677] Stefanie Scherzinger. Build your own SQL-on-Hadoop query engine: a report on a term project in a masterlevel database course. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(2):33–38, June 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2019:RHS

[2678] Marianne Winslett and Vanessa Braganholo. Richard Hipp speaks out on SQLite. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(2):39–46, June 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ailamaki:2019:SRT

[2679] Anastasia Ailamaki, Periklis Chrysogelos, Amol Deshpande, and Tim Kraska. The SIGMOD 2019 Research Track Reviewing System. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(2):47–54, June 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wijsen:2019:FQA

[2680] Jef Wijsen. Foundations of query answering on inconsistent databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(3):6–16, September 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

McCullough:2019:WTS

[2681] B. D. McCullough, Taha Mokfi, and Mahsa Almaeenjad. Wilkinson’s tests and SQL packages. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(3):17–22, September 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barca:2019:EIC

[2682] José María Cavero Barca, Belén Vela Sánchez, and Paloma Cáceres García de Marina. Evaluation of an implementation of cross-row constraints using materialized views. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(3):23–28, September 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Winslett:2019:MFS

[2683] Marianne Winslett and Vanessa Braganholo. Michael Franklin speaks out on data science. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(3):29–35, September 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Palpanas:2019:RFS

[2684] Themis Palpanas and Volker Beckmann. Report on the First and Second Interdisciplinary Time Series Analysis Workshop (ITISA). SIG-MOD Record (ACM Special Interest Group on Management of Data), 48 (3):36–40, September 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kondylakis:2019:RFI

[2685] Haridimos Kondylakis, Kostas Stefanidis, and Praveen Rao. Report on the First International Workshop on Semantic Web Technologies for Health Data Management (SWH 2018). SIG-MOD Record (ACM Special Interest Group on Management of Data), 48 (3):41–44, September 2019. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

CuencaGrau:2020:LDD

[2686] Bernardo Cuenca Grau, Ian Horrocks, Mark Kaminski, Egor V. Kostylev, and Boris Motik. Limit Datalog: a declarative query language for data analysis. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(4):6–17, February 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3385658.3385660.

Zhang:2020:HCS

[2687] Shuhao Zhang, Feng Zhang, Yingjun Wu, Bingsheng He, and Paul Johns. Hardware-conscious stream processing: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(4):18–29, February 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3385658.3385662.

Papadakis:2020:DSA

[2688] George Papadakis, Leonidas Tsekouras, Emmanouil Thanos, George Giannakopoulos, Themis Palpanas, and Manolis Koubarakis. Domain- and structure-agnostic end-to-end entity resolution with JedAI. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(4):30– 36, February 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3385658.
3385664.

Winslett:2020:NAS

[2689] Marianne Winslett and Vanessa Braganholo. Natassa Ailamaki speaks out on how to be a systems researcher and how to manage a large research group. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(4):37–43, February 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3385658.3385666.

Abadi:2020:SRD

[2690] Daniel Abadi, Anastasia Ailamaki, David Andersen, Peter Bailis, Magdalena Balazinska, Philip Bernstein, Peter Boncz, Surajit Chaudhuri, Alvin Cheung, AnHai Doan, Luna Dong, Michael J. Franklin, Juliana Freire, Alon Halevy, Joseph M. Hellerstein, Stratos Idreos, Donald Kossmann, Tim Kraska, Sailesh Krishnamurthy, Volker Markl, Sergey Melnik, Tova Milo, C. Mohan, Thomas Neumann, Beng Chin Ooi, Fatma Ozcan, Jignesh Patel, Andrew Pavlo, Raluca Popa, Raghu Ramakrishnan, Christopher Ré, Michael Stonebraker, and Dan Suciu. The Seattle Report on Database Research. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(4):44–53, February 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3385658.3385668.

Manghi:2020:DLS

[2691] Paolo Manghi, Leonardo Candela, Emma Lazzeri, and Gianmaria Silvello. Digital libraries: Supporting open science. SIGMOD Record (ACM Special Interest Group on Management of Data), 48(4):54–57, February 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3385658.3385669.

Gehrke:2020:TPC

[2692] Johannes Gehrke. Technical perspective: Checking invariant confluence, in whole or in parts. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):6, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3422648.3422650.

Whittaker:2020:CIC

[2693] Michael Whittaker and Joseph M. Hellerstein. Checking invariant confluence, in whole or in parts. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):7–14, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3422648.3422651.

Bernstein:2020:TPC

[2694] Philip A. Bernstein. Technical perspective of concurrent prefix recovery: Performing CPR on a database. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):15, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3422648.3422652.

Prasaad:2020:CPR

[2695] Guna Prasaad, Badrish Chandramouli, and Donald Kossmann. Concurrent prefix recovery: Performing CPR on a database. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):16–23, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3422648.3422653.

Kimelfeld:2020:TPC

[2696] Benny Kimelfeld. Technical perspective: Constant-delay enumeration for nondeterministic document spanners. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):24, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3422648.3422654.

Amarilli:2020:CDE

[2697] Antoine Amarilli, Pierre Bourhis, Stefan Mengel, and Matthias Niewerth. Constant-delay enumeration for nondeterministic document spanners. SIG-MOD Record (ACM Special Interest Group on Management of Data), 49 (1):25–32, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3422648.
3422655.

Getoor:2020:TPD

[2698] Lise Getoor. Technical perspective: Database repair meets algorithmic fairness. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):33, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3422648.3422656.

Salimi:2020:DRM

[2699] Babak Salimi, Bill Howe, and Dan Suciu. Database repair meets algorithmic fairness. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):34–41, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3422648.3422657.

Boehm:2020:TPD

[2700] Matthias Boehm. Technical perspective: Declarative recursive computation on an RDBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):42, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3422648.3422658.

Jankov:2020:DRC

[2701] Dimitrije Jankov, Shangyu Luo, Binhang Yuan, Zhuhua Cai, Jia Zou, Chris Jermaine, and Zekai J. Gao. Declarative recursive computation on an RDBMS: or, why you should use a database for distributed machine learning. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):43–50, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3422648.3422659.

Pichler:2020:TPE

[2702] Reinhard Pichler. Technical perspective: Efficient logspace classes for enumeration, counting, and uniform generation. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):51, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3422648.3422660.

Arenas:2020:ELC

[2703] Marcelo Arenas, Luis Alberto Croquevielle, Rajesh Jayaram, and Cristian Riveros. Efficient logspace classes for enumeration, counting, and uniform generation. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):52–59, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3422648.3422661.

Schelter:2020:TPQ

[2704] Sebastian Schelter. Technical perspective: Query optimization for faster deep CNN explanations. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):60, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3422648.3422662.

Nakandala:2020:QOF

[2705] Supun Nakandala, Arun Kumar, and Yannis Papakonstantinou. Query optimization for faster deep CNN explanations. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):61–68, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3422648.3422663.

Kanza:2020:TPR

[2706] Yaron Kanza. Technical perspective: Revealing every story of data in blockchain systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):69, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3422648.3422664.

Ruan:2020:RES

[2707] Pingcheng Ruan, Tien Tuan Anh Dinh, Qian Lin, Meihui Zhang, Gang Chen, and Beng Chin Ooi. Revealing every story of data in blockchain systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(1):70–77, September 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3422648.3422665.

Barcelo:2020:EPG

[2708] Pablo Barceló, Egor V. Kostylev, Mikaël Monet, Jorge Pérez, Juan L. Reutter, and Juan-Pablo Silva. The expressive power of graph neural networks as a query language. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(2):6–17, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3442322.3442324.

Schwentick:2020:SDC

[2709] Thomas Schwentick, Nils Vortmeier, and Thomas Zeume. Sketches of dynamic complexity. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(2):18–29, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3442322.3442325.

Amer-Yahia:2020:MAM

[2710] Sihem Amer-Yahia, Senjuti Basu Roy, Lei Chen, Atsuyuki Morishima, James Abello Monedero, Pierre Bourhis, François Charoy, Marina Danilevsky, Gautam Das, Gianluca Demartini, Shady Elbassuoni, David Gross-Amblard, Emilie Hoareau, Munenari Inoguchi, Jared Kenworthy, Itaru Kitahara, Dongwon Lee, Yunyao Li, Ria Mae Borromeo, Paolo Papotti, Raghav Rao, Sudeepa Roy, Pierre Senellart, Keishi Tajima, Saravanan Thirumuruganathan, Marion Tommasi, Kazutoshi Umemoto, Andrea Wiggins, and Koichiro Yoshida. Making AI machines work for humans in FoW. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(2):30–35, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL

https://dl.acm.org/doi/10.1145/
3442322.3442327.

Jandre:2020:PCS

[2711] Eduardo Jandre, Bruna Diirr, and Vanessa Braganholo. Provenance in collaborative in silico scientific research: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(2):36–51, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3442322.3442329.

Davidson:2020:SDS

[2712] Susan Davidson. Susan Davidson speaks out on collaborating with other research areas and balancing work and family. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(2):52–58, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3442322.3442331.

Kondylakis:2020:RSI

[2713] Haridimos Kondylakis, Kostas Stefanidis, Praveen Rao, and Dave Parry. Report on the Second International Workshop on Semantic Web Meets Health Data Management (SWH 2019). SIG-MOD Record (ACM Special Interest Group on Management of Data), 49 (2):59–62, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3442322.
                                        3442333.

Hu:2020:MPJ

[2714] Xiao Hu and Ke Yi. Massively parallel join algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(3):6–17, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3444831.3444833.

Hameed:2020:DPS

[2715] Mazhar Hameed and Felix Naumann. Data preparation: a survey of commercial tools. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(3):18–29, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3444831.3444835.

Winslett:2020:GGS

[2716] Marianne Winslett and Vanessa Braganholo. Goetz Graefe speaks out on (not only) query optimization. SIG-MOD Record (ACM Special Interest Group on Management of Data), 49 (3):30–36, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3444831.
3444837.

Jindal:2020:ARL

[2717] Alekh Jindal. Applied research lessons from CloudViews Project. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(3):37–42, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3444831.3444839.

Maier:2020:ASP

[2718] David Maier, Rachel Pottinger, An-Hai Doan, Eduard Dragut, Bill Howe, Joanne Lateulere, John Lateulere, Mostafa Milani, Tilmann Rabl, Dan Suciu, Yufei Tao, Wang-Chiew Tan, and Kristin Tufte. Advice from SIG-MOD/PODS 2020. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(3):43–54, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3444831.3444841.

Stonebraker:2020:PWS

[2719] Michael Stonebraker, Timothy Mattson, Tim Kraska, and Vijay Gadepally. Poly’19 Workshop summary: GDPR. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(3):55–58, December 2020. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3444831.3444842.

Traub:2021:ABT

[2720] Jonas Traub, Zoi Kaoudi, Jorge-Arnulfo Quiané-Ruiz, and Volker Markl. Agora: Bringing together datasets, algorithms, models and more in a unified ecosystem [vision]. SIG-MOD Record (ACM Special Interest Group on Management of Data), 49(4): 6–11, March 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3456859.3456861.

Liu:2021:BMN

[2721] Feilong Liu, Claude Barthels, Spyros Blanas, Hideaki Kimura, and Garret Swart. Beyond MPI: New communication interfaces for database systems and data-intensive applications. SIG-MOD Record (ACM Special Interest Group on Management of Data), 49(4): 12–17, March 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3456859.3456862.

Karumuri:2021:TOD

[2722] Suman Karumuri, Franco Solleza, Stan Zdonik, and Nesime Tatbul. Towards observability data management at scale. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(4):18–23, March 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3456859.3456863.

Artikis:2021:DSF

[2723] Alexander Artikis, Thomas Eiter, Alessandro Margara, and Stijn Vansummeren. Dagstuhl Seminar on the Foundations of Composite Event Recognition. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(4):24–27, March 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/10.
1145/3456859.3456865.

Bonifati:2021:HCO

[2724] Angela Bonifati, Giovanna Guerrini, Carsten Lutz, Wim Martens, Lara Mazilu, Norman W. Paton, Marcos Antonio Vaz Salles, Marc H. Scholl, and Yongluan Zhou. Holding a conference online and live due to Covid-19: Experiences and lessons learned from EDBT/ICDT 2020. SIGMOD Record (ACM Special Interest Group on Management of Data), 49(4):28–32, March 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3456859.3456866.

Cormode:2021:TPF

[2725] Graham Cormode. Technical perspective: a framework for adversarially robust streaming algorithms. SIG-MOD Record (ACM Special Interest Group on Management of Data), 50(1): 5, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471487.

Ben-Eliezer:2021:FAR

[2726] Omri Ben-Eliezer, Rajesh Jayaram, David P. Woodruff, and Eylon Yogev. A framework for adversarially robust streaming algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):6–13, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471488.

Fekete:2021:CCC

[2727] Alan D. Fekete. Chiller: Contentioncentric transaction execution and data partitioning for modern networks (technical perspective). SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):14, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3471485.3471489.

Zamanian:2021:CCC

[2728] Erfan Zamanian, Julian Shun, Carsten Binnig, and Tim Kraska. Chiller: Contention-centric transaction execution and data partitioning for modern networks. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):15–22, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3471485.3471490.

Boncz:2021:TPD

[2729] Peter Boncz. Technical perspective DIAMetrics: Benchmarking query engines at scale. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):23, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3471485.3471491.

Deep:2021:DBQ

[2730] Shaleen Deep, Anja Gruenheid, Kruthi Nagaraj, Hiro Naito, Jeff Naughton, and Stratis Viglas. DIAMetrics: Benchmarking query engines at scale. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):24–31, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3471485.
3471492.

Tao:2021:TPE

[2731] Yufei Tao. Technical perspective of efficient directed densest subgraph discovery. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):32, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3471485.
3471493.

Ma:2021:EDD

[2732] Chenhao Ma, Yixiang Fang, Reynold Cheng, Laks V. S. Lakshmanan, Wenjie Zhang, and Xuemin Lin. Efficient directed densest subgraph discovery. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):33–40, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471494.

Zhang:2021:TPF

[2733] Qin Zhang. Technical perspective:

Fair near neighbor search via sampling. SIGMOD Record (ACM Special Interest Group on Management of Data), 50 (1):41, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3471485.3471495.

Aumuller:2021:FNN

[2734] Martin Aumuller, Sariel Har-Peled, Sepideh Mahabadi, Rasmus Pagh, and Francesco Silvestri. Fair near neighbor search via sampling. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):42–49, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471496.

Howe:2021:TPS

[2735] Bill Howe. Technical perspective: From sketching to natural language: Expressive visual querying for accelerating insight. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):50, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471497.

Siddiqui:2021:SNL

[2736] Tarique Siddiqui, Paul Luh, Zesheng Wang, Karrie Karahalios, and Aditya G. Parameswaran. From sketching to natural language: Expressive visual querying for accelerating insight. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):51–58, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471498.

Zukowski:2021:TPO

[2737] Marcin Zukowski. Technical perspective: Optimistically compressed hash tables & strings in the USSR. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):59, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471499.

Gubner:2021:OCH

[2738] Tim Gubner, Viktor Leis, and Peter Boncz. Optimistically compressed hash tables & strings in the USSR. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):60–67, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471500.

Olteanu:2021:TPP

[2739] Dan Olteanu. Technical perspective: Probabilistic data with continuous distributions. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):68, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471501.

Grohe:2021:PDC

[2740] Martin Grohe, Benjamin Lucien Kaminski, Joost-Pieter Katoen, and Peter Lindner. Probabilistic data with continuous distributions. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):69–76, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471502.

Suciu:2021:TPQ

[2741] Dan Suciu. Technical perspective for:

Query games in databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):77, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471503.

Livshits:2021:QGD

[2742] Ester Livshits, Leopoldo Bertossi, Benny Kimelfeld, and Moshe Sebag. Query games in databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):78–85, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471504.

Ross:2021:TPS

[2743] Kenneth A. Ross. Technical perspective: Scaling dynamic hash tables on real persistent memory. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):86, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3471485.3471505.

Lu:2021:SDH

[2744] Baotong Lu, Xiangpeng Hao, Tianzheng Wang, and Eric Lo. Scaling dynamic hash tables on real persistent memory. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(1):87–94, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3471485.
3471506.

Doleschal:2021:DPC

[2745] Johannes Doleschal, Benny Kimelfeld, and Wim Martens. Database principles and challenges in text analysis. SIG-MOD Record (ACM Special Interest Group on Management of Data), 50(2):

6–17, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3484622.3484624.

Doulkeridis:2021:SBD

[2746] Christos Doulkeridis, Akrivi Vlachou, Nikos Pelekis, and Yannis Theodoridis. A survey on big data processing frameworks for mobility analytics. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(2):18–29, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3484622.3484626.

Valduriez:2021:MRM

[2747] Patrick Valduriez. Making the right move to senior researcher: Some challenges and hints. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(2):30–32, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3484622.3484628.

Athanassoulis:2021:TDT

[2748] Manos Athanassoulis. Transparent data transformation: Can I have my rows and eat my columns too? SIG-MOD Record (ACM Special Interest Group on Management of Data), 50(2): 33–34, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3484622.3484630.

Winslett:2021:SBS

[2749] Marianne Winslett and Vanessa Braganholo. Sourav Bhowmick speaks out on drawing your queries, ethics, and drawing for ethics. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(2):35–40, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3484622.3484632.

Winslett:2021:VLS

[2750] Marianne Winslett and Vanessa Braganholo. Viktor Leis speaks out on concurrency and parallelism on multicore CPUs. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(2):41–43, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3484622.3484633.

Dosso:2021:IRS

[2751] Dennis Dosso, Stefano Ferilli, Paolo Manghi, Antonella Poggi, Giuseppe Serra, and Gianmaria Silvello. Information and research science connecting to digital and library science: Report on the 17th Italian Research Conference on Digital Libraries, IRCDL2021. SIG-MOD Record (ACM Special Interest Group on Management of Data), 50(2): 44–47, June 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3484622.3484635.

Geerts:2021:MQL

[2752] Floris Geerts, Thomas Muñoz, Cristian Riveros, Jan Van den Bussche, and Domagoj Vrgoc. Matrix query languages. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(3):6–19, September 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3503780.3503782.

Jagadish:2021:WBS

[2753] H. V. Jagadish. What becomes a senior researcher. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(3):20–22, September 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3503780.3503784.

Ives:2021:VPS

[2754] Zack Ives, Johannes Gehrke, Jana Giceva, Arun Kumar, and Rachel Pottinger. VLDB panel summary: “The Future of Data(base) Education: Is the Cow Book Dead?”. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(3):23–26, September 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3503780.3503786.

Trummer:2021:DTU

[2755] Immanuel Trummer. Database tuning using natural language processing. SIGMOD Record (ACM Special Interest Group on Management of Data), 50 (3):27–28, September 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3503780.
3503788.

Winslett:2021:JAS

[2756] Marianne Winslett and Vanessa Braganholo. Joy Arulraj speaks out on non-volatile memory database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 50 (3):29–31, September 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3503780.
                                        3503790.

Kondylakis:2021:RTI

[2757] Haridimos Kondylakis, Kostas Stefanidis, and Praveen Rao. Report on the Third International Workshop on Semantic Web Meets Health Data Management (SWH 2020). SIGMOD Record (ACM Special Interest Group on Management of Data), 50(3):32–35, September 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3503780.3503792.

Cormode:2021:CTD

[2758] Graham Cormode. Current trends in data summaries. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(4):6–15, December 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3516431.3516433.

Mansour:2021:FDS

[2759] Essam Mansour, Kavitha Srinivas, and Katja Hose. Federated data science to break down silos [vision]. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(4):16–22, December 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3516431.3516435.

Amer-Yahia:2021:IBE

[2760] Sihem Amer-Yahia, Georgia Koutrika, Martin Braschler, Diego Calvanese, Davide Lanti, Hendrik Lücke-Tieke, Alessandro Mosca, Tarcisio Mendes de Farias, Dimitris Papadopoulos, Yogendra Patil, Guillem Rull, Ellery Smith, Dimitrios Skoutas, Srividya Subramanian, and Kurt Stockinger. INODE: Building an end-to-end data exploration system in practice. SIG-MOD Record (ACM Special Interest Group on Management of Data), 50 (4):23–29, December 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3516431.
3516436.

Bonifati:2021:HIW

[2761] Angela Bonifati, Michael J. Mior, Felix Naumann, and Nele Sina Noack. How inclusive are we? SIGMOD Record (ACM Special Interest Group on Management of Data), 50(4):30–35, December 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3516431.3516438.

Balazinska:2021:CYB

[2762] Magdalena Balazinska. Congratulations! you have become a senior researcher. now what? SIGMOD Record (ACM Special Interest Group on Management of Data), 50(4):36–38, December 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3516431.3516440.

Arulraj:2021:AVA

[2763] Joy Arulraj. Accelerating video analytics. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(4):39–40, December 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3516431.3516442.

Winslett:2021:JFS

[2764] Marianne Winslett and Vanessa Braganholo. Juliana Freire speaks out on reproducibility and hard changes. SIG-MOD Record (ACM Special Interest Group on Management of Data), 50 (4):41–46, December 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3516431.
                                        3516444.

Winslett:2021:HZS

[2765] Marianne Winslett and Vanessa Braganholo. Huanchen Zhang speaks out on memory-efficient search trees. SIG-MOD Record (ACM Special Interest Group on Management of Data), 50 (4):47–49, December 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3516431.
                                        3516445.

Bonnet:2021:VDH

[2766] Philippe Bonnet, Xin Luna Dong, Felix Naumann, and Pinar Tözün. VLDB 2021: Designing a hybrid conference. SIGMOD Record (ACM Special Interest Group on Management of Data), 50(4):50–53, December 2021. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3516431.3516447.

Markl:2022:MLQ

[2767] Volker Markl. Making learned query optimization practical: a technical perspective. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):5, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3542700.
3542702.

Marcus:2022:BML

[2768] Ryan Marcus, Parimarjan Negi, Hongzi Mao, Nesime Tatbul, Mohammad Alizadeh, and Tim Kraska. Bao: Making learned query optimization practical. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):6–13, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542700.3542703.

Alonso:2022:TP

[2769] Gustavo Alonso. Technical perspective.

SIGMOD Record (ACM Special Interest Group on Management of Data), 51 (1):14, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3542700.3542704.

Thostrup:2022:DDF

[2770] Lasse Thostrup, Jan Skrzypczak, Matthias Jasny, Tobias Ziegler, and Carsten Binnig. DFI: The data flow interface for high-speed networks. SIG-MOD Record (ACM Special Interest Group on Management of Data), 51(1): 15–22, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3542700.3542705.

Kemper:2022:TPF

[2771] Alfons Kemper. Technical perspective: FoundationDB: a distributed unbundled transactional key value store. SIGMOD Record (ACM Special Interest Group on Management of Data), 51 (1):23, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3542700.3542706.

Zhou:2022:FDK

[2772] Jingyu Zhou, Meng Xu, Alexander Shraer, Bala Namasivayam, Alex Miller, Evan Tschannen, Steve Atherton, Andrew J. Beamon, Rusty Sears, John Leach, Dave Rosenthal, Xin Dong, Will Wilson, Ben Collins, David Scherer, Alec Grieser, Young Liu, Alvin Moore, Bhaskar Muppana, Xiaoge Su, and Vishesh Yadav. FoundationDB: a distributed key value store. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):24–31, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542700.3542707.

Papotti:2022:TPT

[2773] Paolo Papotti. Technical perspective of TURL: Table understanding through representation learning. SIG-MOD Record (ACM Special Interest Group on Management of Data), 51(1): 32, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3542700.3542708.

Deng:2022:TTU

[2774] Xiang Deng, Huan Sun, Alyssa Lees, You Wu, and Cong Yu. TURL: Table understanding through representation learning. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):33–40, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542700.3542709.

Hogan:2022:TPN

[2775] Aidan Hogan. Technical perspective — no PANE, no gain: Scaling attributed network embedding in a single server. SIGMOD Record (ACM Special Interest Group on Management of Data), 51 (1):41, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3542700.3542710.

Yang:2022:NPN

[2776] Renchi Yang, Jieming Shi, Xiaokui Xiao, Yin Yang, Sourav S. Bhowmick, and Juncheng Liu. No PANE, no gain: Scaling attributed network embedding in a single server. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):42–49, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542700.3542711.

Mamoulis:2022:TPB

[2777] Nikos Mamoulis. Technical perspective: Bipartite matching: What to do in the real world when computing assignment costs dominates finding the optimal assignment. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):50, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542700.3542712.

Abeywickrama:2022:BMW

[2778] Tenindra Abeywickrama, Victor Liang, and Kian-Lee Tan. Bipartite matching: What to do in the real world when computing assignment costs dominates finding the optimal assignment. SIG-MOD Record (ACM Special Interest Group on Management of Data), 51(1): 51–58, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3542700.3542713.

Howe:2022:TPI

[2779] Bill Howe. Technical perspective: Imperative or functional control flow handling: Why not the best of both worlds? SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):59, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3542700.3542714.

Gevay:2022:IFC

[2780] Gábor E. Gévay, Tilmann Rabl, Sebastian Breß, Loránd Madai-Tahy, Jorge-Arnulfo Quiané-Ruiz, and Volker Markl. Imperative or functional control flow handling: Why not the best of both worlds? SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):60–67, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3542700.3542715.

Pagh:2022:TPR

[2781] Rasmus Pagh. Technical perspective: Relative error streaming quantiles.

SIGMOD Record (ACM Special Interest Group on Management of Data), 51 (1):68, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3542700.3542716.

Cormode:2022:RES

[2782] Graham Cormode, Zohar Karnin, Edo Liberty, Justin Thaler, and Pavel Vesely. Relative error streaming quantiles. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):69–76, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542700.3542717.

Yannakakis:2022:TPS

[2783] Mihalis Yannakakis. Technical perspective: Structure and complexity of bag consistency. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):77, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542700.3542718.

Atserias:2022:SCB

[2784] Albert Atserias and Phokion G. Kolaitis. Structure and complexity of bag consistency. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):78–85, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542700.3542719.

Woodruff:2022:TPM

[2785] David P. Woodruff. Technical perspective: Model counting meets distinct elements in a data stream. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):86, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542700.3542720.

Pavan:2022:MCM

[2786] A. Pavan, N. V. Vinodchandran, Arnab Bhattacharyya, and Kuldeep S. Meel. Model counting meets distinct elements in a data stream. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(1):87–94, March 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3542700.3542721.

Khamis:2022:DW

[2787] Mahmoud Abo Khamis, Hung Q. Ngo, Reinhard Pichler, Dan Suciu, and Yisu Remy Wang. Datalog in Wonderland. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):6–17, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3552490.3552492.

Dave:2022:MIR

[2788] Dev Dave, Angelica Celestino, Aparna S. Varde, and Vaibhav Anu. Management of implicit requirements data in large SRS documents: Taxonomy and techniques. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):18–29, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/10.
1145/3552490.3552494.

Psallidas:2022:DST

[2789] Fotis Psallidas, Yiwen Zhu, Bojan Karlas, Jordan Henkel, Matteo Interlandi, Subru Krishnan, Brian Kroth, Venkatesh Emani, Wentao Wu, Ce Zhang, Markus Weimer, Avrilia Floratou, Carlo Curino, and Konstantinos Karanasos. Data science through the looking glass: Analysis of millions of GitHub notebooks and ML.NET pipelines. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):30–37, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/10.
1145/3552490.3552496.

Ghosh:2022:CED

[2790] Dhrubajyoti Ghosh, Peeyush Gupta, Sharad Mehrotra, and Shantanu Sharma. A case for enrichment in data management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):38–43, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3552490.3552497.

Ozsu:2022:RIP

[2791] M. Tamer Özsu. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):44–46, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3552490.3552499.

Maier:2022:DWD

[2792] David Maier. Deciding what not to do. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):47–48, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3552490.3552501.

deWaal:2022:OAC

[2793] Andre de Waal and Ana Echeverri. OpenDS4All: Accelerating the creation of data science curricula at academic institutions. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):49–51, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3552490.3552503.

Fekete:2022:TAD

[2794] Alan D. Fekete and Uwe Röhm. Teaching about data and databases: Why, what, how? SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):52–60, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3552490.3552504.

Appuswamy:2022:TPM

[2795] Raja Appuswamy. Towards passive, migration-free, standardized, longterm database archival. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):61–62, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3552490.3552506.

Cheng:2022:STR

[2796] Reynold Cheng, Chenhao Ma, Xiaodong Li, Yixiang Fang, Ye Liu, Victor Y. L. Wong, Esther Lee, Tai Hing Lam, Sai Yin Ho, Man Ping Wang, Weijie Gong, Wentao Ning, and Ben Kao. The Social Technology and Research (STAR) Lab in the University of Hong Kong. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):63–68, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3552490.3552508.

Amer-Yahia:2022:DIA

[2797] Sihem Amer-Yahia, Yael Amsterdamer, Sourav S. Bhowmick, Angela Bonifati, Philippe Bonnet, Renata Borovica-Gajic, Barbara Catania, Tania Cerquitelli, Silvia Chiusano, Panos K. Chrysanthis, Carlo Curino, Jérôme Darmont, Amr El Abbadi, Avrilia Floratou, Juliana Freire, Alekh Jindal, Vana Kalogeraki, Georgia Koutrika, Arun Kumar, Sujaya Maiyya, Alexandra Meliou, Madhulika Mohanty, Felix Naumann, Nele Sina Noack, Fatma Özcan, Liat Peterfreund, Wenny Rahayu, Wang-Chiew Tan, Yuanyuan Tian, Pinar Tözün, Genoveva Vargas-Solar, Neeraja Yadwadkar, and Meihui Zhang. Diversity and inclusion activities in database conferences: a 2021 report. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):69–73, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/10.
1145/3552490.3552510.

Athanassoulis:2022:AAR

[2798] Manos Athanassoulis, Peter Triantafillou, Raja Appuswamy, Rajesh Bordawekar, Badrish Chandramouli, Xuntao Cheng, Ioana Manolescu, Yannis Papakonstantinou, and Nesime Tatbul. Artifacts availability & reproducibility (VLDB 2021 Round Table). SIGMOD Record (ACM Special Interest Group on Management of Data), 51(2):74–77, June 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3552490.3552511.

Arenas:2022:CAQ

[2799] Marcelo Arenas, Luis Alberto Croquevielle, Rajesh Jayaram, and Cristian Riveros. Counting the answers to a query. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(3):6–17, September 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3572751.3572753.

Azcoitia:2022:SDM

[2800] Santiago Andrés Azcoitia and Nikolaos Laoutaris. A survey of data marketplaces and their business models. SIG-MOD Record (ACM Special Interest Group on Management of Data), 51 (3):18–29, September 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3572751.
                                        3572755.

Kennedy:2022:ROD

[2801] Javen Kennedy, Pranav Subramaniam, Sainyam Galhotra, and Raul Castro Fernandez. Revisiting online data markets in 2022: a seller and buyer perspective. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(3):30–37, September 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3572751.3572757.

Ozcan:2022:RIP

[2802] Fatma Özcan and Yuanyuan Tian.

Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 51 (3):38–39, September 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3572751.
3572759.

Ailamaki:2022:FMC

[2803] Anastasia Ailamaki. The formidable mid-career crisis. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(3):40–42, September 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3572751.3572761.

Kumar:2022:DEU

[2804] Arun Kumar, Alin Deutsch, Amarnath Gupta, Yannis Papakonstantinou, Babak Salimi, and Victor Vianu. Database education at UC San Diego. SIGMOD Record (ACM Special Interest Group on Management of Data), 51 (3):43–46, September 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3572751.
3572763.

Balmau:2022:CML

[2805] Oana Balmau. Characterizing I/O in machine learning with MLPerf storage. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(3):47–48, September 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3572751.3572765.

Jindal:2022:QOS

[2806] Alekh Jindal and Jyoti Leeka. Query optimizer as a service: an idea whose time has come! SIGMOD Record (ACM Special Interest Group on Management of Data), 51(3):49–55, September 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3572751.3572767.

Kumar:2022:VSD

[2807] Arun Kumar, Alon Halevy, and Nesime Tatbul. VLDB scalable data science category: The inaugural year. SIG-MOD Record (ACM Special Interest Group on Management of Data), 51 (3):56–58, September 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3572751.
3572769.

Ketsman:2022:CCD

[2808] Bas Ketsman, Christoph Koch, Frank Neven, and Brecht Vandevoort. Concurrency control for database theorists. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(4):6–17, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3582302.
3582304.

Schlegel:2022:MML

[2809] Marius Schlegel and Kai-Uwe Sattler. Management of machine learning lifecycle artifacts: a survey. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(4):18–35, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3582302.3582306.

Benedikt:2022:PFI

[2810] Michael Benedikt, Fergus Cooper, Stefano Germano, Gabor Gyorkei, Efthymia Tsamoura, Brandon Moore, and Camilo Ortiz. PDQ 2.0: Flexible infrastructure for integrating reasoning and query planning. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(4):36–41, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3582302.3582308.

Rabl:2022:RIP

[2811] Tilmann Rabl. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(4):42–44, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3582302.3582310.

Lakshmanan:2022:MCR

[2812] Laks V. S. Lakshmanan. Midcareer researcher, huh?: What just changed? SIGMOD Record (ACM Special Interest Group on Management of Data), 51(4):45–48, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3582302.3582312.

Aivaloglou:2022:RFI

[2813] Efthimia Aivaloglou, George Fletcher, Michael Liut, and Daphne Miedema. Report on the First International Workshop on Data Systems Education (DataEd ’22). SIGMOD Record (ACM Special Interest Group on Management of Data), 51(4):49–53, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3582302.3582314.

Pirk:2022:CDS

[2814] Holger Pirk. Collaborative data science using scalable homoiconicity. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(4):54–55, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3582302.3582316.

Wu:2022:CWS

[2815] Chenggang Wu. Chenggang Wu speaks out on his ACM SIGMOD Jim Gray Dissertation Award, rejection, believing in your work, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(4):56–59, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3582302.3582318.

Tian:2022:WGD

[2816] Yuanyuan Tian. The world of graph databases from an industry perspective. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(4):60–67, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3582302.3582320.

Perscheid:2022:EPI

[2817] Michael Perscheid, Hasso Plattner, Daniel Ritter, Rainer Schlosser, and Ralf Teusner. Enterprise platform and integration concepts research at HPI. SIGMOD Record (ACM Special Interest Group on Management of Data), 51(4):68–73, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3582302.
3582322.

Bhowmick:2022:HCO

[2818] Sourav S. Bhowmick. How connected are our conference review boards? SIG-MOD Record (ACM Special Interest Group on Management of Data), 51 (4):74–78, December 2022. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3582302.
3582324.

Salem:2023:TPA

[2819] Kenneth Salem. Technical Perspective: Ad hoc transactions: What they are and why we should care. SIG-MOD Record (ACM Special Interest Group on Management of Data), 52 (1):6, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3604437.3604439.

Tang:2023:AHT

[2820] Chuzhe Tang, Zhaoguo Wang, Xiaodong Zhang, Qianmian Yu, Binyu Zang, Haibing Guan, and Haibo Chen. Ad hoc transactions: What they are and why we should care. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):7–15, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3604437.3604440.

Bonifati:2023:TPS

[2821] Angela Bonifati. Technical perspective: Sortledton: a universal graph data structure. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):16, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3604437.3604441.

Fuchs:2023:SUG

[2822] Per Fuchs, Domagoj Margan, and Jana Giceva. Sortledton: a universal graph data structure. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):17–25, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3604437.3604442.

Binnig:2023:TPS

[2823] Carsten Binnig. Technical perspective for skeena: Efficient and consistent cross-engine transactions. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):26, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3604437.3604443.

Zhang:2023:EMC

[2824] Jianqiu Zhang, Kaisong Huang, Tianzheng Wang, and King Lv. Efficiently making cross-engine transactions consistent. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):27–34, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/10.
1145/3604437.3604444.

Fekete:2023:TPW

[2825] Alan D. Fekete. Technical perspective: When is it safe to run a transactional workload under read committed? SIGMOD Record (ACM Special Interest Group on Management of Data), 52 (1):35, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3604437.3604445.

Vandevoort:2023:WIS

[2826] Brecht Vandevoort, Bas Ketsman, Christoph Koch, and Frank Neven. When is it safe to run a transactional workload under Read Committed? SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):36–43, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604446.

Kraska:2023:TPS

[2827] Tim Kraska. Technical perspective for Sherman: a write-optimized distributed B+Tree index on disaggregated memory. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):44, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604447.

Wang:2023:BWO

[2828] Qing Wang, Youyou Lu, and Jiwu Shu. Building write-optimized tree indexes on disaggregated memory. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):45–52, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3604437.3604448.

Vansummeren:2023:TPC

[2829] Stijn Vansummeren. Technical perspective: Conjunctive queries with comparisons. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):53, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL

                                        https://dl.acm.org/doi/10.1145/
                                        3604437.3604449.

Wang:2023:CQC

[2830] Qichen Wang and Ke Yi. Conjunctive queries with comparisons. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):54–62, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3604437.3604450.

Libkin:2023:TPQ

[2831] Leonid Libkin. Technical perspective: Query answers — fewer is faster. SIG-MOD Record (ACM Special Interest Group on Management of Data), 52(1): 63, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3604437.3604451.

Bonifati:2023:TQ

[2832] Angela Bonifati, Stefania Dumbrava, George Fletcher, Jan Hidders, Matthias Hofer, Wim Martens, Filip Murlak, Joshua Shinavier, Slawek Staworko, and Dominik Tomaszuk. Threshold queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):64–73, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/10.
1145/3604437.3604452.

Rudra:2023:TPP

[2833] Atri Rudra. Technical perspective:

(pre-) semirings come to the recursion party. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):74, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604453.

Khamis:2023:CDP

[2834] Mahmoud Abo Khamis, Hung Q. Ngo, Reinhard Pichler, Dan Suciu, and Yisu Remy Wang. Convergence of Datalog over (pre-) semirings. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):75–82, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604454.

Jayaram:2023:TPO

[2835] Rajesh Jayaram. Technical perspective: Optimal algorithms for multiway search on partial orders. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):83, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604455.

Lu:2023:OAP

[2836] Shangqi Lu, Wim Martens, Matthias Niewerth, and Yufei Tao. An optimal algorithm for partial order multiway search. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):84–92, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604456.

Suciu:2023:TPA

[2837] Dan Suciu. Technical perspective: Accurate summary-based cardinality estimation through the lens of cardinality estimation graphs. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):93, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604457.

Chen:2023:ASB

[2838] Jeremy Chen, Yuqing Huang, Mushi Wang, Semih Salihoglu, and Kenneth Salem. Accurate summarybased cardinality estimation through the lens of cardinality estimation graphs. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):94–102, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604458.

Kipf:2023:TPR

[2839] Andreas Kipf. Technical perspective:

Revisiting runtime dynamic optimization for join queries in big data management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):103, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604459.

Pavlopoulou:2023:RRD

[2840] Christina Pavlopoulou, Michael J.

Carey, and Vassilis J. Tsotras. Revisiting runtime dynamic optimization for join queries in big data management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):104–113, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604460.

Cormode:2023:TPR

[2841] Graham Cormode. Technical perspective on ’R2T: Instance-optimal truncation for differentially private query evaluation with foreign keys. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):114, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3604437.3604461.

Dong:2023:RIO

[2842] Wei Dong, Juanru Fang, Ke Yi, Yuchao Tao, and Ashwin Machanavajjhala. R2T: Instance-optimal truncation for differentially private query evaluation with foreign keys. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(1):115– 123, March 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3604437.3604462.

Bertossi:2023:SVD

[2843] Leopoldo Bertossi, Benny Kimelfeld, Ester Livshits, and Mikaël Monet. The Shapley value in database management. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(2):6–17, June 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3615952.
                                        3615954.

Khan:2023:KGQ

[2844] Arijit Khan. Knowledge graphs querying. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(2):18–29, June 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3615952.3615956.

Aboulnaga:2023:RIP

[2845] Ashraf Aboulnaga. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(2):30–31, June 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3615952.3615958.

Luo:2023:MCA

[2846] Qiong Luo. Mid-career academics: What I have learned or wish I had known. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(2):32–35, June 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3615952.3615960.

Crooks:2023:EDS

[2847] Natacha Crooks. Efficient data sharing across trust domains. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(2):36–37, June 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3615952.3615962.

Amer-Yahia:2023:DEI

[2848] Sihem Amer-Yahia, Divyakant Agrawal, Yael Amsterdamer, Sourav S. Bhowmick, Angela Bonifati, Renata Borovica-Gajic, Jesús Camacho-Rodríguez, Barbara Catania, Panos K. Chrysanthis, Carlo Curino, Jérôme Darmont, Gillian Dobbie, Amr El Abbadi, Avrilia Floratou, Juliana Freire, Alekh Jindal, Vana Kalogeraki, Sujaya Maiyya, Alexandra Meliou, Madhulika Mohanty, Behrooz Omidvar-Tehrani, Fatma Özcan, Liat Peterfreund, Wenny Rahayu, Shazia Sadiq, Sana Sellami, Utku Sirin, Wang-Chiew Tan, Bhavani Thuraisingham, Yuanyuan Tian, Pinar Tözün, Genoveva Vargas-Solar, Neeraja Yadwadkar, Victor Zakhary, and Meihui Zhang. Diversity, equity and inclusion activities in database conferences: a 2022 report. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(2):38–42, June 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

URL https://dl.acm.org/doi/10.
1145/3615952.3615964.

DeAngelis:2023:ELL

[2849] Andrea De Angelis, Maurizio Mazzei, Federico Piai, Paolo Merialdo, Giovanni Simonini, Luca Zecchini, Sonia Bergamaschi, Donatella Firmani, Xu Chu, Peng Li, and Renzhi Wu. Experiences and lessons learned from the SIGMOD Entity Resolution Programming Contests. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(2):43–47, June 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3615952.3615965.

Mildenberger:2023:RDS

[2850] Thoralf Mildenberger, Martin Braschler, Andreas Ruckstuhl, Robert Vorburger, and Kurt Stockinger. The role of data scientists in modern enterprises — experience from data science education. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(2):48–52, June 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3615952.
                                        3615966.

Olteanu:2023:RWF

[2851] Dan Olteanu. Report on the workshop on factorized databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(2):53–56, June 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3615952.3615967.

Dong:2023:QEU

[2852] Wei Dong and Ke Yi. Query evaluation under differential privacy. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(3):6–17, September 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3631504.3631506.

Seufitelli:2023:WDD

[2853] Danilo B. Seufitelli, Michele A.

Brandão, Ayane C. A. Fernandes, Kayque M. Siqueira, and Mirella M. Moro. Where do databases and digital forensics meet? A comprehensive survey and taxonomy. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(3):18–29, September 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

org/doi/10.1145/3631504.3631508.

Beedkar:2023:AWU

[2854] Kaustubh Beedkar, Bertty Contreras-Rojas, Haralampos Gavriilidis, Zoi Kaoudi, Volker Markl, Rodrigo Pardo-Meza, and Jorge-Arnulfo Quiané-Ruiz. Apache Wayang: a unified data analytics framework. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(3):30–35, September 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3631504.3631510.

Borovica-Gajic:2023:RIP

[2855] Renata Borovica-Gajic. Reminiscences on influential papers. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(3):36–38, September 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3631504.3631512.

Salihoglu:2023:KDM

[2856] Semih Salihoglu. Kùzu: a database management system for “Beyond Relational” workloads. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(3):39–40, September 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3631504.3631514.

Poppe:2023:PRA

[2857] Olga Poppe, Pablo Castro, Willis Lang, and Jyoti Leeka. Proactive resource allocation policy for Microsoft Azure cognitive search. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(3):41–48, September 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3631504.3631516.

Amer-Yahia:2023:LLM

[2858] Sihem Amer-Yahia, Angela Bonifati, Lei Chen, Guoliang Li, Kyuseok Shim, Jianliang Xu, and Xiaochun Yang. From large language models to databases and back: a discussion on research and education. SIGMOD Record (ACM Special Interest Group on Management of Data), 52(3):49–56, September 2023. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3631504.3631518.

ACM:1984:PTA

[2859] ACM, editor. Proceedings of the Third ACM SIGACT-SIGMOD Symposium on Principles of Database Systems: April 2–4, 1984, Waterloo, Ontario, Canada, ACM SIGMOD Record. ACM Press, New York, NY 10036, USA, 1984. CODEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 A15 1984. ACM order no. 475840.

ACM:1984:SPA

[2860] ACM, editor. SIGMOD ’84: proceedings of annual meeting, Boston, MA, June 18–21, 1984, volume 14(2) of ACM SIGMOD Record. ACM Press, New York, NY 10036, USA, 1984. CO-DEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA76.9.D3 I59 1984.

Yormark:1984:SPA

[2861] Beatrice Yormark, editor. SIGMOD ’84, Proceedings of Annual Meeting, volume 14(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 1984. CO-DEN SRECD8. ISBN 0-89791-128-8. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ????

Anonymous:1988:SIC

[2862] Anonymous, editor. SIGMOD International Conference on Management of Data, volume 17(3) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, September 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1989:ASI

[2863] Anonymous, editor. 1989 ACM SIGMOD International Conference on Management of Data, volume 18(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, June 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1989:WFD

[2864] Anonymous, editor. Workshop on Future Directions in DBMS Research, volume 18(1) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, March 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Garcia-Molina:1990:PAS

[2865] Héctor García-Molina and H. V. Jagadish, editors. Proceedings of the 1990 ACM SIGMOD International Conference on Management of Data, volume 19(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1990:ASI

[2866] Anonymous, editor. 1990 ACM SIGMOD International Conference on Management of Data, volume 19(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1991:ASI

[2867] Anonymous, editor. 1991 ACM SIGMOD International Conference on Management of Data, volume 20(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1993:SAS

[2868] Anonymous, editor. SIGMOD ’93.

1993 ACM SIGMOD. International Conference on Management of Data, volume 22(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, June 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Buneman:1993:PAS

[2869] Peter Buneman and Sushil Jajodia, editors. Proceedings of the 1993 ACM SIGMOD International Conference on Management of Data: Washington, DC, May 26–28, 1993, volume 22(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ????

Anonymous:1994:ASI

[2870] Anonymous, editor. 1994 ACM SIGMOD International Conference on Management of Data, volume 23(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Anonymous:1995:ASI

[2871] Anonymous, editor. 1995 ACM SIGMOD International Conference on Management of Data, volume 24(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, June 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).